### v1

# Learning in Structured Stackelberg Games

Abstract
We study structured Stackelberg games, in which both players (the leader and the follower) observe information about the state of the world at time of play. Importantly, this information may contain information about the follower, which the leader may use when deciding her strategy. Under this setting, we show that no-regret learning is possible if and only if the set of mappings from contexts to follower types that the leader uses to learn is not “too complex”. Specifically, we find that standard learning theoretic measures of complexity do not characterize learnability in our setting and we give a new dimension which does, which we term the Stackelberg-Littlestone dimension. In the distributional setting, we give analogous results by showing that standard complexity measures do not characterize the sample complexity of learning, but a new dimension called the Stackelberg-Natarajan dimension does. We then show that an appropriate empirical risk minimization procedure achieves the corresponding sample complexity.
1 Introduction
Many strategic interactions between a principal and an agent can be cast as Stackelberg games [36]. In a Stackelberg game, the principal (also called the leader) commits to a possibly randomized strategy, before the agent (also called the follower) observes this strategy and best-responds to it. Due to their effectiveness in modeling such leader-follower interactions, Stackelberg games have had significant real-world impact over the past decade, particularly in security domains [35]. For example, in protecting endangered species [38], a wildlife agency determines ranger patrol locations before poachers select targets. Coast guards decide on shoreline patrol routes in advance of illegal fishers’ actions [12] and, similarly, airport security [32] chooses checkpoint screening strategies before smugglers decide which entry points to use.
In many of these real-world settings, the state of the world at time of play can greatly influence the utilities of the players and subsequently the outcome of the game. For example, defending some areas of the airport might be deemed of higher- or lower-importance based on the schedule of flights and the number of passengers on board. Similarly, weather patterns might make it harder to reach animals in certain parts of the park, while breeding seasons might make others more vulnerable thereby affecting their risk levels as targets. When contextual information becomes available over time, using it could help the leader design more effective strategies against the follower. This raises a natural question: given that the state of the world might change in current and future rounds of play, how should the leader in a Stackelberg game design a policy that adapts to new contexts coming in over time?
One way is to study repeated Stackelberg games with side information [24], where at each interaction both the leader and the follower get to observe a context that affects their utilities. As in Balcan et al. [5], in this setting each follower that the leader faces might have different preferences and objectives, i.e. there might be different follower types. Harris et al. [24] study this problem and show that as long as either the followers or the contexts are independently drawn from a fixed and unknown distribution, no-Stackelberg-regret learning is possible. That is, the leader is guaranteed to achieve nearly as much utility as the best fixed policy in hindsight. In the fully adversarial regime, however, regret can grow linearly with the time horizon. Interestingly, this lower bound follows from the adversary’s ability to encode an arbitrarily complex relationship between the sequence of contexts and the follower types.
In this work, we take a more granular view of this impossibility result and ask whether introducing additional structure between the contextual information and the follower types can enhance the leader’s ability to maximize returns, even when both are chosen adversarially to the leader. Namely, there could be correlation between the two, making the context predictive of the type of follower about to appear next. In wildlife protection for example, specific tire tracks might signal the presence of a particular poaching group and market demand trends might be indicative of which animals’ poachers are now more likely to appear. Similarly, in airport security, current affairs often raise or lower the risk of an attack on certain international flights.
Algorithms proposed by Harris et al. [24] and later work [9] are designed to handle changes in utility caused by the context, but crucially they assume no structure in how contexts relate to follower types. If, however, there is a fixed and unknown mapping between contexts and follower types, can the leader learn to exploit it to achieve sublinear regret? In this case, the leader m ust adapt her policy based on which follower types are more likely to appear, given the context and past observations. This poses an additional and possibly more complex learning task for the leader, which we study in this work.
1.1 Our Contributions
In this paper, we revisit the framework of Harris et al. [24] and take a learning theoretic approach to analyzing the leader’s task in predicting the follower type from contexts. Importantly, we find that existing notions of complexity do not capture the difficulty of learning in this setting—we obtain (tight) sample complexity and Stackelberg regret guarantees that are parameterized by new notions of complexity and that directly exploit the structure of Stackelberg games. Concretely, our main contributions in this work are as follows:
  1. 1.
    We formalize online learning in Stackelberg games with side information, under the structural assumption that there exists a fixed, unknown mapping between contexts and follower types. In particular, we study how the leader, given access to a set of possible such mappings, can leverage it to learn an optimal policy. In particular, we study this problem under two different settings: (1) the realizable setting, in which there exists an optimal predictor of the followers’ types that the leader competes with; and (2) the agnostic setting, in which no such assumption is made.
  2. 2.
    We circumvent the impossibility result shown in Harris et al. [24] by providing algorithms with no-regret guarantees under the above structural assumption. We find that our problem is closely related to the problem of multiclass prediction, yet, interestingly, we show that our setting can be easier to learn in. In particular, we show that the (multiclass) Littlestone dimension, which quantifies the difficulty of learning in online multiclass prediction, can give loose mistake bounds in our setting. Motivated by this observation, we introduce a new notion of complexity, the Stackelberg-Littlestone dimension. We show that it characterizes learnability in our setting, and prove it to be at most as large as the multiclass Littlestone dimension. In fact, we give an example in which the latter is infinite, yet the former is zero. We also provide no-regret algorithms both for both the realizable and agnostic settings.
  3. 3.
    We introduce and formalize the study of distributional learning in Stackelberg games with side information. When the leader has access to a set of (context, follower type) examples, we show generalization with PAC-style bounds. Here too, we find that existing measures which characterize the complexity of multiclass prediction, such as the Natarajan dimension, do not characterize the sample complexity of our problem. Analogous to the online setting, we introduce the notion of Stackelberg-Natarajan dimension, with which we achieve an upper bound that is tight up to logarithmic factors.
1.2 Related Work
Stackelberg games. In the one-shot version of a Stackelberg game, Conitzer and Sandholm [16] prove that the problem of computing a Stackelberg equilibrium is NP-hard and they give an algorithm for maximizing the leader’s payoff. Other works focus on relaxing the assumption that all of the game’s parameters are known (e.g. utilities, number of actions) and aim to learn them instead [2, 31].
Our work is closely related to Bayesian Stackelberg games, in which the leader is playing against a single unknown follower type that is drawn from a given set of follower types [16]. This has been studied both in one-shot games, in which the leader must find a single optimal-mixed strategy [16, 30, 37] and in the repeated setting, in which the leader faces the same follower type sequentially and must learn from previous interactions [28, 26].
Learning in Stackelberg games. The problem of learning an optimal strategy for the leader with a single unknown follower type has been studied by Letchford et al. [26] and Blum et al. [11]. These results have been extended by Peng et al. [31], who give faster rates and almost-matching lower bounds. Recently, Bacchiocchi et al. [2] relax some structural assumptions from previous work, including tie-breaking rules for the follower, and revise worst-case sample complexity results of learning the leader’s optimal mixed strategy. Recent research has explored many domains through the framework of Stackelberg games, including strategic classification, where classifiers account for individuals who strategically modify their features [23]; Bayesian persuasion, which models how a leader strategically reveals information to influence a follower’s decisions [25]; interactions with calibrated followers, who update their strategies based on past observations rather than making immediate best responses [21]; and settings with non-myopic followers, who optimize their actions over multiple time steps rather than reacting myopically [20, 15].
Part of our work focuses on learning in online Stackelberg games. Online Stackelberg games with multiple follower types were introduced in Balcan et al. [5]. The authors show that when a series of unknown attacker types is chosen adversarially, playing the well-known Hedge algorithm  [19] over a special set of mixed strategies yields a no-regret algorithm.
Our departing point is the work of Harris et al. [24], in which the authors study the problem of online learning in Stackelberg games with side information. The authors show that when either the sequence of follower types or the sequence of contexts are drawn independently from some fixed, unknown distribution, no-regret learning is indeed possible. However, when both of them are chosen adversarially, regret can scale linearly in the worst case, as the problem can be reduced from online learning linear thresholds. On the other hand, we make a structural assumption about the relationship between contexts and follower types: namely that it belongs to a class of functions that the leader has access to. With this, we show that learning is possible, and our guarantees depend on an appropriate measure of complexity of this function class.
Multiclass learnability. Our work can also be viewed through the lens of multiclass learnability, as it suffices for the leader to learn the hypothesis class of the function mapping the contexts to a (finite) set of follower types. Then the leader can compute the optimal strategy against the predicted follower. Multiclass prediction has been studied extensively both in the ditributional ([29, 17, 13]) and in the online setting ([17, 22]). Indeed we adopt several tools from online multiclass prediction in our work. Interestingly, however, we show that learning in Stackelberg games with contexts is at most as hard as multiclass prediction, and in some cases, can be much easier. For example, in the online setting, we give an example in which the leader can learn to play optimally, under a hypothesis class that has infinite Littlestone dimension.
2 Problem Setting
Notation.
We use $[N]:=\{1,2,\ldots,N\}$ to denote the set of natural numbers up to and including $N$. We use $\Delta(\pazocal{A})$ to denote the probability simplex over a (finite) set $\pazocal{A}$.
Stackelberg games with side information.
Consider a game between a leader and a follower. Before the game starts, some contextual information $\mathbf{z}\in\pazocal{Z}\subseteq\mathbb{R}^{d}$ is revealed to both players. The leader plays first and commits to a (possibly mixed) strategy $\mathbf{x}\in\Delta(\pazocal{A})$, where $\pazocal{A}$ is the set of finite actions available to the leader. The follower best responds to the leader’s strategy, $\mathbf{x}$, and the context, $\mathbf{z}$, by playing some action $a_{f}\in\pazocal{A}_{f}$, where $\pazocal{A}_{f}$ is the finite set of follower actions. [^2] Note that without loss of generality the follower’s best response will always be a pure strategy. We define the follower’s best response as:
$b_{f}(\mathbf{z},\mathbf{x})\in\arg\max_{a_{f}\in\pazocal{A}_{f}}\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u_{f}(\mathbf{z},a_{l},a_{f}).$ (1)
Here, $u_{f}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$ is the utility function of follower $f$. We assume that the follower breaks ties in favor of the leader.
We allow the follower to be one of $K\in\mathbb{N}_{\geq 1}$ follower types in the set $\mathscr{K}=\{f^{(1)},\dots,f^{(K)}\}$. The type of follower characterizes their utility function, $u_{f^{(i)}}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$. We assume that the leader knows the set of possible follower types a priori, but not the realized type. The leader aims to maximize her own utility, $u:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$. By slight abuse of notation, we sometimes use the shorthand
$u(\mathbf{z},\mathbf{x},a_{f})=\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u(\mathbf{z},a_{l},a_{f}).$
We denote a Stackelberg game with side information as a tuple, $\pazocal{G}=(\pazocal{A},\pazocal{A}_{f},\pazocal{Z},u,u_{f^{(1)}},\ldots,u_{f^{(K)}})$, consisting of the leader’s action set, the follower’s action set, the set of possible contexts, as well as the leader’s and followers’ utility functions.
Remark 2.1.
For the remainder of the paper and without loss of generality, we will consider $\mathscr{K}$ to be a set of distinct follower types. We call any two follower types distinct, if the optimal leader’s strategy against them is not the same for every context.
3 Online Learning in Structured Stackelberg Games
Consider an online Stackelberg game in which the leader is faced by a sequence of followers. In each round $t\in[T]$, a context $\mathbf{z}_{t}\in\pazocal{Z}$ is revealed to both players. A follower of type $f_{t}$ arrives, the leader commits to a mixed strategy $\mathbf{x}_{t}\in\Delta(\pazocal{A})$, and the follower best responds to it and the context according to Equation 1. As before, we assume that the leader knows the utilities of all follower types, but the type of follower at time $t$ is not revealed to the leader until after the round is over. In this section, we make no assumptions about either the contexts or the follower types that appear at each timestep. In fact, we handle the case, where they both could be chosen jointly by an adversary.
3.1 Leader’s task in the realizable setting
Consider a set $\pazocal{H}$ (also called a hypothesis class) of candidate functions mapping contexts to follower types, which the leader uses to learn. More formally, for each $h\in\pazocal{H}$, we let $h:\pazocal{Z}\rightarrow\mathscr{K}$. We first consider the realizable setting, where there exists a true mapping $h^{*}\in\pazocal{H}$ such that $h^{*}(\mathbf{z}_{t})=f_{t}$ for all possible $(\mathbf{z}_{t},f_{t})$ pairs. We relax this assumption in Section 3.2. We use $h^{*}$ as a benchmark to measure the leader’s performance over time, and we quantify this via the notion of contextual Stackelberg regret.
Definition 3.1 (Contextual Stackelberg Regret).
Given a sequence of followers $f_{1},\ldots,f_{T}$ and contexts $\mathbf{z}_{1},\dots\mathbf{z}_{T}$, the leader’s contextual Stackelberg regret is given by:
$R(T):=\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t})).$
We let $\pi_{h^{*}}(\mathbf{z})=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h^{*}(\mathbf{z})}(\mathbf{z},\mathbf{x}))$ and we assume that there is a fixed and arbitrary tie-breaking rule over the leader strategies. We let $\mathbf{x}_{t}$ for $t\in[T]$ be the mixed strategy played by the leader at time $t$.
The goal of the leader is to learn a policy $\pi:\pazocal{Z}\to\Delta(\pazocal{A})$ that minimizes her regret, that is, it achieves utility comparable to the utility she would get if she could (1) optimally predict the follower type given the context and (2) play an optimal strategy against that follower type. In the remainder of this section, we will present an algorithm that meets that benchmark. In general, if the regret of an algorithm is sublinear in the time horizon $T$, i.e., $R(T)=o(T)$, we say that the algorithm is a no-regret algorithm.
Remark 3.2.
An important part of our analysis is that the leader be able to compute the optimal strategy for a fixed context and follower type. However, due to the tie-breaking rules of the follower, the leader might not be able to get an exact solution. Previous work [5, 24] shows that the leader can get an arbitrarily small error in the utility by optimizing over an appropriately selected subset of the simplex. For ease of notation, in the remainder of this paper, we will assume that the leader can directly optimize over the entire simplex, without that affecting her regret guarantees. For completeness, we formally state this result in Appendix A.1.
Background on Online Multiclass Learnability.
Our results build off of tools and techniques from multiclass online learning [17]. In this section, we highlight the concepts and definitions from this area that are relevant to our work. We start with the notion of shattering, which captures the following intuition: the more “complex” a set of functions is, the more complex patterns it should be able to fit.
Definition 3.3 ((Multiclass) $\pazocal{H}$-Shattered Littlestone Tree).
Let a Littlestone tree, $T$, be a tree whose internal nodes are instances in $\pazocal{Z}$ and where for each node the node-to-child edges correspond to different labels in a finite set $\mathscr{K}$. We say that $T$ is shattered by a hypothesis class $\pazocal{H}$ if for every root-to-leaf path in the tree that traverses the nodes $\mathbf{z}_{1},\dots,\mathbf{z}_{n}$, there exists a function $h\in\pazocal{H}$ such that the label of every edge $(\mathbf{z}_{i},\mathbf{z}_{i+1})$ is $h(\mathbf{z}_{i})$.
The notion of multiclass Littlestone dimension reflects how the ability of a function class to shatter deeper trees corresponds to greater complexity in the class. [^3] Note that for binary labels, the definition of Littlestone dimension recovers the definition of the well-studied VC-dimension in the case where all nodes on the same depth of a shattered tree are the same. This definition is due to Daniely et al. [17], and corresponds to a natural generalization of the classic definition of the Littlestone dimension for binary classification [27] to the multiclass setting.
Definition 3.4 ((Multiclass) Littlestone Dimension).
The Littlestone dimension of a multiclass hypothesis class $\pazocal{H}$, denoted as $\textnormal{Ldim}(\pazocal{H})$, is the maximal integer $d$, such that there exists a full binary tree of depth $d$ that is shattered by $\pazocal{H}$.
Examples.
The finiteness of the Littlestone dimension of a hypothesis class is necessary for the class to be online learnable [27]. There are several examples of hypothesis classes, which have finite Littlestone dimension and can, therefore, be (online) learned. These include the class of conjunctions, the class of decision lists, and the class of linear thresholds with a margin (see Proposition 17 in [1]).
In Daniely et al. [17], the authors show that the multiclass Littlestone dimension of a hypothesis class characterizes the learnability of the class. In our setting, learning how to optimally predict the follower’s type given a context is an instance of multiclass prediction and is indeed sufficient for learning an optimal policy. Namely, given the context and the predicted follower type, the leader can deterministically find a strategy that maximizes her utility, by solving a sequence of $|\pazocal{A}_{f}|$ linear programs [16]. However, it is unclear a priori whether learning to optimally predict the follower’s type is also necessary. In Theorem 3.5 we show that it is not.
Theorem 3.5.
There exists a Stackelberg game, $\pazocal{G}$, and a hypothesis class, $\pazocal{H}$, such that $\textnormal{Ldim}(\pazocal{H})=\infty$, but the leader’s optimal policy can be determined without learning.
Proof.
Consider the setting where the context space is $\pazocal{Z}=\{1,2\}\times[0,1]$ and we have 4 follower types $\{f^{(1)},f^{(2)},f^{(3)},f^{(4)}\}$. Follower types $f^{(1)}$ and $f^{(2)}$ have fixed utility functions $u_{f^{(1)}},u_{f^{(2)}}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$ such that for all $\mathbf{z}\in\pazocal{Z}$, we have that $\pi_{*}^{(1)}(\mathbf{z})\neq\pi_{*}^{(2)}(\mathbf{z})$, where $\pi_{*}^{(i)}$, refers to the optimal policy against follower $f^{(i)}$. Let follower type $f^{(3)}$ have utility function $u_{f^{(3)}}(\mathbf{z},a_{l},a_{f})=\mathbbm{1}\{\mathbf{z}[1]=1\}\cdot u_{f^{(1)}}(\mathbf{z},a_{l},a_{f})+\mathbbm{1}\{\mathbf{z}[1]=2\}\cdot u_{f^{(2)}}(\mathbf{z},a_{l},a_{f})$ and follower type $f^{(4)}$ have utility function $u_{f^{(4)}}(\mathbf{z},a_{l},a_{f})=\mathbbm{1}\{\mathbf{z}[1]=2\}\cdot u_{f^{(1)}}(\mathbf{z},a_{l},a_{f})+\mathbbm{1}\{\mathbf{z}[1]=1\}\cdot u_{f^{(2)}}(\mathbf{z},a_{l},a_{f})$. Let
$h_{\gamma_{1},\gamma_{2}}(\mathbf{z})=\begin{cases}f^{(1)}\;\;&\text{ if }0\leq\mathbf{z}[2]<\gamma_{1}\\ f^{(2)}\;\;&\text{ if }\gamma_{2}<\mathbf{z}[2]\leq 1\\ f^{(3)}\;\;&\text{ if }(\mathbf{z}[1]=1\text{ and }\gamma_{1}\leq\mathbf{z}[2]<\frac{1}{2})\text{ or }(\mathbf{z}[1]=2\text{ and }\gamma_{2}\leq\mathbf{z}[2]\leq 1/2)\\ f^{(4)}\;\;&\text{ if }(\mathbf{z}[1]=2\text{ and }\gamma_{1}\leq\mathbf{z}[2]<\frac{1}{2})\text{ or }(\mathbf{z}[1]=1\text{ and }\gamma_{2}\leq\mathbf{z}[2]\leq 1/2)\\ \end{cases}$
[…]
Figure 1: Follower type as a function of the context $\mathbf{z}$ in the example of Theorem 3.5. Observe that while $\gamma_{1}$ and $\gamma_{2}$ vary with each hypothesis, the optimal mixed strategy to play is determined by whether $\mathbf{z}[2]\leq\frac{1}{2}$.
Theorem 3.5 shows that the Littlestone dimension does not characterize the complexity of online learning in Stackelberg games with side information. The example constructed in our proof highlights that the Littlestone dimension is blind to how the hypothesis class interacts with the utilities of the players. This observation motivates the definition of a new quantity, which we term the Stakelberg-Littlestone (SL) dimension. In order to define the SL dimension, we first need to introduce the notion of an SL-Tree, which is a natural extension of the Littlestone tree to our setting.
Consider a tree of depth $d$ whose internal nodes are instances in $\pazocal{Z}$ and where the node-to-child edges for a single node, correspond to different labels in a finite set $\mathscr{K}$. The tree $T$ is shattered by a hypothesis class $\pazocal{H}$ if for every root-to-leaf path in the tree that traverses the nodes $\mathbf{z}_{1},\dots,\mathbf{z}_{n}$, there exists a function $h\in\pazocal{H}$ such that the label of every edge $(\mathbf{z}_{i},\mathbf{z}_{i+1})$ is $h(\mathbf{z}_{i})$.
We say that $T$ is shattered by a hypothesis class $\pazocal{H}$ if for every root-to-leaf path in the tree that traverses the nodes $\mathbf{z}_{1},\dots,\mathbf{z}_{n}$, there exists a function $h\in\pazocal{H}$ such that the label of every edge $(\mathbf{z}_{i},\mathbf{z}_{i+1})$ is $h(\mathbf{z}_{i})$.
Definition 3.6 ($\pazocal{H}$-Shattered Stackelberg-Littlestone Tree).
For a Stackelberg game, $\pazocal{G}$, let a Stackelberg-Littlestone tree, $T$, be a tree whose internal nodes are instances in $\pazocal{Z}$ and where for each node, the node-to-child edges correspond to different vectors in $\Delta(\pazocal{A})$. We say that $T$ is shattered by a hypothesis class $\pazocal{H}$ with respect to $\pazocal{G}$ if for every root-to-leaf path in the tree that traverses the nodes $\mathbf{z}_{1},\ldots,\mathbf{z}_{n}$, there exists a hypothesis $h\in\mathscr{H}$, such that the label of every edge $(\mathbf{z}_{i},\mathbf{z}_{i+1})$ is $\mathbf{x}_{i}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{i},\mathbf{x},b_{h(\mathbf{z}_{i})}(\mathbf{z}_{i},\mathbf{x}))$. In the case of ties, we assume that the leader follows a fixed and arbitrary tie-breaking rule.
Equipped with these definitions, we are now ready to introduce the Stackelberg-Littlestone dimension, which characterizes the complexity of online learning in our setting.
Definition 3.7 (Stackelberg-Littlsestone Dimension).
The Stackelberg-Littlestone dimension of a hypothesis class $\mathscr{H}$ with respect to a Stackelberg game, $\pazocal{G}$, denoted as $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$, is the maximal integer $d$, such that there exists a full binary SL tree of depth $d$ that is shattered by $\pazocal{H}$.
3.1.1 Lower bound
We begin by showing that for any online Stackelberg game $\mathscr{G}$, no algorithm can achieve regret smaller than $\Omega(\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H}))$.
Theorem 3.8.
There exists a Stackelberg game, $\pazocal{G}$, such that for any hypothesis class $\pazocal{H}$ any deterministic algorithm will suffer at least $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ regret. Furthermore, any randomized algorithm will suffer at least $\frac{1}{2}\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ regret in expectation.
Proof.
Consider a Stackelberg game, in which for any fixed context, playing the optimal strategy against the true follower type yields utility $1$, whereas playing any other strategy yields utility 00. Concretely, let $\beta_{1},\beta_{2},\ldots,\beta_{K}$ be distinct real numbers and consider the following payoff matrix, where the leader is the row player and the column player is follower type $f^{(i)}$. The leader has 2 pure actions and the follower has 3 pure actions.
$F_{1}$ $F_{2}$ $F_{3}$
$L_{1}$ $0,\frac{1}{\beta_{i}}$ $1,1$ $0,0$
$L_{2}$ $0,0$ $1,1$ $0,\frac{1}{1-\beta_{i}}$
Notice that if the leader puts probability $p>\beta_{i}$ on action $L_{1}$, then the follower will play action $F_{1}$ and the leader gets payoff 00. Similarly, if the leader puts probability $p<\beta_{i}$ on action $L_{1}$ then the follower plays action $F_{3}$ and the leader gets payoff 00. However, if the leader puts probability $p=\beta_{i}$ on action $L_{1}$, then the follower is indifferent and since the follower breaks ties in favor of the leader, she plays action $F_{2}$, giving the leader a payoff of $1$.
(Claim 1.) An adversary can force the learner to make $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes by using the following strategy: (1) select $\mathbf{z}_{t}$ as the next input from an SL-Tree of depth $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ that is shattered by $\pazocal{H}$ and (2) set the optimal mixed strategy, $\mathbf{x}_{t}^{*}$ to be the edge in the shattered tree not chosen by the leader, that is $\mathbf{x}_{t}^{*}\neq\widehat{\mathbf{x}}_{t}$, where $\widehat{\mathbf{x}}_{t}$ is the strategy played by the learner at time $t$. Since the tree is shattered by $\pazocal{H}$, there must exist some $h\in\pazocal{H}$ that is consistent with $(\mathbf{z}_{t},f_{t})_{t=1}^{\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})}$. Therefore, any deterministic learner will make at least $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes.
(Claim 2.) When the learner is randomized, at each timestep $t$, the adversary can select an $\mathbf{x}_{t}^{*}$ such that it minimizes the probability that the leader chooses $\widehat{\mathbf{x}}_{t}=\mathbf{x}_{t}^{*}$. In this case, the probability that the learner will make a mistake is at least 0.5, and equality holds when the leader selects uniformly among two branches of the tree. Summing over each layer of the tree gives us the lower bound on the learner’s expected number of mistakes. ∎
3.1.2 Upper bounds
Having established that the Stackelberg-Littlestone dimension characterizes the difficulty of learning in our setting, in this section we present an algorithm and show that it achieves optimal regret. Our guarantees in this setting are with respect to the leader’s number of mistakes (i.e., timesteps where the utility of the leader is not optimal). By the realizability assumption and the boundedness of the utility function, all mistake bounds imply the same regret bounds.
Warm-up: context-free utilities.
As a warm-up to our main results, consider a simpler setting, in which both the leader and the follower utility functions do not take the context as an input (i.e. they are context-free). This means that the context at each timestep is now only predictive of the follower’s type at that timestep. In Theorem 3.9, we show that we can reduce the problem of learning the optimal policy of the leader to online multiclass prediction. Additionally, the known optimal algorithm in online multiclass prediction (the Standard Optimal Algorithm (SOA)) meets the lower bound shown in Theorem 3.8.
Theorem 3.9.
Consider a Stackelberg game, $\pazocal{G}$, such that the utilities of both the leader and the follower are context-free. Then for any finite class, $\pazocal{H}$, of hypotheses $h:\pazocal{Z}\to\mathscr{K}$, there exists a reduction of this setting to multiclass prediction that runs in $O(\text{poly}(K,|\pazocal{A}_{f}|,|\pazocal{A}|,|\pazocal{H}|))$ time and achieves regret at most $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
Proof.
The reduction is as follows. First, compute the optimal strategy against each type of follower offline. This can be achieved using linear programming and requires $O(\text{poly}(K,|\pazocal{A}_{f}|,|\pazocal{A}|))$ runtime [16]. Then create a (possibly many-to-one) look-up table mapping each follower type $f^{(i)}$ to an optimal strategy, $\mathbf{x}_{*}^{(i)}$, for the leader against that follower. Compute $\pazocal{H}^{\prime}=\{h^{\prime}(\mathbf{z}):=\sum_{i=1}^{K}\mathbf{x}_{*}^{(i)}\cdot\mathbbm{1}\{h(\mathbf{z})=f^{(i)}\}\}_{h\in\pazocal{H}}$, which takes at most $O(\text{poly}(|\pazocal{H}|))$ time. Finally, we run the multiclass Standard Optimal Algorithm (see [17]) on $\pazocal{H}^{\prime}$ considering every optimal strategy for the leader in the look-up table as one label. It remains to show that $\textnormal{Ldim}(\pazocal{H}^{\prime})=\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$. Notice that $\pazocal{H}^{\prime}$ is the set of all hypotheses, mapping each context to an optimal strategy for the leader against every follower type that is consistent with $\pazocal{H}$ for that context. Consider any SL tree and any Littlestone tree such that the context at each corresponding node is the same among them. If the SL tree is shattered by $\pazocal{H}$ then for every root-to-leaf path there must exist a hypothesis in $\pazocal{H}$ that is consistent with the labels along that path. But by definition of $\pazocal{H}^{\prime}$ for that corresponding path in the Littlestone tree there must exist an $h^{\prime}\in\pazocal{H}^{\prime}$ that is consistent with the labelings of that path. With the same argument, the converse is also true. Therefore, the SL tree with maximal depth that is shattered by $\pazocal{H}$ is precisely the generalized Littlestone tree with maximal depth that is shattered by $\pazocal{H}^{\prime}$. ∎
We note that the algorithm described above coincides with Algorithm 1 (presented below), which we prove to be optimal in the general contextual Stackelberg game setting.
Context-dependent utilities.
We now let the utilities of both players to be functions of the context. We present our main algorithm (Algorithm 1), which we show achieves regret that matches the lower bound in Theorem 3.8 for any Stackelberg game. Our approach is inspired by the Standard Optimal Algorithm for online multiclass prediction, with the main difference that our algorithm uses a hypothesis class, $\pazocal{H}$, to optimize directly over leader strategies in the simplex.
Intuitively, our algorithm keeps track of all the “surviving” hypotheses in $\pazocal{H}$, that is, the hypotheses that are consistent with the data seen so far. At each timestep the algorithm observes the context and considers the predicted follower types of all surviving hypotheses. For every such follower type the algorithm computes the optimal leader strategy against them. Then it plays the strategy that maximizes the Stackelberg Littlestone dimension of the set of surviving hypotheses were that to be the optimal strategy for that timestep. If the algorithm makes a mistake, then the set of surviving hypotheses will have strictly smaller Stackelberg-Littlestone dimension than before. In other words, the algorithm learns from this mistake. Notice that the number of possible strategies in the simplex is infinite, yet our algorithm only considers a bounded number of them at each timestep, at most one for each follower type.
Algorithm 1 Stackelberg Standard Optimal Algorithm (SSOA)
1:input: time horizon, $T$, Stackelberg game $\pazocal{G}$, and hypothesis class $\pazocal{H}$, where $h:\pazocal{Z}\to\mathscr{K}$ for every $h\in\pazocal{H}$.
2:Set $V_{0}=\pazocal{H}$
3:for $t\in[T]$ do
4:     Observe $\mathbf{z}_{t}$
5:     Let $\mathbf{x}^{(h,t)}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{h(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))\quad\forall h\in\pazocal{H}$
6:     Let $X_{t}=\{\mathbf{x}^{(h)}:h\in V_{t-1}\}$triangleright Note that $|X_{t}|\leq K$
7:     Let $V_{t}^{(\mathbf{x})}=\{h\in V_{t-1}:\mathbf{x}^{(h,t)}=\mathbf{x}\}\quad\forall\mathbf{x}\in X_{t}$
8:     Play action $\mathbf{x}_{t}=\arg\max_{\mathbf{x}\in X_{t}}\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})$
9:     Observe $f_{t}$
10:     Set $V_{t}=\{h\in V_{t-1}:h(\mathbf{z}_{t})=f_{t}\}$
11:end
Theorem 3.10.
For any Stackelberg game, $\pazocal{G}$ and every hypothesis class, $\pazocal{H}$, the regret of Algorithm 1 is at most $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
Proof.
Let $T_{e}\subseteq[T]$ be the set of time steps on which SSOA errs and let $T_{c}=[T]\setminus T_{e}$ be the set of time steps on which SSOA correctly predicts the optimal strategy against the true follower type. Recall from Remark 3.2 that we can make $\sum_{t\in T_{c}}u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$ arbitrarily small.
Now from the boundedness of the utility function, we also know that $\sum_{t\in T_{e}}u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))\leq|T_{e}|$. It suffices to show that for every time step $t\in T_{e}$ that SSOA makes a mistake, we have that $\textnormal{SLdim}_{\pazocal{G}}(V_{t})\leq\textnormal{SLdim}_{\pazocal{G}}(V_{t+1})-1$. For the sake of contradiction, assume that $\textnormal{SLdim}_{\pazocal{G}}(V_{t})=\textnormal{SLdim}_{\pazocal{G}}(V_{t+1})$. Given that SSOA chooses a strategy $\mathbf{x}$ that maximizes $\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})$ and that SSOA made a mistake at this timestep, there must have been at least two distinct strategies $\mathbf{x},\mathbf{x}^{\prime}$ such that $\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})=\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x}^{\prime})})=\textnormal{SLdim}_{\pazocal{G}}(V_{t})$. However, this would mean that we could construct a full binary SL tree that is shattered by $\pazocal{H}$ and that has depth $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})+1$, which leads to a contradiction.
3.1.3 Relationship with Littlestone dimension
We have so far shown that the Stackelberg-Littlestone dimension characterizes learnability in our setting. We next explore how that relates to multiclass learnability. While, as discussed, learning to predict the follower type from the context is a sufficient condition for learning the leader’s optimal policy, in this section we prove that this is not a necessary condition. In Theorem 3.11, we show that the Littlestone dimension is always at most as large as the Stackelberg Littlestone dimension and in Proposition LABEL:prop:ldim=sldim we show that equality always holds in the special case of two follower types. In Theorem 3.12, we show that an out-of-the-box application of the optimal algorithm for multiclass learning in our setting can result not only in suboptimal number of mistakes, but also in suboptimal utility for the leader. This is in contrast to our proposed algorithm, which achieves optimal utility by accounting for the interaction between the hypothesis class and the Stackelberg game.
Theorem 3.11.
For any hypothesis class $\pazocal{H}$, $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})\leq\textnormal{Ldim}(\pazocal{H})$.
Proof.
If the leader plays SOA, she will make at most $\textnormal{Ldim}(\pazocal{H})$ mistakes. This is because she will predict the follower’s type incorrectly at most $\textnormal{Ldim}(\pazocal{H})$ times, and predicting the follower’s type correctly is a sufficient condition for playing the optimal mixed strategy. Since the SOA makes $\textnormal{Ldim}(\pazocal{H})$ mistakes and we know that any deterministic algorithm will make at least $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes by Theorem 3.8, we can infer that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})\leq\textnormal{Ldim}(\pazocal{H})$. ∎
While it is known that SOA makes at least $\textnormal{Ldim}(\pazocal{H})$ mistakes when learning a hypothesis class, $\pazocal{H}$, a careful reader might notice that this does not immediately show that SOA achieves suboptimal regret in our setting. This is because mispredicting the follower type does not necessarily imply that the optimal strategy against that and the true follower type do not yield the same utility for the leader. In Theorem 3.12, we prove that, indeed, there are instances where SOA will achieve smaller cumulative utility for the leader compared to our proposed algorithm.
Theorem 3.12.
There exists some hypothesis class $\pazocal{H}$ and a Stackelberg game, $\pazocal{G}$, such that running multiclass SOA on $\pazocal{H}$ yields suboptimal utility for the leader.
Proof.
Consider $\pazocal{Z}=\{1,2,3,4\}$ and $\mathscr{K}=\{f^{(1)},f^{(2)},f^{(3)}\}$. We define $\pazocal{H}=\{h_{\sigma}(z):=\mathbbm{1}\{z\neq 4\}\cdot f^{(\sigma(z))}+\mathbbm{1}\{z=4\}\cdot f^{(3)}\}_{\sigma\in\Sigma}\cup\{h,h^{\prime}\}$ , where $\Sigma$ is the set of all permutations of the set $\{1,2,3\}$ and $h,h^{\prime}$ are defined as follows.
$h(z)=\begin{cases}f^{(3)}\;\;&\text{ if }z=1\\ f^{(2)}\;\;&\text{ if }z=2\\ f^{(1)}\;\;&\text{ if }z=3\\ f^{(2)}\;\;&\text{ if }z=4\\ \end{cases}\hskip 56.9055pth^{\prime}(z)=\begin{cases}f^{(3)}\;\;&\text{ if }z=1\\ f^{(1)}\;\;&\text{ if }z=2\\ f^{(2)}\;\;&\text{ if }z=3\\ f^{(2)}\;\;&\text{ if }z=4\\ \end{cases}$
Notice that the Littlestone dimension of this hypothesis class is at least $2$. This follows from the $\pazocal{H}$-shattered tree construction shown in Figure 2.
Figure 2: Construction of an $\pazocal{H}$-shattered Littlestone tree. Each node represents a context in $\pazocal{Z}$ and each edge corresponds to a labeling of the parent node to a follower type.
Consider now a Stackelberg game with utility functions $u^{(1)},u^{(2)},u^{(3)}$ corresponding to follower types $1,2$, and $3$ respectively, defined such that if $z=1$ or $z=2$ then $\pi_{*}^{(1)}(z)=\pi_{*}^{(2)}(z)\neq\pi_{*}^{(3)}(z)$, if $z=3$ then $\pi_{*}^{(1)}(3)=\pi_{*}^{(2)}(3)=\pi_{*}^{(3)}(3)$, and if $z=4$, then $\pi_{*}^{(1)}(4)\neq\pi_{*}^{(2)}(4)=\pi_{*}^{(3)}(4)$. Here, we let $\pi_{*}^{(i)}:\pazocal{Z}\to\Delta(\pazocal{A})$ denote the optimal leader’s policy against follower type $f^{(i)}$. Concretely, consider a Stackelberg game with the leader being the row player and each follower being the column player, as shown in Table 1.
$F_{1}$ $F_{2}$
$L_{1}$ $2,1$ $0,0$
$L_{2}$ $0,0$ $2,0$
(a) Follower type $f^{(1)}$
$F_{1}$ $F_{2}$
$L_{1}$ $2,\mathbbm{1}_{\{z\neq 4\}}$ $0,0$
$L_{2}$ $0,0$ $2,\mathbbm{1}_{\{z=4\}}$
(b) Follower type $f^{(2)}$
$F_{1}$ $F_{2}$
$L_{1}$ $2,\mathbbm{1}_{\{z=3\}}$ $0,0$
$L_{2}$ $0,0$ $2,\mathbbm{1}_{\{z\neq 3\}}$
(c) Follower type $f^{(3)}$
Table 1: Utility tables for the leader (row player) and each follower type (column player)
We now show that the $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=1$. We will consider any possible SL tree. First, note that the context $z=4$ cannot be a node in an $\pazocal{H}$-shattered SL-Tree since for $z=4,\;\pi_{*}^{(2)}(4)=\pi_{*}^{(3)}(4)$, where $f^{(2)}$ and $f^{(3)}$ are the only possible follower types predicted by hypotheses in $\pazocal{H}$ for this context. Similarly, $z=3$ cannot be a node in an $\pazocal{H}$-shattered SL tree, since the leader’s optimal strategy against all follower types is identical ($\pi_{*}^{(1)}(3)=\pi_{*}^{(2)}(3)=\pi_{*}^{(3)}(3)$). So it suffices to consider contexts $1$ and $2$. In Figure 3 we show the two $\pazocal{H}$-shattered SL tree constructions with $z=1$ and $z=2$ as the root nodes respectively. Notice that the maximal depth of a full binary tree is $1$, as desired.
Say that the adversary presents context $1$ followed by context $2$. We go through the SOA outputs per timestep below.
  • ($t=1$) This node has three children corresponding to follower types $f^{(1)},f^{(2)},f^{(3)}$, with $f^{(1)}$ and $f^{(2)}$ leading to a full binary subtree of maximal depth 1 each. To see this, notice that (1) neither $h$ nor $h^{\prime}$ are consistent with the mapping of context $z=1$ to follower type $f^{(1)}$ or $f^{(2)}$, and (2) $\pazocal{H}\setminus\{h,h^{\prime}\}$ consists of all permutations mapping contexts $\{1,2,3\}$ to follower types $\{f^{(1)},f^{(2)},f^{(3)}\}$. Therefore, at most two of these contexts can exist in any root-to-leaf path of any $\pazocal{H}$-shattered tree. The child corresponding to follower type $f^{(3)}$ leads to a full binary subtree of maximal depth at least $2$, as shown in Figure 2. So, the predicted follower type is $f_{1}=f^{(3)}$ and the strategy played is $x_{1}=\pi_{*}^{(3)}(1)$. However, the true follower type is $f_{1}^{*}=f^{(1)}$ (mistake).
  • ($t=2$) This node has two children mapping $z=2$ to follower types $f^{(2)}$ and $f^{(3)}$. No subsequent context can lead to a full-binary tree that is consistent with any of these assignments, so SOA tie-breaks. The adversary in this case can always force a mistake by claiming the correct follower type to be the one not predicted by the algorithm. So if SOA predicts $f_{2}=f^{(2)}$, and plays strategy $x_{2}=\pi_{*}^{(2)}(2)$ the adversary can choose the true follower type to be $f_{2}^{*}=f^{(3)}$ (mistake).
This leads to SOA mispredicting the follower type and getting suboptimal utility in both rounds. Since the Stackelberg Littlestone dimension is $1$ and we have shown that for the above sequence of contexts SOA makes $2$ mistakes, this concludes our proof.
For ease of exposition, we also provide an overview of SSOA for the same series of nodes (see Figure (3(a)) for reference).
  • ($t=1$) This root node has two branches: $\pi_{*}^{(1)}(1)=\pi_{*}^{(2)}(1)$ and $\pi_{*}^{(3)}(1)$. Given the arguments used to prove that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=1$, the former branch leads to a full binary subtree of depth $1$, while the second branch cannot lead to a full binary tree. So SSOA chooses $x_{1}=\pi_{*}^{(1)}(1)=\pi_{*}^{(2)}(1)$ and the true follower type is $f_{1}^{*}=f^{(1)}$ (no mistake).
  • ($t=2$) The context $z=2$ leads to two branches: $\pi_{*}^{(2)}(2)$ and $\pi_{*}^{(3)}(2)$, none of which leads to a complete binary tree. SSOA tie breaks, in which case an adversary can always force a mistake by choosing the correct follower type to be the one not predicted by SSOA (mistake).
Figure 3: $\pazocal{H}$-shattered SL-Tree with (a) $z=1$ as the root node and (b) $z=2$ as the root node.
While we have shown that running multiclass SOA alone can in certain settings achieve suboptimal utility for the leader, in the following theorem we show that our algorithm may be interpreted as running SOA in a transformed space.
Theorem 3.13.
For every online contextual Stackelberg game with $K$ distinct follower types and every hypothesis class $\pazocal{H}$, there exists a hypothesis class $\Pi$ such that running multiclass SOA on $\Pi$ incurs $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes. Moreover, this algorithm coincides with SSOA.
Proof.
Consider a mapping $\zeta:\pazocal{H}\to\Pi$ that maps every $h\in\pazocal{H}$ to a function $\pi_{h}:\pazocal{Z}\to\Delta(\pazocal{A})$ such that:
$\pi_{h}(\mathbf{z}):=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x})).$
Notice that $\zeta$ is a bijection given that each follower type is distinct (see Remark 2.1). Also note that given that learning $\pazocal{H}$ is a realizable problem, so is learning $\Pi$. This is because if there exists $h^{*}\in\pazocal{H}$ such that $\forall t\in[T]\quad h^{*}(\mathbf{z}_{t})=f_{t}$, then by definition there must exist $\pi^{*}\in\Pi$ such that $\forall t\in[T]\quad\pi^{*}(\mathbf{z}_{t})=\mathbf{x}^{*}_{t}$, since $\mathbf{x}^{*}_{t}=\arg\max_{\mathbf{x}\in\Delta(A)}u(\mathbf{z}_{t},\mathbf{x},b_{h^{*}(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))$.
We know that multiclass SOA makes at most $\textnormal{Ldim}(\Pi)$ mistakes for any hypothesis class $\Pi$ with infinite labels (see Theorem 1 in [22]). Therefore, it remains to show that $\textnormal{Ldim}(\Pi)=\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
Consider any SL-Tree for $\pazocal{H}$. Notice that all nodes are contexts and all edges represent optimal leader strategies, so this is a valid Littlestone tree for $\Pi$ as well. Also, for any root-to-leaf path on this tree that is consistent with a function $h\in\pazocal{H}$, there exists $\pi_{h}\in\Pi$ that is consistent with it as well, by definition of $\pi_{h}$. The converse is also true: in any Littlestone tree for $\Pi$ consider any root-to-leaf path, that is consistent with some $\pi_{h}\in\Pi$. Then by $\zeta$ being a bijection, there exists an $h\in\pazocal{H}$ that is also consistent with it. Therefore, a tree is SL-shattered by $\pazocal{H}$ if and only if it is shattered by $\Pi$. This holds for all trees, so it will hold for the maximal-depth trees as well.
3.2 No-regret learning in the agnostic setting
In this section we relax the realizability assumption, and we consider the setting where the sequence of follower types $f_{1},\cdots,f_{T}$ is not necessarily consistent with some hypothesis $h\in\mathscr{H}$. Unlike the realizable setting, here we will not have an absolute benchmark such as the mistake bound. Instead, we measure the performance of our algorithms relative to the best in hindsight hypothesis in the class. We define $h^{*}\in\arg\max_{h\in\mathscr{H}}\sum_{t\in[T]}\mathbbm{1}\{h(\mathbf{z}_{t})=y_{t}\}$ and generalize the notion of regret:
Definition 3.14 ((Agnostic Setting) Contextual Stackelberg Regret).
Given a sequence of followers $f_{1},\dots,f_{T}$ and contexts $\mathbf{z}_{1},\dots\mathbf{z}_{T}$, the leader’s contextual Stackelberg regret in the agnostic setting is given by:
$R(T):=\max_{h\in\mathscr{H}}\sum_{t\in[T]}u(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t})),$
where $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$ is the sequence of mixed strategies played and $\pi_{h}(\mathbf{z}):=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x}))$.
Combining our SSOA algorithm with results and tools from multiclass prediction [17], we get the following regret bound.
Theorem 3.15.
Consider a Stackelberg game, $\pazocal{G}$, and a hypothesis class, $\pazocal{H}$. In the agnostic setting, there exists an algorithm that achieves regret $R(T)\leq\sqrt{\frac{1}{2}\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H})T\log{(TK)}}$.
We use a similar approach to Lemma 14 from [10] to get a (nearly-)matching lower bound.
Theorem 3.16.
Let $\pazocal{H}$ be any hypothesis class. There exists a Stackelberg game, $\pazocal{G}$, such that $\pazocal{H}$ has a finite Stackelberg-Littlestone dimension $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ and such that no algorithm in the agnostic, full-feedback setting can achieve regret smaller than $\Omega\left(\sqrt{\frac{1}{8}\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})T}\right)$.
We include the complete proofs of Theorems 3.15 and 3.16 in Appendix A.3.
4 Distributional Learning
We now shift our focus from online learning to distributional learning. Suppose that there exists an unknown, joint distribution over contexts and follower types. We first study the simpler setting, without contexts, in which we give strong generalization guarantees by exploiting the structure of the leader’s payoff function. When contexts come into play in Section 4.2, we analyze an intuitive approach that uses a black box empirical risk minimizer (ERM) and show that it is suboptimal in our setting. Analogous to our online results, we then give a better sample complexity bound that depends on a new complexity measure.
4.1 Learning without Contexts
Suppose there is a fixed, unknown distribution, $\mathscr{D}$, over the set of possible follower types. We denote with $u(\mathbf{x},b_{f^{(i)}}(\mathbf{x}))$ the leader’s payoff when committing to a (possibly mixed) strategy $\mathbf{x}\in\Delta(\pazocal{A})$ against a follower of type $f^{(i)}$. We use the notation $u_{\mathscr{D}}(\mathbf{x})=\mathbb{E}_{f\sim\mathscr{D}}[u(\mathbf{x},b_{f}(\mathbf{x}))]$ and given a sample $S$ of follower types, we write:
$u_{S}(\mathbf{x})=\frac{1}{|S|}\sum_{f\in S}u(\mathbf{x},b_{f}(\mathbf{x}))$
Given a set $S$ of $m$ samples from $\mathscr{D}$, the goal of the leader is to choose a mixed strategy that fares well in expectation against any new sample from $\mathscr{D}$. In other words, we want to choose $\hat{\mathbf{x}}\in\Delta(\pazocal{A})$ such that with high probability:
$\mathbb{P}_{S}[\max_{\mathbf{x}}u_{\mathscr{D}}(\mathbf{x})-u_{\mathscr{D}}(\hat{\mathbf{x}})>\epsilon]\leq\delta.$
For this, it suffices to show a generalization guarantee, i.e. a bound on the difference between the leader’s expected payoff and the empirical payoff over a set of follower types for any possible mixed strategy.
Definition 4.1.
A generalization guarantee for the leader in a Stackelberg game, $\pazocal{G}$, is a function $\epsilon_{\pazocal{G}}:\mathbb{Z}_{\geq 1}\times(0,1)\to\mathbb{R}_{\geq 0}$ such that for any $\delta\in(0,1)$, any $m\in\mathbb{Z}_{\geq 1}$, and any distribution $\mathscr{D}$ over the follower types, with probability at least $1-\delta$ over the draw of a set $S\sim\mathscr{D}^{m}$ for any $\mathbf{x}\in\Delta(\pazocal{A})$, the difference between the average payoff of $\mathbf{x}$ over $S$ and the expected payoff of $\mathbf{x}$ over $\mathscr{D}$ is at most $\epsilon_{\pazocal{G}}(m,\delta)$:
$\mathbb{P}_{S\sim\mathscr{D}^{m}}[\exists\mathbf{x}\in\Delta(\pazocal{A})\;\;\text{such that}\;\;|u_{\mathscr{D}}(\mathbf{x})-u_{S}(\mathbf{x})|>\epsilon_{\pazocal{G}}(m,\delta)]\leq\delta$
The implication of a generalization guarantee is that it relates the utility of the leader’s strategy that maximizes the average payoff over the training data to the expected utility of the optimal strategy. Formally, this means that for $\hat{x}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u_{S}(\mathbf{x})$ and $x^{*}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u_{\mathscr{D}}(\mathbf{x})$:
$\mathbb{P}_{S\sim\mathscr{D}^{m}}[|u_{\mathscr{D}}(\mathbf{x}^{*})-u_{S}(\hat{\mathbf{x}})|>2\epsilon_{\pazocal{G}}(m,\delta)]\leq\delta.$
Consider the class $\pazocal{P}=\{u_{\mathbf{x}}:\mathscr{K}\to\mathbb{R}\;\;\text{such that}\;\;u_{\mathbf{x}}(f)=u(\mathbf{x},b_{f}(\mathbf{x}))\}$ to be the set of payoff functions for the leader. Each payoff function is parameterized by the leader’s mixed strategy, takes as input the follower’s type, and outputs the leader’s utility. We also define the dual class $\tilde{\pazocal{P}}=\{\tilde{u}_{f}:\Delta(\pazocal{A})\to\mathbb{R}\;\;\text{such that}\;\;\tilde{u}_{f}(\mathbf{x})=u(\mathbf{x},b_{f}(\mathbf{x}))\}$ to be the set of payoff functions for the leader, parameterized by a follower type that take as input the leader’s mixed strategy and output the leader’s utility. Consider the following definition, introduced by [6] and adapted to our setting:
Definition 4.2.
A class $\pazocal{H}$ of hypotheses, $h:\mathscr{K}\to\mathbb{R}$, is $(d,t)$-delineable if:
  1. 1.
    The class $\pazocal{H}$ consists of hypotheses that are parameterized by vectors $\mathbf{x}$ from a set $\pazocal{X}\subseteq\mathbb{R}^{d}$; and
  2. 2.
    For any $f\in\mathscr{K}$, there is a set $\mathscr{J}$ of $t$ hyperplanes such that for any connected component $\pazocal{X}^{\prime}=\pazocal{X}\setminus\mathscr{J}$, the dual function $h_{f}(\mathbf{x})$ is linear over $\pazocal{X}^{\prime}$.
With this definition, we make the following useful proposition.
Proposition 4.3.
The class of leader payoff functions, $\pazocal{P}$ that are parameterized by the leader’s mixed strategy and take as input the follower’s type is $(|\pazocal{A}|,\binom{|\pazocal{A}_{f}|}{2})$-delineable.
Proof.
The statement holds from the following two facts: (1.) Given a fixed follower type, $f$, we have that $\tilde{u}_{f}(\mathbf{x})$ is a piece-wise linear function. (2.) For any fixed follower type, $f$, we have that $u_{f}(\mathbf{x})$ will have at most $|A_{f}|(|A_{f}|-1)$ hyperplanes going through it. Each hyperplane corresponds to a change in best-response strategies for the follower from one pure action to another, so there can be at most $\binom{|\pazocal{A}_{f}|}{2}$ hyperplanes. ∎
Next, we present the definition of pseudo-dimension [33], which is complexity measure that characterizes a hypothesis class and in our setting helps us reason about how difficult the learning task of the leader is. Informally speaking, the more complex the hypothesis class, $\pazocal{P}$, the harder it is for the leader to distinguish between strategies, $\mathbf{x}$, which parameterize the class, and therefore, more samples are needed to learn.
Definition 4.4 (Pseudo-dimension).
Let $\pazocal{F}$ be a class of functions $f:A\to\mathbb{R}$, with an abstract domain, $A$. We say that $y^{(1)},\ldots,y^{(n)}$, witness the shattering of $S=\{x^{(1)},\ldots,x^{(n)}\}\subseteq A$ if for all $T\subseteq S$, there exists a function $f_{T}\in\pazocal{F}$, such that for all $x^{(i)}\in T,\;f_{T}(x^{(i)})\leq y^{(i)}$ and for all $x^{(i)}\not\in T,\;f_{T}(x^{(i)})>y^{(i)}$. We then say that the pseudo-dimension of $\pazocal{F}$, denoted $\textnormal{Pdim}(\pazocal{F})$, is the size of the largest set that can be shattered by $\pazocal{F}$.
Leveraging results from the data-driven algorithm design literature ([4, 8, 7]), such as Lemma 3.10 in [6] and Proposition 4.3 we have the following result:
Theorem 4.5.
The pseudo-dimension of the hypothesis class, $\pazocal{P}$ satisfies $\textnormal{Pdim}(\pazocal{P})\leq 9|\pazocal{A}|\log\left(4|\pazocal{A}||\pazocal{A}_{f}|^{2}\right)$.
So we also get the generalization guarantee from Pollard [33] and Dudley [18] with function:
$\epsilon_{\pazocal{G}}(m,\delta)=O\left(U\sqrt{\frac{|\pazocal{A}|\log(|\pazocal{A}||\pazocal{A}_{f}|^{2})}{m}}+U\sqrt{\frac{\ln(1/\delta)}{m}}\right),$
where $U$ is the maximum payoff value for the leader.
Discussion.
We note that the above generalization result is particularly strong. Informally, it asserts that no matter what the distribution over the follower types is, if the leader has so far encountered a number of followers that scales linearly with the size of her action set and logarithmically with the follower’s action set, it suffices for her to play a strategy that maximizes her average utility over the followers seen so far. This result holds even for an infinite number of follower types and extends prior work on learning in Stackelberg games without contextual information [26].
4.2 Learning with Contexts
We now shift to analyzing the setting where the leader has access to a data set consisting of tuples of contexts and follower types, $S_{m}=\{(\mathbf{z}_{1},f_{1}),\ldots,(\mathbf{z}_{m},f_{m})\}$. Given new contextual information, the leader aims to choose a strategy that maximizes her payoff against a follower, whose type is unknown. Here we introduce the study of the statistical complexity of distributional learning in Stackelberg games with side information.
Consider a distribution $\mathscr{D}$ over $\pazocal{Z}\times\mathscr{K}$. We define the leader’s payoff as a function of a policy $\pi:\pazocal{Z}\to\Delta(\pazocal{A})$ with respect to $\mathscr{D}$ as:
$u_{\mathscr{D}}(\pi)=\mathop{\mathbb{E}}_{(\mathbf{z},f^{(i)})\sim\mathscr{D}}\left[u(\mathbf{z},\pi(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi(\mathbf{z})))\right]$
and the leader’s average payoff on a sample $S_{m}=\{(\mathbf{z}_{1},f_{1}),\ldots,(\mathbf{z}_{m},f_{m})\}\in(\pazocal{Z}\times\mathscr{K})^{m}$ as:
$u_{S_{m}}(\pi)=\frac{1}{m}\sum_{i=1}^{m}\left[u(\mathbf{z}_{i},\pi(\mathbf{z}_{i}),b_{f_{i}}(\mathbf{z}_{i},\pi(\mathbf{z}_{i})))\right].$
The goal of the learner is to deploy an algorithm, $\textsc{Alg}:\cup_{m=0}^{\infty}(\pazocal{Z}\times\mathscr{K})^{m}\to\Delta(A)^{\pazocal{Z}}$, such that given a training sequence $S_{m}=\{(\mathbf{z}_{1},f_{1}),\cdots,(\mathbf{z}_{m},f_{m})\}$, the algorithm outputs a policy $\pi_{\textsc{Alg}}:\pazocal{Z}\to\Delta(A)$ that maximizes $u_{\mathscr{D}}(\pi)$.
Suppose that the leader has access to a set of hypotheses, $\pazocal{H}$, such that each hypothesis in $\pazocal{H}$, maps contexts to follower types ($h:\pazocal{Z}\to\mathscr{K}$). We define the true error function of a hypothesis $h$ with respect to a distribution, $\mathscr{D}$, as:
$\text{Err}_{\mathscr{D}}(h)=\mathbb{P}_{(\mathbf{z},f)\sim\mathscr{D}}\left[h(\mathbf{z})\neq f\right].$
Similarly, we define the empirical error of a hypothesis $h$ on a sample $S_{m}$ as:
$\text{Err}(h)=\frac{1}{m}|\{i\in[m]:h(\mathbf{z}_{i})\neq f_{i}\}|$
and the approximation error of a hypothesis class $\pazocal{H}$ on a distribution $\mathscr{D}$ to be:
$\text{Err}_{\mathscr{D}}(\pazocal{H}):=\inf_{h\in\pazocal{H}}\text{Err}_{\mathscr{D}}(h).$
Consider an optimal hypothesis, $h^{*}\in\pazocal{H}$, such that given any context, $h^{*}$ always predicts the follower type correctly, i.e. $\text{Err}_{\mathscr{D}}(h^{*})=0$. Since we are in the realizable setting, for the remainder of this section we will overload notation and use $\mathscr{D}$ to refer to the marginal distribution over the context space $\pazocal{Z}$.
We now define the Natarajan dimension, a measure of complexity of a multiclass hypothesis class (introduced by  Natarajan [29]), which will be useful in our setting. Informally, the Natarajan dimension measures the ability of a class of functions, whose image is a finite set, to fit complex patterns. [^4] Note that for binary labels, the definition of Natarajan dimension recovers the well-studied VC-dimension of a hypothesis class.
Definition 4.6 (Natarajan Dimension).
Let $\pazocal{H}\subseteq\mathscr{K}^{\pazocal{Z}}$ be a hypothesis class and let $S\subseteq\pazocal{Z}$. We say that $\pazocal{H}$ N-shatters $S$ if there exist $\phi_{1},\phi_{2}:S\to\mathscr{K}$ such that $\forall\mathbf{z}\in\pazocal{Z}\;\phi_{1}(\mathbf{z})\neq\phi_{2}(\mathbf{z})$ and for any $T\subseteq S$ there exists a $g\in\pazocal{H}$ such that:
$\forall\mathbf{z}\in T,\;g(\mathbf{z})=\phi_{1}(\mathbf{z})\quad\text{and}\quad\forall\mathbf{z}\in S\setminus T,\;g(\mathbf{z})=\phi_{2}(\mathbf{z}).$
The Natarajan dimension of $\pazocal{H}$, denoted $d_{N}(\pazocal{H})$, is the maximal cardinality of a set that is shattered by $\pazocal{H}$.
We define the realizable sample complexity of a Stackelberg-learning algorithm Alg for a hypothesis class $\pazocal{H}$ to be the function $m_{\textsc{Alg},\pazocal{H}}^{r}$ such that for every $\epsilon,\delta>0$, $m_{\textsc{Alg},\pazocal{H}}^{r}$ is the minimal integer such that for every $m\geq m_{\textsc{Alg},\pazocal{H}}^{r}$ and every distribution $\mathscr{D}$ over $\pazocal{Z}$, we have:
$\mathbb{P}_{S_{m}}\left(u_{\mathscr{D}}(\textsc{Alg}(S_{m}))<u_{\mathscr{D}}(\pi_{h^{*}})-\epsilon\right)\leq\delta.$ (2)
where $\pi_{h^{*}}(\mathbf{z}):=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h^{*}(\mathbf{z})}(\mathbf{z},\mathbf{x}))$.
4.2.1 Vanilla ERM learner
We define an Empirical Risk Minimization (ERM) learner for the hypothesis class $\pazocal{H}$ as an algorithm, such that for any sample $S_{m}$ returns a function that minimizes the empirical error of that function relative to any other function in $\pazocal{H}$. Formally, we say:
$h_{ERM}$ $\in\arg\min_{h\in\pazocal{H}}\frac{1}{m}\sum_{i=1}^{m}\mathbbm{1}\{h(\mathbf{z}_{i})\neq f_{i}\}.$ (3)
We now show that the finiteness of the Natarajan dimension is a sufficient condition for learning the optimal policy for the leader in structured Stackelberg games. This can be achieved by using a simple ERM learner as in (3).
Proposition 4.7.
Consider an algorithm $\textsc{Alg}_{\text{ERM}}:\cup_{n=0}^{\infty}(\pazocal{Z}\times\mathscr{K})^{n}\to\Delta(\pazocal{A})^{\pazocal{Z}}$ that outputs the policy:
$\pi_{\text{ERM}}(\mathbf{z})$ $\in\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h_{ERM}(\mathbf{z})}(\mathbf{z},\mathbf{x})),$
where $h_{ERM}$ satisfies equation (3). Then for any such $h_{\text{ERM}}$, there exists constant $C>0$, such that for every $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$,
$m^{r}_{\pazocal{A}_{\text{ERM}}}\leq C\left(\frac{d_{N}(\pazocal{H})\ln(K)\ln(\frac{1}{\epsilon})+\ln(\frac{1}{\delta})}{\epsilon}\right)$
We present the proof of Proposition 4.7 in Appendix B.
4.2.2 Stackelberg ERM learner
In this section, we show how to improve upon the sample complexity bound shown in Proposition 4.7, by leveraging the interaction between the Stackelberg game and the hypothesis class. Consider the following hypothesis class of possible leader policies:
$\Pi_{\pazocal{H},\pazocal{G}}=\Bigl{\{}\pi_{h}:\pazocal{Z}\to\Delta(\pazocal{A})|h\in\pazocal{H},\pi_{h}(\mathbf{z})=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x}))\Bigl{\}}$
and the following variation of an ERM rule.
Definition 4.8 (Stackelberg ERM).
Consider a hypothesis class $\pazocal{H}\subseteq\mathscr{K}^{\pazocal{Z}}$ and a Stackelberg game, $\pazocal{G}$. We define the Stackelberg ERM learning algorithm, $\textsc{Alg}_{\text{S-ERM}}:\cup_{n=0}^{\infty}(\pazocal{Z}\times\mathscr{K})^{n}\to\Delta(\pazocal{A})^{\pazocal{Z}}$, such that given a training sequence $S_{m}=\{(\mathbf{z}_{1},f_{1}),\cdots,(\mathbf{z}_{m},f_{m})\}$, outputs the policy $\pi_{\text{S-ERM}}:\pazocal{Z}^{m}\to\Delta(\pazocal{A})$ that satisfies:
$\pi_{\text{S-ERM}}\in\arg\min_{\pi\in\Pi_{\pazocal{H},\pazocal{G}}}\frac{1}{m}\sum_{i=1}^{m}\mathbbm{1}\{\pi(\mathbf{z}_{i})\neq\mathbf{x}_{i}^{*}\},$ (4)
where $\mathbf{x}_{i}^{*}=\arg\max_{x\in\Delta(\pazocal{A})}\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{i},\mathbf{x},b_{f_{i}}(\mathbf{z}_{i},\mathbf{x}))$.
We define the following measure of complexity of a hypothesis class with respect to a Stackelberg game.
Definition 4.9 (Stackelberg-Natarajan Dimension).
Consider a hypothesis class $\pazocal{H}\subseteq\mathscr{K}^{\pazocal{Z}}$, a Stackelberg game, $\pazocal{G}$, and let $S\subseteq\pazocal{Z}$. We say that $\pazocal{H}$ SN-shatters $S$ with respect to game $\pazocal{G}$, if there exist $\phi_{1},\phi_{2}:S\to\Delta(\pazocal{A})$ such that $\forall\mathbf{z}\in\pazocal{Z}\;\phi_{1}(\mathbf{z})\neq\phi_{2}(\mathbf{z})$ and for any $T\subseteq S$ there exists a $g\in\pazocal{H}$ such that:
$\forall\mathbf{z}\in T,\;\pi_{g}(\mathbf{z})=\phi_{1}(\mathbf{z})\quad\text{and}\quad\forall\mathbf{z}\in S\setminus T,\;\pi_{g}(\mathbf{z})=\phi_{2}(\mathbf{z}).$
The Stackelberg-Natarajan dimension of $\pazocal{H}$ with respect to game $\pazocal{G}$, denoted as $d_{SN,\pazocal{G}}(\pazocal{H})$, is the maximal cardinality of a set that is SN-shattered by $\pazocal{H}$.
Theorem 4.10.
For any hypothesis class $\pazocal{H}$, $d_{SN,\pazocal{G}}(\pazocal{H})\leq d_{N}(\pazocal{H})$. Additionally, there exists a hypothesis class and a Stackelberg game with distinct follower types for which the strict inequality holds.
Proof.
(Claim 1.) Consider a hypothesis class, $\pazocal{H}$. It suffices to show that for any set $S\subseteq\pazocal{Z}$ SN-shattered by $\pazocal{H}$, $S$ is also N-shattered by $\pazocal{H}$. This holds because for any pair of points $\mathbf{z}_{1},\mathbf{z}_{2}\in S$, for which there are two different leader strategies consistent with some $g\in\pazocal{H}$, (and assuming a fixed tie-breaking rule) then it must be that $g(\mathbf{z}_{1})\neq g(\mathbf{z}_{2})$, i.e. that function predicts two different follower types for these contexts.
(Claim 2.) Consider the set $\pazocal{Z}=\{\mathbf{z}_{1},\mathbf{z}_{2}\}$ and the set of follower types $\{f^{(0)},f^{(1)},f^{(2)}\}$. Also consider the following hypothesis class $\pazocal{H}=\{h_{i}\}_{i\in[4]}$, where $h_{1}(\mathbf{z}_{1})=h_{4}(\mathbf{z}_{1})=f^{(0)}$, $h_{2}(\mathbf{z}_{1})=h_{3}(\mathbf{z}_{1})=f^{(1)}$, $h_{1}(\mathbf{z}_{2})=h_{3}(\mathbf{z}_{2})=f^{(2)}$, and $h_{2}(\mathbf{z}_{2})=h_{4}(\mathbf{z}_{2})=f^{(1)}$.
It is easy to verify that for $\phi_{1},\phi_{2}$, where $\phi_{1}(\mathbf{z}_{1})=f^{(0)},\;\phi_{1}(\mathbf{z}_{2})=f^{(1)},\;\phi_{2}(\mathbf{z}_{1})=f^{(1)}$, and $\phi_{2}(\mathbf{z}_{2})=f^{(2)}$, $\pazocal{H}$ shatters the set $\{\mathbf{z}_{1},\mathbf{z}_{2}\}$, therefore, $d_{N}(\pazocal{H})=2$.
Now consider a Stackelberg game, where for all contexts $\mathbf{z}$, $\pi_{*}^{(0)}(\mathbf{z})\neq\pi_{*}^{(2)}(\mathbf{z})$, while $\pi_{*}^{(1)}(\mathbf{z}_{1})=\pi_{*}^{(0)}(\mathbf{z}_{1})$, and $\pi_{*}^{(1)}(\mathbf{z}_{2})=\pi_{*}^{(2)}(\mathbf{z}_{2})$, where $\pi_{*}^{(i)}:\pazocal{Z}\to\Delta(\pazocal{A})$, denotes the optimal policy of the leader against a follower type $f^{(i)}$. Then notice that we always have $\pi_{h^{*}}(\mathbf{z}_{1})=\pi_{*}^{(0)}(\mathbf{z}_{1})$ and $\pi_{h^{*}}(\mathbf{z}_{2})=\pi_{*}^{(2)}(\mathbf{z}_{2})$. Therefore, $d_{SN,\pazocal{G}}(\pazocal{H})=0$. ∎
As in the online setting, depending on the hypothesis class and the Stackelberg game at hand, the gap between the Stackelberg-Natarajan dimension and the Natarajan dimension can be arbitrarily large. We illustrate this via the following example.
Example 4.11.
Consider the hypothesis class of conjunctions, where each hypothesis is defined over a set of Boolean variables, $\pazocal{Z}=\{z_{1},\cdots,z_{n}\}$, such that $\pazocal{H}_{\text{conj}}=\{h:\{0,1\}^{n}\to\{0,1\}|h(z)=\land_{i\in S}z_{i}\land_{j\in T}z_{j},S,T\subseteq[n],S\cap T=\emptyset\}$. Since each hypothesis outputs in binary, this hypothesis class could naturally be considered in a Stackelberg game with two follower types, $f^{(0)},f^{(1)}$. The Natarajan dimension in this case reduces to the well-studied notion of the VC-dimension, which is known to be equal to $n$ for $\pazocal{H}_{\text{conj}}$. At the same time, the $d_{SN,\pazocal{G}}(\pazocal{H}_{\text{conj}})$ will be upper bounded by the size of the set $\{z\in\{0,1\}^{n}|\pi_{*}^{(0)}(z)\neq\pi_{*}^{(1)}(z)\}$, which could be as small as $1$ for two distinct follower types.
Theorem 4.12.
Consider the Stackelberg ERM policy, $\pi_{\text{S-ERM}}$, as in Definition 4.9 and let $h_{\text{S-ERM}}$ be the corresponding hypothesis in $\pazocal{H}$. Then universally for any such policy, there exists constant $C$, such that for every $\pazocal{H}\subseteq\mathscr{K}^{\pazocal{Z}}$ and distribution $\mathscr{D}$,
$m^{r}_{\pazocal{A}_{\text{ERM}}}\leq C\frac{1}{\epsilon}\left(d_{SN}(\pazocal{H})\log\left(\frac{d_{SN}(\pazocal{H})\cdot K}{\epsilon}\right)+\log\left(\frac{1}{\delta}\right)\right)$
Proof Sketch: The proof follows from a similar argument as that of the sample complexity bounds of multiclass learning [17]. The only part of the proof that does not apply here is Natarajan’s lemma. We prove an appropriate generalization in Lemma 4.13. The full proof of Theorem 4.12 is presented in Appendix B.
Lemma 4.13.
$|\Pi_{\pazocal{H},\pazocal{G}}|\leq|\pazocal{Z}|^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$.
We present the complete proof of Lemma 4.13 in Appendix B.
Remark 4.14.
It is worth noting that even when there is only one context (i.e. $\pazocal{Z}$ is a singleton), our results in this section are incomparable with those in Section 4.1. This is because in this section, the goal of the leader is to learn a deterministic mapping from contexts to follower types, assuming that one exists. So, if the contexts are the same between all examples in the training set, the follower types corresponding to these examples will also be the same. In contrast, the context-free setting we consider in Section 4.1 allows for a distribution over follower types. Thus the two bounds are generally incomparable.
5 Conclusion
In this work we study structured Stackelberg games, where contextual information is predictive of the follower’s type. The follower type that the leader is playing against is unknown, while the context that carries information about the state of the world is presented to both players. We show that the leader can use the unknown, fixed mapping from contexts to follower types to inform her strategy. In the online setting, our algorithms learn the optimal policy for the leader and achieve tight no-regret bounds. These bounds scale with a new notion of statistical complexity that directly exploits the structure of Stackelberg games. We show similar results in the distributional setting, where contexts and follower types are drawn from a (joint) distribution.
An interesting direction for future research is to analyze structured Stackelberg games in the beyond worst-case analysis framework. It has recently been shown that smooth online learning is as easy as distributional learning, with regret bounds scaling with the VC dimension of the hypothesis class. Extending these findings to multiclass learning and structured Stackelberg games might lead to improved results both from a statistical and computational point of view.
Appendix A Details and Proofs from Section 3
A.1 A technical subtlety
An important component of our analysis is that the leader must be able to compute the optimal strategy for a given context and follower type. For this, we state that for any context $\mathbf{z}\in\pazocal{Z}$ the leader incurs negligible regret by restricting herself to a subset $\mathscr{E}_{\mathbf{z}}\subseteq\Delta(\pazocal{A})$ which is finite and computable. To formally state our result we present the notion of a contextual best response region, which was first introduced in [24] and which we repeat here for completeness.
Definition A.1 (Contextual Follower Best-Response Region).
For context $\mathbf{z}$, follower type $f^{(i)}$, and follower action $a_{f}\in\pazocal{A}_{f}$, let $X_{\mathbf{z}}(f^{(i)},a_{f})\subseteq\pazocal{X}$ denote the leader’s mixed strategies such that a follower of type $f^{(i)}$ best-responds to all $x\in X_{\mathbf{z}}(f^{(i)},a_{f})$ by playing action $a_{f}$, that is, $X_{\mathbf{z}}(f^{(i)},a_{f})=\{\mathbf{x}\in\pazocal{X}:b_{f^{(i)}}(\mathbf{z},\mathbf{x})=a_{f}\}$
Notice that any possible follower best-response region is (1) a subset of $\Delta(\pazocal{A})$ and (2) the intersection of finitely-many halfspaces. Therefore, every $X_{\mathbf{z}}(f^{(i)},a_{f})$ is convex and bounded. However, the leader may not be able to optimize over the extreme points of a best-response region, as it may not be closed due to the followers’ tie-breaking.
Definition A.2 ($\delta$-approximate extreme points).
Fix a context $\mathbf{z}\in\pazocal{Z}$ and consider the set of all non-empty contextual best-response regions. For $\delta>0$, $\mathscr{E}_{\mathbf{z}}(\delta)$ is the set of leader mixed strategies such that for all best-response functions $\sigma$ and any $\mathbf{x}\in\Delta(\pazocal{A})$ that is an extreme point of cl$(\pazocal{X}_{\mathbf{z}}(\sigma))$, $\mathbf{x}\in\mathscr{E}_{\mathbf{z}}(\delta)$ if $\mathbf{x}\in\pazocal{X}_{\mathbf{z}}(\sigma)$. Otherwise there is some $\mathbf{x}^{\prime}\in\mathscr{E}_{\mathbf{z}}(\delta)$ such that $\mathbf{x}^{\prime}\in\pazocal{X}_{\mathbf{z}}(\sigma)$ and $\|\mathbf{x}^{\prime}-\mathbf{x}\|_{1}\leq\delta$.
Notice that for sufficiently small $\delta$ i.e. $\delta=\mathscr{O}(\frac{1}{T})$, we get that the additional regret of the leader by choosing actions over $\mathscr{E}^{(i)}$ rather than $X(f^{(i)},a_{f})$ is negligible. Therefore, we use the shorthand $\mathscr{E}^{(i)}:=\mathscr{E}^{(i)}(\delta)$ for the rest of the section. The following lemma from [24] formalizes this notion.
Lemma A.3.
For any sequence of followers $f_{1},\ldots f_{T}$ and any leader policy $\pi$, there exists a policy $\pi^{(\mathscr{E})}:\pazocal{Z}\rightarrow\cup_{\mathbf{z}\in\pazocal{Z}}\mathscr{E}_{\mathbf{z}}$ that, when given context $\mathbf{z}$, plays a mixed strategy in $\mathscr{E}_{\mathbf{z}}$ and guarantees that $\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\pi^{(\mathscr{E})}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi^{(\mathscr{E})}(\mathbf{z}_{t})))\leq 1$. Moreover, the same result holds in expectation over any distribution over follower types.
Given Lemma A.3 and for ease of notation, in this paper we assume that the leader will always set the parameter $\delta$, such that it does not affect our regret bounds.
A.2 Using multiclass SOA as a black-box algorithm
Consider the following algorithm.
Algorithm 2 Learning with full-feedback and realizable follower types
1:for $t\in[T]$ do
2:     Observe $\mathbf{z}_{t}$
3:     Evaluate $\hat{f}_{t}=h_{t}(\mathbf{z}_{t})\leftarrow\text{SOA}(\mathbf{z}_{t})$
4:     Commit to $\mathbf{x}_{t}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{\hat{f}_{t}}(\mathbf{z}_{t},\mathbf{x}))$
5:     Observe follower type $f_{t}$
6:end
Here SOA refers to the online multiclass learning algorithm proposed by [17].
A.3 Online Learning in the Agnostic Setting
A.3.1 Upper bound
Algorithmic procedure.
Consider the “learning with expert advice” setting where the leader at each timestep $t$ would get to observe not only the context $\mathbf{z}_{t}$ but also $N$ advices, $\{\phi_{t}^{(1)},\cdots,\phi_{t}^{(N)}\}\in\Delta(\pazocal{A})^{N}$. Then we can denote the cumulative loss of expert $i$ at timestep $T$ as:
$L_{T,i}=-\sum_{t=1}^{T}u(\mathbf{z}_{t},\phi_{t}^{(i)},b_{f_{t}}(\mathbf{z}_{t},\phi_{t}^{(i)})).$
For an appropriately defined set of experts we can formalize the following algorithm.
Algorithm 3 Learning in the full-feedback, agnostic setting
1:input: time horizon $T$
2:Set $\eta=\sqrt{8\ln(N)/T}$
3:for $t\in[T]$ do
4:     Observe $\mathbf{z}_{t}$
5:     Receive advices $\{\phi_{t}^{(1)},\cdots,\phi_{t}^{(N)}\}\in\Delta(A)^{N}$
6:     Commit to $\mathbf{x}_{t}=\phi_{t}^{(i)}$ with probability proportional to $\exp(-\eta L_{t-1,i})$
7:     Receive utility $u(\mathbf{z}_{t},a_{l},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$, where $a_{l}$ is drawn from $\mathbf{x}_{t}$
8:     Observe follower type $f_{t}$
9:end
See 3.15
Proof.
First consider the case where we construct an expert for every $h\in\mathscr{H}$ and keep track of their cumulative loss $L_{t,h}$ at each timestep $t$. In our case, for each $h\in\pazocal{H}$, we can construct an expert $i$ that will predict the type, $f_{t}^{(i)}=h(\mathbf{z}_{t})$ at that timestep and based on that evaluate:
$\phi_{t}^{(i)}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u(\mathbf{z}_{t},a_{l},b_{f_{t}^{(i)}}(\mathbf{z}_{t},\mathbf{x})).$
From the regret analysis of the Hedge algorithm (see [14] Chapter 2), we know that:
$\sum_{t\in[T]}\mathbb{E}\left[u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))\right]\geq\max_{i\in[N]}\sum_{t=1}^{T}u(\mathbf{z}_{t},\phi_{t}^{(i)},b_{f_{t}}(\mathbf{z}_{t},\phi_{t}^{i}))-\sqrt{\frac{1}{2}\ln(|\mathscr{H}|)T},$
where the expectation is taken over the randomness in the Hedge algorithm.
Now it remains to construct a smaller set of experts such that the upper bound on the regret depends on the $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$, instead of $\ln(|\pazocal{H}|)$. For some time horizon $T$, we define $S_{T}=\{S\subset[T]:|S|\leq\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})\}$.
For each $S\in S_{T}$ and $g:S\rightarrow\Delta(\pazocal{A})$, we construct an expert, $E_{S,g}$. Expert $E_{S,g}$ immitates an instance of the SSOA algorithm, where SSOA errs at all timesteps $t\in S$ and for each timestep $t\in S$, the optimal strategy for the leader at the timestep is given by $g(t)$. We show the implementation of each expert in Algorithm 4 below.
Algorithm 4 Implementation of expert $E_{S,g}$
1:input: time horizon $T$
2:Set $V_{0}=\pazocal{H}$
3:for $t\in[T]$ do
4:     Observe $\mathbf{z}_{t}$
5:     Let $\mathbf{x}^{(h,t)}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{h(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))\quad\forall h\in\pazocal{H}$
6:     Let $X_{t}=\{\mathbf{x}^{(h)}:h\in V_{t-1}\}$triangleright $|X_{t}|\leq K$
7:     Let $V_{t}^{(\mathbf{x})}=\{h\in V_{t-1}:\mathbf{x}^{(h,t)}=\mathbf{x}\}\quad\forall\mathbf{x}\in X_{t}$
8:     Play action $l_{t}=\arg\max_{\mathbf{x}\in X_{t}}\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})$
9:     Set $l_{t}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{\mathbf{z}_{t}\rightarrow k})$
10:     if $t\in S$ then
11:         Predict $g(t)$
12:         $V_{t+1}\leftarrow V_{t}^{\mathbf{z}_{t}\rightarrow g(t)}$
13:     else
14:         Predict $l_{t}$
15:         $V_{t+1}\leftarrow V_{t}^{\mathbf{z}_{t}\rightarrow l_{t}}$
16:end
Notice that for each possible timestep $t\in T$, a single context $\mathbf{z}_{t}$ is presented, so there are at most $K$ possible optimal strategies for the leader, one for each follower type. So, the number of experts that can be constructed in this way is:
$\sum_{i=0}^{\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H})}\binom{T}{i}(K-1)^{i}\leq(TK)^{\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H})}.$
So using the above regret upper bound with this set of experts we get that Algorithm 3 achieves utility at least:
$\max_{S,g}-L_{S,g,T}-\sqrt{\frac{1}{2}\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H})T\ln(TK)},$
where $-L_{S,g,T}$ is the accumulated utility of expert $E_{S,g}$ up to timestep $T$. It remains to show that $\max_{S,g}-L_{S,g,T}\geq\max_{h\in\mathscr{H}}-L_{h,T}$. Notice that for any $h\in\mathscr{H}$ and for any set of timesteps $S$ that SSOA errs in, there exists an expert in the constructed set of experts such that $L_{S,g,T}=L_{h,T}$ when executed on the sequence $(\mathbf{z}_{1},h(\mathbf{z}_{1})),\cdots,(\mathbf{z}_{T},h(\mathbf{z}_{T}))$.
A.3.2 Lower bound
See 3.16
Proof.
We let $d=\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$. For simplicity assume that $T$ is an integer multiple of $d$, say $T=kd$ for $k\in\mathbb{N}$. Consider a tree, $(\mathbf{v}_{1},\mathbf{v}_{1},\ldots,\mathbf{v}_{2^{d}-1})$, of depth, $d$, that is SL-shattered by $\pazocal{H}$ such that each node $\mathbf{v}_{i}\in\pazocal{Z}$ has two edges (each leading to a child) corresponding to two strategies for the leader $\mathbf{x}_{i,1},\mathbf{x}_{i,2}\in\Delta(\pazocal{A})$, such that $\mathbf{x}_{i,1}\neq\mathbf{x}_{i,2}$. By definition, we have that for each $i$, there exist at least two functions $h_{i,1},h_{i,2}\in\pazocal{H}$ such that at context $\mathbf{v}_{i}$, they predict follower types $f_{1}=h_{i,1}(\mathbf{v}_{i})$ and $f_{2}=h_{i,2}(\mathbf{v}_{i})$, against which the strategies $\mathbf{x}_{i,1}$ and respectively $\mathbf{x}_{i,2}$ are optimal. In other words, we construct a Stackelberg game, playing the optimal strategy against the mispredicted follower type gives utility 00, whereas playing the optimal strategy against the correct follower type gives utility $1$ (for an explicit construction see proof of Theorem 3.8). Then, we have that for every $i$:
$u(\mathbf{v}_{i},\mathbf{x}_{i,1},b_{f_{1}}(\mathbf{v}_{i},\mathbf{x}_{i,1}))-u(\mathbf{v}_{i},\mathbf{x}_{i,2},b_{f_{1}}(\mathbf{v}_{i},\mathbf{x}_{i,2}))=1$
and similarly,
$u(\mathbf{v}_{i},\mathbf{x}_{i,2},b_{f_{2}}(\mathbf{v}_{i},\mathbf{x}_{i,2}))-u(\mathbf{v}_{i},\mathbf{x}_{i,1},b_{f_{2}}(\mathbf{v}_{i},\mathbf{x}_{i,1}))=1.$
We will now construct a sequence $(\mathbf{z}_{1},\mathbf{x}_{1}),(\mathbf{z}_{2},\mathbf{x}_{2}),\ldots,(\mathbf{z}_{T},\mathbf{x}_{T})$, which we will show forces any algorithm to meet the lower bound in the theorem statement. From the SL-shattered tree, we follow a root-to-leaf path, say $(\tilde{\mathbf{z}}_{1},\tilde{\mathbf{x}}_{1}),\ldots,(\tilde{\mathbf{z}}_{d},\tilde{\mathbf{x}}_{d})$. For each node $\tilde{\mathbf{z}}_{i}$ on the path, where $i\in[d]$, we construct a set of $k$ pairs of contexts and follower types $\{(\mathbf{z}_{(i-1)k+1}^{a},\mathbf{x}_{(i-1)k+1}^{a}),\ldots,(\mathbf{z}_{ik}^{a},\mathbf{x}_{ik}^{a})\}$ as follows.
  • We set $\mathbf{z}_{(i-1)k+1}^{a}=\ldots=\mathbf{z}_{ik}^{a}=\tilde{\mathbf{z}}_{i}$.
  • We draw each $\mathbf{x}_{t}^{a}$ for $t\in\{(i-1)k+1,\ldots,ik\}$, independently and uniformly at random from the two optimal strategies corresponding to the edges leading to the two children of $\mathbf{z}_{i}$. We denote the follower type and optimal strategy for each child, $f_{A,i},\mathbf{x}_{A,i}$ and $f_{B,i},\mathbf{x}_{B,i}$ respectively.
For each set of $k$ examples, let $T_{i}=\{(i-1)k+1,\ldots,ik\}$ be the set of indeces of the examples corresponding to $\tilde{\mathbf{z}}_{i}$. It suffices to lower bound:
$\max_{h\in\pazocal{H}}\sum_{t\in T_{i}}u(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$
For ease of notation, we denote $h^{*}=\arg\max_{h\in\pazocal{H}}\sum_{t\in T_{i}}u(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t})))$, $\;u_{h^{*},t}=u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))$, $u_{t}=u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$, and $u_{t}^{*}=u(\mathbf{z}_{t},\pi_{f_{t}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{f_{t}}(\mathbf{z}_{t})))$, where $\pi_{f_{t}}(\mathbf{z}_{t})$ corresponds to the optimal policy against the true follower type and context at time $t$. So equivalently, we want to lower bound:
$\sum_{t\in T_{i}}u_{h^{*},t}-u_{t}=\sum_{t\in T_{i}}(u_{t}^{*}-u_{t})-(u_{t}^{*}-u_{h^{*},t}).$
Let $r_{i}$ be the number of timesteps in $T_{i}$, where $f_{A,i}$ appears in the example. Then:
$u_{t}^{*}-u_{h^{*},t}=\begin{cases}k-r\quad\text{if}\;r\geq k/2\\ r\hskip 28.45274pt\text{if}\;r<k/2\end{cases}$
Therefore, we get that $k/2-\sum_{t\in T_{i}}u_{t}^{*}-u_{h^{*},t}=|r-k/2|$. By taking an expectation over the randomness in the $f_{t}$’s, we get that:
$k/2-\mathbb{E}\left[\sum_{t\in T_{i}}u_{t}^{*}-u_{h^{*},t}\right]=\mathbb{E}\left[|r-k/2|\right]\geq\sqrt{k/8},$
where the last inequality holds by Khinchine’s inequality (see Lemma A.9 in [14]). Now it suffices to sum over all $i\in[d]$ to get:
$\frac{dk}{2}-\mathbb{E}\left[\sum_{t\in[T]}u_{t}^{*}-u_{h^{*},t}\right]=\sum_{i\in[d]}\left(\frac{k}{2}-\mathbb{E}\left[\sum_{t\in T_{i}}u_{t}^{*}-u_{h^{*},t}\right]\right)\geq d\sqrt{k/8}$
Finally, notice that $dk/2=T/2=\sum_{t\in[T]}(u_{t}^{*}-u_{t})$ for any algorithm, which concludes the proof. ∎
Appendix B Details and Proofs from Section 4
See 4.7
Proof.
We have that:
$\mathbb{P}_{S_{m}}\left[u_{\mathscr{D}}(\pi_{h^{*}})-u_{\mathscr{D}}(\pi_{\text{ERM}})\geq\epsilon\right]$
$\leq$ $\mathbb{P}_{S_{m}}\left[\mathop{\mathbb{E}}_{\mathbf{z}\sim\mathscr{D}}\left[u(\mathbf{z},\pi_{h^{*}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{h^{*}}(\mathbf{z})))-u(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z})))\right]>\epsilon\right]$
By the boundedness of the utility function we have that the expectation above lies within $[-1,1]$. Therefore, we get:
$\leq$ $\mathbb{P}_{S_{m}}\left[\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}\left[u(\mathbf{z},\pi_{h^{*}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{h^{*}}(\mathbf{z})))-u(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z})))>0\right]>\epsilon\right]$
But getting less utility than the optimal predictor in the class, implies that our prediction of the follower’s type is incorrect. So,
$\leq$ $\mathbb{P}_{S_{m}}\left[\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}\left[h_{*}(\mathbf{z})\neq h_{\text{ERM}}(\mathbf{z})\right]>\epsilon\right]$
But from Theorem 6 in [17], we know that for any $\epsilon,\delta>0$, there exists constant $C_{2}$ such that for any $m\geq C_{2}\left(\frac{d_{N}(\pazocal{H}))\ln(K)\ln(\frac{1}{\epsilon})+\ln(\frac{1}{\delta})}{\epsilon}\right)$, we have that for any $h_{\text{ERM}}$ that satisfies equation (3):
$\mathbb{P}_{S_{m}}\left(\text{Err}_{\mathscr{D}}(h_{\text{ERM}})>\text{Err}_{\mathscr{D}}(\pazocal{H})+\epsilon\right)\leq\delta$
Since we are in the realizable setting, we know that $\text{Err}_{\mathscr{D}}(\pazocal{H})=0$. By definition, we have that $\text{Err}_{\mathscr{D}}(h_{\text{ERM}})>\epsilon\iff\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}\left[h_{\text{ERM}}(\mathbf{z})\neq h^{*}(\mathbf{z})\right]>\epsilon$, which concludes our proof. ∎
B.1 Proof of Theorem 4.12
See 4.12
Proof.
Note that the following proof follows parts of Balcan [3], which we reproduce for completeness. By definition, it suffices to show that for any $\epsilon,\delta\in(0,1)$ if we draw a sample $S$ from $\mathscr{D}$ of size:
$m>C\cdot\frac{1}{\epsilon}\Bigr{[}d_{SN,\pazocal{G}}(\pazocal{H})\log_{2}\Bigl{(}\frac{d_{SN,\pazocal{G}}(\pazocal{H})\cdot K}{\epsilon}\Bigl{)}+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}\Bigr{]}$
for some large enough constant $C$, then with probability at least $1-\delta$,
$\mathbb{P}_{S_{m}}\left(u_{\mathscr{D}}(\pi_{\text{S-ERM}})\geq u_{\mathscr{D}}(\pi_{h^{*}})-\epsilon\right)$ (5)
where $\pi_{h^{*}}(\mathbf{z}):=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h^{*}(\mathbf{z})}(\mathbf{z},\mathbf{x}))$. In other words, it suffices to prove that all the policies in $\Pi_{\pazocal{H},\pazocal{G}}$ with $u_{\mathscr{D}}(\pi_{h})<u_{\mathscr{D}}(\pi_{h^{*}})-\epsilon$ have low utility in the training sample too, that is, $u_{S}(\pi_{h})\leq u_{S}(\pi_{h^{*}})$.
It suffices then to bound the following bad event:
$B:\exists\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}\text{ with }u_{S}(\pi_{h})=u_{S}(\pi_{h^{*}})\text{ but }u_{D}(\pi_{h})<u_{D}(\pi_{h^{*}})-\epsilon.$
Let us denote the training sample by $S=\{\mathbf{z}_{1},\mathbf{z}_{2},\ldots,\mathbf{z}_{m}\}$. Now suppose $S^{\prime}=\{\mathbf{z}^{\prime}_{1},\mathbf{z}^{\prime}_{2},\ldots,\mathbf{z}^{\prime}_{m}\}$ is another sample drawn i.i.d. from $\mathscr{D}$ (a ghost sample). We define the event:
$B^{\prime}:\exists\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}\text{ with }u_{S}(\pi_{h})=u_{S}(\pi_{h^{*}})\text{ but }u_{S^{\prime}}(\pi_{h})<u_{S^{\prime}}(\pi_{h^{*}})-\epsilon/2.$
Lemma B.1.
If $m>\frac{8}{\epsilon}$, then $\mathbb{P}[B]\leq 2\mathbb{P}[B^{\prime}]$.
Proof.
Suppose that we know that there exists an $\pi_{h}$ that is consistent with $S$ but $L_{\mathscr{D}}(\pi_{h}):=u_{D}(\pi_{h^{*}})-u_{D}(\pi_{h})>\epsilon$. Since $S^{\prime}$ is drawn i.i.d. from $\mathscr{D}$, $\mathbb{E}[L_{S^{\prime}}(\pi_{h})]\geq\epsilon m$. By a Chernoff bound, $\mathbb{P}[L_{S^{\prime}}(\pi_{h})\leq\epsilon m/2]<e^{-m\epsilon/8}\leq\frac{1}{2}$ for $m>\frac{8}{\epsilon}$. This implies that:
$\frac{\mathbb{P}[B^{\prime}]}{\mathbb{P}[B]}\geq\frac{\mathbb{P}[B^{\prime}\cap B]}{\mathbb{P}[B]}=\mathbb{P}[B^{\prime}|B]\geq\frac{1}{2},$
so $\mathbb{P}[B]\leq 2\mathbb{P}[B^{\prime}]$. ∎
Thus, it suffices to bound $\mathbb{P}[B^{\prime}]$ (this probability is over the choices of $S$ and $S^{\prime}$). Given two samples $S$ and $S^{\prime}$, consider the random process SwapR:
  • For $i$ from $1$ to $m$, flip a fair coin.
  • If heads, swap $\mathbf{z}_{i}$ and $\mathbf{z}^{\prime}_{i}$, else do nothing.
Let $T$ and $T^{\prime}$ denote the resulting collections.
Lemma B.2.
Suppose $S$ and $S^{\prime}$ are chosen according to $\mathscr{D}$ and SwapR is applied. Then $T$ and $T^{\prime}$ are identically distributed to $S$ and $S^{\prime}$.
Now, we define the event:
$B^{\prime\prime}:\exists\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}\text{ with }L_{T}(\pi_{h})=0\text{ but }L_{T^{\prime}}(\pi_{h})>\epsilon/2.$
By Claim B.2, $\mathbb{P}[B^{\prime\prime}]=\mathbb{P}[B^{\prime}]$ so, it suffices to upper bound $\mathbb{P}[B^{\prime\prime}]$. Notice that the first probability is over the choice of $S$, $S^{\prime}$, and the randomness of SwapR, while the second is over the choice of $S$ and $S^{\prime}$.
Lemma B.3.
Fix $h\in\pazocal{H}$. Then $\mathbb{P}[L_{T}(\pi_{h})=0\land L_{T^{\prime}}(\pi_{h})>\epsilon/2|S,S^{\prime}]\leq 2^{-\epsilon m/2}.$
Proof.
Consider:
$\pi_{h}(\mathbf{z}_{1}),\pi_{h}(\mathbf{z}_{2}),\ldots,\pi_{h}(\mathbf{z}_{m})$
$\pi_{h}(\mathbf{z}^{\prime}_{1}),\pi_{h}(\mathbf{z}^{\prime}_{2}),\ldots,\pi_{h}(\mathbf{z}^{\prime}_{m}).$
Notice the following: (1.) If there is a column, with both predictions wrong, then $L_{T}(\pi_{h})=0$ can never happen, so we are done. (2.) Similarly, if more than $(1-\epsilon/2)m$ columns have both predictions correct, $L_{T^{\prime}}(\pi_{h})>\epsilon/2$ cannot happen.
So, there must be at least $r\geq\epsilon m/2$ columns with one correct and one wrong prediction. In this case, SwapR must always ensure that the correct prediction remains on top, which happens with probability at most $2^{-r}\leq 2^{-\epsilon m/2}$. ∎
We are now ready to upper bound $\mathbb{P}[B^{\prime\prime}]$.
Lemma B.4.
$\mathbb{P}[B^{\prime\prime}]\leq\Pi_{\pazocal{H},\pazocal{G}}[2m]\cdot 2^{-\epsilon m/2}$.
Proof.
By definition, we have that:
$\mathbb{P}[B^{\prime\prime}]$ $=\mathop{\mathbb{E}}_{S,S^{\prime}}[\mathbb{P}_{\texttt{SwapR}}[\exists\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}},\;L_{T}(h)=0\land L_{T^{\prime}}(h)\geq m\epsilon/2|S,S^{\prime}]]$
$=\mathop{\mathbb{E}}_{S,S^{\prime}}[\mathbb{P}_{\texttt{SwapR}}[\exists\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}[S\cup S^{\prime}],\;L_{T}(h)=0\land L_{T^{\prime}}(h)\geq m\epsilon/2|S,S^{\prime}]].$
Using the union bound, we get that:
$\mathbb{P}[B^{\prime\prime}]$ $\leq\mathop{\mathbb{E}}_{S,S^{\prime}}\Bigr{[}\sum_{\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}[S\cup S^{\prime}]}\mathbb{P}_{\texttt{SwapR}}[\;L_{T}(\pi_{h})=0\land L_{T^{\prime}}(\pi_{h})\geq m\epsilon/2|S,S^{\prime}]\Bigr{]}$
$\leq\Pi_{\pazocal{H},\pazocal{G}}[2m]2^{-\epsilon m/2}$
Combining all these, we get that $\mathbb{P}[B]\leq\delta$, whenever $2\Pi_{\pazocal{H},\pazocal{G}}[2m]2^{-\epsilon m/2}\leq\delta$, which proves that:
$m>\frac{2}{\epsilon}\Bigr{[}\log_{2}(2\Pi_{\pazocal{H},\pazocal{G}}[2m])+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}\Bigr{]}$
It remains to show that $\Pi_{\pazocal{H},\pazocal{G}}\leq d_{SN,\pazocal{G}}(\pazocal{H})\ln(K)$. We prove this using the Lemma 4.13, which states that $\Pi_{\pazocal{H},\pazocal{G}}[m]\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$. So we get that:
$m$ $>\frac{2}{\epsilon}\Bigr{[}\log_{2}(2\cdot(2m)^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})})+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}\Bigr{]}$
$=\frac{2}{\epsilon}\Bigr{[}1+d_{SN,\pazocal{G}}(\pazocal{H})\log_{2}(m)+d_{SN,\pazocal{G}}(\pazocal{H})+2d_{SN,\pazocal{G}}(\pazocal{H})\log(K)+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}\Bigr{]}$
Rearranging, we obtain the requirement that:
$m>\frac{2d_{SN,\pazocal{G}}(\pazocal{H})}{\epsilon}\Bigr{[}\log_{2}(m)\Bigr{]}+\frac{2}{\epsilon}\Bigr{[}2d_{SN,\pazocal{G}}(\pazocal{H})\log(K)+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}+d_{SN,\pazocal{G}}(\pazocal{H})+1\Bigr{]}$
Using lemma B.5, a sufficient condition for the above to hold is that:
$m>\frac{8d_{SN,\pazocal{G}}(\pazocal{H})}{\epsilon}\Bigr{[}\log_{2}\Bigl{(}\frac{4d_{SN,\pazocal{G}}(\pazocal{H})}{\epsilon}\Bigl{)}\Bigr{]}+\frac{4}{\epsilon}\Bigr{[}2d_{SN,\pazocal{G}}(\pazocal{H})\log(K)+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}+d_{SN,\pazocal{G}}(\pazocal{H})+1\Bigr{]}$
And a sufficient condition for the above is:
$m>\frac{8d_{SN,\pazocal{G}}(\pazocal{H})}{\epsilon}\Bigr{[}\log_{2}\Bigl{(}\frac{4d_{SN,\pazocal{G}}(\pazocal{H})\cdot K}{\epsilon}\Bigl{)}\Bigr{]}+\frac{4}{\epsilon}\Bigr{[}\log\Bigl{(}\frac{1}{\delta}\Bigl{)}+d_{SN,\pazocal{G}}(\pazocal{H})+1\Bigr{]}$
The above holds if there exists a constant $C$, large enough such that:
$m>C\cdot\frac{1}{\epsilon}\Bigr{[}d_{SN,\pazocal{G}}(\pazocal{H})\log_{2}\Bigl{(}\frac{d_{SN,\pazocal{G}}(\pazocal{H})\cdot K}{\epsilon}\Bigl{)}+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}\Bigr{]}.$
Lemma B.5 (Shalev-Shwartz and Ben-David [34], Lemma A.2).
Let $a\geq 1$ and $b>0$. Then $x\geq 4a\log(2a)+2b$ implies that $x\geq a\log(x)+b$.
B.2 Proof of Lemma 4.13
See 4.13
Proof.
We will show that for $|Z|=m$:
$|\Pi_{\pazocal{H},\pazocal{G}}|$ |$\leq|\{B\subseteq\pazocal{Z}:\pazocal{H}\;\text{SN-shatters}\;B\}|\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$ (6)
$\leq\sum_{i=1}^{d_{SN,\pazocal{G}}(\pazocal{H})}\binom{m}{i}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$ (7)
$\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$ (8)
Notice that equation (7) holds by enumerating all possible subsets of $\pazocal{Z}$ that can be SN-shattered by $\pazocal{H}$ up to $d_{SN,\pazocal{G}}(\pazocal{H})$, which is the largest size subset that $\pazocal{H}$ can shatter. So it suffices to prove equation (6). We proceed by induction on the size of $|\pazocal{Z}|$.
Base Case: For $|\pazocal{Z}|=1$, we have that $|\Pi_{\pazocal{H},\pazocal{G}}|\leq K$ so (6) holds.
Inductive Hypothesis: For $|\pazocal{Z}|=m$ assume that (6) holds.
Inductive Step: We will show (6) for $|S|=m+1$. Consider the set $\pazocal{Z}=\{\mathbf{z}_{0},\mathbf{z}_{1},\cdots,\mathbf{z}_{m}\}$ and we define the following notation:
$\Lambda^{(i)}=\{\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}|\pi_{h}(\mathbf{z}_{0})=i\}$
$\Lambda^{(i,j)}=\{\pi_{h}\in\Lambda^{(i)}|\exists\pi_{h}^{\prime}\in\Lambda^{(j)}\;\text{s.t.}\;\pi_{h}(\mathbf{z})=\pi_{h}^{\prime}(\mathbf{z})\;\forall\mathbf{z}\in\pazocal{Z}\setminus\{\mathbf{z}_{0}\}\}$
For $i,j\in[K]$ also denote $\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}$ to be all the policies $\pi_{f}:\pazocal{Z}\setminus\{\mathbf{z}_{0}\}\to[K]$ that can be extended to a policy in $\Pi_{\pazocal{H},\pazocal{G}}$ both by defining $\pi_{f}(\mathbf{z}_{0})=i$ and by defining $\pi_{f}(\mathbf{z}_{0})=j$. Then we claim the following:
$|\Pi_{\pazocal{H},\pazocal{G}}|\leq|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}|\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|,$ (9)
where $|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|$ denotes the restriction of $\Pi_{\pazocal{H},\pazocal{G}}$ to the set $\pazocal{Z}\setminus\{\mathbf{z}_{0}\}$.
(Proof of equation (9)). Consider the $\ell$’th labeling of elements in $\pazocal{Z}\setminus\{\mathbf{z}_{0}\}$ that is consistent with a policy in $\Pi_{\pazocal{H},\pazocal{G}}$. Let $\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}$ denote the policies in $\Pi_{\pazocal{H},\pazocal{G}}$ that are consistent with that labeling. Then we have that:
$|\Pi_{\pazocal{H},\pazocal{G}}|$ $=\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}\mathbbm{1}\{\exists\pi\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\}$
$=\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}1+\left(\sum_{i=1}^{K}\mathbbm{1}\{\exists\pi\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\}\right)-1$
$=|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}\mathbbm{1}\{\exists\pi\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\}-1$
$\leq|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}-1+\sum_{j=1}^{K}\mathbbm{1}\{\exists\pi,\pi^{\prime}\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\wedge\pi^{\prime}(\mathbf{z}_{0})=j\}$
$=|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}\left(\mathbbm{1}\{\exists\pi\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\}-1\right)$
$\hskip 85.00012pt+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}\mathbbm{1}\{\exists\pi,\pi^{\prime}\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\wedge\pi^{\prime}(\mathbf{z}_{0})=j\}$
$\leq|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}\mathbbm{1}\{\exists\pi,\pi^{\prime}\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\wedge\pi^{\prime}(\mathbf{z}_{0})=j\}$
$=|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}|\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|.$ (10)
We proceed with the induction. From the inductive hypothesis we have that:
$|\displaystyle|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|$$\leq|\{B\subseteq\pazocal{Z}\setminus\{\mathbf{z}_{0}\}:\pazocal{H}\;\text{SN-shatters}\;B\}|\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$
$\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})},$ (11)
where the last inequality holds from equation (8). Also, notice that each $\Lambda^{(i,j)}$ SN-shatters $\mathbf{z}_{0}$ by definition. So $\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{x_{0}\}}$ cannot shatter a set larger than $d_{SN,\pazocal{G}}(\pazocal{H})-1$. Therefore, by the inductive hypothesis we get that:
$|\displaystyle|\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{x_{0}\}}|$$\leq|\{B\subseteq\pazocal{Z}\setminus\{\mathbf{z}_{0}\}:\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}\;\text{SN-shatters}\;B\}|\cdot K^{2(d_{SN,\pazocal{G}}(\pazocal{H})-1)}$
$\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})-1}\cdot K^{2(d_{SN,\pazocal{G}}(\pazocal{H})-1)}$ (12)
Combining inequalities (10), (11), and (12), we get that:
$|\Pi_{\pazocal{H},\pazocal{G}}|$ $\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}+\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}m^{d_{SN,\pazocal{G}}(\pazocal{H})-1}\cdot K^{2(d_{SN,\pazocal{G}}(\pazocal{H})-1)}$
$=m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}+m^{d_{SN,\pazocal{G}}(\pazocal{H})-1}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$
$=m^{d_{SN,\pazocal{G}}(\pazocal{H})-1}K^{2d_{SN,\pazocal{G}}(\pazocal{H})}(m+1)$
$\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$

### v2

@@
 Abstract
-We study structured Stackelberg games, in which both players (the leader and the follower) observe information about the state of the world at time of play.
-Importantly, this information may contain information about the follower, which the leader may use when deciding her strategy.
-Under this setting, we show that no-regret learning is possible if and only if the set of mappings from contexts to follower types that the leader uses to learn is not “too complex”.
-Specifically, we find that standard learning theoretic measures of complexity do not characterize learnability in our setting and we give a new dimension which does, which we term the Stackelberg-Littlestone dimension.
-In the distributional setting, we give analogous results by showing that standard complexity measures do not characterize the sample complexity of learning, but a new dimension called the Stackelberg-Natarajan dimension does.
-We then show that an appropriate empirical risk minimization procedure achieves the corresponding sample complexity.
+We study structured Stackelberg games, in which both players (the leader and the follower) observe contextual information about the state of the world at time of play.
+The leader plays against one of a finite number of followers, but the follower’s type is not known until after the game has ended.
+Importantly, we assume a fixed relationship between the contextual information and the follower’s type, thereby allowing the leader to leverage this additional structure when deciding her strategy.
+Under this setting, we find that standard learning theoretic measures of complexity do not characterize the difficulty of the leader’s learning task.
+Instead, we introduce a new notion of dimension, the Stackelberg-Littlestone dimension, which we show characterizes the instance-optimal regret of the leader in the online setting.
+Based on this, we also provide a provably optimal learning algorithm.
+We extend our results to the distributional setting, where we use two new notions of dimension, the $\gamma$-Stackelberg-Natarajan dimension and $\gamma$-Stackelberg-Graph dimension.
+We prove that these control the sample complexity lower and upper bounds respectively, and we design a simple, improper algorithm that achieves the upper bound.
 1 Introduction
@@
 In a Stackelberg game, the principal (also called the leader) commits to a possibly randomized strategy, before the agent (also called the follower) observes this strategy and best-responds to it.
-Due to their effectiveness in modeling such leader-follower interactions, Stackelberg games have had significant real-world impact over the past decade, particularly in security domains [35].
-For example, in protecting endangered species [38], a wildlife agency determines ranger patrol locations before poachers select targets.
-Coast guards decide on shoreline patrol routes in advance of illegal fishers’ actions [12] and, similarly, airport security [32] chooses checkpoint screening strategies before smugglers decide which entry points to use.
-In many of these real-world settings, the state of the world at time of play can greatly influence the utilities of the players and subsequently the outcome of the game.
-For example, defending some areas of the airport might be deemed of higher- or lower-importance based on the schedule of flights and the number of passengers on board.
-Similarly, weather patterns might make it harder to reach animals in certain parts of the park, while breeding seasons might make others more vulnerable thereby affecting their risk levels as targets.
-When contextual information becomes available over time, using it could help the leader design more effective strategies against the follower.
-This raises a natural question: given that the state of the world might change in current and future rounds of play, how should the leader in a Stackelberg game design a policy that adapts to new contexts coming in over time?
-One way is to study repeated Stackelberg games with side information [24], where at each interaction both the leader and the follower get to observe a context that affects their utilities.
-As in Balcan et al.
-[5], in this setting each follower that the leader faces might have different preferences and objectives, i.e.
-there might be different follower types.
+Stackelberg games have had significant real-world impact over the past decade, particularly in security domains [37].
+For example, when protecting endangered species [39], a wildlife agency determines ranger patrol locations before poachers select targets.
+Coast guards decide on shoreline patrol routes in advance of illegal fishers’ actions [17] and, airport security [33] chooses checkpoint screening strategies before smugglers decide which entry points to use.
+In many of these settings, the state of the world at time of play can greatly influence the utilities of the players and subsequently the outcome of the game.
+For example, patrolling some areas of the airport might be deemed of higher- or lower-importance based on the schedule of flights and the number of passengers on board.
+Similarly, weather patterns may make it harder to reach animals in certain parts of the park, while migration patterns might make others more vulnerable.
+When contextual information becomes available over time, leveraging it will generally help the leader deploy more effective strategies.
+This raises a natural question: given that the state of the world might evolve over time, how should the leader in a Stackelberg game design a strategy that adapts to new information coming in over time?
+Previous work [27, 13] has formalized such settings as repeated Stackelberg games with side information, where at each interaction both players observe contextual information that affects their utilities.
+In this setting, followers in different rounds may have different types that determine their payoff for each outcome of the game.
 Harris et al.
-[24] study this problem and show that as long as either the followers or the contexts are independently drawn from a fixed and unknown distribution, no-Stackelberg-regret learning is possible.
+[27] show that as long as either the sequence of followers or the sequence of contexts are drawn i.i.d.
+from a fixed distribution, no-regret learning is possible.
 That is, the leader is guaranteed to achieve nearly as much utility as the best fixed policy in hindsight.
-In the fully adversarial regime, however, regret can grow linearly with the time horizon.
+In the fully adversarial regime, however, they show that regret can grow linearly with the time horizon.
 Interestingly, this lower bound follows from the adversary’s ability to encode an arbitrarily complex relationship between the sequence of contexts and the follower types.
-In this work, we take a more granular view of this impossibility result and ask whether introducing additional structure between the contextual information and the follower types can enhance the leader’s ability to maximize returns, even when both are chosen adversarially to the leader.
-Namely, there could be correlation between the two, making the context predictive of the type of follower about to appear next.
-In wildlife protection for example, specific tire tracks might signal the presence of a particular poaching group and market demand trends might be indicative of which animals’ poachers are now more likely to appear.
-Similarly, in airport security, current affairs often raise or lower the risk of an attack on certain international flights.
-Algorithms proposed by Harris et al.
-[24] and later work [9] are designed to handle changes in utility caused by the context, but crucially they assume no structure in how contexts relate to follower types.
-If, however, there is a fixed and unknown mapping between contexts and follower types, can the leader learn to exploit it to achieve sublinear regret?
-In this case, the leader m ust adapt her policy based on which follower types are more likely to appear, given the context and past observations.
-This poses an additional and possibly more complex learning task for the leader, which we study in this work.
+Naturally, there are many cases in which the relationship between the contextual information and the follower about to appear next is not arbitrarily complex, but instead exhibits some structure.
+In fact, in many settings the contextual information can be predictive of the follower’s type.
+For example, in wildlife protection, tire tracks might signal the presence of poachers in a particular area of the park, and market demand trends might be indicative of which animals’ poachers are now more likely to appear.
+Similarly in airport and border security, current affairs might make it more or less likely that certain smuggler groups are active.
+Motivated by this observation, we study how the leader can learn effectively when there is additional structure between the contextual information and the follower types, thereby circumventing the previous impossibility result.
 1.1 Our Contributions
-In this paper, we revisit the framework of Harris et al.
+We revisit the framework of Harris et al.
 [27] and take a learning theoretic approach to analyzing the leader’s task in predicting the follower type from contexts.
-Importantly, we find that existing notions of complexity do not capture the difficulty of learning in this setting—we obtain (tight) sample complexity and Stackelberg regret guarantees that are parameterized by new notions of complexity and that directly exploit the structure of Stackelberg games.
-Concretely, our main contributions in this work are as follows:
-  1. 1.
-    We formalize online learning in Stackelberg games with side information, under the structural assumption that there exists a fixed, unknown mapping between contexts and follower types.
-In particular, we study how the leader, given access to a set of possible such mappings, can leverage it to learn an optimal policy.
-In particular, we study this problem under two different settings: (1) the realizable setting, in which there exists an optimal predictor of the followers’ types that the leader competes with; and (2) the agnostic setting, in which no such assumption is made.
-  2. 2.
-    We circumvent the impossibility result shown in Harris et al.
-[24] by providing algorithms with no-regret guarantees under the above structural assumption.
-We find that our problem is closely related to the problem of multiclass prediction, yet, interestingly, we show that our setting can be easier to learn in.
-In particular, we show that the (multiclass) Littlestone dimension, which quantifies the difficulty of learning in online multiclass prediction, can give loose mistake bounds in our setting.
-Motivated by this observation, we introduce a new notion of complexity, the Stackelberg-Littlestone dimension.
-We show that it characterizes learnability in our setting, and prove it to be at most as large as the multiclass Littlestone dimension.
-In fact, we give an example in which the latter is infinite, yet the former is zero.
-We also provide no-regret algorithms both for both the realizable and agnostic settings.
-  3. 3.
-    We introduce and formalize the study of distributional learning in Stackelberg games with side information.
-When the leader has access to a set of (context, follower type) examples, we show generalization with PAC-style bounds.
-Here too, we find that existing measures which characterize the complexity of multiclass prediction, such as the Natarajan dimension, do not characterize the sample complexity of our problem.
-Analogous to the online setting, we introduce the notion of Stackelberg-Natarajan dimension, with which we achieve an upper bound that is tight up to logarithmic factors.
+Importantly, we find that existing notions of complexity do not capture the difficulty of learning in this setting—we obtain (tight) Stackelberg regret and sample complexity guarantees that are parameterized by new complexity measures and that directly exploit the structure of Stackelberg games.
+Concretely, our main contributions in this work are as follows.
+Instance-optimal Stackelberg regret bounds.
+We initiate the study of online learning in Stackelberg games with side information, under the structural assumption that there exists a fixed, unknown mapping between contexts and follower types.
+In particular, we study how the leader, given access to a predefined hypothesis class, can leverage it to learn.
+We study the realizable setting, under which the class includes an optimal predictor of the followers’ types.
+Then the leader’s task is to learn an optimal policy on the fly while achieving as high cumulative utility as possible.
+Naturally, the difficulty of the leader’s learning task scales with the complexity of the hypothesis class.
+In the learning theory literature, the (multiclass) Littlestone dimension [30, 22] is used to characterize the complexity of a class in settings with a finite set of labels (which in structured Stackelberg games corresponds to the follower types).
+Interestingly, we prove that the Littlestone dimension can give arbitrarily loose regret bounds (Theorem˜3.5).
+Our construction leverages the fact that the Littlestone dimension is completely blind to the utility space defined by the Stackelberg game, which ultimately determines the performance of the leader.
+Based on these insights, we introduce a new notion of complexity, the Stackelberg-Littlestone dimension (Definitions˜3.6 and 3.7, which directly exploits the interaction between the hypothesis class and the utility space of the game.
+We base our definition on shattered trees, which are the workhorse of classic online learning.
+The challenge in our setting relies on having a finite set of follower types, yet a continuous utility function, which requires a much more nuanced approach to constructing the trees.
+We show that the Stackelberg-Littlestone dimension exactly characterizes the difficulty of learning in our setting (Theorem˜3.8), and in particular gives instance-optimal regret bounds.
+That is, for any hypothesis class and any Stackelberg game, the Stackelberg-Littlestone dimension identifies exactly the optimal regret.
+The proof of our upper bound Theorem˜3.9 is constructive (Algorithm˜1), which yields a complete characterization of learnability in online structured Stackelberg games.
+PAC learning guarantees.
+We introduce and formalize the study of distributional learning in structured Stackelberg games (Section˜4).
+The leader has access to a set of (context, follower type) examples, where the contexts are drawn from some fixed, unknown distribution and the follower types are determined by a true hypothesis in a given hypothesis class.
+Upon the draw of a new context, the leader’s goal is to deploy a strategy such as to maximize her expected payoff.
+We show generalization guarantees with Probably Approximately Correct (PAC)-style bounds.
+We introduce the notion of $\gamma$-Stackelberg-Natarajan dimension (Definitions˜4.9 and 4.10) and $\gamma$-Stackelberg-Graph dimension (Definitions˜4.12 and 4.13), which respectively control the lower and upper bounds on the sample complexity (Theorems˜4.11 and 4.14).
+Our definitions are inspired by analogous notions in classification and regression [31, 3, 21], yet we give a tailored approach to what constitutes a “mistake”.
+This leverages the discreteness of the set of follower types to determine how much hypotheses “differ” in a minmax sense.
+We give a simple, improper algorithm (Algorithm˜2) that with high probability achieves (almost) optimal expected utility with as many samples as our sample complexity upper bound.
+We also analyze a simpler, context-free setting, in which the training set is only over the follower types.
+Using tools from data-driven algorithm design, we show strong generalization guarantees Theorem˜4.5.
+Optimal strategy misprediction bounds.
+In the Appendix, we study structured Stackelberg games under a relaxed benchmark that seeks to minimize the total rounds in which the leader does not play the optimal strategy against the realized follower type.
+We give results for both the online (Appendix˜B) and distributional setting (Section˜B.2), and in both cases, we show stronger guarantees than those achieved by applying existing approaches from learning theory.
+These results are both conceptually and quantitatively superseded by the ones we present in the main body and we include them for completeness and further insight.
+We summarize our results in  Table˜1.
+Setting: Online Distributional
+Dimension Definitions˜3.6 and 3.7 Definitions˜4.10 and 4.13
+Lower Bound Theorem˜3.8 Theorem˜4.11
+Upper Bound Theorem˜3.9 Theorem˜4.14
+Algorithm Algorithm˜1 Algorithm˜2
+Table 1: Summary of our main results.
 1.2 Related Work
-Stackelberg games.
-In the one-shot version of a Stackelberg game, Conitzer and Sandholm [16] prove that the problem of computing a Stackelberg equilibrium is NP-hard and they give an algorithm for maximizing the leader’s payoff.
-Other works focus on relaxing the assumption that all of the game’s parameters are known (e.g.
-utilities, number of actions) and aim to learn them instead [2, 31].
-Our work is closely related to Bayesian Stackelberg games, in which the leader is playing against a single unknown follower type that is drawn from a given set of follower types [16].
-This has been studied both in one-shot games, in which the leader must find a single optimal-mixed strategy [16, 30, 37] and in the repeated setting, in which the leader faces the same follower type sequentially and must learn from previous interactions [28, 26].
 Learning in Stackelberg games.
@@
 [4] relax some structural assumptions from previous work, including tie-breaking rules for the follower, and revise worst-case sample complexity results of learning the leader’s optimal mixed strategy.
-Recent research has explored many domains through the framework of Stackelberg games, including strategic classification, where classifiers account for individuals who strategically modify their features [23]; Bayesian persuasion, which models how a leader strategically reveals information to influence a follower’s decisions [25]; interactions with calibrated followers, who update their strategies based on past observations rather than making immediate best responses [21]; and settings with non-myopic followers, who optimize their actions over multiple time steps rather than reacting myopically [20, 15].
-Part of our work focuses on learning in online Stackelberg games.
+Our main focus is on online learning in Stackelberg games.
 Online Stackelberg games with multiple follower types were introduced in Balcan et al.
 [7].
-The authors show that when a series of unknown attacker types is chosen adversarially, playing the well-known Hedge algorithm  [19] over a special set of mixed strategies yields a no-regret algorithm.
+The authors show that when a series of unknown attacker types is chosen adversarially, playing the well-known Hedge algorithm  [25] over a carefully selected set of mixed strategies yields a no-regret algorithm.
 Our departing point is the work of Harris et al.
@@
 The authors show that when either the sequence of follower types or the sequence of contexts are drawn independently from some fixed, unknown distribution, no-regret learning is indeed possible.
-However, when both of them are chosen adversarially, regret can scale linearly in the worst case, as the problem can be reduced from online learning linear thresholds.
-On the other hand, we make a structural assumption about the relationship between contexts and follower types: namely that it belongs to a class of functions that the leader has access to.
-With this, we show that learning is possible, and our guarantees depend on an appropriate measure of complexity of this function class.
+Follow-up work [13] obtains better regret guarantees under bandit feedback.
+However, when both sequences are chosen adversarially, regret can scale linearly in the worst case, as the problem can be reduced from online learning linear thresholds.
+In contrast to this line of work, we circumvent this impossibility result by making an additional structural assumption about the relationship between contexts and follower types—namely that it belongs to a class of functions that the leader has access to.
+Under this assumption, we show that online learning is indeed possible, and our guarantees depend on an appropriate measure of complexity of this function class.
 Multiclass learnability.
@@
 Multiclass prediction has been studied extensively both in the ditributional ([31, 22, 18]) and in the online setting ([22, 26]).
-Indeed we adopt several tools from online multiclass prediction in our work.
-Interestingly, however, we show that learning in Stackelberg games with contexts is at most as hard as multiclass prediction, and in some cases, can be much easier.
-For example, in the online setting, we give an example in which the leader can learn to play optimally, under a hypothesis class that has infinite Littlestone dimension.
+Interestingly, we show that learning in Stackelberg games with contexts is at most as hard as multiclass prediction, and in some cases, can be much easier.
+For example, we give a problem instance the leader can learn to play optimally under a hypothesis class that has infinite Littlestone dimension.
+Online Learning beyond Multiclass Prediction.
+There has been tremendous interest in understanding online learnability in many settings of interest.
+Notably, in online realizable regression Daskalakis and Golowich [23] designed a proper algorithm that achieved near-optimal cumulative absolute loss with respect to the sequential fat-shattering dimension of the hypothesis class.
+Our work is more closely related to Attias et al.
+[3], who introduce the online (resp.
+$\gamma$-Graph) dimension of a function class and show that it tightly characterizes online regression (resp.
+distributional learning) in the realizable setting up to a factor of 2.
+While our setting also considers utility functions with continuous image sets, the dimensions we introduce explicitly exploit the discrete structure of the Stackelberg game that arise from the finite set of follower types.
+Another recent line of work that captures incentives in online learning is online strategic classification.
+In this setting, agents can modify their features according to a directed graph over the feature space, before providing these features to the learner.
+The authors use a new combinatorial measure, the Strategic Littlestone Dimension, to achieve instance optimal mistake bounds.
+While seemingly similar, this measure of complexity is fundamentally distinct from the one we use to derive our results.
+Learning piecewise Lipschitz functions.
+Given the piecewise linear structure of the leader’s utility function, our work also connects to the line of work on learning functions that are piecewise Lipschitz [8, 36, 10, 11, 9].
+We draw from this literature in part of our study that explores a simpler distributional setting without contexts.
+The well-structured payoff function of the leader against a fixed follower type allows us to adopt tools from Balcan et al.
+[9] and get better generalization guarantees compared to prior work that studied a similar setting [29].
 2 Problem Setting
 Notation.
-We use $[N]:=\{1,2,\ldots,N\}$ to denote the set of natural numbers up to and including $N$. We use $\Delta(\pazocal{A})$ to denote the probability simplex over a (finite) set $\pazocal{A}$.
-Stackelberg games with side information.
-Consider a game between a leader and a follower.
+We use $[N]:=\{1,2,\ldots,N\}$ to denote the set of natural numbers up to and including $N$.
+We use $\Delta(\pazocal{A})$ to denote the probability simplex over a (finite) set $\pazocal{A}$.
+For a given string we use inclusive slice indexing such that $x[i:j]$ are the elements of string $x$ from position $i$ to $j$.
+Dropping either $i$ or $j$ means that the slice extends from the beginning of the string or to the end, respectively.
+We use negative indexing to denote the elements of a string starting at the end i.e., $x[-1]$ is the last element of $x$.
+We consider a Stackelberg game between a leader and a follower.
 Before the game starts, some contextual information $\mathbf{z}\in\pazocal{Z}\subseteq\mathbb{R}^{d}$ is revealed to both players.
 The leader plays first and commits to a (possibly mixed) strategy $\mathbf{x}\in\Delta(\pazocal{A})$, where $\pazocal{A}$ is the set of finite actions available to the leader.
-The follower best responds to the leader’s strategy, $\mathbf{x}$, and the context, $\mathbf{z}$, by playing some action $a_{f}\in\pazocal{A}_{f}$, where $\pazocal{A}_{f}$ is the finite set of follower actions.
-[^2] Note that without loss of generality the follower’s best response will always be a pure strategy.
-We define the follower’s best response as:
+The follower best-responds to the leader’s strategy and the context by playing some action $a_{f}\in\pazocal{A}_{f}$, where $\pazocal{A}_{f}$ is the finite set of follower actions.
+[^2] Without loss of generality the follower’s best-response will always be a pure strategy.
+We define the follower’s best-response as
 $b_{f}(\mathbf{z},\mathbf{x})\in\arg\max_{a_{f}\in\pazocal{A}_{f}}\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u_{f}(\mathbf{z},a_{l},a_{f}).$ (1)
-Here, $u_{f}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$ is the utility function of follower $f$. We assume that the follower breaks ties in favor of the leader.
-We allow the follower to be one of $K\in\mathbb{N}_{\geq 1}$ follower types in the set $\mathscr{K}=\{f^{(1)},\dots,f^{(K)}\}$.
+Here, $u_{f}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$ is the utility function of follower $f$.
+In case of ties, we assume that there is a fixed and known ordering over follower actions that determines the follower’s tie-breaking rule.
+We allow the follower to be one of $K\in\mathbb{N}_{\geq 1}$ follower types in the set $\{f^{(1)},\dots,f^{(K)}\}$.
 The type of follower characterizes their utility function, $u_{f^{(i)}}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$.
-We assume that the leader knows the set of possible follower types a priori, but not the realized type.
-The leader aims to maximize her own utility, $u:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$.
-By slight abuse of notation, we sometimes use the shorthand
-$u(\mathbf{z},\mathbf{x},a_{f})=\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u(\mathbf{z},a_{l},a_{f}).$
-We denote a Stackelberg game with side information as a tuple, $\pazocal{G}=(\pazocal{A},\pazocal{A}_{f},\pazocal{Z},u,u_{f^{(1)}},\ldots,u_{f^{(K)}})$, consisting of the leader’s action set, the follower’s action set, the set of possible contexts, as well as the leader’s and followers’ utility functions.
+We assume that the leader knows the set of possible follower types a priori, but not the realized type and aims to maximize her own utility, $u:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$.
+By slight abuse of notation, we sometimes use the shorthand $u(\mathbf{z},\mathbf{x},a_{f})=\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u(\mathbf{z},a_{l},a_{f}).$
+We will denote a Stackelberg game as a tuple, $\pazocal{G}=\left(\pazocal{A},\pazocal{A}_{f},\pazocal{Z},u,u_{f^{(1)}},\ldots,u_{f^{(K)}}\right)$, consisting of the leader action set, the follower action set, the set of possible contexts, as well as the leader and followers’ utility functions.
 Remark 2.1.
-For the remainder of the paper and without loss of generality, we will consider $\mathscr{K}$ to be a set of distinct follower types.
+For the remainder of the paper and without loss of generality, we will consider $[K]$ to be a set of distinct follower types.
 We call any two follower types distinct, if the optimal leader’s strategy against them is not the same for every context.
 3 Online Learning in Structured Stackelberg Games
-Consider an online Stackelberg game in which the leader is faced by a sequence of followers.
+In the online setting, the leader is faced with a sequence of followers.
 In each round $t\in[T]$, a context $\mathbf{z}_{t}\in\pazocal{Z}$ is revealed to both players.
-A follower of type $f_{t}$ arrives, the leader commits to a mixed strategy $\mathbf{x}_{t}\in\Delta(\pazocal{A})$, and the follower best responds to it and the context according to Equation 1.
-As before, we assume that the leader knows the utilities of all follower types, but the type of follower at time $t$ is not revealed to the leader until after the round is over.
-In this section, we make no assumptions about either the contexts or the follower types that appear at each timestep.
-In fact, we handle the case, where they both could be chosen jointly by an adversary.
-3.1 Leader’s task in the realizable setting
-Consider a set $\pazocal{H}$ (also called a hypothesis class) of candidate functions mapping contexts to follower types, which the leader uses to learn.
-More formally, for each $h\in\pazocal{H}$, we let $h:\pazocal{Z}\rightarrow\mathscr{K}$.
-We first consider the realizable setting, where there exists a true mapping $h^{*}\in\pazocal{H}$ such that $h^{*}(\mathbf{z}_{t})=f_{t}$ for all possible $(\mathbf{z}_{t},f_{t})$ pairs.
-We relax this assumption in Section 3.2.
-We use $h^{*}$ as a benchmark to measure the leader’s performance over time, and we quantify this via the notion of contextual Stackelberg regret.
+A follower of type $f_{t}$ arrives, the leader commits to a mixed strategy $\mathbf{x}_{t}\in\Delta(\pazocal{A})$, and the follower best-responds according to Equation˜1.
+We assume that the leader knows the utilities of all follower types, but the type of follower at time $t$ is not revealed to the leader until after the round is over.
+We make no assumptions about how the sequence of contexts is chosen, i.e., we allow them to be chosen adversarially and adaptively to the leader’s sequence of play.
+3.1 Leader’s objective
+Let $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$ be a set of candidate functions (also called a hypothesis class) mapping contexts to follower types, which the leader uses to learn.
+We consider the realizable setting, where there exists a true mapping $h^{*}\in\pazocal{H}$ such that $h^{*}(\mathbf{z}_{t})=f_{t}$ for all possible $(\mathbf{z}_{t},f_{t})$ pairs.
+We use $h^{*}$ as a benchmark for the leader’s performance over time, which we quantify via the notion of contextual Stackelberg regret.
 Definition 3.1 (Contextual Stackelberg Regret).
-Given a sequence of followers $f_{1},\ldots,f_{T}$ and contexts $\mathbf{z}_{1},\dots\mathbf{z}_{T}$, the leader’s contextual Stackelberg regret is given by:
-$R(T):=\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t})).$
-We let $\pi_{h^{*}}(\mathbf{z})=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h^{*}(\mathbf{z})}(\mathbf{z},\mathbf{x}))$ and we assume that there is a fixed and arbitrary tie-breaking rule over the leader strategies.
-We let $\mathbf{x}_{t}$ for $t\in[T]$ be the mixed strategy played by the leader at time $t$.
-The goal of the leader is to learn a policy $\pi:\pazocal{Z}\to\Delta(\pazocal{A})$ that minimizes her regret, that is, it achieves utility comparable to the utility she would get if she could (1) optimally predict the follower type given the context and (2) play an optimal strategy against that follower type.
-In the remainder of this section, we will present an algorithm that meets that benchmark.
-In general, if the regret of an algorithm is sublinear in the time horizon $T$, i.e., $R(T)=o(T)$, we say that the algorithm is a no-regret algorithm.
+Given a sequence of pairs of contexts and followers $\{(\mathbf{z}_{i},f_{i})\}_{t=1}^{T}$ the leader’s regret is
+$R(T):=\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t})),$
+where $\pi_{h^{*}}(\mathbf{z})=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h^{*}(\mathbf{z})}(\mathbf{z},\mathbf{x}))$ and $\mathbf{x}_{t}\in\Delta(\pazocal{A})$ is the leader’s strategy at timestep $t$.
+The goal of the leader is to minimize her regret, that is, to achieve utility comparable to the utility she would have gotten if she could (1) optimally predict the type of follower given the context and (2) play an optimal strategy against that type of follower.
+In Section˜3 we present an algorithm that is optimal with respect to this benchmark.
 Remark 3.2.
-An important part of our analysis is that the leader be able to compute the optimal strategy for a fixed context and follower type.
+An important part of our analysis relies on the leader being able to compute the optimal strategy for any fixed context and follower type.
 However, due to the tie-breaking rules of the follower, the leader might not be able to get an exact solution.
@@
 For ease of notation, in the remainder of this paper, we will assume that the leader can directly optimize over the entire simplex, without that affecting her regret guarantees.
-For completeness, we formally state this result in Appendix A.1.
-Background on Online Multiclass Learnability.
-Our results build off of tools and techniques from multiclass online learning [17].
-In this section, we highlight the concepts and definitions from this area that are relevant to our work.
-We start with the notion of shattering, which captures the following intuition: the more “complex” a set of functions is, the more complex patterns it should be able to fit.
-Definition 3.3 ((Multiclass) $\pazocal{H}$-Shattered Littlestone Tree).
-Let a Littlestone tree, $T$, be a tree whose internal nodes are instances in $\pazocal{Z}$ and where for each node the node-to-child edges correspond to different labels in a finite set $\mathscr{K}$.
-We say that $T$ is shattered by a hypothesis class $\pazocal{H}$ if for every root-to-leaf path in the tree that traverses the nodes $\mathbf{z}_{1},\dots,\mathbf{z}_{n}$, there exists a function $h\in\pazocal{H}$ such that the label of every edge $(\mathbf{z}_{i},\mathbf{z}_{i+1})$ is $h(\mathbf{z}_{i})$.
+For completeness, we formally state this result in Section˜A.1.
+3.2 Littlestone Dimension
+In this section, we highlight the concepts and definitions from multiclass online learning [22] that are relevant to our work.
+We start with the notion of shattering, which captures the following intuition: the richer a set of functions is, the more complex patterns it should be able to fit.
+Definition 3.3 (Multiclass $\pazocal{H}$-Shattered Littlestone Tree).
+[…]
 The notion of multiclass Littlestone dimension reflects how the ability of a function class to shatter deeper trees corresponds to greater complexity in the class.
-[^3] Note that for binary labels, the definition of Littlestone dimension recovers the definition of the well-studied VC-dimension in the case where all nodes on the same depth of a shattered tree are the same.
+[^3] We note that for binary labels, the Littlestone dimension recovers the definition of the well-studied VC-dimension in the case where all nodes at the same depth of a shattered tree are the same.
 This definition is due to Daniely et al.
-[17], and corresponds to a natural generalization of the classic definition of the Littlestone dimension for binary classification [27] to the multiclass setting.
-Definition 3.4 ((Multiclass) Littlestone Dimension).
+[22] and it generalizes the classic definition of the Littlestone dimension for binary classification [30] to the multiclass setting.
+Definition 3.4 (Multiclass Littlestone Dimension).
 The Littlestone dimension of a multiclass hypothesis class $\pazocal{H}$, denoted as $\textnormal{Ldim}(\pazocal{H})$, is the maximal integer $d$, such that there exists a full binary tree of depth $d$ that is shattered by $\pazocal{H}$.
@@
 The finiteness of the Littlestone dimension of a hypothesis class is necessary for the class to be online learnable [30].
-There are several examples of hypothesis classes, which have finite Littlestone dimension and can, therefore, be (online) learned.
-These include the class of conjunctions, the class of decision lists, and the class of linear thresholds with a margin (see Proposition 17 in [1]).
+There are several examples of hypothesis classes that have finite Littlestone dimension and can, therefore, be (online) learned.
+These include the class of conjunctions, the class of decision lists, the class of linear separators with a margin (cf.
+Proposition 17 in [1]), and in general anything learnable in the mistake bound model.
 In Daniely et al.
@@
 Namely, given the context and the predicted follower type, the leader can deterministically find a strategy that maximizes her utility, by solving a sequence of $|\pazocal{A}_{f}|$ linear programs [20].
-However, it is unclear a priori whether learning to optimally predict the follower’s type is also necessary.
-In Theorem 3.5 we show that it is not.
+However, it is a priori unclear whether learning to optimally predict the follower’s type is also necessary.
+In  Theorem˜3.5 we show that the Littlestone dimension does not characterize the complexity of online learning in Stackelberg games with side information.
+The full proof can be found in Section˜A.2.
 Theorem 3.5.
 There exists a Stackelberg game, $\pazocal{G}$, and a hypothesis class, $\pazocal{H}$, such that $\textnormal{Ldim}(\pazocal{H})=\infty$, but the leader’s optimal policy can be determined without learning.
+Proof sketch.
+Our proof captures the following intuition: in a structured Stackelberg game, the follower types can be completely distinct from the perspective of the leader (i.e., any one irreducible to another), and the hypothesis class can be arbitrarily hard to learn.
+Yet, to truly know how difficult the leader’s learning task is, one has to account for how the parameters of the game interact with the hypothesis class.
+Our construction uses distinct follower types and the hypothesis class indicates that there is a true threshold that splits the context domain in two parts, each of which is predictive of a different follower type.
+Learning that true threshold can be arbitrarily hard ($\textnormal{Ldim}(\pazocal{H})=\infty$), yet unnecessary if the two follower types happen to induce the same optimal strategy for the leader near the threshold.
+This result highlights that the Littlestone dimension is completely blind to how the hypothesis class interacts with the utilities of the players, which motivates a new measure of complexity.
+3.3 Stackelberg-Littlestone Dimension
+In this section we provide an instance optimal learner for structured Stackelberg games. We start with the definition of node-weighted shattered trees.
+[…]
+Definition 3.6 (Shattered Stackelberg-Littlestone Tree).
+[…]
+  • (Consistency) $\pazocal{H}$ is consistent with every path in $\pazocal{T}_{d}$ i.e., for any $s\in S_{d}$, there exists $h\in\pazocal{H}$ such that for all $i\in\{\varnothing,1,\dots,d-1\}$ the label of edge $(\mathbf{z}_{s[:i]},\mathbf{z}_{s[:i+1]})$ is $h(\mathbf{z}_{s[:i]})$.
+  • (Max-branch property) $\pazocal{T}_{d}$ satisfies the max-branch property, i.e., every node, $\mathbf{z}_{s}$, has children $\mathbf{z}_{si}$ for all $i\in[K]$ where $\pazocal{H}$ is consistent with path $si$.
+  • (Weights) each node, $\mathbf{z}_{s}\in\mathscr{T}_{d}$ has a weight, $\rho_{s}$ defined as follows: if $\text{len}(s)=d+1$ then $\rho_{s}=0$, else
+    $\rho_{s}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{sj\in S_{d}}\biggl{[}r(\mathbf{z},\mathbf{x},f^{(j)})+\rho_{sj}\biggr{]},$
+    where $r(\mathbf{z},\mathbf{x},j):=u(\mathbf{z},\pi_{h^{*}}(\mathbf{z}),b_{f^{(j)}}(\mathbf{z},\pi_{h^{*}}(\mathbf{z})))-u(\mathbf{z},\mathbf{x},b_{j}(\mathbf{z},\mathbf{x}))$ is the leader’s instantaneous regret, given context $\mathbf{z}$, mixed strategy $\mathbf{x}$, and follower type $j$ [^4] The infimum might not be achieved due to the followers’ tie breaking rule.
+We refer the reader to Section A.1 for further details.
+Definition 3.7 (Stackelberg-Littlestone dimension).
+The Stackelberg-Littlestone dimension of a hypothesis class $\mathscr{H}$ with respect to a Stackelberg game, $\pazocal{G}$, denoted as $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$, is the supremum root node weight value among all finite node-weighted trees that are shattered by hypothesis class $\pazocal{H}$ under Stackelberg game $\pazocal{G}$.
+Note that taking the supremum is necessary since it is possible that no finite-depth tree achieves the optimal value.
+We begin by showing that no deterministic algorithm can achieve regret lower than the Stackelberg-Littlestone dimension.
+Theorem 3.8.
+For any hypothesis class, $\pazocal{H}$, any Stackelberg game $\pazocal{G}$, and any deterministic algorithm, there exists an adversarial sequence of contexts and follower types such that the learner suffers at least $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})-\epsilon$ regret, for any $\epsilon>0$.
+Proof.
+We prove by induction on the depth $d$ of shattered trees that across $d$ timesteps the learner can be forced to incur regret at least as much as the weight of the tree’s root node.
+Since the Stackelberg-Littlestone dimension is defined as the supremum root node weight over all finite-depth trees, proving this suffices to show the desired claim.
+In case the supremum in the definition of $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ is not achieved by any finite tree, there will always be an additive error of $\epsilon$ that will get arbitrarily small as $d\to\infty$.
+(Base Case) We consider a shattered tree of depth 111. Then
+$\rho_{\varnothing}$ $=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})$
+$=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl{[}u(\mathbf{z}_{\varnothing},\pi_{h^{*}}(\mathbf{z}_{\varnothing}),b_{f^{(i)}}(\mathbf{z}_{\varnothing},\pi_{h^{*}}(\mathbf{z}_{\varnothing})))-u(\mathbf{z}_{\varnothing},\mathbf{x},b_{f^{(i)}}(\mathbf{z}_{\varnothing},\mathbf{x}))\biggr{]}.$
+So $\rho_{\varnothing}$ corresponds to the min-max value of the one-shot Stackelberg game and therefore, to a lower bound on the instantaneous regret of the learner.
+(Inductive step) We assume that for any tree of depth $d$ that is shattered by $\pazocal{H}$ under $\pazocal{G}$, the weight of the root node lower bounds the learner’s cumulative regret across $d$ timesteps.
+We now consider a tree $\mathscr{T}_{d+1}$ of depth $d+1$ that is shattered by $\pazocal{H}$ under Stackelberg game $\pazocal{G}$.
+We will prove that for such a tree, the adversary can guarantee regret at least $\rho_{\varnothing}$, where $\mathbf{z}_{\varnothing}$ is the root node of this tree.
+By Definition˜3.6, we know that $\rho_{\varnothing}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl{[}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})+\rho_{i}\biggr{]}$.
+Let $\mathbf{x}_{*}\in\Delta(\pazocal{A})$ and $i_{*}\in[K]$ be the respective arginf and argmax values in the above expression.
+If the learner chooses to play $\mathbf{x}_{*}$ at this round, then by definition the adversary can guarantee instantaneous regret equal to $r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})$ by selecting the true follower type to be $i_{*}$ (which is realizable by $\pazocal{H}$ by definition of a shattered tree).
+If this happens, then by the inductive hypothesis the adversary can guarantee regret $\rho_{i_{*}}$ in the next $d$ timesteps and regret $\rho_{\varnothing}$ overall.
+We will now prove the same even when the learner plays a strategy different from $\mathbf{x}_{*}$.
+Say that there exists $\tilde{\mathbf{x}}\in\Delta(\pazocal{A})$ such that the learner can guarantee instantaneous regret $r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(\tilde{i})})<r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})$, where $\tilde{i}=\textnormal{argmax}_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl{[}r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(i)})+\rho_{i}\biggr{]}$.
+There are two cases.
+(Case 1) Suppose $\tilde{i}=i_{*}$. Then we know that
+$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl{[}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})+\rho_{i}\biggr{]}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\biggl{[}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i_{*})})+\rho_{i_{*}}\biggr{]}.$
+Since $\mathbf{x}_{*}$ achieves the infimum value
+$r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})+\rho_{i_{*}}$ $\leq r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(i_{*})})+\rho_{i_{*}}\;\;\forall\tilde{\mathbf{x}}\in\Delta(\pazocal{A})$
+$\implies r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})$ $\leq r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(i_{*})})\;\;\forall\tilde{\mathbf{x}}\in\Delta(\pazocal{A}),$
+which leads to a contradiction.
+(Case 2) Suppose $\tilde{i}\neq i_{*}$. Then we have that
+$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl{[}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})+\rho_{i})\biggr{]}\leq\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl{[}r(\mathbf{z}_{\varnothing},\mathbf{x}^{\prime},f^{(i)})+\rho_{i})\biggr{]}\;\;\forall\mathbf{x}^{\prime}\in\Delta(\pazocal{A})$
+Taking $\mathbf{x}^{\prime}=\tilde{\mathbf{x}}$, we get
+$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl{[}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})+\rho_{i})\biggr{]}$ $\leq\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl{[}r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(i)})+\rho_{i})\biggr{]}$
+$\implies r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})+\rho_{i_{*}}$ $\leq r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(\tilde{i})})+\rho_{\tilde{i}}$
+$\implies r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})-r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(\tilde{i})})$ $\leq\rho_{\tilde{i}}-\rho_{i_{*}}$
+We denote $\ell:=r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})-r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(\tilde{i})})$ and by definition of $\tilde{\mathbf{x}}$ we know that $\ell>0$.
+We get that $\rho_{\tilde{i}}\geq\rho_{i_{*}}+\ell$.
+Now by the boundedness of the utility function, we know that if $\rho_{i_{*}}+\ell\geq d$, then the above inequality leads to a contradiction (the learner cannot incur strictly more than $d$ regret over $d$ timesteps).
+Otherwise, we have a subtree with root node weight at least $\rho_{i_{*}}+\ell$ that is shattered by the induced hypothesis space $\pazocal{H}^{(\mathbf{z}_{\varnothing}\to f^{(\tilde{i})})}$.
+By the inductive hypothesis we know that the adversary can force at least $\rho_{i_{*}}+\ell$ regret over the remaining $d$ timesteps and therefore at least $r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(\tilde{i})})+\rho_{i_{*}}+\ell=r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})+\rho_{i_{*}}=\rho_{\varnothing}$ regret over $d+1$ timesteps, as desired.
+∎
+3.4 Stackelberg Standard Optimal Algorithm
+In this section, we present an algorithm that achieves instance-optimal regret that is equal to the Stackelberg-Littlestone dimension.
+We call Algorithm˜1 the Stackelberg Standard Optimal Algorithm (SSOA), which maintains a similar approach to the Standard Optimal Algorithm (SOA) from online multiclass prediction.
+Specifically, it keeps track of the subset of hypotheses that are consistent with the history so far (˜8) and plays the mixed strategy that minimizes the worst-case possible instantaneous regret “regularized” by the Stackelberg-Littlestone dimension of the induced hypothesis class (˜6).
+The latter works as a proxy to the difficulty of the learning task of the leader that would remain after the current round is over.
+Algorithm 1 Stackelberg Standard Optimal Algorithm (SSOA)
+1:input: time horizon, $T$, Stackelberg game $\pazocal{G}$, and hypothesis class $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$
+2:Set $V_{1}=\pazocal{H}$
+3:for $t\in[T]$ do
+4:  Observe $\mathbf{z}_{t}$
+5:  Let $u_{*}^{(i,\mathbf{z}_{t})}:=\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{f^{(i)}}(\mathbf{z}_{t},\mathbf{x}))$ for all $i\in V_{t}(\mathbf{z}_{t})$
+6:  Play $\mathbf{x}_{t}=\arg\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in V_{t}(\mathbf{z}_{t})}\left[u_{*}^{(j,\mathbf{z}_{t})}-u(\mathbf{z}_{t},\mathbf{x},b_{f^{(j)}}(\mathbf{z}_{t},\mathbf{x}))+\textnormal{SLdim}_{\pazocal{G}}\left(V_{t}^{(\mathbf{z}_{t}\to j)}\right)\right]$
+7:  Observe $f_{t}$
+8:  Set $V_{t+1}=V_{t}^{(\mathbf{z}_{t}\to f_{t})}$
+9:end
+In Theorem˜3.9, we prove the optimality of our algorithm.
+Our proof follows a high-level idea similar to that of the Standard Optimal Algorithm (SOA) from online multiclass learning: with each mistake, the learner’s task becomes easier, as the Littlestone dimension decreases.
+Similarly here we show that when the learner incurs positive instantaneous regret, the Stackelberg Littlestone dimension decreases by at least this amount.
+Theorem 3.9.
+For any Stackelberg game, $\pazocal{G}$, and every hypothesis class, $\pazocal{H}$, the regret of Algorithm˜1 is at most $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
+Proof.
+We will show that at each timestep $t$, if the learner incurred instantaneous regret $r_{t}=r(\mathbf{z}_{t},\mathbf{x}_{t},f_{t})$, then the the Stackelberg-Littlestone dimension must have decreased by at least $r_{t}$.
+In other words, $\textnormal{SLdim}_{\pazocal{G}}(V_{t+1})\leq\textnormal{SLdim}_{\pazocal{G}}(V_{t})-r_{t}$, where $V_{t}$ is the set of hypotheses consistent with the history up to timestep $t$.
+We have that the regret at timestep $t$ is
+$r_{t}$ $=u(\mathbf{z}_{t},\mathbf{x}_{*}^{(i,\mathbf{z}_{t})},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{*}^{(i,\mathbf{z}_{t})})-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))+(\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{z}_{t}\to f_{t})})-\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{z}_{t}\to f_{t})}))$
+$\leq\max_{i\in V_{t}(\mathbf{z}_{t})}\left[u(\mathbf{z}_{t},\mathbf{x}_{*}^{(i,\mathbf{z}_{t})},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{*}^{(i,\mathbf{z}_{t})})-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))+\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{z}_{t}\to i)})\right]-\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{z}_{t}\to f_{t})})$
+$=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in V_{t}(\mathbf{z}_{t})}\left[u(\mathbf{z}_{t},\mathbf{x}_{*}^{(i,\mathbf{z}_{t})},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{*}^{(i,\mathbf{z}_{t})})-u(\mathbf{z}_{t},\mathbf{x},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}))+\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{z}_{t}\to i)})\right]-\textnormal{SLdim}_{\pazocal{G}}(V_{t+1})$ (by definition of Algorithm 1)
+$\leq\sup_{\mathbf{z}\in\pazocal{Z}}\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in V_{t}(\mathbf{z})}\left[u(\mathbf{z},\mathbf{x}_{*}^{(i,\mathbf{z})},b_{f_{t}}(\mathbf{z},\mathbf{x}_{*}^{(i,\mathbf{z})})-u(\mathbf{z},\mathbf{x},b_{f_{t}}(\mathbf{z},\mathbf{x}))+\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{z}\to i)})\right]-\textnormal{SLdim}_{\pazocal{G}}(V_{t+1})$
+$=\textnormal{SLdim}_{\pazocal{G}}(V_{t})-\textnormal{SLdim}_{\pazocal{G}}(V_{t+1}),$
+where $\mathbf{x}_{*}^{(i,\mathbf{z})}$ is the leader’s optimal strategy against follower $f^{(i)}$ for context $\mathbf{z}$. ∎
+3.5 Relationship with Littlestone dimension
+We have so far shown that the Stackelberg-Littlestone dimension characterizes learnability in our setting.
+We next explore how that relates to multiclass learnability.
+While, as discussed, learning to predict the follower type from the context is sufficient to learn the leader’s optimal policy, in this section we prove that this is not a necessary condition.
+We start by showing that the Littlestone dimension upper bounds the Stackelberg Littlestone dimension.
+The proof of Lemma˜3.10 can be found in Section˜A.4.
+Lemma 3.10.
+For any hypothesis class $\pazocal{H}$ and Stackelberg game, $\pazocal{G}$, $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})\leq\textnormal{Ldim}(\pazocal{H})$.
+While it is known that any learner makes at least $\textnormal{Ldim}(\pazocal{H})$ mispredictions when learning a hypothesis class, $\pazocal{H}$, a careful reader might notice that this does not immediately show that the optimal multiclass learning algorithm (SOA) [22] achieves suboptimal regret in our setting.
+This is because mispredicting the follower type does not necessarily imply that the optimal strategy against the mispredicted and the true follower type do not yield the same utility for the leader.
+An illustrative example of this is our game construction in Theorem˜3.5, which shows that any learner can mispredict the follower type ad infinitum, yet the leader will achieve optimal cumulative utility regardless.
+In Theorem˜3.11, we prove that, indeed, there are instances where SOA will achieve smaller cumulative utility for the leader compared to our proposed algorithm.
+[^5] In fact, there are cases in which SOA yields strictly lower utility even against the algorithm that minimizes policy mispredictions.
+We prove this result in Lemma B.6.
+Theorem 3.11.
+There exists some hypothesis class $\pazocal{H}$ and a Stackelberg game, $\pazocal{G}$, such that running multiclass SOA on $\pazocal{H}$ yields suboptimal utility for the leader.
+Proof.
+[…]
+$h(z)=\begin{cases}f^{(3)}\;\;&\text{ if }z=1\\ f^{(2)}\;\;&\text{ if }z=2\\ f^{(1)}\;\;&\text{ if }z=3\\ f^{(2)}\;\;&\text{ if }z=4\\ \end{cases}\hskip 56.9055pth^{\prime}(z)=\begin{cases}f^{(3)}\;\;&\text{ if }z=1\\ f^{(1)}\;\;&\text{ if }z=2\\ f^{(2)}\;\;&\text{ if }z=3\\ f^{(2)}\;\;&\text{ if }z=4\\ \end{cases}$
+Notice that the Littlestone dimension of this hypothesis class is at least 222. This follows from the $\pazocal{H}$-shattered tree construction shown in Figure˜1.
+Figure 1: Construction of an $\pazocal{H}$-shattered Littlestone tree. Each node represents a context in $\pazocal{Z}$ and each edge corresponds to a follower type.
+Consider now a Stackelberg game with utility functions $u_{f^{(1)}},u_{f^{(2)}},u_{f^{(3)}}$, defined such that if $z=1$ or $z=2$ then $\pi_{*}^{(1)}(z)=\pi_{*}^{(2)}(z)\neq\pi_{*}^{(3)}(z)$, if $z=3$ then $\pi_{*}^{(1)}(3)=\pi_{*}^{(2)}(3)=\pi_{*}^{(3)}(3)$, and if $z=4$, then $\pi_{*}^{(1)}(4)\neq\pi_{*}^{(2)}(4)=\pi_{*}^{(3)}(4)$.
+Here, we let $\pi_{*}^{(i)}:\pazocal{Z}\to\Delta(\pazocal{A})$ denote the optimal leader’s policy against follower type $f^{(i)}$.
+Concretely, consider a Stackelberg game with the leader being the row player and each follower being the column player, as shown in Table˜2.
+$F_{1}$ $F_{2}$
+$L_{1}$ 1,1/21,1/21 , 1 / 2 $,0$ 0
+$L_{2}$ $,0$ 0 1,01,01 , 0
+(a) Follower type $f^{(1)}$
+$F_{1}$ $F_{2}$
+$L_{1}$ $1,\mathbbm{1}_{\{z\neq 4\}}/2$ $,0$ 0
+$L_{2}$ $,0$ 0 $1,\mathbbm{1}_{\{z=4\}}/2$
+(b) Follower type $f^{(2)}$
+$F_{1}$ $F_{2}$
+$L_{1}$ $1,\mathbbm{1}_{\{z=3\}}/2$ $,0$ 0
+$L_{2}$ $,0$ 0 $1,\mathbbm{1}_{\{z\neq 3\}}/2$
+(c) Follower type $f^{(3)}$
+Table 2: Utility tables for the leader (row player) and each follower type (column player)
+We now show that the $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=3/4$.
+We will consider any possible Stackelberg Littlestone tree.
+First, note that without loss of generality we can consider trees that do not include the context $z=4$ since for $z=4,\;\pi_{*}^{(2)}(4)=\pi_{*}^{(3)}(4)$, where $f^{(2)}$ and $f^{(3)}$ are the only possible follower types predicted by hypotheses in $\pazocal{H}$ for this context.
+Similarly, for $z=3$, since the leader’s optimal strategy against all follower types is identical ($\pi_{*}^{(1)}(3)=\pi_{*}^{(2)}(3)=\pi_{*}^{(3)}(3)$).
+So it suffices to consider contexts 111 and 222.
+In Figure˜2 we show the two shattered Stackelberg-Littlestone tree constructions with $z=1$ and $z=2$ as the root nodes respectively.
+We start with the tree in Figure˜2(a).
+For $z=2$, its not hard to show that $\inf_{\mathbf{x}}\max_{j\in\{1,3\}}r(\mathbf{x},2,f^{(j)})=\inf_{\mathbf{x}}\max_{j\in\{2,3\}}r(\mathbf{x},2,f^{(j)})=1/2$, achieved at $\mathbf{x}=(1/2,1/2)$.
+For the root node $z=1$, we have that $\inf_{\mathbf{x}}\max\{r(\mathbf{x},1,f^{(1)})+1/2,\;r(\mathbf{x},1,f^{(2)})+1/2,\;r(\mathbf{x},1,f^{(3)})\}=3/4$ achieved at $\mathbf{x}=(3/4,1/4)$.
+Therefore, the weight of the root node is 3/43/43 / 4.
+A similar analysis shows that the root node weight of the SL-tree in Figure˜2(b) is 3/43/43 / 4.
+Therefore, $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ (and also the worst-case cumulative regret) equals 3/43/43 / 4.
+We consider the case, where the adversary present context 111 followed by context 222. We go through the SOA outputs per timestep below.
+[…]
+  • ($t=2$) This node has two children mapping $z=2$ to follower types $f^{(2)}$ and $f^{(3)}$.
+No subsequent context can lead to a full-binary tree that is consistent with any of these assignments, so SOA tie-breaks.
+The adversary in this case can always force a mistake by claiming the correct follower type to be the one not predicted by the algorithm.
+So if SOA predicts $f_{2}=f^{(2)}$, and plays strategy $x_{2}=\pi_{*}^{(2)}(2)$ the adversary can choose the true follower type to be $f_{2}^{*}=f^{(3)}$ (instantaneous regret equals 1).
+This leads to SOA mispredicting the follower type and getting suboptimal utility in both rounds.
+Since the Stackelberg Littlestone dimension is 3/43/43 / 4 and we have shown that for the above sequence of contexts running SOA yields regret 222, this concludes our proof.
+Figure 2: $\pazocal{H}$-shattered SL-Tree with (a) $z=1$ as the root node and (b) $z=2$ as the root node.
+4 Distributional Learning in Structured Stackelberg Games
+We now shift our focus from online learning to distributional learning.
+Consider an unknown, joint distribution over contexts and follower types, from which the leader receives samples drawn i.i.d.
+The leader’s objective is to design her strategy in such a way to maximize the utility she will get on a new sample drawn from the same distribution.
+We first study the simpler setting, without contexts, in which we give strong generalization guarantees by exploiting the structure of the leader’s payoff function.
+When contexts come into play in Section˜4.2, we prove necessary and sufficient conditions for PAC learnability.
+4.1 Learning without Contexts
+Suppose there is a fixed, unknown distribution, $\mathscr{D}$, over the set of possible follower types.
+We denote with $u(\mathbf{x},b_{f^{(i)}}(\mathbf{x}))$ the leader’s payoff when committing to a (possibly mixed) strategy $\mathbf{x}\in\Delta(\pazocal{A})$ against a follower of type $f^{(i)}$.
+We use the notation $u_{\mathscr{D}}(\mathbf{x})=\mathbb{E}_{f\sim\mathscr{D}}[u(\mathbf{x},b_{f}(\mathbf{x}))]$ and given a sample $S=\{f_{1},\dots,f_{m}\}$ of follower types, we write $u_{S}(\mathbf{x})=\frac{1}{m}\sum_{f\in S}u(\mathbf{x},b_{f}(\mathbf{x}))$.
+Given such a sample, the goal of the leader is to choose a mixed strategy that fares well in expectation against any new sample from $\mathscr{D}$.
+Concretely, we want to choose $\hat{\mathbf{x}}\in\Delta(\pazocal{A})$ such that
+$\mathbb{P}_{S}\left[\sup_{\mathbf{x}}u_{\mathscr{D}}(\mathbf{x})-u_{\mathscr{D}}(\hat{\mathbf{x}})>\epsilon\right]\leq\delta.$
+To achieve this, it suffices to show a generalization guarantee i.e., a bound on the difference between the leader’s expected payoff and the empirical payoff over a set of follower types for any possible mixed strategy.
+Definition 4.1 (Generalization guarantee).
+A generalization guarantee for the leader in a Stackelberg game, $\pazocal{G}$, is a function $\epsilon_{\pazocal{G}}:\mathbb{Z}_{\geq 1}\times(0,1)\to\mathbb{R}_{\geq 0}$ such that for any $\delta\in(0,1)$, any $m\in\mathbb{Z}_{\geq 1}$, and any distribution $\mathscr{D}$ over the follower types, with probability at least $1-\delta$ over the draw of a set $S\sim\mathscr{D}^{m}$ for any $\mathbf{x}\in\Delta(\pazocal{A})$, the difference between the average payoff of $\mathbf{x}$ over $S$ and the expected payoff of $\mathbf{x}$ over $\mathscr{D}$ is at most $\epsilon_{\pazocal{G}}(m,\delta)$:
+$\mathbb{P}_{S\sim\mathscr{D}^{m}}\left[\exists\mathbf{x}\in\Delta(\pazocal{A})\;\;\text{such that}\;\;|u_{\mathscr{D}}(\mathbf{x})-u_{S}(\mathbf{x})|>\epsilon_{\pazocal{G}}(m,\delta)\right]\leq\delta$
+This implies a relationship between the utility of the leader’s strategy maximizing the average payoff over the training data to the expected utility of the optimal strategy.
+Formally, for $\hat{\mathbf{x}}\in\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u_{S}(\mathbf{x})$ and $\mathbf{x}^{*}\in\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u_{\mathscr{D}}(\mathbf{x})$
+$\mathbb{P}_{S\sim\mathscr{D}^{m}}\left[|u_{\mathscr{D}}(\mathbf{x}^{*})-u_{S}(\hat{\mathbf{x}})|>2\epsilon_{\pazocal{G}}(m,\delta)\right]\leq\delta.$
+We let $\pazocal{P}=\{u_{\mathbf{x}}:[K]\to\mathbb{R}\;\;\text{such that}\;\;u_{\mathbf{x}}(f)=u(\mathbf{x},b_{f}(\mathbf{x}))\}$ be the set of payoff functions for the leader.
+Each payoff function is parameterized by the leader’s mixed strategy, takes as input the follower’s type, and outputs the leader’s utility.
+We also define the dual class $\tilde{\pazocal{P}}=\{\tilde{u}_{f}:\Delta(\pazocal{A})\to\mathbb{R}\;\;\text{such that}\;\;\tilde{u}_{f}(\mathbf{x})=u(\mathbf{x},b_{f}(\mathbf{x}))\}$.
+Consider the following definition, introduced by Balcan et al.
+[9] and adapted to our setting.
+Definition 4.2.
+[…]
+  • The class $\pazocal{H}$ consists of hypotheses that are parameterized by vectors $\mathbf{x}$ from a set $\pazocal{X}\subseteq\mathbb{R}^{d}$,
+  • For any $f\in[K]$, there is a set $\mathscr{J}$ of $t$ hyperplanes such that for any connected component $\pazocal{X}^{\prime}=\pazocal{X}\setminus\mathscr{J}$, the dual function $h_{f}(\mathbf{x})$ is linear over $\pazocal{X}^{\prime}$.
+We can now show that the primal class $\pazocal{P}$ is deliniable, which is a property we will use to show a generalization guarantee.
+Lemma 4.3.
+The class of leader payoff functions, $\pazocal{P}$ that are parameterized by the leader’s mixed strategy and take as input the follower’s type is $(|\pazocal{A}|,\binom{|\pazocal{A}_{f}|}{2})$-delineable.
+Proof.
+The statement holds from the following two facts.
+First, given a fixed follower type, $f$, we have that $\tilde{u}_{f}(\mathbf{x})$ is a piecewise linear function.
+Second, for any fixed follower type, $f$, we have that $u_{f}(\mathbf{x})$ will have at most $|A_{f}|(|A_{f}|-1)$ hyperplanes going through it.
+Each hyperplane corresponds to a change in the follower’s best-response strategies, from one pure action to another.
+Therefore, there can be at most $\binom{|\pazocal{A}_{f}|}{2}$ hyperplanes.
+∎
+Next, we present the definition of pseudo-dimension [34], which is complexity measure that characterizes a hypothesis class and in our setting helps us reason about how difficult the learning task of the leader is.
+Informally speaking, the more complex the hypothesis class, $\pazocal{P}$, the harder it is for the leader to distinguish between strategies, $\mathbf{x}$, which parameterize the class, and therefore, more samples are needed to learn.
+Definition 4.4 (Pseudo-dimension).
+Let $\pazocal{F}$ be a class of functions $f:A\to\mathbb{R}$, with an abstract domain, $A$.
+We say that $y^{(1)},\ldots,y^{(n)}$, witness the shattering of $S=\{x^{(1)},\ldots,x^{(n)}\}\subseteq A$ if for all $T\subseteq S$, there exists a function $f_{T}\in\pazocal{F}$, such that for all $x^{(i)}\in T,\;f_{T}(x^{(i)})\leq y^{(i)}$ and for all $x^{(i)}\not\in T,\;f_{T}(x^{(i)})>y^{(i)}$.
+We then say that the pseudo-dimension of $\pazocal{F}$, denoted $\textnormal{Pdim}(\pazocal{F})$, is the size of the largest set that can be shattered by $\pazocal{F}$.
+Using Lemma˜4.3 and results from the data-driven algorithm design literature [6, 12, 14], we immediately get the following bound on the pseudo-dimension.
+Theorem 4.5 (Adaptation of Lemma 3.10 in Balcan et al. [9]).
+$\textnormal{Pdim}(\pazocal{P})\leq 9|\pazocal{A}|\log\left(4|\pazocal{A}||\pazocal{A}_{f}|^{2}\right)$.
+It is known that the finiteness of the pseudo-dimension of a function class is a sufficient condition to show a generalization guarantee for the class [34, 24].
+In our case, the above bound implies a generalization guarantee with function
+$\epsilon_{\pazocal{G}}(m,\delta)=O\left(\sqrt{\frac{|\pazocal{A}|\log(|\pazocal{A}||\pazocal{A}_{f}|^{2})}{m}}+\sqrt{\frac{\ln(1/\delta)}{m}}\right).$
+Discussion.
+We note that the above generalization result is particularly strong.
+Informally, it asserts that for any distribution over the follower types, if the leader has so far encountered a number of followers that scales linearly with the size of her action set and logarithmically with the follower’s action set, it suffices for her to play a strategy that maximizes her average utility over the followers seen so far.
+This result holds even for an infinite number of follower types and extends prior work on learning in Stackelberg games without contextual information [29].
+4.2 Learning with Contexts: Problem Setting and Definitions
+In this section, we introduce the study of the statistical complexity of distributional learning in structured Stackelberg games.
+The leader has access to a data set consisting of tuples of contexts and follower types drawn from a fixed and unknown distribution.
+Given new contextual information drawn from the same distribution, the leader aims to choose a strategy that maximizes her payoff against a follower whose type is unknown.
+As in the online setting, this comes with the promise that there exists a function, $h^{*}$, in a class, $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$, that perfectly predicts each follower type given the context.
+We denote a learning algorithm for distributional learning in structured Stackelberg games as a mapping $A:(\pazocal{Z}\times[K])^{*}\to\Delta(\pazocal{A}){Z}$.
+Given a context, $\mathbf{z}$, we write the leader’s loss with respect to $h^{*}$, when playing strategy $\hat{\mathbf{x}}\in\Delta(\pazocal{A})$ as
+$r(\mathbf{z},\hat{\mathbf{x}},f^{(h^{*}(\mathbf{z}))})=\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u\left(\mathbf{z},\mathbf{x},b_{f^{(h^{*}(\mathbf{z}))}}(\mathbf{z},\mathbf{x})\right)-u\left(\mathbf{z},\hat{\mathbf{x}},b_{f^{(h^{*}(\mathbf{z}))}}(\mathbf{z},\hat{\mathbf{x}})\right)$
+We handle two variants of the notion of sample complexity, which will be central in showing our results. We begin with the classic PAC sample complexity, which we adapt to our setting.
+Definition 4.6 (PAC Sample complexity).
+The PAC sample complexity $m(\pazocal{H},\pazocal{G};\epsilon,\delta)$ of a function class $\pazocal{H}$ with respect to a Stackelberg game $\pazocal{G}$ is defined as $m(\pazocal{H},\pazocal{G};\epsilon,\delta)=\inf_{A}m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta)$, where the infimum is over all possible learning algorithms.
+With $m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta)$ we denote the minimal integer such that for any $m\geq m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta)$, every distribution $\mathscr{D}$ on $\pazocal{Z}$, and any true target hypothesis $h^{*}\in\pazocal{H}$, the expected loss of $A$,
+$\mathop{\mathbb{E}}_{\mathbf{z}\sim\mathscr{D}}\left[r(\mathbf{z},A(S)(\mathbf{z}),f^{(h^{*}(\mathbf{z}))})\right],$
+is at most $\epsilon$ with probability $1-\delta$ over the set $S=\{(\mathbf{z}_{i},f^{(h^{*}(\mathbf{z}_{i}))})\}_{i\in[m]}$, where $\forall i\in[m]\;\mathbf{z}_{i}\sim\mathscr{D}$.
+Definition 4.7 (PAC Cut-off sample complexity).
+The PAC sample complexity $m(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$ of a function class $\pazocal{H}$ with respect to a Stackelberg game $\pazocal{G}$ is defined as $m(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)=\inf_{A}m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$, where $m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$ is the minimal integer such that for any $m\geq m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$, every distribution $\mathscr{D}$ on $\pazocal{Z}$, and any true target hypothesis $h^{*}\in\pazocal{H}$, the expected cut-off loss of algorithm $A$,
+$\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}[r(\mathbf{z},A(S)(\mathbf{z}),f^{(h^{*}(\mathbf{z}))})>\gamma],$
+is at most $\epsilon$ with probability $1-\delta$ over the set $S=\{(\mathbf{z}_{i},f^{(h^{*}(\mathbf{z}_{i}))})\}_{i\in[m]}$, where $\forall i\in[m]\;\mathbf{z}_{i}\sim\mathscr{D}$.
+For the remainder of this section our results will be with respect to the PAC cut-off sample complexity.
+Any bound with respect to this benchmark immediately implies a bound on the PAC sample complexity as well.
+It is not hard to generalize the result of Lemma˜4.8 to our setting, so we omit the proof.
+Lemma 4.8 (Lemma 1 in [3]).
+For every $\epsilon,\delta\in(0,1)^{2}$ and every $\pazocal{H}\subseteq[0,1]^{\pazocal{Z}}$, it holds that
+$m(\pazocal{H};\sqrt{\epsilon},\delta,\sqrt{\epsilon})\leq m(\pazocal{H};\epsilon,\delta)\leq m(\pazocal{H},\epsilon/2,\delta,\epsilon/2).$
+4.3 Lower bound
+In this section, we introduce the $\gamma$-valued Stackelberg-Natarajan dimension, a new measure of complexity that characterizes a structured Stackelberg game and, as we will see, controls the lower bound of the sample complexity.
+Indeed, we prove that the finiteness of this measure is a necessary condition for learnability in our setting.
+This definition is inspired by the classic Natarajan dimension from multiclass prediction [31, 21], and more recently, the $\gamma$-Natarajan dimension from realizable regression [3].
+We start with introducing the notion of a $\gamma$-SN-shattered set.
+Informally, a set of contexts is $\gamma$-SN-shattered if $\pazocal{H}$ can fit complex patterns of follower types on the set.
+Further, these follower types must be sufficiently different, such that if the type is chosen adversarially among them, the leader incurs loss at least $\gamma$ regardless of her strategy.
+Definition 4.9 ($\gamma$-SN-Shattered Set).
+Let $\pazocal{H}:[K]^{\pazocal{Z}}$ and $\pazocal{G}$ be a Stackelberg game.
+We say that a set $S\in\pazocal{Z}^{n}$ is $\gamma$-SN-shattered by $\pazocal{H}$ with respect to $\pazocal{G}$ if there exist $g_{0},g_{1}:\pazocal{Z}\to[K]$ such that
+$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{g_{0}(\mathbf{z}_{i}),g_{1}(\mathbf{z}_{i})\}}r(\mathbf{z}_{i},\mathbf{x},f^{(j)})>\gamma\;\;\forall\mathbf{z}_{i}\in S$
+and for all $b\in\{0,1\}^{n}$ there exists an $h_{b}\in\pazocal{H}$ such that $h_{b}(\mathbf{z}_{i})=\begin{cases}g_{0}(\mathbf{z}_{i})\;\;\text{if}\;\;b_{i}=0\\ g_{1}(\mathbf{z}_{i})\;\;\text{if}\;\;b_{i}=1\end{cases}$.
+Definition 4.10 ($\gamma$-valued Stackelberg-Natarajan Dimension).
+The $\gamma$-valued Stackelberg-Natarajan dimension, denoted as $\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$, is the cardinality of the largest set that is $\gamma$-SN-shattered by $\pazocal{H}$ with respect to $\pazocal{G}$.
+Before stating our main result of this section, we point out that it is important that the two functions, $g_{0}$ and $g_{1}$ be $\gamma$-far from each other in a minmax sense.
+This allows us to prove a lower bound that holds for any learning algorithm.
+Theorem 4.11.
+Let $A$ be any learning algorithm and $\epsilon,\delta,\gamma\in(0,1)^{3}$ such that $\delta<\epsilon$, then
+$m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)\geq\Omega\left(\frac{\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})+\log(1/\delta)}{\epsilon}\right).$
+Proof.
+Let $n_{0}=\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$.
+Then we know that for any $1\leq n\leq n_{0}$ there exists a set $S_{n}\in\pazocal{Z}^{n}$ and functions $g_{0},g_{1}:\pazocal{Z}\to\Delta(\pazocal{A})$ such that for any set $T\subseteq S$ there exists an $h$ that agrees with $g_{0}$ on all $\mathbf{z}\in T$ and with $g_{1}$ for any $\mathbf{z}\in S\setminus T$.
+We will show that for a given learning algorithm $A:(\pazocal{Z}\times[K])^{*}\to\Delta(\pazocal{A}){Z}$, there exists a distribution $\mathscr{D}$ on $\pazocal{Z}$ such that the algorithm requires at least $C_{0}\frac{\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})+\log(\frac{1}{\delta})}{\epsilon}$ samples.
+To do that, we will construct a distribution with support the $n_{0}$-sized $\gamma$-shattered set $S$.
+We take $\delta<1/15$ and define the distribution $\mathscr{D}$ on $S$ as follows.
+$\mathbb{P}[\mathbf{z}_{1}]=1-16\epsilon,\;\;\;\;\mathbb{P}[\mathbf{z}_{i}]=\frac{16\epsilon}{d-1}\;\forall i\in\{2,\dots,d-1\}.$
+We draw a set $X$ of $m=\frac{d-1}{64\epsilon}$ i.i.d.
+samples from this distribution.
+We let $A$ be an algorithm that takes $X$ as the input and then produces a mapping $A(X):\pazocal{Z}\to\Delta(\pazocal{A})$.
+We define
+$\text{er}_{\mathscr{D},\gamma,h}(A(X)):=\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}[r(\mathbf{z},A(X)(\mathbf{z}),f^{(h(\mathbf{z}))})>\gamma].$
+It suffices to establish that there exists $h^{*}\in\pazocal{H}$ such that $\mathbb{P}_{X\sim\mathscr{D}^{m}}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right]>1/15$.
+First consider the event
+$B:|\{\mathbf{x}\in X:\mathbf{x}\in\{\mathbf{z}_{2},\dots,\mathbf{z}_{n}\}\}|\leq(d-1)/2.$
+Using Markov’s inequality, it is easy to show that $\mathbb{P}[B]\geq 1/2$.
+Let $U=\{\mathbf{z}_{2},\dots,\mathbf{z}_{n}\}\setminus X$ be the points excluding $\mathbf{z}_{1}$ that are unseen by the algorithm through the training set.
+Given $B$ we know that $|U|\geq\frac{d-1}{2}$.
+For the remainder of this section we will focus on
+$\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X)):=\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}[r(\mathbf{z},A(X)(\mathbf{z}),f^{(h(\mathbf{z}))})>\gamma\land\mathbf{z}\in\{\mathbf{z}_{2},\dots,\mathbf{z}_{n_{0}}\}],$
+which trivially lower bounds $\text{er}_{\mathscr{D},\gamma,h}(A(X))$.
+Also, we consider the following way to choose a true hypothesis $h$.
+First, we choose vector $b$ uniformly at random from the set $\{0,1\}^{n_{0}}$.
+We then set $h(\mathbf{z}_{i})=g_{0}(\mathbf{z}_{i})$ for all $b_{i}=0$ and $h(\mathbf{z}_{i})=g_{1}(\mathbf{z}_{i})$ for all $b_{i}=1$.
+We note that choosing a random $h$ in this way is equivalent to flipping a fair coin for each point in $S$ to determine its label.
+Since for any point $\mathbf{z}_{i}\in U$ both labels $g_{0}(\mathbf{z}_{i})$ and $g_{1}(\mathbf{z}_{i})$ are consistent with the training set $X$ (by definition of an SN-shattered set), we have that $A(X)$ is independent of the labeling of points in $U$.
+We also know that if we condition on the point drawn from $\mathscr{D}$ being some unseen point $\mathbf{u}\in U$, we get that for any strategy $A(X)(\mathbf{u})\in\Delta(\pazocal{A})$ of the learner
+$\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\geq\begin{cases}1\;\;\text{if}\;h(\mathbf{u})=\arg\max_{j\in\{g_{0}(\mathbf{u}),g_{1}(\mathbf{u})\}}r(\mathbf{u},A(X)(\mathbf{u}),f^{(j)})\\ 0\;\;\text{otherwise}.\end{cases}$
+Since we are randomly choosing $h(\mathbf{z})$ to be $g_{0}(\mathbf{z}_{i})$ or $g_{1}(\mathbf{z}_{i})$, we have that the two conditions above happen with probability 1/21/21 / 2 each.
+Given that there are at least $(d-1)/2$ unseen points, each sampled with probability $\frac{16\epsilon}{d-1}$, and each forcing the learner to make a $\gamma$-error with probability 1/21/21 / 2, we get that
+$\mathop{\mathbb{E}}_{h,X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\;|\;B\right]>\frac{1}{2}\cdot\frac{16\epsilon}{d-1}\cdot\frac{(d-1)}{2}=4\epsilon.$
+From this analysis and the probability of $B$, we get that
+$\mathop{\mathbb{E}}_{h,X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\right]$ $=\mathop{\mathbb{E}}_{h,X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\;|\;B\right]\cdot\mathbb{P}[B]$
+$>4\epsilon\cdot\frac{1}{2}=2\epsilon.$
+So there must exists some $h^{*}$ such that $\mathop{\mathbb{E}}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h^{*}}(A(X))\right]>2\epsilon$.
+We take $h^{*}$ as the target concept and show that $A$ is likely to produce high error rate.
+We know that
+$\mathbb{P}_{X}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right]\cdot\mathop{\mathbb{E}}_{X}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))\;|\;\text{er}_{\mathscr{D},\gamma,h}(A(X))>\epsilon\right]$
+$+(1-\mathbb{P}_{X}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right])\cdot\mathop{\mathbb{E}}_{X}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))\;|\;\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))\leq\epsilon\right]>2\epsilon$
+Given that for any algorithm, $A$, it holds that $\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\leq\mathbb{P}[\mathbf{z}\in\{\mathbf{z}_{2},\dots,\mathbf{z}_{n}\}]=16\epsilon$ we get
+$\mathop{\mathbb{E}}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\;|\;\text{er}_{\mathscr{D},\gamma,h}(A(X))>\epsilon\right]\leq 16\epsilon\;\;\text{for any}\;\;h$
+which implies that for $h^{*}$
+$\mathbb{P}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right]\cdot 16\epsilon+(1-\mathbb{P}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right])\cdot\epsilon>2\epsilon$
+$\iff\mathbb{P}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right]>1/15.$
+Therefore, the learner needs at least $m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)\geq\frac{d-1}{64\epsilon}$ samples.
+We now consider the case, where $X$ contains only $\mathbf{z}_{1}$.
+This occurs with probability $(1-16\epsilon)^{m}\geq e^{-32m\epsilon}$, which is greater than $\delta$, when $m\leq\frac{\log(1/\delta)}{32\epsilon}$.
+Therefore we get that for any algorithm $A$:
+$m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)\geq\max\biggl{\{}\frac{d-1}{64\epsilon},\frac{\log(1/\delta)}{16\epsilon}\biggr{\}}=C_{0}\frac{\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})+\log(\frac{1}{\delta})}{\epsilon}$
+Lastly, we consider the case, where $\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})=\infty$.
+Then for any shattered set of size $n\in\mathbb{N}$, we can repeat the argument above, which would show that the cut-off sample complexity of any algorithm is at least $\Omega\left(\frac{n+\log(\frac{1}{\delta})}{\epsilon}\right)$, thus establishing that $m(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)=\infty$.
+∎
+4.4 Upper bound
+In this section, we introduce the $\gamma$-valued Stackelberg-Graph dimension, which, as we will see, is a sufficient condition for distributional learning in structured Stackelberg games.
+It is also a natural relaxation of the $\gamma$-Natarajan dimension and an adaptation of similar dimensions from classification [31, 21] and regression [3] to our setting.
+Definition 4.12 ($\gamma$-SG-Shattered Set).
+Let $\pazocal{H}:[K]^{\pazocal{Z}}$ and $\pazocal{G}$ be a Stackelberg game.
+We say that a sample $S\in\pazocal{Z}^{n}$ is shattered by $\pazocal{H}$ with respect to $\pazocal{G}$ if there exists a function $g:\pazocal{Z}\to[K]$ such that for all $b\in\{0,1\}^{n}$ there exists an $h_{b}\in\pazocal{H}$ such that:
+  • $h_{b}(\mathbf{z}_{i})=g(i)\;\;\forall i\in[n]$ such that $b_{i}=0$, and
+  • $\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{g(\mathbf{z}_{i}),h_{b}(\mathbf{z}_{i})\}}r(\mathbf{z}_{i},\mathbf{x},f^{(j)})\geq\gamma\;\;\forall i\in[n]$ such that $b_{i}=1$.
+Definition 4.13 ($\gamma$-valued Stackelberg-Graph Dimension).
+The $\gamma$-valued Stackelberg-Graph dimension, denoted as $\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$ is the cardinality of the largest set that is $\gamma$-shattered by $\pazocal{H}$ with respect to $\pazocal{G}$.
+We are not ready to show our main result in this section, which shows that the $\gamma$-Graph dimension upper bounds the PAC sample complexity in our setting.
+We design the following simple, improper algorithm that achieves the corresponding bound.
+Algorithm 2 Stackelberg Distributional Learner ($\mathfrak{L}^{*}$)
+1:input: sample $S=\{(\mathbf{z}_{1},f_{1}),\dots,(\mathbf{z}_{n},f_{n})\}$ that is realizable by $\pazocal{H}$ and test point $\mathbf{z}\in\pazocal{Z}$
+2:Consider the subset of hypotheses consistent with the sample, $\pazocal{H}_{c}=\pazocal{H}|_{S}$.
+3:Consider the set of followers predicted by $\pazocal{H}_{c}$ on test point $\mathbf{z}$, $F=\pazocal{H}_{c}(\mathbf{z})$
+4:Compute and return $\mathbf{x}^{*}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in F}r(\mathbf{z},\mathbf{x},f^{(i)})$
+5:end
+Theorem 4.14.
+For any class $\pazocal{H}:[K]^{\pazocal{Z}}$, any Stackelberg game, $\pazocal{G}$, and for any $\epsilon,\delta,\gamma\in(0,1)$, there exists some constant $C_{1}$ such that Algorithm˜2 achieves:
+$m_{\mathfrak{L}^{*}}^{r}(\pazocal{H},\epsilon,\delta,\gamma)\leq C_{1}\frac{\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})\log(1/\epsilon)+\log(1/\delta)}{\epsilon}.$
+Proof.
+We work with the cut-off loss problem for some parameters $(\epsilon,\delta,\gamma)\in(0,1)^{3}$. We define the cut-off loss
+$\text{er}_{\mathscr{D},\gamma}(h):=\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}\left[\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}),h^{*}(\mathbf{z})\}}r(\mathbf{z},\mathbf{x},f^{(j)})>\gamma\right]$
+and the empirical cut-off loss for a dataset $S\in(\pazocal{Z}\times[K])^{n}$
+$\hat{\text{er}}_{S,\gamma}(h):=\frac{1}{n}\sum_{(\mathbf{z}_{i},f_{i})\in S}\mathbbm{1}\biggl{\{}\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}),h^{*}(\mathbf{z})\}}r(\mathbf{z},\mathbf{x},f^{(j)})>\gamma\biggr{\}}$
+and the empirical loss
+$\hat{\text{er}}_{S}(h):=\frac{1}{n}\sum_{(\mathbf{z}_{i},f_{i})\in S}\mathbbm{1}\biggl{\{}h(\mathbf{z}_{i})\neq h^{\prime}(\mathbf{z}_{i})\biggr{\}}.$
+For $\pazocal{M}=\pazocal{Z}\times[K]$ we are interested in the following set
+$Q_{\epsilon,\gamma}=\{M\in\pazocal{M}^{n}:\exists h\in\pazocal{H}:\hat{\text{er}}_{M}(h)=0,\text{er}_{\mathscr{D},\gamma}(h)>\epsilon\}$
+Now using a classic symmetrization argument [35, 3], we can upper bound the probability of $Q_{\epsilon,\gamma}$ with that of the following set
+$R_{\epsilon,\gamma}=\{(N,S)\in\pazocal{M}^{n}\times\pazocal{M}^{n}:\exists h\in\pazocal{H}:\hat{\text{er}}_{N}(h)=0,\hat{\text{er}}_{S,\gamma}(h)\geq\epsilon/2\}.$
+The idea here is to bound the probability that there exists a hypothesis yielding maximum utility for the leader over the sample, but is $\gamma$-far from the optimal hypothesis, by the probability that this hypothesis has poor performance on another sample.
+More concretely, it can be shown that $\mathscr{D}^{n}(Q_{\epsilon,\gamma})\leq 2\mathscr{D}^{2n}(R_{\epsilon,\gamma})$ for $n\geq c/\epsilon$, for some constant $c$.
+To see this, we first write
+$\mathscr{D}^{2n}(R_{\epsilon,\gamma})=\int_{Q_{\epsilon,\gamma}}\mathscr{D}^{n}(\{S\in\pazocal{M}^{n}:\exists h\in\pazocal{H}:\hat{\text{er}}_{N}(h)=0,\hat{\text{er}}_{S,\gamma}(h)\geq\epsilon/2\})d\mathscr{D}^{n}(N)$
+For some $N\in Q_{\epsilon,\gamma}$ consider some $h_{N}\in\pazocal{H}$ that satisfies $\hat{\text{er}}_{N,\gamma}(h_{N})=0$ and $\text{er}_{\mathscr{D},\gamma}(h_{N})>\epsilon$.
+We show that $\mathscr{D}(\hat{\text{er}}_{S,\gamma}(h_{S})\geq\epsilon/2)\geq 1/2$, from which the claim follows straightforwardly.
+We know that $\text{er}_{\mathscr{D},\gamma}(h_{N})>\epsilon$, and therefore $n\cdot\hat{\text{er}}_{S,\gamma}(h_{N})\leq n\cdot\hat{\text{er}}_{S}(h_{N})$ follows a binomial distribution with probability of success at least $\epsilon$.
+We note that the inequality above follows since mispredicting the follower type is a necessary condition for suboptimal utility.
+Using a Chernoff bound, we can show that
+$\mathscr{D}^{n}(S:\hat{\text{er}}_{S,\gamma}(h_{N})<\epsilon/2)\leq e^{-n\epsilon/8}$
+and for $n=c/\epsilon$, we get the desired result.
+The next step is to use a random swap argument to upper bound $\mathscr{D}^{2n}(R_{\epsilon,\gamma})$ with something that is independent of $\mathscr{D}$.
+We let $\Gamma_{n}$ denote the set of all permutations on $\{1,\dots,2n\}$ that potentially swap the $i$’th and $n+i$’th elements for all $i\in[n]$.
+So for all $\sigma\in\Gamma_{n}$ and $i\in[n]$, we have that either $\sigma(i)=i,\sigma(n+i)=n+i$ or $\sigma(i)=n+i,\sigma(n+i)=i$.
+From Lemma 3 in Attias et al.
+[3] (cf.
+also [2]), we get that for some $\pazocal{M}=\pazocal{Z}\times[K]$ and distribution $\mathscr{D}$ on $V$:
+$\mathscr{D}^{2n}(R_{\epsilon,\gamma})=\mathop{\mathbb{E}}_{Z\sim\mathscr{D}^{2n}}\mathbb{P}_{\sigma\sim\text{Unif}(\Gamma_{n})}[\sigma(V)\in R_{\epsilon,\gamma}]\leq\max_{Z\in\pazocal{Z}^{2n}}\mathbb{P}_{\sigma\sim\text{Unif}(\Gamma_{n})}[\sigma(V)\in R_{\epsilon,\gamma}].$
+The last step of the proof, requires reducing to learning a partial binary hypothesis class and using a disambiguation argument to turn the partial class into a total one, as in Attias et al.
+[3].
+Concretely, we start by fixing a set $V\in\pazocal{M}^{2n}$, where $v_{i}=(\mathbf{z}_{i},h^{*}(\mathbf{z}_{i}))$ and a set $S=\{\mathbf{z}_{1},\dots,\mathbf{z}_{2n}\}$.
+For the projection, $\pazocal{H}(S)$ of the hypothesis class on the set, one can define the partial binary hypothesis class
+$\pazocal{H}:=\left\{h^{\prime}\in\{0,1,\star\}^{2n}:\exists h\in\pazocal{H}|_{S}:\forall i\in[2n]\;\;h^{\prime}(\mathbf{z}_{i})=\begin{cases}0\;\;\text{if}\;\;h(\mathbf{z}_{i})=h^{*}(\mathbf{z}_{i})\\ 1\;\;\text{if}\;\;\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}),h^{*}(\mathbf{z})\}}r(\mathbf{z},\mathbf{x},f^{(j)})>\gamma\\ \star\;\;\text{if}\;\;\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}),h^{*}(\mathbf{z})\}}r(\mathbf{z},\mathbf{x},f^{(j)})\leq\gamma\end{cases}\right\}$
+It is easy to verify that $\textnormal{VCdim}(\pazocal{H})\leq\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$.
+We can now turn this partial concept class into a total concept class of size $n^{O(\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})\log(n))}$ using the following disambiguation argument.
+Lemma 4.15 ([1]).
+Let $\pazocal{H}^{\prime}$ be a partial concept class on a finite space $S$, with $\textnormal{VCdim}(\pazocal{H}^{\prime})=d$.
+Then there exists a disambiguation $\tilde{\pazocal{H}}^{\prime}$ of $\pazocal{H}^{\prime}$ of size $|\tilde{\pazocal{H}}^{\prime}|=|S|^{O(d\log|S|)}$.
+We have that $\sigma(V)\in R_{\epsilon,\gamma}$ iff there exists some $h\in\pazocal{H}$ such that $h(\mathbf{z}_{\sigma(i)})=h^{*}(\mathbf{z}_{\sigma(i)})\;\;\;\forall\;i\in[n]$ and
+$\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\biggl{\{}\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}_{\sigma(n+i)}),h^{*}(\mathbf{z}_{\sigma(n+i)})\}}r(\mathbf{z}_{\sigma(n+i)},\mathbf{x},f^{(j)})>\gamma\biggr{\}}\geq\epsilon/2.$
+We know that if there exists such an $h$ that is correct on the first $n$ points but $\gamma$-far on at least $\epsilon n/2$ of the last $n$ points, there exists a function in $\tilde{\pazocal{H}}^{\prime}$, that is 0 on the first $n$ points and 111 on the points on which $h$ is $\gamma$-far.
+We also know that if for some $\sigma\in\Gamma_{n}$, $\tilde{h}^{\prime}\in\pazocal{H}^{\prime}$ is a witness that $\sigma(V)\in R_{\epsilon,\gamma}$, then at least one of $\tilde{h}^{\prime}(i),\tilde{h}^{\prime}(i)$ must be 0, and at least $n\epsilon/2$ elements must be non-zero.
+Thus for a permutation drawn uniformly at random, the probability that all 1’s are on the last n elements is at most $2^{-n\epsilon/2}$.
+Also since $|\tilde{\pazocal{H}}^{\prime}|$ is bounded we an use a union bound that yields $\mathbb{P}_{\sigma\sim\text{Unif}(\Gamma_{n})}[\sigma(V)\in R_{\epsilon,\gamma}]\leq n^{O(\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})\log(n))}2^{-n\epsilon/2}$.
+Given that $V\in\pazocal{M}^{2n}$, we can take a $\max$ over all $V\in\pazocal{M}^{2n}$, which gives the desired result.
+∎
+Remark 4.16.
+It is worth noting that even when there is only one context (i.e.
+$\pazocal{Z}$ is a singleton), our results in this section are incomparable with those in Section˜4.1.
+This is because in this section, the goal of the leader is to learn given a deterministic mapping from contexts to follower types.
+So, if the contexts are the same between all examples in the training set, the follower types corresponding to these examples will also be the same.
+In contrast, the context-free setting we consider in Section˜4.1 allows for a distribution over follower types.
+Thus the two bounds are generally incomparable.
+5 Computational Complexity Considerations
+Computational complexity has been a longstanding concern in the analysis of Stackelberg games.
+In the simplest setting, where a leader interacts with a single known follower type Conitzer and Sandholm [20] have shown that computing the optimal mixed strategy for the leader can be achieved in polynomial time using linear programming.
+However, in more complex Stackelberg games such as Bayesian Stackelberg games, Stackelberg security games, and even 3-player Stackelberg games computing the leader’s optimal strategy has been shown to be NP-hard [20, 28].
+Therefore computational considerations have so far been largely ignored.
+In our setting, there are several interesting cases of structured Stackelberg games for which the leader’s optimal strategy can be learned in polynomial time with provable (albeit not optimal) regret guarantees.
+Observation.
+In a structured Stackelberg game with hypothesis class $\pazocal{H}$ the leader’s optimal policy can be learned in polynomial time if $\pazocal{H}$ is efficiently online learnable.
+We give a high-level overview of this approach.
+First, we consider a learning algorithm for the hypothesis class $\pazocal{H}$.
+The leader can use this algorithm to predict the follower’s type and based on that compute the optimal mixed strategy against that type as if her prediction were correct.
+As shown by Conitzer and Sandholm [20], for a fixed and known follower type, computing the leader’s optimal mixed strategy can be achieved in polynomial time using linear programming.
+The learner will make a finite number of follower type mispredictions before finding $h^{*}$, and at this point will play optimally against every follower type that appears in every subsequent round.
+Therefore, for any hypothesis class in the realizable setting that is efficiently online- (and therefore also PAC-) learnable we can establish a computationally efficient algorithm for learning the leader’s optimal policy.
+We now present one such example of a hypothesis class that can be learned efficiently when used to predict the followers’ types.
+Example.
+[…]
+Theorem 5.1.
+The class of multiclass decision lists is efficiently online learnable.
+We present the full proof of Theorem˜5.1 and the corresponding algorithm in Section˜A.5.
+6 Conclusion and Future Directions
+We initiate the study of structured Stackelberg games, where contextual information is predictive of the follower’s type.
+Our main results include a complete characterization of instance-optimal regret bounds in the online setting and PAC sample complexity guarantees in the distributional setting.
+In this section, we discuss concrete open problems and future directions.
+Computationally efficient algorithms.
+An interesting avenue for future research is to consider the computational complexity of structured Stackelberg game algorithms.
+Although we have shown how to achieve optimal regret bounds, our proposed algorithm is not computationally efficient.
+As we discuss in Section˜5, our model opens the door for efficient, yet so far sub-optimal learning algorithms in certain settings.
+It would be interesting to further explore whether the structure introduced in our work can lead to both optimal and efficient algorithms or if there is an underlying tradeoff between computational complexity and tight regret guarantees.
+Regret guarantees for the agnostic setting.
+Our work focuses on structured Stackelberg games in the realizable setting, where the leader is given a class of hypotheses with the promise that it includes the optimal predictor.
+The agnostic learning setting is a natural relaxation of this assumption.
+While we provide some results (cf.
+Section˜B.1.4) for this setting, our bounds are not instance optimal.
+Providing a complete characterization for the agnostic case, analogously to the realizable one we study here, remains an avenue for future research.
+It is plausible that an entirely new complexity measure will be required here, which would show a clear separation between our setting and multiclass classification, where the Littlestone dimension characterizes both the realizable and the agnostic setting.
+Beyond worst-case analysis frameworks.
+Another direction for future research is to analyze structured Stackelberg games in the beyond worst-case analysis framework.
+It has recently been shown that smooth online learning is as easy as distributional learning, with regret bounds scaling with the VC dimension of the hypothesis class.
+Extending these findings to multiclass learning and structured Stackelberg games might lead to improved results both from a statistical and computational point of view.
+Appendix A Details and Omitted Proofs
+A.1 A technical subtlety
+An important component of our analysis is that the leader must be able to compute the optimal strategy for a given context and follower type.
+For this, we state that for any context $\mathbf{z}\in\pazocal{Z}$ the leader incurs negligible regret by restricting herself to a subset $\mathscr{E}_{\mathbf{z}}\subseteq\Delta(\pazocal{A})$ which is finite and computable.
+To formally state our result we present the notion of a contextual best-response region, which was first introduced in [27] and which we repeat here for completeness.
+Definition A.1 (Contextual Follower Best-Response Region).
+For context $\mathbf{z}$, follower type $f^{(i)}$, and follower action $a_{f}\in\pazocal{A}_{f}$, let $X_{\mathbf{z}}(f^{(i)},a_{f})\subseteq\pazocal{X}$ denote the leader’s mixed strategies such that a follower of type $f^{(i)}$ best-responds to all $x\in X_{\mathbf{z}}(f^{(i)},a_{f})$ by playing action $a_{f}$, that is, $X_{\mathbf{z}}(f^{(i)},a_{f})=\{\mathbf{x}\in\pazocal{X}:b_{f^{(i)}}(\mathbf{z},\mathbf{x})=a_{f}\}$
+Notice that any possible follower best-response region is (1) a subset of $\Delta(\pazocal{A})$ and (2) the intersection of finitely-many halfspaces.
+Therefore, every $X_{\mathbf{z}}(f^{(i)},a_{f})$ is convex and bounded.
+However, the leader may not be able to optimize over the extreme points of a best-response region, as it may not be closed due to the followers’ tie-breaking.
+Definition A.2 ($\delta$-approximate extreme points).
+[…]
+Notice that for sufficiently small $\delta$ i.e.
+$\delta=\mathscr{O}(\frac{1}{T})$, we get that the additional regret of the leader by choosing actions over $\mathscr{E}^{(i)}$ rather than $X(f^{(i)},a_{f})$ is negligible.
+Therefore, we use the shorthand $\mathscr{E}^{(i)}:=\mathscr{E}^{(i)}(\delta)$ for the rest of the section.
+The following lemma from [27] formalizes this notion.
+Lemma A.3.
+For any sequence of followers $f_{1},\ldots f_{T}$ and any leader policy $\pi$, there exists a policy $\pi^{(\mathscr{E})}:\pazocal{Z}\rightarrow\cup_{\mathbf{z}\in\pazocal{Z}}\mathscr{E}_{\mathbf{z}}$ that, when given context $\mathbf{z}$, plays a mixed strategy in $\mathscr{E}_{\mathbf{z}}$ and guarantees that $\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\pi^{(\mathscr{E})}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi^{(\mathscr{E})}(\mathbf{z}_{t})))\leq 1$.
+Moreover, the same result holds in expectation over any distribution over follower types.
+Given Lemma˜A.3 and for ease of notation, in this paper we assume that the leader will always set the parameter $\delta$, such that it does not affect our regret bounds.
+A.2 Proof of Theorem˜3.5
+See 3.5
 Proof.
@@
 Let follower type $f^{(3)}$ have utility function $u_{f^{(3)}}(\mathbf{z},a_{l},a_{f})=\mathbbm{1}\{\mathbf{z}[1]=1\}\cdot u_{f^{(1)}}(\mathbf{z},a_{l},a_{f})+\mathbbm{1}\{\mathbf{z}[1]=2\}\cdot u_{f^{(2)}}(\mathbf{z},a_{l},a_{f})$ and follower type $f^{(4)}$ have utility function $u_{f^{(4)}}(\mathbf{z},a_{l},a_{f})=\mathbbm{1}\{\mathbf{z}[1]=2\}\cdot u_{f^{(1)}}(\mathbf{z},a_{l},a_{f})+\mathbbm{1}\{\mathbf{z}[1]=1\}\cdot u_{f^{(2)}}(\mathbf{z},a_{l},a_{f})$.
+Concretely, one can consider followers with the following utility functions.
+The leader in this case is the row player and the follower is the column player.
+$F_{1}$ $F_{2}$
+$L_{1}$ 1,0.51,0.51 , 0.5 0.5,00.5,00.5 , 0
+$L_{2}$ 0.25,10.25,10.25 , 1 0.75,00.75,00.75 , 0
+(a) Follower type $f^{(1)}$
+$F_{1}$ $F_{2}$
+$L_{1}$ 1,01,01 , 0 $,0.5$ 0.5
+$L_{2}$ 0.25,00.25,00.25 , 0 0.75,10.75,10.75 , 1
+(b) Follower type $f^{(2)}$
+$F_{1}$ $F_{2}$
+$L_{1}$ $1,|\mathbf{z}[1]-2|$ $,0.5$ 0.5
+$L_{2}$ 0.25,0.50.25,0.50.25 , 0.5 $0.75,|\mathbf{z}[1]-1|$
+(c) Follower type $f^{(3)}$
+$F_{1}$ $F_{2}$
+$L_{1}$ $1,|\mathbf{z}[1]-1|$ $,0.5$ 0.5
+$L_{2}$ 0.25,0.50.25,0.50.25 , 0.5 $0.75,|\mathbf{z}[1]-2|$
+(d) Follower type $f^{(4)}$
+Table 3: Utility tables for Theorem˜3.5
 Let
-$h_{\gamma_{1},\gamma_{2}}(\mathbf{z})=\begin{cases}f^{(1)}\;\;&\text{ if }0\leq\mathbf{z}[2]<\gamma_{1}\\ f^{(2)}\;\;&\text{ if }\gamma_{2}<\mathbf{z}[2]\leq 1\\ f^{(3)}\;\;&\text{ if }(\mathbf{z}[1]=1\text{ and }\gamma_{1}\leq\mathbf{z}[2]<\frac{1}{2})\text{ or }(\mathbf{z}[1]=2\text{ and }\gamma_{2}\leq\mathbf{z}[2]\leq 1/2)\\ f^{(4)}\;\;&\text{ if }(\mathbf{z}[1]=2\text{ and }\gamma_{1}\leq\mathbf{z}[2]<\frac{1}{2})\text{ or }(\mathbf{z}[1]=1\text{ and }\gamma_{2}\leq\mathbf{z}[2]\leq 1/2)\\ \end{cases}$
+$h_{\gamma_{1},\gamma_{2}}(\mathbf{z})=\begin{cases}f^{(1)}\;\;&\text{ if }0\leq\mathbf{z}[2]<\gamma_{1}\\ f^{(2)}\;\;&\text{ if }\gamma_{2}<\mathbf{z}[2]\leq 1\\ f^{(3)}\;\;&\text{ if }(\mathbf{z}[1]=1\text{ and }\gamma_{1}\leq\mathbf{z}[2]<0.5)\text{ or }(\mathbf{z}[1]=2\text{ and }\gamma_{2}\leq\mathbf{z}[2]\leq 0.5)\\ f^{(4)}\;\;&\text{ if }(\mathbf{z}[1]=2\text{ and }\gamma_{1}\leq\mathbf{z}[2]<0.5)\text{ or }(\mathbf{z}[1]=1\text{ and }\gamma_{2}\leq\mathbf{z}[2]\leq 0.5)\\ \end{cases}$
 […]
-Figure 1: Follower type as a function of the context $\mathbf{z}$ in the example of Theorem 3.5.
-Observe that while $\gamma_{1}$ and $\gamma_{2}$ vary with each hypothesis, the optimal mixed strategy to play is determined by whether $\mathbf{z}[2]\leq\frac{1}{2}$.
-Theorem 3.5 shows that the Littlestone dimension does not characterize the complexity of online learning in Stackelberg games with side information.
-The example constructed in our proof highlights that the Littlestone dimension is blind to how the hypothesis class interacts with the utilities of the players.
-This observation motivates the definition of a new quantity, which we term the Stakelberg-Littlestone (SL) dimension.
-In order to define the SL dimension, we first need to introduce the notion of an SL-Tree, which is a natural extension of the Littlestone tree to our setting.
-Consider a tree of depth $d$ whose internal nodes are instances in $\pazocal{Z}$ and where the node-to-child edges for a single node, correspond to different labels in a finite set $\mathscr{K}$.
-The tree $T$ is shattered by a hypothesis class $\pazocal{H}$ if for every root-to-leaf path in the tree that traverses the nodes $\mathbf{z}_{1},\dots,\mathbf{z}_{n}$, there exists a function $h\in\pazocal{H}$ such that the label of every edge $(\mathbf{z}_{i},\mathbf{z}_{i+1})$ is $h(\mathbf{z}_{i})$.
-We say that $T$ is shattered by a hypothesis class $\pazocal{H}$ if for every root-to-leaf path in the tree that traverses the nodes $\mathbf{z}_{1},\dots,\mathbf{z}_{n}$, there exists a function $h\in\pazocal{H}$ such that the label of every edge $(\mathbf{z}_{i},\mathbf{z}_{i+1})$ is $h(\mathbf{z}_{i})$.
-Definition 3.6 ($\pazocal{H}$-Shattered Stackelberg-Littlestone Tree).
-For a Stackelberg game, $\pazocal{G}$, let a Stackelberg-Littlestone tree, $T$, be a tree whose internal nodes are instances in $\pazocal{Z}$ and where for each node, the node-to-child edges correspond to different vectors in $\Delta(\pazocal{A})$.
+Figure 3: Follower type as a function of the context in the construction of Theorem˜3.5.
+While $\gamma_{1}$ and $\gamma_{2}$ vary with each hypothesis, the leader’s optimal strategy only depends on whether $\mathbf{z}[2]<0.5$.
+A.3 Alternate definition of the Stackelberg Littlestone dimension
+In this section we provide an alternate definition of the Stackelberg Littlestone dimension that can be defined for infinite trees.
+In Theorem˜A.6 we prove that for finite trees, this is equivalent to Definition˜3.7.
+Definition A.4 ($\kappa$-shattered edge-weighted tree).
+[…]
+  1. 1.
+    (Consistency) $\pazocal{H}$ is consistent with every path in $\pazocal{T}_{d}$ i.e., for any $s\in S_{d}$, there exists $h\in\pazocal{H}$ such that for all $i\in\{\varnothing,1,\dots,d-1\}$ the label of edge $(\mathbf{z}_{s[:i]},\mathbf{z}_{s[:i+1]})$ is $h(\mathbf{z}_{s[:i]})$.
+  2. 2.
+    (Max-branch property) $\pazocal{T}_{d}$ satisfies the max-branch property, i.e., every node, $\mathbf{z}_{s}$, has children $\mathbf{z}_{si}$ for all $i\in[K]$ where $\pazocal{H}$ is consistent with path $si$.
+  3. 3.
+    (Weights) for each node, $\mathbf{z}_{s}\in\pazocal{Z}$ with children $\{\mathbf{z}_{si}\}_{i\in M\subseteq[K]}$ the set of node-to-child edge weights, $\{\nu_{s}^{(i)}\}_{i\in M\subseteq[K]}$, is such that for every $\mathbf{x}\in\Delta(\pazocal{A})$ there exists a follower type $i$ such that $r(\mathbf{z},\mathbf{x},f^{(i)})\geq\nu_{s}^{(i)}$.
+  4. 4.
+    $\inf_{s\in S_{d}}\biggl{[}\sum_{i=0}^{d}\nu_{s[:i]}^{(s[i+1])}\biggr{]}=\kappa$
+Definition A.5 (Stackelberg-Littlestone dimension).
+The Stackelberg-Littlestone dimension of a hypothesis class $\mathscr{H}$ with respect to a Stackelberg game, $\pazocal{G}$, denoted as $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$, is the supremum value of $\kappa$ among all edge-weighted finite trees that are $\kappa$-shattered by hypothesis class $\pazocal{H}$ under Stackelberg game $\pazocal{G}$.
+Theorem A.6.
+Definition˜3.7 and Definition˜A.5 are equivalent for finite trees.
+Proof.
+It suffices to show that any fixed finite-depth tree can be $\kappa$-shattered under Definition˜A.4 if and only if it can be shattered under Definition˜3.6 with root node weight $\kappa$.
+We proceed by induction on the depth of shattered trees.
+(Base Case) Consider a tree, $\pazocal{T}_{1}$, of depth 111 with root node $\mathbf{z}_{\varnothing}$.
+We will argue that the root node weight specified by Definition˜3.6 is equal to the minimum edge weight of the maximum set of valid edge weights under Definition˜A.4 and Definition˜A.5.
+It suffices to show that
+$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})=\max_{N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}}\min_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\{\nu_{\varnothing}^{(i)}:\nu_{\varnothing}^{(i)}\in N\}$
+where $\pazocal{N}_{\mathbf{z}_{\varnothing}}$ is the collection of all sets $N=\{\nu_{\varnothing}^{(i)}\}_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}$ with the property that for all $\mathbf{x}\in\Delta(\pazocal{A})$, there exists an $i\in\pazocal{H}(\mathbf{z}_{\varnothing})$ such that $r(\mathbf{z},\mathbf{x},f^{(i)})\geq\nu_{\varnothing}^{(i)}$.
+We will first show that $\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\max_{N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}}\min_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\{\nu_{\varnothing}^{(i)}:\nu_{\varnothing}^{(i)}\in N\}$.
+Consider any set of valid edge-weights $N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}$.
+By definition we know that for every $\mathbf{x}\in\Delta(\pazocal{A})$, $\exists i\in\pazocal{H}(\mathbf{z}_{\varnothing})$ such that $r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\nu_{\varnothing}^{(i)}$.
+This means that
+$\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\min_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\{\nu_{\varnothing}^{(i)}:\nu_{\varnothing}^{(i)}\in N\}\;\;\;\forall\mathbf{x}\in\Delta(\pazocal{A})\;\;\forall N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}.$
+Considering the $\mathbf{x}\in\Delta(\pazocal{A})$ that minimizes the LHS and an $N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}$ that maximizes the RHS, we get
+$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\max_{N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}}\min_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\{\nu_{\varnothing}^{(i)}:\nu_{\varnothing}^{(i)}\in N\}.$
+We will now show that $\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\leq\max_{N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}}\min_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\{\nu_{\varnothing}^{(i)}:\nu_{\varnothing}^{(i)}\in N\}$.
+Let us denote $\mathbf{x}^{*}=\arg\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})$, and $i^{*}=\arg\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing}})r(\mathbf{z}_{\varnothing},\mathbf{x}^{*},f^{(i)})$.
+Let $N^{*}=\{\nu_{\varnothing}^{(i)}:=r(\mathbf{z}_{\varnothing},x^{*},f^{(i^{*})})\}_{i\in[K]}$.
+We argue that $N^{*}\in\pazocal{N}_{\mathbf{z}_{\varnothing}}$, i.e., that
+$\forall\mathbf{x}\in\Delta(\pazocal{A})\;\;\exists i\in\pazocal{H}(\mathbf{z}_{\varnothing})\;\;\text{such that}\;\;r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq r(\mathbf{z},\mathbf{x}^{*},f^{(i^{*})})$
+and it suffices to argue that $\forall\mathbf{x}\in\Delta(\pazocal{A})$ if $i_{\mathbf{x}}=\textnormal{argmax}_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})$, then $\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})$, which holds trivially.
+(Inductive Step) Say that the two dimensions are equivalent for any tree of depth $d-1$.
+We will show that they are also equivalent for any tree of depth $d$.
+Consider such a tree, $\pazocal{T}_{d}$, with root node weight $\mathbf{z}_{\varnothing}$ and children $\{\mathbf{z}_{0i}\}_{i\in[K]}$.
+We want to show that the root node weight $\rho_{\varnothing}$ defined as in Definition˜3.6 is equal to the weight of the minimum path in the tree, where the edge weights are defined as in Definition˜A.4.
+Concretely, we want to show that
+$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in[K]}\biggl{[}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(j)})+\rho_{\varnothing j}\biggr{]}$ $=\sup_{\pazocal{N}_{d}}\inf_{s\in S_{d}}\Biggl{[}\sum_{i=0}^{d}\nu_{s[:i]}^{(s[i+1])}\Biggr{]},$
+where $\pazocal{N}_{d}$ is the collection of sets of all possible edge weights, one for each node in the tree of depth $d$. This holds if and only if
+$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in[K]}\biggl{[}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(j)})+\rho_{\varnothing j}\biggr{]}$ $=\max_{\pazocal{N}_{0}}\min_{i\in[K]}\Biggl{[}\nu_{\varnothing}^{(s[1])}+\max_{\pazocal{N}_{d-1}}\inf_{s\in S_{d-1}}\sum_{i=1}^{d}\nu_{s[:i]}^{(s[i+1])}\Biggr{]}$
+We know that for any $j\in[K]$ and $s\in S_{d}$ such that $s[1]=j$, $\rho_{\varnothing j}=\max_{\pazocal{N}_{d-1}}\inf_{s\in S_{d-1}}\sum_{i=1}^{d}\nu_{s[:i]}^{(s[i+1])}$ by the inductive hypothesis.
+Therefore, using an argument similar to the base case, the equation above holds.
+∎
+A.4 Proof of Lemma˜3.10
+See 3.10
+Proof.
+We assume for the sake of contradiction that there exists a hypothesis class with $\textnormal{Ldim}(\pazocal{H})=d$ and a Stackelberg game $\pazocal{G}$, such that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})>d$.
+According to Definition˜A.5, this implies that there exists a shattered tree, whose minimum-weighted root-to-leaf path has weight strictly greater than $d$.
+This implies that all paths in the tree have cumulative weight strictly greater than $d$.
+Given that the weight of each edge in the tree is at most 111, the length of the minimum path (and therefore of all other paths in the tree) must be of length at least $d+1$.
+This implies that there exists a tree of depth $d+1$ that is shattered by $\pazocal{H}$, which leads to a contradiction.
+∎
+A.5 Proof of Theorem˜5.1
+Algorithm 3 Multiclass Decision Lists Learning Algorithm
+[…]
+2:Define the set of rules $R=\{(\ell_{i},k_{i})\;\text{such that}\;\ell_{i}\in\{x_{i},\bar{x}_{i}\},x_{i}\in\pazocal{Z},\;k_{i}\in[K]\}$
+3:Set $\tilde{h}=[R]$
+4:for $t=1,2,3,\ldots$ do
+5:  Observe $\mathbf{z}_{t}$
+6:  for  each set $S=\{(\ell_{1},k_{1}),\ldots\}\in\tilde{h}$ do
+7:   if  there exists $(\ell_{i},k_{i})\in S$ such that $\ell_{i}=\mathbf{z}[i]$ then
+8:     Output label $\tilde{k}_{t}=k_{i}$
+9:     Observe the true label $h^{*}(\mathbf{z})$
+10:     if $\tilde{k}_{t}\neq h^{*}(\mathbf{z})$ then
+11:      Set $M=\{(\ell_{i},k_{i})\in S\;|\;\ell_{i}=\mathbf{z}[i]\land k_{i}\neq k_{t}\}$
+12:      Replace $S$ in $\tilde{h}$ with $S\setminus M$ followed by $M$
+13:     else
+14:      Break
+15:   else
+16:     Continue
+17:end
+See 5.1
+Proof.
+[…]
+Appendix B Optimal Strategy Misprediction Bounds
+In the main body of our work we include examples that highlight that existing tools and techniques from multiclass prediction would yield loose regret bounds in our setting (cf.
+proofs of Theorems˜3.5 and 3.11).
+A common theme that emerges from these is that for two distinct follower types, $i,j\in[K]$, the optimal policies $\pi_{*}^{(i)},\pi_{*}^{(j)}:\pazocal{Z}\to\Delta(\pazocal{A})$ for each follower type respectively might coincide for a subset of the context space, that is, they yield the same optimal strategy for the leader.
+Naturally, this prompts us to consider the goal of minimizing the number of timesteps in which the leader mispredicts the optimal strategy.
+Would we get tighter regret bounds if the leader played optimally against this benchmark?
+We answer this in the affirmative, by giving a full characterization of learning in both the online and distributional learning settings.
+While our approaches here perform better than using off-the-shelf classification algorithms, the results we present in the main body strictly supersede them.
+We find this to be for two reasons: (1) what counts as a “misprediction” of the optimal strategy is sensitive to the leader’s tie-breaking rule, and (2) among the suboptimal strategies, some yield higher utility than others, yet each misprediction counts as either adding loss of 1 or 0 to the total.
+B.1 Online Learning
+In this section, we present no-regret learning algorithms for online learning in structured Stackelberg games, that are designed to minimize the number of optimal strategy mispredictions that the leader makes in a time horizon $T$.
+In general, if the regret of an algorithm is sublinear in the time horizon $T$, i.e., $R(T)=o(T)$, we say that the algorithm is a no-regret algorithm.
+We first introduce the notion of a Mistake Stackelberg Littlestone- (MSL-)Tree, which is a natural extension of the Littlestone tree to our setting.
+Definition B.1 ($\pazocal{H}$-Shattered MSL Tree).
+For a Stackelberg game, $\pazocal{G}$, let a MSL tree, $T$, be a tree whose internal nodes are instances in $\pazocal{Z}$ and where for each node, the node-to-child edges correspond to different vectors in $\Delta(\pazocal{A})$.
 We say that $T$ is shattered by a hypothesis class $\pazocal{H}$ with respect to $\pazocal{G}$ if for every root-to-leaf path in the tree that traverses the nodes $\mathbf{z}_{1},\ldots,\mathbf{z}_{n}$, there exists a hypothesis $h\in\mathscr{H}$, such that the label of every edge $(\mathbf{z}_{i},\mathbf{z}_{i+1})$ is $\mathbf{x}_{i}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{i},\mathbf{x},b_{h(\mathbf{z}_{i})}(\mathbf{z}_{i},\mathbf{x}))$.
 In the case of ties, we assume that the leader follows a fixed and arbitrary tie-breaking rule.
-Equipped with these definitions, we are now ready to introduce the Stackelberg-Littlestone dimension, which characterizes the complexity of online learning in our setting.
-Definition 3.7 (Stackelberg-Littlsestone Dimension).
-The Stackelberg-Littlestone dimension of a hypothesis class $\mathscr{H}$ with respect to a Stackelberg game, $\pazocal{G}$, denoted as $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$, is the maximal integer $d$, such that there exists a full binary SL tree of depth $d$ that is shattered by $\pazocal{H}$.
-3.1.1 Lower bound
-We begin by showing that for any online Stackelberg game $\mathscr{G}$, no algorithm can achieve regret smaller than $\Omega(\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H}))$.
-Theorem 3.8.
-There exists a Stackelberg game, $\pazocal{G}$, such that for any hypothesis class $\pazocal{H}$ any deterministic algorithm will suffer at least $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ regret.
-Furthermore, any randomized algorithm will suffer at least $\frac{1}{2}\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ regret in expectation.
+Equipped with this definition, we are now ready to introduce the MSL dimension, which characterizes the complexity of online learning when the leader only cares about minimizing the number of mistakes.
+Definition B.2 (MSL Dimension).
+The MSL dimension of a hypothesis class $\mathscr{H}$ with respect to a Stackelberg game, $\pazocal{G}$, denoted as $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$, is the maximal integer $d$, such that there exists a full binary MSL tree of depth $d$ that is shattered by $\pazocal{H}$.
+B.1.1 Lower bound
+We begin by showing that for any online Stackelberg game $\mathscr{G}$, no algorithm can make fewer than $\Omega(\textnormal{MSLdim}_{\pazocal{G}}(\mathscr{H}))$ mistakes.
+Theorem B.3.
+There exists a Stackelberg game, $\pazocal{G}$, such that for any hypothesis class $\pazocal{H}$ any deterministic algorithm will make at least $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes.
+Furthermore, any randomized algorithm will make at least $\frac{1}{2}\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes in expectation.
 Proof.
-Consider a Stackelberg game, in which for any fixed context, playing the optimal strategy against the true follower type yields utility $1$, whereas playing any other strategy yields utility 00.
-Concretely, let $\beta_{1},\beta_{2},\ldots,\beta_{K}$ be distinct real numbers and consider the following payoff matrix, where the leader is the row player and the column player is follower type $f^{(i)}$.
+Consider a Stackelberg game, in which for any fixed context, playing the optimal strategy against the true follower type yields utility 111, whereas playing any other strategy yields utility 0.
+In other words, the leader gets utility 1 if and only if she correctly predicts the follower type and plays optimally against them.
+Concretely, consider the following game.
+Let $\beta_{1},\beta_{2},\ldots,\beta_{K}$ be distinct, positive real numbers and consider the following payoff matrix, where the leader is the row player and the column player is follower type $f^{(i)}$.
 The leader has 2 pure actions and the follower has 3 pure actions.
+Let $\epsilon\in(0,\min_{i}\beta_{i})$ be a small real number.
+Notice that if the leader puts probability $p>\beta_{i}+\epsilon$ on action $L_{1}$, then the follower will play action $F_{1}$ and the leader gets payoff 0.
+Similarly, if the leader puts probability $p<\beta_{i}-\epsilon$ on action $L_{1}$ then the follower plays action $F_{3}$ and the leader gets payoff 0.
+To guarantee that the leader gets utility 1, the leader must play action $L_{1}$ with probability $p\in(\beta_{i}-\epsilon,\beta_{i}+\epsilon)$.
+Then it is always beneficial for the follower to play $F_{2}$ (this holds no matter how the follower tie-breaks).
+So as long as the intervals $(\beta_{1}-\epsilon,\beta_{1}+\epsilon),(\beta_{2}-\epsilon,\beta_{2}+\epsilon),\ldots,(\beta_{K}-\epsilon,\beta_{K}+\epsilon)$ do not overlap, the leader will get utility 1 if and only if she correctly predicts the follower type she is playing against and plays optimally against them.
 $F_{1}$ $F_{2}$ $F_{3}$
-$L_{1}$ $0,\frac{1}{\beta_{i}}$ $1,1$ $0,0$
-$L_{2}$ $0,0$ $1,1$ $0,\frac{1}{1-\beta_{i}}$
-Notice that if the leader puts probability $p>\beta_{i}$ on action $L_{1}$, then the follower will play action $F_{1}$ and the leader gets payoff 00.
-Similarly, if the leader puts probability $p<\beta_{i}$ on action $L_{1}$ then the follower plays action $F_{3}$ and the leader gets payoff 00.
-However, if the leader puts probability $p=\beta_{i}$ on action $L_{1}$, then the follower is indifferent and since the follower breaks ties in favor of the leader, she plays action $F_{2}$, giving the leader a payoff of $1$.
-(Claim 1.) An adversary can force the learner to make $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes by using the following strategy: (1) select $\mathbf{z}_{t}$ as the next input from an SL-Tree of depth $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ that is shattered by $\pazocal{H}$ and (2) set the optimal mixed strategy, $\mathbf{x}_{t}^{*}$ to be the edge in the shattered tree not chosen by the leader, that is $\mathbf{x}_{t}^{*}\neq\widehat{\mathbf{x}}_{t}$, where $\widehat{\mathbf{x}}_{t}$ is the strategy played by the learner at time $t$.
-Since the tree is shattered by $\pazocal{H}$, there must exist some $h\in\pazocal{H}$ that is consistent with $(\mathbf{z}_{t},f_{t})_{t=1}^{\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})}$.
-Therefore, any deterministic learner will make at least $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes.
+$L_{1}$ $0,\frac{1}{\beta_{i}+\epsilon}$ $,1$ 1 $,0$ 0
+$L_{2}$ $,0$ 0 $,1$ 1 $0,\frac{1}{1-\beta_{i}+\epsilon}$
+Table 4: Game construction for Theorem˜B.3
+(Claim 1.) An adversary can force the learner to make $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes by using the following strategy: (1) select $\mathbf{z}_{t}$ as the next input from an MSL-Tree of depth $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ that is shattered by $\pazocal{H}$ and (2) set the optimal mixed strategy, $\mathbf{x}_{t}^{*}$ to be the edge in the shattered tree not chosen by the leader, that is $\mathbf{x}_{t}^{*}\neq\widehat{\mathbf{x}}_{t}$, where $\widehat{\mathbf{x}}_{t}$ is the strategy played by the learner at time $t$.
+Since the tree is shattered by $\pazocal{H}$, there must exist some $h\in\pazocal{H}$ that is consistent with $(\mathbf{z}_{t},f_{t})_{t=1}^{\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})}$.
+Therefore, any deterministic learner will make at least $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes.
 (Claim 2.) When the learner is randomized, at each timestep $t$, the adversary can select an $\mathbf{x}_{t}^{*}$ such that it minimizes the probability that the leader chooses $\widehat{\mathbf{x}}_{t}=\mathbf{x}_{t}^{*}$.
@@
 ∎
-3.1.2 Upper bounds
-Having established that the Stackelberg-Littlestone dimension characterizes the difficulty of learning in our setting, in this section we present an algorithm and show that it achieves optimal regret.
-Our guarantees in this setting are with respect to the leader’s number of mistakes (i.e., timesteps where the utility of the leader is not optimal).
-By the realizability assumption and the boundedness of the utility function, all mistake bounds imply the same regret bounds.
+B.1.2 Upper bounds
+Having established that the MSL dimension lower-bounds the number of mispredictions of the optimal strategy for any learner, in this section we present an algorithm and show that it achieves the optimal mistake bound.
 Warm-up: context-free utilities.
-As a warm-up to our main results, consider a simpler setting, in which both the leader and the follower utility functions do not take the context as an input (i.e.
+As a warm-up, consider a simpler setting, in which both the leader and the follower utility functions do not take the context as an input (i.e.
 they are context-free).
 This means that the context at each timestep is now only predictive of the follower’s type at that timestep.
-In Theorem 3.9, we show that we can reduce the problem of learning the optimal policy of the leader to online multiclass prediction.
-Additionally, the known optimal algorithm in online multiclass prediction (the Standard Optimal Algorithm (SOA)) meets the lower bound shown in Theorem 3.8.
-Theorem 3.9.
+In Theorem˜B.4, we show that we can reduce the problem of learning the optimal policy of the leader to online multiclass prediction.
+Additionally, the known optimal algorithm in online multiclass prediction (the Standard Optimal Algorithm (SOA)) meets the lower bound shown in Theorem˜B.3.
+Theorem B.4.
 Consider a Stackelberg game, $\pazocal{G}$, such that the utilities of both the leader and the follower are context-free.
-Then for any finite class, $\pazocal{H}$, of hypotheses $h:\pazocal{Z}\to\mathscr{K}$, there exists a reduction of this setting to multiclass prediction that runs in $O(\text{poly}(K,|\pazocal{A}_{f}|,|\pazocal{A}|,|\pazocal{H}|))$ time and achieves regret at most $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
+Then for any finite class, $\pazocal{H}$, of hypotheses $h:\pazocal{Z}\to[K]$, there exists a reduction of this setting to multiclass prediction that runs in $O(\text{poly}(K,|\pazocal{A}_{f}|,|\pazocal{A}|,|\pazocal{H}|))$ time and makes at most $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes.
 Proof.
@@
 Finally, we run the multiclass Standard Optimal Algorithm (see [22]) on $\pazocal{H}^{\prime}$ considering every optimal strategy for the leader in the look-up table as one label.
-It remains to show that $\textnormal{Ldim}(\pazocal{H}^{\prime})=\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
+It remains to show that $\textnormal{Ldim}(\pazocal{H}^{\prime})=\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$.
 Notice that $\pazocal{H}^{\prime}$ is the set of all hypotheses, mapping each context to an optimal strategy for the leader against every follower type that is consistent with $\pazocal{H}$ for that context.
-Consider any SL tree and any Littlestone tree such that the context at each corresponding node is the same among them.
-If the SL tree is shattered by $\pazocal{H}$ then for every root-to-leaf path there must exist a hypothesis in $\pazocal{H}$ that is consistent with the labels along that path.
+Consider any MSL tree and any Littlestone tree such that the context at each corresponding node is the same among them.
+If the MSL tree is shattered by $\pazocal{H}$ then for every root-to-leaf path there must exist a hypothesis in $\pazocal{H}$ that is consistent with the labels along that path.
 But by definition of $\pazocal{H}^{\prime}$ for that corresponding path in the Littlestone tree there must exist an $h^{\prime}\in\pazocal{H}^{\prime}$ that is consistent with the labelings of that path.
 With the same argument, the converse is also true.
-Therefore, the SL tree with maximal depth that is shattered by $\pazocal{H}$ is precisely the generalized Littlestone tree with maximal depth that is shattered by $\pazocal{H}^{\prime}$.
+Therefore, the MSL tree with maximal depth that is shattered by $\pazocal{H}$ is precisely the generalized Littlestone tree with maximal depth that is shattered by $\pazocal{H}^{\prime}$.
 ∎
-We note that the algorithm described above coincides with Algorithm 1 (presented below), which we prove to be optimal in the general contextual Stackelberg game setting.
+We note that the algorithm described above coincides with Algorithm˜4 (presented below), which we prove to be optimal in the general contextual Stackelberg game setting.
 Context-dependent utilities.
 We now let the utilities of both players to be functions of the context.
-We present our main algorithm (Algorithm 1), which we show achieves regret that matches the lower bound in Theorem 3.8 for any Stackelberg game.
+We present our main algorithm (Algorithm˜4), which we show achieves regret that matches the lower bound in Theorem˜B.3 for any Stackelberg game.
 Our approach is inspired by the Standard Optimal Algorithm for online multiclass prediction, with the main difference that our algorithm uses a hypothesis class, $\pazocal{H}$, to optimize directly over leader strategies in the simplex.
@@
 For every such follower type the algorithm computes the optimal leader strategy against them.
-Then it plays the strategy that maximizes the Stackelberg Littlestone dimension of the set of surviving hypotheses were that to be the optimal strategy for that timestep.
-If the algorithm makes a mistake, then the set of surviving hypotheses will have strictly smaller Stackelberg-Littlestone dimension than before.
+Then it plays the strategy that maximizes the MSL dimension of the set of surviving hypotheses were that to be the optimal strategy for that timestep.
+If the algorithm makes a mistake, then the set of surviving hypotheses will have strictly smaller MSL dimension than before.
 In other words, the algorithm learns from this mistake.
 Notice that the number of possible strategies in the simplex is infinite, yet our algorithm only considers a bounded number of them at each timestep, at most one for each follower type.
-Algorithm 1 Stackelberg Standard Optimal Algorithm (SSOA)
-1:input: time horizon, $T$, Stackelberg game $\pazocal{G}$, and hypothesis class $\pazocal{H}$, where $h:\pazocal{Z}\to\mathscr{K}$ for every $h\in\pazocal{H}$.
+Algorithm 4 Mistake-Stackelberg Standard Optimal Algorithm (MSSOA)
+1:Input: time horizon, $T$, Stackelberg game $\pazocal{G}$, and hypothesis class $\pazocal{H}$, where $h:\pazocal{Z}\to[K]$ for every $h\in\pazocal{H}$.
 2:Set $V_{0}=\pazocal{H}$
@@
 4:  Observe $\mathbf{z}_{t}$
-5:     Let $\mathbf{x}^{(h,t)}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{h(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))\quad\forall h\in\pazocal{H}$
+5:  Let $\mathbf{x}^{(h,t)}=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{h(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))\quad\forall h\in\pazocal{H}$
 6:  Let $X_{t}=\{\mathbf{x}^{(h)}:h\in V_{t-1}\}$triangleright Note that $|X_{t}|\leq K$
 7:  Let $V_{t}^{(\mathbf{x})}=\{h\in V_{t-1}:\mathbf{x}^{(h,t)}=\mathbf{x}\}\quad\forall\mathbf{x}\in X_{t}$
-8:     Play action $\mathbf{x}_{t}=\arg\max_{\mathbf{x}\in X_{t}}\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})$
+8:  Play action $\mathbf{x}_{t}=\arg\max_{\mathbf{x}\in X_{t}}\textnormal{MSLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})$
 9:  Observe $f_{t}$
 10:  Set $V_{t}=\{h\in V_{t-1}:h(\mathbf{z}_{t})=f_{t}\}$
-11:end
-Theorem 3.10.
-For any Stackelberg game, $\pazocal{G}$ and every hypothesis class, $\pazocal{H}$, the regret of Algorithm 1 is at most $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
+Theorem B.5.
+For any Stackelberg game, $\pazocal{G}$ and every hypothesis class, $\pazocal{H}$, Algorithm˜4 makes at most $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes.
 Proof.
-Let $T_{e}\subseteq[T]$ be the set of time steps on which SSOA errs and let $T_{c}=[T]\setminus T_{e}$ be the set of time steps on which SSOA correctly predicts the optimal strategy against the true follower type.
-Recall from Remark 3.2 that we can make $\sum_{t\in T_{c}}u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$ arbitrarily small.
+Let $T_{e}\subseteq[T]$ be the set of time steps on which MSSOA errs and let $T_{c}=[T]\setminus T_{e}$ be the set of time steps on which MSSOA correctly predicts the optimal strategy against the true follower type.
+Recall from Remark˜3.2 that we can make $\sum_{t\in T_{c}}u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$ arbitrarily small.
 Now from the boundedness of the utility function, we also know that $\sum_{t\in T_{e}}u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))\leq|T_{e}|$.
-It suffices to show that for every time step $t\in T_{e}$ that SSOA makes a mistake, we have that $\textnormal{SLdim}_{\pazocal{G}}(V_{t})\leq\textnormal{SLdim}_{\pazocal{G}}(V_{t+1})-1$.
-For the sake of contradiction, assume that $\textnormal{SLdim}_{\pazocal{G}}(V_{t})=\textnormal{SLdim}_{\pazocal{G}}(V_{t+1})$.
-Given that SSOA chooses a strategy $\mathbf{x}$ that maximizes $\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})$ and that SSOA made a mistake at this timestep, there must have been at least two distinct strategies $\mathbf{x},\mathbf{x}^{\prime}$ such that $\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})=\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x}^{\prime})})=\textnormal{SLdim}_{\pazocal{G}}(V_{t})$.
-However, this would mean that we could construct a full binary SL tree that is shattered by $\pazocal{H}$ and that has depth $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})+1$, which leads to a contradiction.
-3.1.3 Relationship with Littlestone dimension
-We have so far shown that the Stackelberg-Littlestone dimension characterizes learnability in our setting.
-We next explore how that relates to multiclass learnability.
-While, as discussed, learning to predict the follower type from the context is a sufficient condition for learning the leader’s optimal policy, in this section we prove that this is not a necessary condition.
-In Theorem 3.11, we show that the Littlestone dimension is always at most as large as the Stackelberg Littlestone dimension and in Proposition LABEL:prop:ldim=sldim we show that equality always holds in the special case of two follower types.
-In Theorem 3.12, we show that an out-of-the-box application of the optimal algorithm for multiclass learning in our setting can result not only in suboptimal number of mistakes, but also in suboptimal utility for the leader.
-This is in contrast to our proposed algorithm, which achieves optimal utility by accounting for the interaction between the hypothesis class and the Stackelberg game.
-Theorem 3.11.
-For any hypothesis class $\pazocal{H}$, $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})\leq\textnormal{Ldim}(\pazocal{H})$.
+It suffices to show that for every time step $t\in T_{e}$ that MSSOA makes a mistake, we have that $\textnormal{MSLdim}_{\pazocal{G}}(V_{t+1})\leq\textnormal{MSLdim}_{\pazocal{G}}(V_{t})-1$.
+For the sake of contradiction, assume that $\textnormal{MSLdim}_{\pazocal{G}}(V_{t})=\textnormal{MSLdim}_{\pazocal{G}}(V_{t+1})$.
+Given that MSSOA chooses a strategy $\mathbf{x}$ that maximizes $\textnormal{MSLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})$ and that MSSOA made a mistake at this timestep, there must have been at least two distinct strategies $\mathbf{x},\mathbf{x}^{\prime}$ such that $\textnormal{MSLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})=\textnormal{MSLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x}^{\prime})})=\textnormal{MSLdim}_{\pazocal{G}}(V_{t})$.
+However, this would mean that we could construct a full binary MSL tree that is shattered by $\pazocal{H}$ and that has depth $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})+1$, which leads to a contradiction.
+B.1.3 Using multiclass SOA as a black-box algorithm
+We consider the following algorithm that uses SOA [22] to predict the follower’s type given the context.
+Then proceeds to play the optimal strategy against that follower type as if the prediction were correct.
+Algorithm 5 Learning with full-feedback and realizable follower types
+1:for $t\in[T]$ do
+2:  Observe $\mathbf{z}_{t}$
+3:  Evaluate $\hat{f}_{t}=h_{t}(\mathbf{z}_{t})\leftarrow\text{SOA}(\mathbf{z}_{t})$
+4:  Commit to $\mathbf{x}_{t}=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{\hat{f}_{t}}(\mathbf{z}_{t},\mathbf{x}))$
+5:  Observe follower type $f_{t}$
+6:end
+Lemma B.6.
+There exists a hypothesis class $\pazocal{H}$ and a Stackelberg game $\pazocal{G}$ such that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})<\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})<\textnormal{Ldim}(\pazocal{H})$.
+Further, SOA achieves strictly larger regret than MSSOA, which in turn achieves strictly larger regret than SSOA.
 Proof.
-If the leader plays SOA, she will make at most $\textnormal{Ldim}(\pazocal{H})$ mistakes.
-This is because she will predict the follower’s type incorrectly at most $\textnormal{Ldim}(\pazocal{H})$ times, and predicting the follower’s type correctly is a sufficient condition for playing the optimal mixed strategy.
-Since the SOA makes $\textnormal{Ldim}(\pazocal{H})$ mistakes and we know that any deterministic algorithm will make at least $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes by Theorem 3.8, we can infer that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})\leq\textnormal{Ldim}(\pazocal{H})$.
-∎
-While it is known that SOA makes at least $\textnormal{Ldim}(\pazocal{H})$ mistakes when learning a hypothesis class, $\pazocal{H}$, a careful reader might notice that this does not immediately show that SOA achieves suboptimal regret in our setting.
-This is because mispredicting the follower type does not necessarily imply that the optimal strategy against that and the true follower type do not yield the same utility for the leader.
-In Theorem 3.12, we prove that, indeed, there are instances where SOA will achieve smaller cumulative utility for the leader compared to our proposed algorithm.
-Theorem 3.12.
-There exists some hypothesis class $\pazocal{H}$ and a Stackelberg game, $\pazocal{G}$, such that running multiclass SOA on $\pazocal{H}$ yields suboptimal utility for the leader.
+We revisit the game presented in the proof of Theorem˜3.11.
+We have shown that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=3/4$ and $\textnormal{Ldim}(\pazocal{H})=2$.
+We will now show that $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})=1$.
+We consider any possible MSL tree.
+First, note that the context $z=4$ cannot be a node in an $\pazocal{H}$-shattered MSL-Tree since for $z=4,\;\pi_{*}^{(2)}(4)=\pi_{*}^{(3)}(4)$, where $f^{(2)}$ and $f^{(3)}$ are the only possible follower types predicted by hypotheses in $\pazocal{H}$ for this context.
+Similarly, $z=3$ cannot be a node in an $\pazocal{H}$-shattered SL tree, since the leader’s optimal strategy against all follower types is identical ($\pi_{*}^{(1)}(3)=\pi_{*}^{(2)}(3)=\pi_{*}^{(3)}(3)$).
+So it suffices to consider contexts 111 and 222.
+In Figure˜4 we show the two $\pazocal{H}$-shattered MSL tree constructions with $z=1$ and $z=2$ as the root nodes respectively.
+Notice that the maximal depth of a full binary tree is 111, as desired.
+We now prove the regret values.
+First, since $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ tightly characterizes the regret of SSOA, we know that the regret of SSOA is 3/43/43 / 4.
+Also, from Theorem˜3.11 we immediately get that the cumulative regret of SOA equals 2 when the sequence of contexts is 111, followed by 222.
+We now proceed with the regret of MSSOA for the same sequence.
+  • ($t=1$) This root node has two branches: $\pi_{*}^{(1)}(1)=\pi_{*}^{(2)}(1)$ and $\pi_{*}^{(3)}(1)$.
+Given the arguments used to prove that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=1$, the former branch leads to a full binary subtree of depth 111, while the second branch cannot lead to a full binary tree.
+So SSOA chooses $x_{1}=\pi_{*}^{(1)}(1)=\pi_{*}^{(2)}(1)$ and the true follower type is $f_{1}^{*}=f^{(1)}$ (no instantaneous regret).
+  • ($t=2$) The context $z=2$ leads to two branches: $\pi_{*}^{(2)}(2)$ and $\pi_{*}^{(3)}(2)$, none of which leads to a complete binary tree.
+SSOA tie breaks, in which case an adversary can always force a mistake by choosing the correct follower type to be the one not predicted by MSSOA (thereby incurring instantaneous regret 1).
+Figure 4: $\pazocal{H}$-shattered MSL-Tree with (a) $z=1$ and (b) $z=2$ as the root node.
+Theorem B.7.
+For every online contextual Stackelberg game with $K$ distinct follower types and every hypothesis class $\pazocal{H}$, there exists a hypothesis class $\Pi$ such that running multiclass SOA on $\Pi$ incurs $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes.
+Moreover, this algorithm coincides with MSSOA.
 Proof.
-Consider $\pazocal{Z}=\{1,2,3,4\}$ and $\mathscr{K}=\{f^{(1)},f^{(2)},f^{(3)}\}$.
-We define $\pazocal{H}=\{h_{\sigma}(z):=\mathbbm{1}\{z\neq 4\}\cdot f^{(\sigma(z))}+\mathbbm{1}\{z=4\}\cdot f^{(3)}\}_{\sigma\in\Sigma}\cup\{h,h^{\prime}\}$ , where $\Sigma$ is the set of all permutations of the set $\{1,2,3\}$ and $h,h^{\prime}$ are defined as follows.
-$h(z)=\begin{cases}f^{(3)}\;\;&\text{ if }z=1\\ f^{(2)}\;\;&\text{ if }z=2\\ f^{(1)}\;\;&\text{ if }z=3\\ f^{(2)}\;\;&\text{ if }z=4\\ \end{cases}\hskip 56.9055pth^{\prime}(z)=\begin{cases}f^{(3)}\;\;&\text{ if }z=1\\ f^{(1)}\;\;&\text{ if }z=2\\ f^{(2)}\;\;&\text{ if }z=3\\ f^{(2)}\;\;&\text{ if }z=4\\ \end{cases}$
-Notice that the Littlestone dimension of this hypothesis class is at least $2$. This follows from the $\pazocal{H}$-shattered tree construction shown in Figure 2.
-Figure 2: Construction of an $\pazocal{H}$-shattered Littlestone tree. Each node represents a context in $\pazocal{Z}$ and each edge corresponds to a labeling of the parent node to a follower type.
-Consider now a Stackelberg game with utility functions $u^{(1)},u^{(2)},u^{(3)}$ corresponding to follower types $1,2$, and $3$ respectively, defined such that if $z=1$ or $z=2$ then $\pi_{*}^{(1)}(z)=\pi_{*}^{(2)}(z)\neq\pi_{*}^{(3)}(z)$, if $z=3$ then $\pi_{*}^{(1)}(3)=\pi_{*}^{(2)}(3)=\pi_{*}^{(3)}(3)$, and if $z=4$, then $\pi_{*}^{(1)}(4)\neq\pi_{*}^{(2)}(4)=\pi_{*}^{(3)}(4)$.
-Here, we let $\pi_{*}^{(i)}:\pazocal{Z}\to\Delta(\pazocal{A})$ denote the optimal leader’s policy against follower type $f^{(i)}$.
-Concretely, consider a Stackelberg game with the leader being the row player and each follower being the column player, as shown in Table 1.
-$F_{1}$ $F_{2}$
-$L_{1}$ $2,1$ $0,0$
-$L_{2}$ $0,0$ $2,0$
-(a) Follower type $f^{(1)}$
-$F_{1}$ $F_{2}$
-$L_{1}$ $2,\mathbbm{1}_{\{z\neq 4\}}$ $0,0$
-$L_{2}$ $0,0$ $2,\mathbbm{1}_{\{z=4\}}$
-(b) Follower type $f^{(2)}$
-$F_{1}$ $F_{2}$
-$L_{1}$ $2,\mathbbm{1}_{\{z=3\}}$ $0,0$
-$L_{2}$ $0,0$ $2,\mathbbm{1}_{\{z\neq 3\}}$
-(c) Follower type $f^{(3)}$
-Table 1: Utility tables for the leader (row player) and each follower type (column player)
-We now show that the $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=1$.
-We will consider any possible SL tree.
-First, note that the context $z=4$ cannot be a node in an $\pazocal{H}$-shattered SL-Tree since for $z=4,\;\pi_{*}^{(2)}(4)=\pi_{*}^{(3)}(4)$, where $f^{(2)}$ and $f^{(3)}$ are the only possible follower types predicted by hypotheses in $\pazocal{H}$ for this context.
-Similarly, $z=3$ cannot be a node in an $\pazocal{H}$-shattered SL tree, since the leader’s optimal strategy against all follower types is identical ($\pi_{*}^{(1)}(3)=\pi_{*}^{(2)}(3)=\pi_{*}^{(3)}(3)$).
-So it suffices to consider contexts $1$ and $2$.
-In Figure 3 we show the two $\pazocal{H}$-shattered SL tree constructions with $z=1$ and $z=2$ as the root nodes respectively.
-Notice that the maximal depth of a full binary tree is $1$, as desired.
-Say that the adversary presents context $1$ followed by context $2$. We go through the SOA outputs per timestep below.
-  • ($t=1$) This node has three children corresponding to follower types $f^{(1)},f^{(2)},f^{(3)}$, with $f^{(1)}$ and $f^{(2)}$ leading to a full binary subtree of maximal depth 1 each.
-To see this, notice that (1) neither $h$ nor $h^{\prime}$ are consistent with the mapping of context $z=1$ to follower type $f^{(1)}$ or $f^{(2)}$, and (2) $\pazocal{H}\setminus\{h,h^{\prime}\}$ consists of all permutations mapping contexts $\{1,2,3\}$ to follower types $\{f^{(1)},f^{(2)},f^{(3)}\}$.
-Therefore, at most two of these contexts can exist in any root-to-leaf path of any $\pazocal{H}$-shattered tree.
-The child corresponding to follower type $f^{(3)}$ leads to a full binary subtree of maximal depth at least $2$, as shown in Figure 2.
-So, the predicted follower type is $f_{1}=f^{(3)}$ and the strategy played is $x_{1}=\pi_{*}^{(3)}(1)$.
-However, the true follower type is $f_{1}^{*}=f^{(1)}$ (mistake).
-  • ($t=2$) This node has two children mapping $z=2$ to follower types $f^{(2)}$ and $f^{(3)}$.
-No subsequent context can lead to a full-binary tree that is consistent with any of these assignments, so SOA tie-breaks.
-The adversary in this case can always force a mistake by claiming the correct follower type to be the one not predicted by the algorithm.
-So if SOA predicts $f_{2}=f^{(2)}$, and plays strategy $x_{2}=\pi_{*}^{(2)}(2)$ the adversary can choose the true follower type to be $f_{2}^{*}=f^{(3)}$ (mistake).
-This leads to SOA mispredicting the follower type and getting suboptimal utility in both rounds.
-Since the Stackelberg Littlestone dimension is $1$ and we have shown that for the above sequence of contexts SOA makes $2$ mistakes, this concludes our proof.
-For ease of exposition, we also provide an overview of SSOA for the same series of nodes (see Figure (3(a)) for reference).
-  • ($t=1$) This root node has two branches: $\pi_{*}^{(1)}(1)=\pi_{*}^{(2)}(1)$ and $\pi_{*}^{(3)}(1)$.
-Given the arguments used to prove that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=1$, the former branch leads to a full binary subtree of depth $1$, while the second branch cannot lead to a full binary tree.
-So SSOA chooses $x_{1}=\pi_{*}^{(1)}(1)=\pi_{*}^{(2)}(1)$ and the true follower type is $f_{1}^{*}=f^{(1)}$ (no mistake).
-  • ($t=2$) The context $z=2$ leads to two branches: $\pi_{*}^{(2)}(2)$ and $\pi_{*}^{(3)}(2)$, none of which leads to a complete binary tree.
-SSOA tie breaks, in which case an adversary can always force a mistake by choosing the correct follower type to be the one not predicted by SSOA (mistake).
-Figure 3: $\pazocal{H}$-shattered SL-Tree with (a) $z=1$ as the root node and (b) $z=2$ as the root node.
-While we have shown that running multiclass SOA alone can in certain settings achieve suboptimal utility for the leader, in the following theorem we show that our algorithm may be interpreted as running SOA in a transformed space.
-Theorem 3.13.
-For every online contextual Stackelberg game with $K$ distinct follower types and every hypothesis class $\pazocal{H}$, there exists a hypothesis class $\Pi$ such that running multiclass SOA on $\Pi$ incurs $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes.
-Moreover, this algorithm coincides with SSOA.
-Proof.
-Consider a mapping $\zeta:\pazocal{H}\to\Pi$ that maps every $h\in\pazocal{H}$ to a function $\pi_{h}:\pazocal{Z}\to\Delta(\pazocal{A})$ such that:
-$\pi_{h}(\mathbf{z}):=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x})).$
-Notice that $\zeta$ is a bijection given that each follower type is distinct (see Remark 2.1).
+Consider a mapping $\zeta:\pazocal{H}\to\Pi$ that maps every $h\in\pazocal{H}$ to a function $\pi_{h}:\pazocal{Z}\to\Delta(\pazocal{A})$ such that $\pi_{h}(\mathbf{z}):=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x})).$
+Notice that $\zeta$ is a bijection given that each follower type is distinct (cf.
+Remark˜2.1).
 Also note that given that learning $\pazocal{H}$ is a realizable problem, so is learning $\Pi$.
-This is because if there exists $h^{*}\in\pazocal{H}$ such that $\forall t\in[T]\quad h^{*}(\mathbf{z}_{t})=f_{t}$, then by definition there must exist $\pi^{*}\in\Pi$ such that $\forall t\in[T]\quad\pi^{*}(\mathbf{z}_{t})=\mathbf{x}^{*}_{t}$, since $\mathbf{x}^{*}_{t}=\arg\max_{\mathbf{x}\in\Delta(A)}u(\mathbf{z}_{t},\mathbf{x},b_{h^{*}(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))$.
-We know that multiclass SOA makes at most $\textnormal{Ldim}(\Pi)$ mistakes for any hypothesis class $\Pi$ with infinite labels (see Theorem 1 in [22]).
-Therefore, it remains to show that $\textnormal{Ldim}(\Pi)=\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
-Consider any SL-Tree for $\pazocal{H}$.
+This is because if there exists $h^{*}\in\pazocal{H}$ such that $\forall t\in[T]\quad h^{*}(\mathbf{z}_{t})=f_{t}$, then by definition there must exist $\pi^{*}\in\Pi$ such that $\forall t\in[T]\quad\pi^{*}(\mathbf{z}_{t})=\mathbf{x}^{*}_{t}$, since $\mathbf{x}^{*}_{t}=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{h^{*}(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))$.
+We know that multiclass SOA makes at most $\textnormal{Ldim}(\Pi)$ mistakes for any hypothesis class $\Pi$ with infinite labels (cf.
+Theorem 1 in [26]).
+Therefore, it remains to show that $\textnormal{Ldim}(\Pi)=\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$.
+Consider any MSL-Tree for $\pazocal{H}$.
 Notice that all nodes are contexts and all edges represent optimal leader strategies, so this is a valid Littlestone tree for $\Pi$ as well.
@@
 Then by $\zeta$ being a bijection, there exists an $h\in\pazocal{H}$ that is also consistent with it.
-Therefore, a tree is SL-shattered by $\pazocal{H}$ if and only if it is shattered by $\Pi$.
+Therefore, a tree is MSL-shattered by $\pazocal{H}$ if and only if it is shattered by $\Pi$.
 This holds for all trees, so it will hold for the maximal-depth trees as well.
-3.2 No-regret learning in the agnostic setting
+B.1.4 No-regret learning in the agnostic setting
 In this section we relax the realizability assumption, and we consider the setting where the sequence of follower types $f_{1},\cdots,f_{T}$ is not necessarily consistent with some hypothesis $h\in\mathscr{H}$.
@@
 We define $h^{*}\in\arg\max_{h\in\mathscr{H}}\sum_{t\in[T]}\mathbbm{1}\{h(\mathbf{z}_{t})=y_{t}\}$ and generalize the notion of regret:
-Definition 3.14 ((Agnostic Setting) Contextual Stackelberg Regret).
+Definition B.8 ((Agnostic Setting) Contextual Stackelberg Regret).
 Given a sequence of followers $f_{1},\dots,f_{T}$ and contexts $\mathbf{z}_{1},\dots\mathbf{z}_{T}$, the leader’s contextual Stackelberg regret in the agnostic setting is given by:
 $R(T):=\max_{h\in\mathscr{H}}\sum_{t\in[T]}u(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t})),$
-where $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$ is the sequence of mixed strategies played and $\pi_{h}(\mathbf{z}):=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x}))$.
-Combining our SSOA algorithm with results and tools from multiclass prediction [17], we get the following regret bound.
-Theorem 3.15.
+where $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$ is the sequence of mixed strategies played and $\pi_{h}(\mathbf{z}):=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x}))$.
+Algorithmic procedure.
+Consider the “learning with expert advice” setting where the leader at each timestep $t$ would get to observe not only the context $\mathbf{z}_{t}$ but also $N$ advices, $\{\phi_{t}^{(1)},\cdots,\phi_{t}^{(N)}\}\in\Delta(\pazocal{A})^{N}$.
+Then we can denote the cumulative loss of expert $i$ at timestep $T$ as:
+$L_{T,i}=-\sum_{t=1}^{T}u(\mathbf{z}_{t},\phi_{t}^{(i)},b_{f_{t}}(\mathbf{z}_{t},\phi_{t}^{(i)})).$
+For an appropriately defined set of experts we can formalize the following algorithm.
+Algorithm 6 Learning in the full-feedback, agnostic setting
+1:input: time horizon $T$
+2:Set $\eta=\sqrt{8\ln(N)/T}$
+3:for $t\in[T]$ do
+4:  Observe $\mathbf{z}_{t}$
+5:  Receive advices $\{\phi_{t}^{(1)},\cdots,\phi_{t}^{(N)}\}\in\Delta(A)^{N}$
+6:  Commit to $\mathbf{x}_{t}=\phi_{t}^{(i)}$ with probability proportional to $\exp(-\eta L_{t-1,i})$
+7:  Receive utility $u(\mathbf{z}_{t},a_{l},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$, where $a_{l}$ is drawn from $\mathbf{x}_{t}$
+8:  Observe follower type $f_{t}$
+9:end
+Combining our MSSOA algorithm with the above results and tools from multiclass prediction [22], we get the following regret bound.
+Theorem B.9.
 Consider a Stackelberg game, $\pazocal{G}$, and a hypothesis class, $\pazocal{H}$.
-In the agnostic setting, there exists an algorithm that achieves regret $R(T)\leq\sqrt{\frac{1}{2}\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H})T\log{(TK)}}$.
+In the agnostic setting, there exists an algorithm that achieves regret $R(T)\leq\sqrt{\frac{1}{2}\textnormal{MSLdim}_{\pazocal{G}}(\mathscr{H})T\log{(TK)}}$.
+Proof.
+First consider the case where we construct an expert for every $h\in\mathscr{H}$ and keep track of their cumulative loss $L_{t,h}$ at each timestep $t$.
+In our case, for each $h\in\pazocal{H}$, we can construct an expert $i$ that will predict the type, $f_{t}^{(i)}=h(\mathbf{z}_{t})$ at that timestep and based on that evaluate
+$\phi_{t}^{(i)}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u(\mathbf{z}_{t},a_{l},b_{f_{t}^{(i)}}(\mathbf{z}_{t},\mathbf{x})).$
+From the regret analysis of the Hedge algorithm (see [19] Chapter 2), we know that
+$\sum_{t\in[T]}\mathbb{E}\left[u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))\right]\geq\max_{i\in[N]}\sum_{t=1}^{T}u(\mathbf{z}_{t},\phi_{t}^{(i)},b_{f_{t}}(\mathbf{z}_{t},\phi_{t}^{i}))-\sqrt{\frac{1}{2}\ln(|\mathscr{H}|)T},$
+where the expectation is taken over the randomness in the Hedge algorithm.
+Now it remains to construct a smaller set of experts such that the upper bound on the regret depends on the $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$, instead of $\ln(|\pazocal{H}|)$.
+For some time horizon $T$, we define $S_{T}=\{S\subset[T]:|S|\leq\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})\}$.
+For each $S\in S_{T}$ and $g:S\rightarrow\Delta(\pazocal{A})$, we construct an expert, $E_{S,g}$.
+Expert $E_{S,g}$ immitates an instance of the MSSOA algorithm, where MSSOA errs at all timesteps $t\in S$ and for each timestep $t\in S$, the optimal strategy for the leader at the timestep is given by $g(t)$.
+We show the implementation of each expert in Algorithm˜7 below.
+Algorithm 7 Implementation of expert $E_{S,g}$
+1:input: time horizon $T$
+2:Set $V_{0}=\pazocal{H}$
+3:for $t\in[T]$ do
+4:  Observe $\mathbf{z}_{t}$
+5:  Let $\mathbf{x}^{(h,t)}=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{h(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))\quad\forall h\in\pazocal{H}$
+6:  Let $X_{t}=\{\mathbf{x}^{(h)}:h\in V_{t-1}\}$triangleright $|X_{t}|\leq K$
+7:  Let $V_{t}^{(\mathbf{x})}=\{h\in V_{t-1}:\mathbf{x}^{(h,t)}=\mathbf{x}\}\quad\forall\mathbf{x}\in X_{t}$
+8:  Play action $l_{t}=\arg\max_{\mathbf{x}\in X_{t}}\textnormal{MSLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})$
+9:  Set $l_{t}=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}\textnormal{MSLdim}_{\pazocal{G}}(V_{t}^{\mathbf{z}_{t}\rightarrow k})$
+10:  if $t\in S$ then
+11:   Predict $g(t)$
+12:   $V_{t+1}\leftarrow V_{t}^{\mathbf{z}_{t}\rightarrow g(t)}$
+13:  else
+14:   Predict $l_{t}$
+15:   $V_{t+1}\leftarrow V_{t}^{\mathbf{z}_{t}\rightarrow l_{t}}$
+16:end
+Notice that for each possible timestep $t\in T$, a single context $\mathbf{z}_{t}$ is presented, so there are at most $K$ possible optimal strategies for the leader, one for each follower type.
+So, the number of experts that can be constructed in this way is $\sum_{i=0}^{\textnormal{MSLdim}_{\pazocal{G}}(\mathscr{H})}\binom{T}{i}(K-1)^{i}\leq(TK)^{\textnormal{MSLdim}_{\pazocal{G}}(\mathscr{H})}$.
+So using the above regret upper bound with this set of experts we get that Algorithm˜6 achieves utility at least $\max_{S,g}-L_{S,g,T}-\sqrt{\frac{1}{2}\textnormal{MSLdim}_{\pazocal{G}}(\mathscr{H})T\ln(TK)}$, where $-L_{S,g,T}$ is the accumulated utility of expert $E_{S,g}$ up to timestep $T$.
+It remains to show that $\max_{S,g}-L_{S,g,T}\geq\max_{h\in\mathscr{H}}-L_{h,T}$.
+Notice that for any $h\in\mathscr{H}$ and for any set of timesteps $S$ that MSSOA errs in, there exists an expert in the constructed set of experts such that $L_{S,g,T}=L_{h,T}$ when executed on the sequence $(\mathbf{z}_{1},h(\mathbf{z}_{1})),\cdots,(\mathbf{z}_{T},h(\mathbf{z}_{T}))$.
 We use a similar approach to Lemma 14 from [15] to get a (nearly-)matching lower bound.
-Theorem 3.16.
+Theorem B.10.
 Let $\pazocal{H}$ be any hypothesis class.
-There exists a Stackelberg game, $\pazocal{G}$, such that $\pazocal{H}$ has a finite Stackelberg-Littlestone dimension $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ and such that no algorithm in the agnostic, full-feedback setting can achieve regret smaller than $\Omega\left(\sqrt{\frac{1}{8}\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})T}\right)$.
-We include the complete proofs of Theorems 3.15 and 3.16 in Appendix A.3.
-4 Distributional Learning
-We now shift our focus from online learning to distributional learning.
-Suppose that there exists an unknown, joint distribution over contexts and follower types.
-We first study the simpler setting, without contexts, in which we give strong generalization guarantees by exploiting the structure of the leader’s payoff function.
-When contexts come into play in Section 4.2, we analyze an intuitive approach that uses a black box empirical risk minimizer (ERM) and show that it is suboptimal in our setting.
-Analogous to our online results, we then give a better sample complexity bound that depends on a new complexity measure.
-4.1 Learning without Contexts
-Suppose there is a fixed, unknown distribution, $\mathscr{D}$, over the set of possible follower types.
-We denote with $u(\mathbf{x},b_{f^{(i)}}(\mathbf{x}))$ the leader’s payoff when committing to a (possibly mixed) strategy $\mathbf{x}\in\Delta(\pazocal{A})$ against a follower of type $f^{(i)}$.
-We use the notation $u_{\mathscr{D}}(\mathbf{x})=\mathbb{E}_{f\sim\mathscr{D}}[u(\mathbf{x},b_{f}(\mathbf{x}))]$ and given a sample $S$ of follower types, we write:
-$u_{S}(\mathbf{x})=\frac{1}{|S|}\sum_{f\in S}u(\mathbf{x},b_{f}(\mathbf{x}))$
-Given a set $S$ of $m$ samples from $\mathscr{D}$, the goal of the leader is to choose a mixed strategy that fares well in expectation against any new sample from $\mathscr{D}$.
-In other words, we want to choose $\hat{\mathbf{x}}\in\Delta(\pazocal{A})$ such that with high probability:
-$\mathbb{P}_{S}[\max_{\mathbf{x}}u_{\mathscr{D}}(\mathbf{x})-u_{\mathscr{D}}(\hat{\mathbf{x}})>\epsilon]\leq\delta.$
-For this, it suffices to show a generalization guarantee, i.e.
-a bound on the difference between the leader’s expected payoff and the empirical payoff over a set of follower types for any possible mixed strategy.
-Definition 4.1.
-A generalization guarantee for the leader in a Stackelberg game, $\pazocal{G}$, is a function $\epsilon_{\pazocal{G}}:\mathbb{Z}_{\geq 1}\times(0,1)\to\mathbb{R}_{\geq 0}$ such that for any $\delta\in(0,1)$, any $m\in\mathbb{Z}_{\geq 1}$, and any distribution $\mathscr{D}$ over the follower types, with probability at least $1-\delta$ over the draw of a set $S\sim\mathscr{D}^{m}$ for any $\mathbf{x}\in\Delta(\pazocal{A})$, the difference between the average payoff of $\mathbf{x}$ over $S$ and the expected payoff of $\mathbf{x}$ over $\mathscr{D}$ is at most $\epsilon_{\pazocal{G}}(m,\delta)$:
-$\mathbb{P}_{S\sim\mathscr{D}^{m}}[\exists\mathbf{x}\in\Delta(\pazocal{A})\;\;\text{such that}\;\;|u_{\mathscr{D}}(\mathbf{x})-u_{S}(\mathbf{x})|>\epsilon_{\pazocal{G}}(m,\delta)]\leq\delta$
-The implication of a generalization guarantee is that it relates the utility of the leader’s strategy that maximizes the average payoff over the training data to the expected utility of the optimal strategy.
-Formally, this means that for $\hat{x}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u_{S}(\mathbf{x})$ and $x^{*}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u_{\mathscr{D}}(\mathbf{x})$:
-$\mathbb{P}_{S\sim\mathscr{D}^{m}}[|u_{\mathscr{D}}(\mathbf{x}^{*})-u_{S}(\hat{\mathbf{x}})|>2\epsilon_{\pazocal{G}}(m,\delta)]\leq\delta.$
-Consider the class $\pazocal{P}=\{u_{\mathbf{x}}:\mathscr{K}\to\mathbb{R}\;\;\text{such that}\;\;u_{\mathbf{x}}(f)=u(\mathbf{x},b_{f}(\mathbf{x}))\}$ to be the set of payoff functions for the leader.
-Each payoff function is parameterized by the leader’s mixed strategy, takes as input the follower’s type, and outputs the leader’s utility.
-We also define the dual class $\tilde{\pazocal{P}}=\{\tilde{u}_{f}:\Delta(\pazocal{A})\to\mathbb{R}\;\;\text{such that}\;\;\tilde{u}_{f}(\mathbf{x})=u(\mathbf{x},b_{f}(\mathbf{x}))\}$ to be the set of payoff functions for the leader, parameterized by a follower type that take as input the leader’s mixed strategy and output the leader’s utility.
-Consider the following definition, introduced by [6] and adapted to our setting:
-Definition 4.2.
-A class $\pazocal{H}$ of hypotheses, $h:\mathscr{K}\to\mathbb{R}$, is $(d,t)$-delineable if:
-  1. 1.
-    The class $\pazocal{H}$ consists of hypotheses that are parameterized by vectors $\mathbf{x}$ from a set $\pazocal{X}\subseteq\mathbb{R}^{d}$; and
-  2. 2.
-    For any $f\in\mathscr{K}$, there is a set $\mathscr{J}$ of $t$ hyperplanes such that for any connected component $\pazocal{X}^{\prime}=\pazocal{X}\setminus\mathscr{J}$, the dual function $h_{f}(\mathbf{x})$ is linear over $\pazocal{X}^{\prime}$.
-With this definition, we make the following useful proposition.
-Proposition 4.3.
-The class of leader payoff functions, $\pazocal{P}$ that are parameterized by the leader’s mixed strategy and take as input the follower’s type is $(|\pazocal{A}|,\binom{|\pazocal{A}_{f}|}{2})$-delineable.
+There exists a Stackelberg game, $\pazocal{G}$, such that $\pazocal{H}$ has a finite MSL dimension $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ and such that no algorithm in the agnostic, full-feedback setting can achieve regret smaller than $\Omega\left(\sqrt{\frac{1}{8}\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})T}\right)$.
 Proof.
-The statement holds from the following two facts: (1.) Given a fixed follower type, $f$, we have that $\tilde{u}_{f}(\mathbf{x})$ is a piece-wise linear function.
-(2.) For any fixed follower type, $f$, we have that $u_{f}(\mathbf{x})$ will have at most $|A_{f}|(|A_{f}|-1)$ hyperplanes going through it.
-Each hyperplane corresponds to a change in best-response strategies for the follower from one pure action to another, so there can be at most $\binom{|\pazocal{A}_{f}|}{2}$ hyperplanes.
-∎
-Next, we present the definition of pseudo-dimension [33], which is complexity measure that characterizes a hypothesis class and in our setting helps us reason about how difficult the learning task of the leader is.
-Informally speaking, the more complex the hypothesis class, $\pazocal{P}$, the harder it is for the leader to distinguish between strategies, $\mathbf{x}$, which parameterize the class, and therefore, more samples are needed to learn.
-Definition 4.4 (Pseudo-dimension).
-Let $\pazocal{F}$ be a class of functions $f:A\to\mathbb{R}$, with an abstract domain, $A$.
-We say that $y^{(1)},\ldots,y^{(n)}$, witness the shattering of $S=\{x^{(1)},\ldots,x^{(n)}\}\subseteq A$ if for all $T\subseteq S$, there exists a function $f_{T}\in\pazocal{F}$, such that for all $x^{(i)}\in T,\;f_{T}(x^{(i)})\leq y^{(i)}$ and for all $x^{(i)}\not\in T,\;f_{T}(x^{(i)})>y^{(i)}$.
-We then say that the pseudo-dimension of $\pazocal{F}$, denoted $\textnormal{Pdim}(\pazocal{F})$, is the size of the largest set that can be shattered by $\pazocal{F}$.
-Leveraging results from the data-driven algorithm design literature ([4, 8, 7]), such as Lemma 3.10 in [6] and Proposition 4.3 we have the following result:
-Theorem 4.5.
-The pseudo-dimension of the hypothesis class, $\pazocal{P}$ satisfies $\textnormal{Pdim}(\pazocal{P})\leq 9|\pazocal{A}|\log\left(4|\pazocal{A}||\pazocal{A}_{f}|^{2}\right)$.
-So we also get the generalization guarantee from Pollard [33] and Dudley [18] with function:
-$\epsilon_{\pazocal{G}}(m,\delta)=O\left(U\sqrt{\frac{|\pazocal{A}|\log(|\pazocal{A}||\pazocal{A}_{f}|^{2})}{m}}+U\sqrt{\frac{\ln(1/\delta)}{m}}\right),$
-where $U$ is the maximum payoff value for the leader.
-Discussion.
-We note that the above generalization result is particularly strong.
-Informally, it asserts that no matter what the distribution over the follower types is, if the leader has so far encountered a number of followers that scales linearly with the size of her action set and logarithmically with the follower’s action set, it suffices for her to play a strategy that maximizes her average utility over the followers seen so far.
-This result holds even for an infinite number of follower types and extends prior work on learning in Stackelberg games without contextual information [26].
-4.2 Learning with Contexts
-We now shift to analyzing the setting where the leader has access to a data set consisting of tuples of contexts and follower types, $S_{m}=\{(\mathbf{z}_{1},f_{1}),\ldots,(\mathbf{z}_{m},f_{m})\}$.
-Given new contextual information, the leader aims to choose a strategy that maximizes her payoff against a follower, whose type is unknown.
-Here we introduce the study of the statistical complexity of distributional learning in Stackelberg games with side information.
-Consider a distribution $\mathscr{D}$ over $\pazocal{Z}\times\mathscr{K}$.
-We define the leader’s payoff as a function of a policy $\pi:\pazocal{Z}\to\Delta(\pazocal{A})$ with respect to $\mathscr{D}$ as:
+We let $d=\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$.
+For simplicity assume that $T$ is an integer multiple of $d$, say $T=kd$ for $k\in\mathbb{N}$.
+Consider a tree, $(\mathbf{v}_{1},\mathbf{v}_{1},\ldots,\mathbf{v}_{2^{d}-1})$, of depth, $d$, that is SL-shattered by $\pazocal{H}$ such that each node $\mathbf{v}_{i}\in\pazocal{Z}$ has two edges (each leading to a child) corresponding to two strategies for the leader $\mathbf{x}_{i,1},\mathbf{x}_{i,2}\in\Delta(\pazocal{A})$, such that $\mathbf{x}_{i,1}\neq\mathbf{x}_{i,2}$.
+By definition, we have that for each $i$, there exist at least two functions $h_{i,1},h_{i,2}\in\pazocal{H}$ such that at context $\mathbf{v}_{i}$, they predict follower types $f_{1}=h_{i,1}(\mathbf{v}_{i})$ and $f_{2}=h_{i,2}(\mathbf{v}_{i})$, against which the strategies $\mathbf{x}_{i,1}$ and respectively $\mathbf{x}_{i,2}$ are optimal.
+In other words, we construct a Stackelberg game, playing the optimal strategy against the mispredicted follower type gives utility 0, whereas playing the optimal strategy against the correct follower type gives utility 111 (for an explicit construction see proof of Theorem˜B.3).
+Then, we have that for every $i$
+$u(\mathbf{v}_{i},\mathbf{x}_{i,1},b_{f_{1}}(\mathbf{v}_{i},\mathbf{x}_{i,1}))-u(\mathbf{v}_{i},\mathbf{x}_{i,2},b_{f_{1}}(\mathbf{v}_{i},\mathbf{x}_{i,2}))=1$
+and similarly,
+$u(\mathbf{v}_{i},\mathbf{x}_{i,2},b_{f_{2}}(\mathbf{v}_{i},\mathbf{x}_{i,2}))-u(\mathbf{v}_{i},\mathbf{x}_{i,1},b_{f_{2}}(\mathbf{v}_{i},\mathbf{x}_{i,1}))=1.$
+We will now construct a sequence $(\mathbf{z}_{1},\mathbf{x}_{1}),(\mathbf{z}_{2},\mathbf{x}_{2}),\ldots,(\mathbf{z}_{T},\mathbf{x}_{T})$, which we will show forces any algorithm to meet the lower bound in the theorem statement.
+From the MSL-shattered tree, we follow a root-to-leaf path, say $(\tilde{\mathbf{z}}_{1},\tilde{\mathbf{x}}_{1}),\ldots,(\tilde{\mathbf{z}}_{d},\tilde{\mathbf{x}}_{d})$.
+For each node $\tilde{\mathbf{z}}_{i}$ on the path, where $i\in[d]$, we construct a set of $k$ pairs of contexts and follower types $\{(\mathbf{z}_{(i-1)k+1}^{a},\mathbf{x}_{(i-1)k+1}^{a}),\ldots,(\mathbf{z}_{ik}^{a},\mathbf{x}_{ik}^{a})\}$ as follows.
+  • We set $\mathbf{z}_{(i-1)k+1}^{a}=\ldots=\mathbf{z}_{ik}^{a}=\tilde{\mathbf{z}}_{i}$.
+  • We draw each $\mathbf{x}_{t}^{a}$ for $t\in\{(i-1)k+1,\ldots,ik\}$, independently and uniformly at random from the two optimal strategies corresponding to the edges leading to the two children of $\mathbf{z}_{i}$.
+We denote the follower type and optimal strategy for each child, $f_{A,i},\mathbf{x}_{A,i}$ and $f_{B,i},\mathbf{x}_{B,i}$ respectively.
+For each set of $k$ examples, let $T_{i}=\{(i-1)k+1,\ldots,ik\}$ be the set of indeces of the examples corresponding to $\tilde{\mathbf{z}}_{i}$. It suffices to lower bound:
+$\max_{h\in\pazocal{H}}\sum_{t\in T_{i}}u(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$
+For ease of notation, we denote $h^{*}=\arg\max_{h\in\pazocal{H}}\sum_{t\in T_{i}}u(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t})))$, $\;u_{h^{*},t}=u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))$, $u_{t}=u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$, and $u_{t}^{*}=u(\mathbf{z}_{t},\pi_{f_{t}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{f_{t}}(\mathbf{z}_{t})))$, where $\pi_{f_{t}}(\mathbf{z}_{t})$ corresponds to the optimal policy against the true follower type and context at time $t$.
+So equivalently, we want to lower bound
+$\sum_{t\in T_{i}}u_{h^{*},t}-u_{t}=\sum_{t\in T_{i}}(u_{t}^{*}-u_{t})-(u_{t}^{*}-u_{h^{*},t}).$
+Let $r_{i}$ be the number of timesteps in $T_{i}$, where $f_{A,i}$ appears in the example. Then
+$u_{t}^{*}-u_{h^{*},t}=\begin{cases}k-r\quad\text{if}\;r\geq k/2\\ r\hskip 28.45274pt\text{if}\;r<k/2\end{cases}$
+Therefore, we get that $k/2-\sum_{t\in T_{i}}u_{t}^{*}-u_{h^{*},t}=|r-k/2|$. By taking an expectation over the randomness in the $f_{t}$’s, we get that
+$k/2-\mathbb{E}\left[\sum_{t\in T_{i}}u_{t}^{*}-u_{h^{*},t}\right]=\mathbb{E}\left[|r-k/2|\right]\geq\sqrt{k/8},$
+where the last inequality holds by Khinchine’s inequality (see Lemma A.9 in [19]). Now it suffices to sum over all $i\in[d]$ to get
+$\frac{dk}{2}-\mathbb{E}\left[\sum_{t\in[T]}u_{t}^{*}-u_{h^{*},t}\right]=\sum_{i\in[d]}\left(\frac{k}{2}-\mathbb{E}\left[\sum_{t\in T_{i}}u_{t}^{*}-u_{h^{*},t}\right]\right)\geq d\sqrt{k/8}$
+Finally, notice that $dk/2=T/2=\sum_{t\in[T]}(u_{t}^{*}-u_{t})$ for any algorithm, which concludes the proof. ∎
+B.2 Distributional Learning
+We now give similar results in the distributional setting, in which the leader has access to a batch of training data. We start with some key definitions.
+Consider a distribution $\mathscr{D}$ over $\pazocal{Z}\times[K]$. We define the leader’s payoff as a function of a policy $\pi:\pazocal{Z}\to\Delta(\pazocal{A})$ with respect to $\mathscr{D}$ as
 $u_{\mathscr{D}}(\pi)=\mathop{\mathbb{E}}_{(\mathbf{z},f^{(i)})\sim\mathscr{D}}\left[u(\mathbf{z},\pi(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi(\mathbf{z})))\right]$
-and the leader’s average payoff on a sample $S_{m}=\{(\mathbf{z}_{1},f_{1}),\ldots,(\mathbf{z}_{m},f_{m})\}\in(\pazocal{Z}\times\mathscr{K})^{m}$ as:
+and the leader’s average payoff on a sample $S_{m}=\{(\mathbf{z}_{1},f_{1}),\ldots,(\mathbf{z}_{m},f_{m})\}\in(\pazocal{Z}\times[K])^{m}$ as
 $u_{S_{m}}(\pi)=\frac{1}{m}\sum_{i=1}^{m}\left[u(\mathbf{z}_{i},\pi(\mathbf{z}_{i}),b_{f_{i}}(\mathbf{z}_{i},\pi(\mathbf{z}_{i})))\right].$
-The goal of the learner is to deploy an algorithm, $\textsc{Alg}:\cup_{m=0}^{\infty}(\pazocal{Z}\times\mathscr{K})^{m}\to\Delta(A)^{\pazocal{Z}}$, such that given a training sequence $S_{m}=\{(\mathbf{z}_{1},f_{1}),\cdots,(\mathbf{z}_{m},f_{m})\}$, the algorithm outputs a policy $\pi_{\textsc{Alg}}:\pazocal{Z}\to\Delta(A)$ that maximizes $u_{\mathscr{D}}(\pi)$.
-Suppose that the leader has access to a set of hypotheses, $\pazocal{H}$, such that each hypothesis in $\pazocal{H}$, maps contexts to follower types ($h:\pazocal{Z}\to\mathscr{K}$).
-We define the true error function of a hypothesis $h$ with respect to a distribution, $\mathscr{D}$, as:
+The goal of the learner is to deploy an algorithm, $\textsc{Alg}:\cup_{m=0}^{\infty}(\pazocal{Z}\times[K])^{m}\to\Delta(A)^{\pazocal{Z}}$, such that given a training sequence $S_{m}=\{(\mathbf{z}_{1},f_{1}),\cdots,(\mathbf{z}_{m},f_{m})\}$, the algorithm outputs a policy $\pi_{\textsc{Alg}}:\pazocal{Z}\to\Delta(A)$ that maximizes $u_{\mathscr{D}}(\pi)$.
+Suppose that the leader has access to a set of hypotheses, $\pazocal{H}$, such that each hypothesis in $\pazocal{H}$, maps contexts to follower types ($h:\pazocal{Z}\to[K]$).
+We define the true error function of a hypothesis $h$ with respect to a distribution, $\mathscr{D}$, as
 $\text{Err}_{\mathscr{D}}(h)=\mathbb{P}_{(\mathbf{z},f)\sim\mathscr{D}}\left[h(\mathbf{z})\neq f\right].$
-Similarly, we define the empirical error of a hypothesis $h$ on a sample $S_{m}$ as:
+Similarly, we define the empirical error of a hypothesis $h$ on a sample $S_{m}$ as
 $\text{Err}(h)=\frac{1}{m}|\{i\in[m]:h(\mathbf{z}_{i})\neq f_{i}\}|$
-and the approximation error of a hypothesis class $\pazocal{H}$ on a distribution $\mathscr{D}$ to be:
+and the approximation error of a hypothesis class $\pazocal{H}$ on a distribution $\mathscr{D}$ to be
 $\text{Err}_{\mathscr{D}}(\pazocal{H}):=\inf_{h\in\pazocal{H}}\text{Err}_{\mathscr{D}}(h).$
+Remark B.11.
+For the remainder of this section, we will assume that for any context, $\mathbf{z}$, and follower type $f^{(i)}$ there exists a fixed and known tie-breaking rule over the leader’s optimal strategy against $f^{(i)}$ given $\mathbf{z}$.
 Consider an optimal hypothesis, $h^{*}\in\pazocal{H}$, such that given any context, $h^{*}$ always predicts the follower type correctly, i.e.
@@
 [^6] Note that for binary labels, the definition of Natarajan dimension recovers the well-studied VC-dimension of a hypothesis class.
-Definition 4.6 (Natarajan Dimension).
-Let $\pazocal{H}\subseteq\mathscr{K}^{\pazocal{Z}}$ be a hypothesis class and let $S\subseteq\pazocal{Z}$.
-We say that $\pazocal{H}$ N-shatters $S$ if there exist $\phi_{1},\phi_{2}:S\to\mathscr{K}$ such that $\forall\mathbf{z}\in\pazocal{Z}\;\phi_{1}(\mathbf{z})\neq\phi_{2}(\mathbf{z})$ and for any $T\subseteq S$ there exists a $g\in\pazocal{H}$ such that:
+Definition B.12 (Natarajan Dimension).
+Let $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$ be a hypothesis class and let $S\subseteq\pazocal{Z}$.
+We say that $\pazocal{H}$ N-shatters $S$ if there exist $\phi_{1},\phi_{2}:S\to[K]$ such that $\forall\mathbf{z}\in\pazocal{Z}\;\;\phi_{1}(\mathbf{z})\neq\phi_{2}(\mathbf{z})$ and for any $T\subseteq S$ there exists a $g\in\pazocal{H}$ such that
 $\forall\mathbf{z}\in T,\;g(\mathbf{z})=\phi_{1}(\mathbf{z})\quad\text{and}\quad\forall\mathbf{z}\in S\setminus T,\;g(\mathbf{z})=\phi_{2}(\mathbf{z}).$
 The Natarajan dimension of $\pazocal{H}$, denoted $d_{N}(\pazocal{H})$, is the maximal cardinality of a set that is shattered by $\pazocal{H}$.
-We define the realizable sample complexity of a Stackelberg-learning algorithm Alg for a hypothesis class $\pazocal{H}$ to be the function $m_{\textsc{Alg},\pazocal{H}}^{r}$ such that for every $\epsilon,\delta>0$, $m_{\textsc{Alg},\pazocal{H}}^{r}$ is the minimal integer such that for every $m\geq m_{\textsc{Alg},\pazocal{H}}^{r}$ and every distribution $\mathscr{D}$ over $\pazocal{Z}$, we have:
+We define the realizable sample complexity of a Stackelberg-learning algorithm Alg for a hypothesis class $\pazocal{H}$ to be the function $m_{\textsc{Alg},\pazocal{H}}^{r}$ such that for every $\epsilon,\delta>0$, $m_{\textsc{Alg},\pazocal{H}}^{r}$ is the minimal integer such that for every $m\geq m_{\textsc{Alg},\pazocal{H}}^{r}$ and every distribution $\mathscr{D}$ over $\pazocal{Z}$, we have
 $\mathbb{P}_{S_{m}}\left(u_{\mathscr{D}}(\textsc{Alg}(S_{m}))<u_{\mathscr{D}}(\pi_{h^{*}})-\epsilon\right)\leq\delta.$ (2)
 where $\pi_{h^{*}}(\mathbf{z}):=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h^{*}(\mathbf{z})}(\mathbf{z},\mathbf{x}))$.
-4.2.1 Vanilla ERM learner
+B.2.1 Vanilla ERM learner
 We define an Empirical Risk Minimization (ERM) learner for the hypothesis class $\pazocal{H}$ as an algorithm, such that for any sample $S_{m}$ returns a function that minimizes the empirical error of that function relative to any other function in $\pazocal{H}$.
-Formally, we say:
+Formally, we say
 $h_{ERM}$ $\in\arg\min_{h\in\pazocal{H}}\frac{1}{m}\sum_{i=1}^{m}\mathbbm{1}\{h(\mathbf{z}_{i})\neq f_{i}\}.$ (3)
@@
 This can be achieved by using a simple ERM learner as in (3).
-Proposition 4.7.
-Consider an algorithm $\textsc{Alg}_{\text{ERM}}:\cup_{n=0}^{\infty}(\pazocal{Z}\times\mathscr{K})^{n}\to\Delta(\pazocal{A})^{\pazocal{Z}}$ that outputs the policy:
+Proposition B.13.
+Consider an algorithm $\textsc{Alg}_{\text{ERM}}:\cup_{n=0}^{\infty}(\pazocal{Z}\times[K])^{n}\to\Delta(\pazocal{A})^{\pazocal{Z}}$ that outputs the policy
 $\pi_{\text{ERM}}(\mathbf{z})$ $\in\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h_{ERM}(\mathbf{z})}(\mathbf{z},\mathbf{x})),$
-where $h_{ERM}$ satisfies equation (3). Then for any such $h_{\text{ERM}}$, there exists constant $C>0$, such that for every $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$,
+where $h_{ERM}$ satisfies Equation˜3. Then for any such $h_{\text{ERM}}$, there exists constant $C>0$, such that for every $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$,
 $m^{r}_{\pazocal{A}_{\text{ERM}}}\leq C\left(\frac{d_{N}(\pazocal{H})\ln(K)\ln(\frac{1}{\epsilon})+\ln(\frac{1}{\delta})}{\epsilon}\right)$
-We present the proof of Proposition 4.7 in Appendix B.
-4.2.2 Stackelberg ERM learner
-In this section, we show how to improve upon the sample complexity bound shown in Proposition 4.7, by leveraging the interaction between the Stackelberg game and the hypothesis class.
-Consider the following hypothesis class of possible leader policies:
-$\Pi_{\pazocal{H},\pazocal{G}}=\Bigl{\{}\pi_{h}:\pazocal{Z}\to\Delta(\pazocal{A})|h\in\pazocal{H},\pi_{h}(\mathbf{z})=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x}))\Bigl{\}}$
+Proof.
+We have that
+$\mathbb{P}_{S_{m}}\left[u_{\mathscr{D}}(\pi_{h^{*}})-u_{\mathscr{D}}(\pi_{\text{ERM}})\geq\epsilon\right]$
+$\leq$ $\mathbb{P}_{S_{m}}\left[\mathop{\mathbb{E}}_{\mathbf{z}\sim\mathscr{D}}\left[u(\mathbf{z},\pi_{h^{*}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{h^{*}}(\mathbf{z})))-u(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z})))\right]>\epsilon\right]$
+By the boundedness of the utility function we have that the expectation above lies within [1,1][-1,1][ - 1 , 1 ]. Therefore, we get
+$\leq$ $\mathbb{P}_{S_{m}}\left[\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}\left[u(\mathbf{z},\pi_{h^{*}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{h^{*}}(\mathbf{z})))-u(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z})))>0\right]>\epsilon\right]$
+But getting less utility than the optimal predictor in the class, implies that our prediction of the follower’s type is incorrect. So,
+$\leq$ $\mathbb{P}_{S_{m}}\left[\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}\left[h_{*}(\mathbf{z})\neq h_{\text{ERM}}(\mathbf{z})\right]>\epsilon\right]$
+But from Theorem 6 in [22], we know that for any $\epsilon,\delta>0$, there exists constant $C_{2}$ such that for any $m\geq C_{2}\left(\frac{d_{N}(\pazocal{H})\ln(K)\ln(\frac{1}{\epsilon})+\ln(\frac{1}{\delta})}{\epsilon}\right)$, we have that for any $h_{\text{ERM}}$ that satisfies Equation˜3
+$\mathbb{P}_{S_{m}}\left(\text{Err}_{\mathscr{D}}(h_{\text{ERM}})>\text{Err}_{\mathscr{D}}(\pazocal{H})+\epsilon\right)\leq\delta$
+Since we are in the realizable setting, we know that $\text{Err}_{\mathscr{D}}(\pazocal{H})=0$.
+By definition, we have that $\text{Err}_{\mathscr{D}}(h_{\text{ERM}})>\epsilon\iff\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}\left[h_{\text{ERM}}(\mathbf{z})\neq h^{*}(\mathbf{z})\right]>\epsilon$, which concludes our proof.
+∎
+B.2.2 Stackelberg ERM learner
+In this section, we show how to improve upon the sample complexity bound shown in Proposition˜B.13, by leveraging the interaction between the Stackelberg game and the hypothesis class.
+Consider the following hypothesis class of possible leader policies
+$\Pi_{\pazocal{H},\pazocal{G}}=\Bigl{\{}\pi_{h}:\pazocal{Z}\to\Delta(\pazocal{A})|h\in\pazocal{H},\pi_{h}(\mathbf{z})\in\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x}))\Bigl{\}}$
 and the following variation of an ERM rule.
-Definition 4.8 (Stackelberg ERM).
-Consider a hypothesis class $\pazocal{H}\subseteq\mathscr{K}^{\pazocal{Z}}$ and a Stackelberg game, $\pazocal{G}$.
-We define the Stackelberg ERM learning algorithm, $\textsc{Alg}_{\text{S-ERM}}:\cup_{n=0}^{\infty}(\pazocal{Z}\times\mathscr{K})^{n}\to\Delta(\pazocal{A})^{\pazocal{Z}}$, such that given a training sequence $S_{m}=\{(\mathbf{z}_{1},f_{1}),\cdots,(\mathbf{z}_{m},f_{m})\}$, outputs the policy $\pi_{\text{S-ERM}}:\pazocal{Z}^{m}\to\Delta(\pazocal{A})$ that satisfies:
-$\pi_{\text{S-ERM}}\in\arg\min_{\pi\in\Pi_{\pazocal{H},\pazocal{G}}}\frac{1}{m}\sum_{i=1}^{m}\mathbbm{1}\{\pi(\mathbf{z}_{i})\neq\mathbf{x}_{i}^{*}\},$ (4)
-where $\mathbf{x}_{i}^{*}=\arg\max_{x\in\Delta(\pazocal{A})}\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{i},\mathbf{x},b_{f_{i}}(\mathbf{z}_{i},\mathbf{x}))$.
+Definition B.14 (Stackelberg ERM).
+Consider a hypothesis class $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$ and a Stackelberg game, $\pazocal{G}$.
+We define the Stackelberg ERM learning algorithm, $\textsc{Alg}_{\text{S-ERM}}:\cup_{n=0}^{\infty}(\pazocal{Z}\times[K])^{n}\to\Delta(\pazocal{A})^{\pazocal{Z}}$, such that given a training sequence $S_{m}=\{(\mathbf{z}_{1},f_{1}),\cdots,(\mathbf{z}_{m},f_{m})\}$, outputs the policy $\pi_{\text{S-ERM}}:\pazocal{Z}^{m}\to\Delta(\pazocal{A})$ that satisfies
+$\pi_{\text{S-ERM}}\in\arg\max_{\pi\in\Pi_{\pazocal{H},\pazocal{G}}}\frac{1}{m}\sum_{i=1}^{m}u(\mathbf{z}_{i},\pi(\mathbf{z}_{i}),b_{f_{i}}(\mathbf{z}_{i},\pi(\mathbf{z}_{i})).$ (4)
 We define the following measure of complexity of a hypothesis class with respect to a Stackelberg game.
-Definition 4.9 (Stackelberg-Natarajan Dimension).
-Consider a hypothesis class $\pazocal{H}\subseteq\mathscr{K}^{\pazocal{Z}}$, a Stackelberg game, $\pazocal{G}$, and let $S\subseteq\pazocal{Z}$.
-We say that $\pazocal{H}$ SN-shatters $S$ with respect to game $\pazocal{G}$, if there exist $\phi_{1},\phi_{2}:S\to\Delta(\pazocal{A})$ such that $\forall\mathbf{z}\in\pazocal{Z}\;\phi_{1}(\mathbf{z})\neq\phi_{2}(\mathbf{z})$ and for any $T\subseteq S$ there exists a $g\in\pazocal{H}$ such that:
+Definition B.15 (Stackelberg-Natarajan Dimension).
+Consider a hypothesis class $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$, a Stackelberg game, $\pazocal{G}$, and let $S\subseteq\pazocal{Z}$.
+We say that $\pazocal{H}$ SN-shatters $S$ with respect to game $\pazocal{G}$, if there exist $\phi_{1},\phi_{2}:S\to\Delta(\pazocal{A})$ such that $\forall\mathbf{z}\in\pazocal{Z}\;\phi_{1}(\mathbf{z})\neq\phi_{2}(\mathbf{z})$ and for any $T\subseteq S$ there exists a $g\in\pazocal{H}$ such that
 $\forall\mathbf{z}\in T,\;\pi_{g}(\mathbf{z})=\phi_{1}(\mathbf{z})\quad\text{and}\quad\forall\mathbf{z}\in S\setminus T,\;\pi_{g}(\mathbf{z})=\phi_{2}(\mathbf{z}).$
 The Stackelberg-Natarajan dimension of $\pazocal{H}$ with respect to game $\pazocal{G}$, denoted as $d_{SN,\pazocal{G}}(\pazocal{H})$, is the maximal cardinality of a set that is SN-shattered by $\pazocal{H}$.
-Theorem 4.10.
+Theorem B.16.
 For any hypothesis class $\pazocal{H}$, $d_{SN,\pazocal{G}}(\pazocal{H})\leq d_{N}(\pazocal{H})$.
@@
 We illustrate this via the following example.
-Example 4.11.
+Example B.17.
 Consider the hypothesis class of conjunctions, where each hypothesis is defined over a set of Boolean variables, $\pazocal{Z}=\{z_{1},\cdots,z_{n}\}$, such that $\pazocal{H}_{\text{conj}}=\{h:\{0,1\}^{n}\to\{0,1\}|h(z)=\land_{i\in S}z_{i}\land_{j\in T}z_{j},S,T\subseteq[n],S\cap T=\emptyset\}$.
@@
 The Natarajan dimension in this case reduces to the well-studied notion of the VC-dimension, which is known to be equal to $n$ for $\pazocal{H}_{\text{conj}}$.
-At the same time, the $d_{SN,\pazocal{G}}(\pazocal{H}_{\text{conj}})$ will be upper bounded by the size of the set $\{z\in\{0,1\}^{n}|\pi_{*}^{(0)}(z)\neq\pi_{*}^{(1)}(z)\}$, which could be as small as $1$ for two distinct follower types.
-Theorem 4.12.
-Consider the Stackelberg ERM policy, $\pi_{\text{S-ERM}}$, as in Definition 4.9 and let $h_{\text{S-ERM}}$ be the corresponding hypothesis in $\pazocal{H}$.
-Then universally for any such policy, there exists constant $C$, such that for every $\pazocal{H}\subseteq\mathscr{K}^{\pazocal{Z}}$ and distribution $\mathscr{D}$,
+At the same time, the $d_{SN,\pazocal{G}}(\pazocal{H}_{\text{conj}})$ will be upper bounded by the size of the set $\{z\in\{0,1\}^{n}|\pi_{*}^{(0)}(z)\neq\pi_{*}^{(1)}(z)\}$, which could be as small as 111 for two distinct follower types.
+Theorem B.18.
+Consider the Stackelberg ERM policy, $\pi_{\text{S-ERM}}$, as in Definition˜B.15 and let $h_{\text{S-ERM}}$ be the corresponding hypothesis in $\pazocal{H}$.
+Then universally for any such policy, there exists constant $C$, such that for every $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$ and distribution $\mathscr{D}$,
 $m^{r}_{\pazocal{A}_{\text{ERM}}}\leq C\frac{1}{\epsilon}\left(d_{SN}(\pazocal{H})\log\left(\frac{d_{SN}(\pazocal{H})\cdot K}{\epsilon}\right)+\log\left(\frac{1}{\delta}\right)\right)$
@@
 The only part of the proof that does not apply here is Natarajan’s lemma.
-We prove an appropriate generalization in Lemma 4.13.
-The full proof of Theorem 4.12 is presented in Appendix B.
-Lemma 4.13.
-$|\Pi_{\pazocal{H},\pazocal{G}}|\leq|\pazocal{Z}|^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$.
-We present the complete proof of Lemma 4.13 in Appendix B.
-Remark 4.14.
-It is worth noting that even when there is only one context (i.e.
-$\pazocal{Z}$ is a singleton), our results in this section are incomparable with those in Section 4.1.
-This is because in this section, the goal of the leader is to learn a deterministic mapping from contexts to follower types, assuming that one exists.
-So, if the contexts are the same between all examples in the training set, the follower types corresponding to these examples will also be the same.
-In contrast, the context-free setting we consider in Section 4.1 allows for a distribution over follower types.
-Thus the two bounds are generally incomparable.
-5 Conclusion
-In this work we study structured Stackelberg games, where contextual information is predictive of the follower’s type.
-The follower type that the leader is playing against is unknown, while the context that carries information about the state of the world is presented to both players.
-We show that the leader can use the unknown, fixed mapping from contexts to follower types to inform her strategy.
-In the online setting, our algorithms learn the optimal policy for the leader and achieve tight no-regret bounds.
-These bounds scale with a new notion of statistical complexity that directly exploits the structure of Stackelberg games.
-We show similar results in the distributional setting, where contexts and follower types are drawn from a (joint) distribution.
-An interesting direction for future research is to analyze structured Stackelberg games in the beyond worst-case analysis framework.
-It has recently been shown that smooth online learning is as easy as distributional learning, with regret bounds scaling with the VC dimension of the hypothesis class.
-Extending these findings to multiclass learning and structured Stackelberg games might lead to improved results both from a statistical and computational point of view.
-Appendix A Details and Proofs from Section 3
-A.1 A technical subtlety
-An important component of our analysis is that the leader must be able to compute the optimal strategy for a given context and follower type.
-For this, we state that for any context $\mathbf{z}\in\pazocal{Z}$ the leader incurs negligible regret by restricting herself to a subset $\mathscr{E}_{\mathbf{z}}\subseteq\Delta(\pazocal{A})$ which is finite and computable.
-To formally state our result we present the notion of a contextual best response region, which was first introduced in [24] and which we repeat here for completeness.
-Definition A.1 (Contextual Follower Best-Response Region).
-For context $\mathbf{z}$, follower type $f^{(i)}$, and follower action $a_{f}\in\pazocal{A}_{f}$, let $X_{\mathbf{z}}(f^{(i)},a_{f})\subseteq\pazocal{X}$ denote the leader’s mixed strategies such that a follower of type $f^{(i)}$ best-responds to all $x\in X_{\mathbf{z}}(f^{(i)},a_{f})$ by playing action $a_{f}$, that is, $X_{\mathbf{z}}(f^{(i)},a_{f})=\{\mathbf{x}\in\pazocal{X}:b_{f^{(i)}}(\mathbf{z},\mathbf{x})=a_{f}\}$
-Notice that any possible follower best-response region is (1) a subset of $\Delta(\pazocal{A})$ and (2) the intersection of finitely-many halfspaces.
-Therefore, every $X_{\mathbf{z}}(f^{(i)},a_{f})$ is convex and bounded.
-However, the leader may not be able to optimize over the extreme points of a best-response region, as it may not be closed due to the followers’ tie-breaking.
-Definition A.2 ($\delta$-approximate extreme points).
-Fix a context $\mathbf{z}\in\pazocal{Z}$ and consider the set of all non-empty contextual best-response regions.
-For $\delta>0$, $\mathscr{E}_{\mathbf{z}}(\delta)$ is the set of leader mixed strategies such that for all best-response functions $\sigma$ and any $\mathbf{x}\in\Delta(\pazocal{A})$ that is an extreme point of cl$(\pazocal{X}_{\mathbf{z}}(\sigma))$, $\mathbf{x}\in\mathscr{E}_{\mathbf{z}}(\delta)$ if $\mathbf{x}\in\pazocal{X}_{\mathbf{z}}(\sigma)$.
-Otherwise there is some $\mathbf{x}^{\prime}\in\mathscr{E}_{\mathbf{z}}(\delta)$ such that $\mathbf{x}^{\prime}\in\pazocal{X}_{\mathbf{z}}(\sigma)$ and $\|\mathbf{x}^{\prime}-\mathbf{x}\|_{1}\leq\delta$.
-Notice that for sufficiently small $\delta$ i.e.
-$\delta=\mathscr{O}(\frac{1}{T})$, we get that the additional regret of the leader by choosing actions over $\mathscr{E}^{(i)}$ rather than $X(f^{(i)},a_{f})$ is negligible.
-Therefore, we use the shorthand $\mathscr{E}^{(i)}:=\mathscr{E}^{(i)}(\delta)$ for the rest of the section.
-The following lemma from [24] formalizes this notion.
-Lemma A.3.
-For any sequence of followers $f_{1},\ldots f_{T}$ and any leader policy $\pi$, there exists a policy $\pi^{(\mathscr{E})}:\pazocal{Z}\rightarrow\cup_{\mathbf{z}\in\pazocal{Z}}\mathscr{E}_{\mathbf{z}}$ that, when given context $\mathbf{z}$, plays a mixed strategy in $\mathscr{E}_{\mathbf{z}}$ and guarantees that $\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\pi^{(\mathscr{E})}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi^{(\mathscr{E})}(\mathbf{z}_{t})))\leq 1$.
-Moreover, the same result holds in expectation over any distribution over follower types.
-Given Lemma A.3 and for ease of notation, in this paper we assume that the leader will always set the parameter $\delta$, such that it does not affect our regret bounds.
-A.2 Using multiclass SOA as a black-box algorithm
-Consider the following algorithm.
-Algorithm 2 Learning with full-feedback and realizable follower types
-1:for $t\in[T]$ do
-2:     Observe $\mathbf{z}_{t}$
-3:     Evaluate $\hat{f}_{t}=h_{t}(\mathbf{z}_{t})\leftarrow\text{SOA}(\mathbf{z}_{t})$
-4:     Commit to $\mathbf{x}_{t}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{\hat{f}_{t}}(\mathbf{z}_{t},\mathbf{x}))$
-5:     Observe follower type $f_{t}$
-6:end
-Here SOA refers to the online multiclass learning algorithm proposed by [17].
-A.3 Online Learning in the Agnostic Setting
-A.3.1 Upper bound
-Algorithmic procedure.
-Consider the “learning with expert advice” setting where the leader at each timestep $t$ would get to observe not only the context $\mathbf{z}_{t}$ but also $N$ advices, $\{\phi_{t}^{(1)},\cdots,\phi_{t}^{(N)}\}\in\Delta(\pazocal{A})^{N}$.
-Then we can denote the cumulative loss of expert $i$ at timestep $T$ as:
-$L_{T,i}=-\sum_{t=1}^{T}u(\mathbf{z}_{t},\phi_{t}^{(i)},b_{f_{t}}(\mathbf{z}_{t},\phi_{t}^{(i)})).$
-For an appropriately defined set of experts we can formalize the following algorithm.
-Algorithm 3 Learning in the full-feedback, agnostic setting
-1:input: time horizon $T$
-2:Set $\eta=\sqrt{8\ln(N)/T}$
-3:for $t\in[T]$ do
-4:     Observe $\mathbf{z}_{t}$
-5:     Receive advices $\{\phi_{t}^{(1)},\cdots,\phi_{t}^{(N)}\}\in\Delta(A)^{N}$
-6:     Commit to $\mathbf{x}_{t}=\phi_{t}^{(i)}$ with probability proportional to $\exp(-\eta L_{t-1,i})$
-7:     Receive utility $u(\mathbf{z}_{t},a_{l},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$, where $a_{l}$ is drawn from $\mathbf{x}_{t}$
-8:     Observe follower type $f_{t}$
-9:end
-See 3.15
-Proof.
-First consider the case where we construct an expert for every $h\in\mathscr{H}$ and keep track of their cumulative loss $L_{t,h}$ at each timestep $t$.
-In our case, for each $h\in\pazocal{H}$, we can construct an expert $i$ that will predict the type, $f_{t}^{(i)}=h(\mathbf{z}_{t})$ at that timestep and based on that evaluate:
-$\phi_{t}^{(i)}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u(\mathbf{z}_{t},a_{l},b_{f_{t}^{(i)}}(\mathbf{z}_{t},\mathbf{x})).$
-From the regret analysis of the Hedge algorithm (see [14] Chapter 2), we know that:
-$\sum_{t\in[T]}\mathbb{E}\left[u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))\right]\geq\max_{i\in[N]}\sum_{t=1}^{T}u(\mathbf{z}_{t},\phi_{t}^{(i)},b_{f_{t}}(\mathbf{z}_{t},\phi_{t}^{i}))-\sqrt{\frac{1}{2}\ln(|\mathscr{H}|)T},$
-where the expectation is taken over the randomness in the Hedge algorithm.
-Now it remains to construct a smaller set of experts such that the upper bound on the regret depends on the $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$, instead of $\ln(|\pazocal{H}|)$.
-For some time horizon $T$, we define $S_{T}=\{S\subset[T]:|S|\leq\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})\}$.
-For each $S\in S_{T}$ and $g:S\rightarrow\Delta(\pazocal{A})$, we construct an expert, $E_{S,g}$.
-Expert $E_{S,g}$ immitates an instance of the SSOA algorithm, where SSOA errs at all timesteps $t\in S$ and for each timestep $t\in S$, the optimal strategy for the leader at the timestep is given by $g(t)$.
-We show the implementation of each expert in Algorithm 4 below.
-Algorithm 4 Implementation of expert $E_{S,g}$
-1:input: time horizon $T$
-2:Set $V_{0}=\pazocal{H}$
-3:for $t\in[T]$ do
-4:     Observe $\mathbf{z}_{t}$
-5:     Let $\mathbf{x}^{(h,t)}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{h(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))\quad\forall h\in\pazocal{H}$
-6:     Let $X_{t}=\{\mathbf{x}^{(h)}:h\in V_{t-1}\}$triangleright $|X_{t}|\leq K$
-7:     Let $V_{t}^{(\mathbf{x})}=\{h\in V_{t-1}:\mathbf{x}^{(h,t)}=\mathbf{x}\}\quad\forall\mathbf{x}\in X_{t}$
-8:     Play action $l_{t}=\arg\max_{\mathbf{x}\in X_{t}}\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})$
-9:     Set $l_{t}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}\textnormal{SLdim}_{\pazocal{G}}(V_{t}^{\mathbf{z}_{t}\rightarrow k})$
-10:     if $t\in S$ then
-11:         Predict $g(t)$
-12:         $V_{t+1}\leftarrow V_{t}^{\mathbf{z}_{t}\rightarrow g(t)}$
-13:     else
-14:         Predict $l_{t}$
-15:         $V_{t+1}\leftarrow V_{t}^{\mathbf{z}_{t}\rightarrow l_{t}}$
-16:end
-Notice that for each possible timestep $t\in T$, a single context $\mathbf{z}_{t}$ is presented, so there are at most $K$ possible optimal strategies for the leader, one for each follower type.
-So, the number of experts that can be constructed in this way is:
-$\sum_{i=0}^{\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H})}\binom{T}{i}(K-1)^{i}\leq(TK)^{\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H})}.$
-So using the above regret upper bound with this set of experts we get that Algorithm 3 achieves utility at least:
-$\max_{S,g}-L_{S,g,T}-\sqrt{\frac{1}{2}\textnormal{SLdim}_{\pazocal{G}}(\mathscr{H})T\ln(TK)},$
-where $-L_{S,g,T}$ is the accumulated utility of expert $E_{S,g}$ up to timestep $T$.
-It remains to show that $\max_{S,g}-L_{S,g,T}\geq\max_{h\in\mathscr{H}}-L_{h,T}$.
-Notice that for any $h\in\mathscr{H}$ and for any set of timesteps $S$ that SSOA errs in, there exists an expert in the constructed set of experts such that $L_{S,g,T}=L_{h,T}$ when executed on the sequence $(\mathbf{z}_{1},h(\mathbf{z}_{1})),\cdots,(\mathbf{z}_{T},h(\mathbf{z}_{T}))$.
-A.3.2 Lower bound
-See 3.16
-Proof.
-We let $d=\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
-For simplicity assume that $T$ is an integer multiple of $d$, say $T=kd$ for $k\in\mathbb{N}$.
-Consider a tree, $(\mathbf{v}_{1},\mathbf{v}_{1},\ldots,\mathbf{v}_{2^{d}-1})$, of depth, $d$, that is SL-shattered by $\pazocal{H}$ such that each node $\mathbf{v}_{i}\in\pazocal{Z}$ has two edges (each leading to a child) corresponding to two strategies for the leader $\mathbf{x}_{i,1},\mathbf{x}_{i,2}\in\Delta(\pazocal{A})$, such that $\mathbf{x}_{i,1}\neq\mathbf{x}_{i,2}$.
-By definition, we have that for each $i$, there exist at least two functions $h_{i,1},h_{i,2}\in\pazocal{H}$ such that at context $\mathbf{v}_{i}$, they predict follower types $f_{1}=h_{i,1}(\mathbf{v}_{i})$ and $f_{2}=h_{i,2}(\mathbf{v}_{i})$, against which the strategies $\mathbf{x}_{i,1}$ and respectively $\mathbf{x}_{i,2}$ are optimal.
-In other words, we construct a Stackelberg game, playing the optimal strategy against the mispredicted follower type gives utility 00, whereas playing the optimal strategy against the correct follower type gives utility $1$ (for an explicit construction see proof of Theorem 3.8).
-Then, we have that for every $i$:
-$u(\mathbf{v}_{i},\mathbf{x}_{i,1},b_{f_{1}}(\mathbf{v}_{i},\mathbf{x}_{i,1}))-u(\mathbf{v}_{i},\mathbf{x}_{i,2},b_{f_{1}}(\mathbf{v}_{i},\mathbf{x}_{i,2}))=1$
-and similarly,
-$u(\mathbf{v}_{i},\mathbf{x}_{i,2},b_{f_{2}}(\mathbf{v}_{i},\mathbf{x}_{i,2}))-u(\mathbf{v}_{i},\mathbf{x}_{i,1},b_{f_{2}}(\mathbf{v}_{i},\mathbf{x}_{i,1}))=1.$
-We will now construct a sequence $(\mathbf{z}_{1},\mathbf{x}_{1}),(\mathbf{z}_{2},\mathbf{x}_{2}),\ldots,(\mathbf{z}_{T},\mathbf{x}_{T})$, which we will show forces any algorithm to meet the lower bound in the theorem statement.
-From the SL-shattered tree, we follow a root-to-leaf path, say $(\tilde{\mathbf{z}}_{1},\tilde{\mathbf{x}}_{1}),\ldots,(\tilde{\mathbf{z}}_{d},\tilde{\mathbf{x}}_{d})$.
-For each node $\tilde{\mathbf{z}}_{i}$ on the path, where $i\in[d]$, we construct a set of $k$ pairs of contexts and follower types $\{(\mathbf{z}_{(i-1)k+1}^{a},\mathbf{x}_{(i-1)k+1}^{a}),\ldots,(\mathbf{z}_{ik}^{a},\mathbf{x}_{ik}^{a})\}$ as follows.
-  • We set $\mathbf{z}_{(i-1)k+1}^{a}=\ldots=\mathbf{z}_{ik}^{a}=\tilde{\mathbf{z}}_{i}$.
-  • We draw each $\mathbf{x}_{t}^{a}$ for $t\in\{(i-1)k+1,\ldots,ik\}$, independently and uniformly at random from the two optimal strategies corresponding to the edges leading to the two children of $\mathbf{z}_{i}$.
-We denote the follower type and optimal strategy for each child, $f_{A,i},\mathbf{x}_{A,i}$ and $f_{B,i},\mathbf{x}_{B,i}$ respectively.
-For each set of $k$ examples, let $T_{i}=\{(i-1)k+1,\ldots,ik\}$ be the set of indeces of the examples corresponding to $\tilde{\mathbf{z}}_{i}$. It suffices to lower bound:
-$\max_{h\in\pazocal{H}}\sum_{t\in T_{i}}u(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$
-For ease of notation, we denote $h^{*}=\arg\max_{h\in\pazocal{H}}\sum_{t\in T_{i}}u(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t})))$, $\;u_{h^{*},t}=u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))$, $u_{t}=u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$, and $u_{t}^{*}=u(\mathbf{z}_{t},\pi_{f_{t}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{f_{t}}(\mathbf{z}_{t})))$, where $\pi_{f_{t}}(\mathbf{z}_{t})$ corresponds to the optimal policy against the true follower type and context at time $t$.
-So equivalently, we want to lower bound:
-$\sum_{t\in T_{i}}u_{h^{*},t}-u_{t}=\sum_{t\in T_{i}}(u_{t}^{*}-u_{t})-(u_{t}^{*}-u_{h^{*},t}).$
-Let $r_{i}$ be the number of timesteps in $T_{i}$, where $f_{A,i}$ appears in the example. Then:
-$u_{t}^{*}-u_{h^{*},t}=\begin{cases}k-r\quad\text{if}\;r\geq k/2\\ r\hskip 28.45274pt\text{if}\;r<k/2\end{cases}$
-Therefore, we get that $k/2-\sum_{t\in T_{i}}u_{t}^{*}-u_{h^{*},t}=|r-k/2|$. By taking an expectation over the randomness in the $f_{t}$’s, we get that:
-$k/2-\mathbb{E}\left[\sum_{t\in T_{i}}u_{t}^{*}-u_{h^{*},t}\right]=\mathbb{E}\left[|r-k/2|\right]\geq\sqrt{k/8},$
-where the last inequality holds by Khinchine’s inequality (see Lemma A.9 in [14]). Now it suffices to sum over all $i\in[d]$ to get:
-$\frac{dk}{2}-\mathbb{E}\left[\sum_{t\in[T]}u_{t}^{*}-u_{h^{*},t}\right]=\sum_{i\in[d]}\left(\frac{k}{2}-\mathbb{E}\left[\sum_{t\in T_{i}}u_{t}^{*}-u_{h^{*},t}\right]\right)\geq d\sqrt{k/8}$
-Finally, notice that $dk/2=T/2=\sum_{t\in[T]}(u_{t}^{*}-u_{t})$ for any algorithm, which concludes the proof. ∎
-Appendix B Details and Proofs from Section 4
-See 4.7
-Proof.
-We have that:
-$\mathbb{P}_{S_{m}}\left[u_{\mathscr{D}}(\pi_{h^{*}})-u_{\mathscr{D}}(\pi_{\text{ERM}})\geq\epsilon\right]$
-$\leq$ $\mathbb{P}_{S_{m}}\left[\mathop{\mathbb{E}}_{\mathbf{z}\sim\mathscr{D}}\left[u(\mathbf{z},\pi_{h^{*}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{h^{*}}(\mathbf{z})))-u(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z})))\right]>\epsilon\right]$
-By the boundedness of the utility function we have that the expectation above lies within $[-1,1]$. Therefore, we get:
-$\leq$ $\mathbb{P}_{S_{m}}\left[\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}\left[u(\mathbf{z},\pi_{h^{*}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{h^{*}}(\mathbf{z})))-u(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z})))>0\right]>\epsilon\right]$
-But getting less utility than the optimal predictor in the class, implies that our prediction of the follower’s type is incorrect. So,
-$\leq$ $\mathbb{P}_{S_{m}}\left[\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}\left[h_{*}(\mathbf{z})\neq h_{\text{ERM}}(\mathbf{z})\right]>\epsilon\right]$
-But from Theorem 6 in [17], we know that for any $\epsilon,\delta>0$, there exists constant $C_{2}$ such that for any $m\geq C_{2}\left(\frac{d_{N}(\pazocal{H}))\ln(K)\ln(\frac{1}{\epsilon})+\ln(\frac{1}{\delta})}{\epsilon}\right)$, we have that for any $h_{\text{ERM}}$ that satisfies equation (3):
-$\mathbb{P}_{S_{m}}\left(\text{Err}_{\mathscr{D}}(h_{\text{ERM}})>\text{Err}_{\mathscr{D}}(\pazocal{H})+\epsilon\right)\leq\delta$
-Since we are in the realizable setting, we know that $\text{Err}_{\mathscr{D}}(\pazocal{H})=0$.
-By definition, we have that $\text{Err}_{\mathscr{D}}(h_{\text{ERM}})>\epsilon\iff\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}\left[h_{\text{ERM}}(\mathbf{z})\neq h^{*}(\mathbf{z})\right]>\epsilon$, which concludes our proof.
-∎
-B.1 Proof of Theorem 4.12
-See 4.12
+We prove an appropriate generalization in Lemma˜B.24.
 Proof.
@@
 for some large enough constant $C$, then with probability at least $1-\delta$,
-$\mathbb{P}_{S_{m}}\left(u_{\mathscr{D}}(\pi_{\text{S-ERM}})\geq u_{\mathscr{D}}(\pi_{h^{*}})-\epsilon\right)$ (5)
-where $\pi_{h^{*}}(\mathbf{z}):=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h^{*}(\mathbf{z})}(\mathbf{z},\mathbf{x}))$.
-In other words, it suffices to prove that all the policies in $\Pi_{\pazocal{H},\pazocal{G}}$ with $u_{\mathscr{D}}(\pi_{h})<u_{\mathscr{D}}(\pi_{h^{*}})-\epsilon$ have low utility in the training sample too, that is, $u_{S}(\pi_{h})\leq u_{S}(\pi_{h^{*}})$.
+$\mathbb{P}_{S_{m}}\left(u_{\mathscr{D}}(\pi_{\text{S-ERM}})\geq u_{\mathscr{D}}(\pi_{h^{*}})-\epsilon\right).$ (5)
+In other words, it suffices to prove that all the policies in $\Pi_{\pazocal{H},\pazocal{G}}$ with $u_{\mathscr{D}}(\pi_{h})<u_{\mathscr{D}}(\pi_{h^{*}})-\epsilon$ have low utility in the training sample too, that is, $u_{S}(\pi_{h})<u_{S}(\pi_{h^{*}})$.
 It suffices then to bound the following bad event:
@@
 Thus, it suffices to bound $\mathbb{P}[B^{\prime}]$ (this probability is over the choices of $S$ and $S^{\prime}$). Given two samples $S$ and $S^{\prime}$, consider the random process SwapR:
-  • For $i$ from $1$ to $m$, flip a fair coin.
+  • For $i$ from 111 to $m$, flip a fair coin.
   • If heads, swap $\mathbf{z}_{i}$ and $\mathbf{z}^{\prime}_{i}$, else do nothing.
@@
 $B^{\prime\prime}:\exists\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}\text{ with }L_{T}(\pi_{h})=0\text{ but }L_{T^{\prime}}(\pi_{h})>\epsilon/2.$
-By Claim B.2, $\mathbb{P}[B^{\prime\prime}]=\mathbb{P}[B^{\prime}]$ so, it suffices to upper bound $\mathbb{P}[B^{\prime\prime}]$.
+By Lemma˜B.20, $\mathbb{P}[B^{\prime\prime}]=\mathbb{P}[B^{\prime}]$ so, it suffices to upper bound $\mathbb{P}[B^{\prime\prime}]$.
 Notice that the first probability is over the choice of $S$, $S^{\prime}$, and the randomness of SwapR, while the second is over the choice of $S$ and $S^{\prime}$.
@@
 Proof.
-Consider:
+We consider
 $\pi_{h}(\mathbf{z}_{1}),\pi_{h}(\mathbf{z}_{2}),\ldots,\pi_{h}(\mathbf{z}_{m})$
@@
 Proof.
-By definition, we have that:
+By definition, we have that
 $\mathbb{P}[B^{\prime\prime}]$ $=\mathop{\mathbb{E}}_{S,S^{\prime}}[\mathbb{P}_{\texttt{SwapR}}[\exists\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}},\;L_{T}(h)=0\land L_{T^{\prime}}(h)\geq m\epsilon/2|S,S^{\prime}]]$
 $=\mathop{\mathbb{E}}_{S,S^{\prime}}[\mathbb{P}_{\texttt{SwapR}}[\exists\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}[S\cup S^{\prime}],\;L_{T}(h)=0\land L_{T^{\prime}}(h)\geq m\epsilon/2|S,S^{\prime}]].$
-Using the union bound, we get that:
+Using the union bound, we get that
 $\mathbb{P}[B^{\prime\prime}]$ $\leq\mathop{\mathbb{E}}_{S,S^{\prime}}\Bigr{[}\sum_{\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}[S\cup S^{\prime}]}\mathbb{P}_{\texttt{SwapR}}[\;L_{T}(\pi_{h})=0\land L_{T^{\prime}}(\pi_{h})\geq m\epsilon/2|S,S^{\prime}]\Bigr{]}$
 $\leq\Pi_{\pazocal{H},\pazocal{G}}[2m]2^{-\epsilon m/2}$
-Combining all these, we get that $\mathbb{P}[B]\leq\delta$, whenever $2\Pi_{\pazocal{H},\pazocal{G}}[2m]2^{-\epsilon m/2}\leq\delta$, which proves that:
+Combining all these, we get that $\mathbb{P}[B]\leq\delta$, whenever $2\Pi_{\pazocal{H},\pazocal{G}}[2m]2^{-\epsilon m/2}\leq\delta$, which proves that
 $m>\frac{2}{\epsilon}\Bigr{[}\log_{2}(2\Pi_{\pazocal{H},\pazocal{G}}[2m])+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}\Bigr{]}$
 It remains to show that $\Pi_{\pazocal{H},\pazocal{G}}\leq d_{SN,\pazocal{G}}(\pazocal{H})\ln(K)$.
-We prove this using the Lemma 4.13, which states that $\Pi_{\pazocal{H},\pazocal{G}}[m]\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$.
-So we get that:
+We prove this using Lemma˜B.24, which states that $\Pi_{\pazocal{H},\pazocal{G}}[m]\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$.
+So we get that
 $m$ $>\frac{2}{\epsilon}\Bigr{[}\log_{2}(2\cdot(2m)^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})})+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}\Bigr{]}$
 $=\frac{2}{\epsilon}\Bigr{[}1+d_{SN,\pazocal{G}}(\pazocal{H})\log_{2}(m)+d_{SN,\pazocal{G}}(\pazocal{H})+2d_{SN,\pazocal{G}}(\pazocal{H})\log(K)+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}\Bigr{]}$
-Rearranging, we obtain the requirement that:
+Rearranging, we obtain the requirement that
 $m>\frac{2d_{SN,\pazocal{G}}(\pazocal{H})}{\epsilon}\Bigr{[}\log_{2}(m)\Bigr{]}+\frac{2}{\epsilon}\Bigr{[}2d_{SN,\pazocal{G}}(\pazocal{H})\log(K)+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}+d_{SN,\pazocal{G}}(\pazocal{H})+1\Bigr{]}$
-Using lemma B.5, a sufficient condition for the above to hold is that:
+Using Lemma˜B.23, a sufficient condition for the above to hold is that
 $m>\frac{8d_{SN,\pazocal{G}}(\pazocal{H})}{\epsilon}\Bigr{[}\log_{2}\Bigl{(}\frac{4d_{SN,\pazocal{G}}(\pazocal{H})}{\epsilon}\Bigl{)}\Bigr{]}+\frac{4}{\epsilon}\Bigr{[}2d_{SN,\pazocal{G}}(\pazocal{H})\log(K)+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}+d_{SN,\pazocal{G}}(\pazocal{H})+1\Bigr{]}$
-And a sufficient condition for the above is:
+And a sufficient condition for the above is
 $m>\frac{8d_{SN,\pazocal{G}}(\pazocal{H})}{\epsilon}\Bigr{[}\log_{2}\Bigl{(}\frac{4d_{SN,\pazocal{G}}(\pazocal{H})\cdot K}{\epsilon}\Bigl{)}\Bigr{]}+\frac{4}{\epsilon}\Bigr{[}\log\Bigl{(}\frac{1}{\delta}\Bigl{)}+d_{SN,\pazocal{G}}(\pazocal{H})+1\Bigr{]}$
-The above holds if there exists a constant $C$, large enough such that:
+The above holds if there exists a constant $C$, large enough such that
 $m>C\cdot\frac{1}{\epsilon}\Bigr{[}d_{SN,\pazocal{G}}(\pazocal{H})\log_{2}\Bigl{(}\frac{d_{SN,\pazocal{G}}(\pazocal{H})\cdot K}{\epsilon}\Bigl{)}+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}\Bigr{]}.$
@@
 Let $a\geq 1$ and $b>0$. Then $x\geq 4a\log(2a)+2b$ implies that $x\geq a\log(x)+b$.
-B.2 Proof of Lemma 4.13
-See 4.13
+Lemma B.24.
+For a given class $\pazocal{H}$ and Stackelberg game $\pazocal{G}$, let $d_{SN,\pazocal{G}}(\pazocal{H})=d$.
+Then for a set of contexts $S=\{\mathbf{z}_{1},\dots,\mathbf{z}_{|S|}\}$, let $\Pi_{\pazocal{H},\pazocal{G}}(S)=\{(\pi(\mathbf{z}_{1}),\dots,\pi(\mathbf{z}_{|S|}))\;\;|\;\;\pi\in\Pi_{\pazocal{H}},\pazocal{G}\}$.
+It holds that
+$|\Pi_{\pazocal{H},\pazocal{G}}(S)|\leq|S|^{d}\cdot K^{2d}.$ (6)
 Proof.
-We will show that for $|Z|=m$:
-$|\Pi_{\pazocal{H},\pazocal{G}}|$ |$\leq|\{B\subseteq\pazocal{Z}:\pazocal{H}\;\text{SN-shatters}\;B\}|\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$ (6)
-$\leq\sum_{i=1}^{d_{SN,\pazocal{G}}(\pazocal{H})}\binom{m}{i}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$ (7)
-$\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$ (8)
-Notice that equation (7) holds by enumerating all possible subsets of $\pazocal{Z}$ that can be SN-shattered by $\pazocal{H}$ up to $d_{SN,\pazocal{G}}(\pazocal{H})$, which is the largest size subset that $\pazocal{H}$ can shatter.
-So it suffices to prove equation (6).
-We proceed by induction on the size of $|\pazocal{Z}|$.
-Base Case: For $|\pazocal{Z}|=1$, we have that $|\Pi_{\pazocal{H},\pazocal{G}}|\leq K$ so (6) holds.
-Inductive Hypothesis: For $|\pazocal{Z}|=m$ assume that (6) holds.
-Inductive Step: We will show (6) for $|S|=m+1$. Consider the set $\pazocal{Z}=\{\mathbf{z}_{0},\mathbf{z}_{1},\cdots,\mathbf{z}_{m}\}$ and we define the following notation:
-$\Lambda^{(i)}=\{\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}|\pi_{h}(\mathbf{z}_{0})=i\}$
-$\Lambda^{(i,j)}=\{\pi_{h}\in\Lambda^{(i)}|\exists\pi_{h}^{\prime}\in\Lambda^{(j)}\;\text{s.t.}\;\pi_{h}(\mathbf{z})=\pi_{h}^{\prime}(\mathbf{z})\;\forall\mathbf{z}\in\pazocal{Z}\setminus\{\mathbf{z}_{0}\}\}$
-For $i,j\in[K]$ also denote $\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}$ to be all the policies $\pi_{f}:\pazocal{Z}\setminus\{\mathbf{z}_{0}\}\to[K]$ that can be extended to a policy in $\Pi_{\pazocal{H},\pazocal{G}}$ both by defining $\pi_{f}(\mathbf{z}_{0})=i$ and by defining $\pi_{f}(\mathbf{z}_{0})=j$.
-Then we claim the following:
-$|\Pi_{\pazocal{H},\pazocal{G}}|\leq|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}|\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|,$ (9)
-where $|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|$ denotes the restriction of $\Pi_{\pazocal{H},\pazocal{G}}$ to the set $\pazocal{Z}\setminus\{\mathbf{z}_{0}\}$.
-(Proof of equation (9)).
-Consider the $\ell$’th labeling of elements in $\pazocal{Z}\setminus\{\mathbf{z}_{0}\}$ that is consistent with a policy in $\Pi_{\pazocal{H},\pazocal{G}}$.
-Let $\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}$ denote the policies in $\Pi_{\pazocal{H},\pazocal{G}}$ that are consistent with that labeling.
-Then we have that:
-$|\Pi_{\pazocal{H},\pazocal{G}}|$ $=\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}\mathbbm{1}\{\exists\pi\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\}$
-$=\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}1+\left(\sum_{i=1}^{K}\mathbbm{1}\{\exists\pi\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\}\right)-1$
-$=|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}\mathbbm{1}\{\exists\pi\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\}-1$
-$\leq|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}-1+\sum_{j=1}^{K}\mathbbm{1}\{\exists\pi,\pi^{\prime}\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\wedge\pi^{\prime}(\mathbf{z}_{0})=j\}$
-$=|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}\left(\mathbbm{1}\{\exists\pi\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\}-1\right)$
-$\hskip 85.00012pt+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}\mathbbm{1}\{\exists\pi,\pi^{\prime}\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\wedge\pi^{\prime}(\mathbf{z}_{0})=j\}$
-$\leq|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{\ell=1}^{|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|}\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}\mathbbm{1}\{\exists\pi,\pi^{\prime}\in\Pi_{\pazocal{H},\pazocal{G}}^{(\ell)}:\pi(\mathbf{z}_{0})=i\wedge\pi^{\prime}(\mathbf{z}_{0})=j\}$
-$=|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|+\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}|\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|.$ (10)
-We proceed with the induction. From the inductive hypothesis we have that:
-$|\displaystyle|\Pi_{\pazocal{H},\pazocal{G}}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}|$$\leq|\{B\subseteq\pazocal{Z}\setminus\{\mathbf{z}_{0}\}:\pazocal{H}\;\text{SN-shatters}\;B\}|\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$
-$\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})},$ (11)
-where the last inequality holds from equation (8).
-Also, notice that each $\Lambda^{(i,j)}$ SN-shatters $\mathbf{z}_{0}$ by definition.
-So $\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{x_{0}\}}$ cannot shatter a set larger than $d_{SN,\pazocal{G}}(\pazocal{H})-1$.
-Therefore, by the inductive hypothesis we get that:
-$|\displaystyle|\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{x_{0}\}}|$$\leq|\{B\subseteq\pazocal{Z}\setminus\{\mathbf{z}_{0}\}:\Lambda^{(i,j)}|_{\pazocal{Z}\setminus\{\mathbf{z}_{0}\}}\;\text{SN-shatters}\;B\}|\cdot K^{2(d_{SN,\pazocal{G}}(\pazocal{H})-1)}$
-$\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})-1}\cdot K^{2(d_{SN,\pazocal{G}}(\pazocal{H})-1)}$ (12)
-Combining inequalities (10), (11), and (12), we get that:
-$|\Pi_{\pazocal{H},\pazocal{G}}|$ $\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}+\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}m^{d_{SN,\pazocal{G}}(\pazocal{H})-1}\cdot K^{2(d_{SN,\pazocal{G}}(\pazocal{H})-1)}$
-$=m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}+m^{d_{SN,\pazocal{G}}(\pazocal{H})-1}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$
-$=m^{d_{SN,\pazocal{G}}(\pazocal{H})-1}K^{2d_{SN,\pazocal{G}}(\pazocal{H})}(m+1)$
-$\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$
+[…]
+We will now show Equation˜6 for $|S|=m+1$.
+First we fix a set $S=\{\mathbf{z}_{0},\mathbf{z}_{1},\cdots,\mathbf{z}_{m}\}$.
+For simplicity and without loss of generality we write $\Pi_{\pazocal{H},\pazocal{G}}(S)$ as $\Pi_{\pazocal{H},\pazocal{G}}$.
+We now define the following
+$\Lambda^{(i)}=\{\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}|\pi_{h}(\mathbf{z}_{0})=\mathbf{x}_{*}^{(i)}\}$
+$\Lambda^{(i,j)}=\{\pi_{h}\in\Lambda^{(i)}|\exists\pi_{h}^{\prime}\in\Lambda^{(j)}\;\text{s.t.}\;\pi_{h}(\mathbf{z})=\pi_{h}^{\prime}(\mathbf{z})\;\forall\mathbf{z}\in\pazocal{Z}\},$
+where $\mathbf{x}_{*}^{(i)}$ is the leader’s optimal strategy against $f^{(i)}$, given context $\mathbf{z}_{0}$. We also define
+$\Pi_{0}=\Pi_{\pazocal{H},\pazocal{G}}\setminus\left(\cup_{i\neq j}\Lambda^{(i,j)}\right).$
+Then it immediately follows that
+$|\Pi_{\pazocal{H},\pazocal{G}}(S)|\leq|\Pi_{0}(S)|+\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}|\Lambda^{(i,j)}(S)|.$ (7)
+By definition we know that each policy in $\Pi_{0}$, produces a distinct set of optimal strategies on the contexts in $S\setminus\{\mathbf{z}_{0}\}$.
+Therefore, $|\Pi_{0}(S)|=|\Pi_{0}(S\setminus\{\mathbf{z}_{0}\})|$.
+Also, $\Pi_{0}\subset\Pi_{\pazocal{H}},\pazocal{G}$ so $\Pi_{0}$ cannot SN-shatter a set of size greater than $d$.
+Therefore, from the inductive hypothesis we get that
+$|\Pi_{0}(S)|=|\Pi_{0}(S\setminus\{\mathbf{z}_{0}\})|\leq m^{d}\cdot K^{2d}.$ (8)
+Next, we argue that each $\Lambda^{(i,j)}$ cannot SN-shatter a set of size greater than $d-1$.
+To see this, we let $L=\{\tilde{\mathbf{z}}_{1},\dots,\tilde{\mathbf{z}}_{|}L|\}$ be a set that $\Lambda^{(i,j)}$ can SN-shatter.
+Then $L$ cannot include $\mathbf{z}_{0}$, since $\pi(\mathbf{z}_{0})=\mathbf{x}_{*}^{(i)}$ for all $\pi\in\Lambda^{(i,j)}$.
+However, $\Pi_{\pazocal{H},\pazocal{G}}$ can SN-shatter $L\cup\{\mathbf{z}_{0}\}$.
+So $|L|\leq d-1$.
+Therefore, by the inductive hypothesis we get
+$|\Lambda^{(i,j)}(S)|\leq m^{d-1}\cdot K^{2(d-1)}$ (9)
+Combining inequalities (7), (8), and (9), we get that
+$|\Pi_{\pazocal{H},\pazocal{G}}(S)|$ $\leq m^{d}\cdot K^{2d}+\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}m^{d-1}\cdot K^{2(d-1)}$
+$\leq m^{d}\cdot K^{2d}+m^{d-1}\cdot K^{2d}$
+$=m^{d-1}K^{2d}(m+1)$
+$\leq(m+1)^{d}K^{2d}$

### v3

@@
 Abstract
-We study structured Stackelberg games, in which both players (the leader and the follower) observe contextual information about the state of the world at time of play.
-The leader plays against one of a finite number of followers, but the follower’s type is not known until after the game has ended.
-Importantly, we assume a fixed relationship between the contextual information and the follower’s type, thereby allowing the leader to leverage this additional structure when deciding her strategy.
-Under this setting, we find that standard learning theoretic measures of complexity do not characterize the difficulty of the leader’s learning task.
-Instead, we introduce a new notion of dimension, the Stackelberg-Littlestone dimension, which we show characterizes the instance-optimal regret of the leader in the online setting.
-Based on this, we also provide a provably optimal learning algorithm.
-We extend our results to the distributional setting, where we use two new notions of dimension, the $\gamma$-Stackelberg-Natarajan dimension and $\gamma$-Stackelberg-Graph dimension.
-We prove that these control the sample complexity lower and upper bounds respectively, and we design a simple, improper algorithm that achieves the upper bound.
+We initiate the study of structured Stackelberg games, a novel form of strategic interaction between a leader and a follower where contextual information can be predictive of the follower’s (unknown) type.
+Motivated by applications such as security games and AI safety, we show how this additional structure can help the leader learn a utility-maximizing policy in both the online and distributional settings.
+In the online setting, we first prove that standard learning-theoretic measures of complexity do not characterize the difficulty of the leader’s learning task.
+Remarkably, we find that there exists a learning-theoretic measure of complexity, analogous to the Littlestone dimension in online classification, that tightly characterizes the leader’s instance-optimal regret.
+We term this the Stackelberg-Littlestone dimension, and leverage it to provide a provably optimal online learning algorithm.
+In the distributional setting, we provide analogous results by showing that two new dimensions control the sample complexity upper- and lower-bound.
 1 Introduction
-Many strategic interactions between a principal and an agent can be cast as Stackelberg games [38].
-In a Stackelberg game, the principal (also called the leader) commits to a possibly randomized strategy, before the agent (also called the follower) observes this strategy and best-responds to it.
-Stackelberg games have had significant real-world impact over the past decade, particularly in security domains [37].
-For example, when protecting endangered species [39], a wildlife agency determines ranger patrol locations before poachers select targets.
-Coast guards decide on shoreline patrol routes in advance of illegal fishers’ actions [17] and, airport security [33] chooses checkpoint screening strategies before smugglers decide which entry points to use.
-In many of these settings, the state of the world at time of play can greatly influence the utilities of the players and subsequently the outcome of the game.
-For example, patrolling some areas of the airport might be deemed of higher- or lower-importance based on the schedule of flights and the number of passengers on board.
-Similarly, weather patterns may make it harder to reach animals in certain parts of the park, while migration patterns might make others more vulnerable.
-When contextual information becomes available over time, leveraging it will generally help the leader deploy more effective strategies.
-This raises a natural question: given that the state of the world might evolve over time, how should the leader in a Stackelberg game design a strategy that adapts to new information coming in over time?
-Previous work [27, 13] has formalized such settings as repeated Stackelberg games with side information, where at each interaction both players observe contextual information that affects their utilities.
-In this setting, followers in different rounds may have different types that determine their payoff for each outcome of the game.
-Harris et al.
-[27] show that as long as either the sequence of followers or the sequence of contexts are drawn i.i.d.
-from a fixed distribution, no-regret learning is possible.
-That is, the leader is guaranteed to achieve nearly as much utility as the best fixed policy in hindsight.
-In the fully adversarial regime, however, they show that regret can grow linearly with the time horizon.
-Interestingly, this lower bound follows from the adversary’s ability to encode an arbitrarily complex relationship between the sequence of contexts and the follower types.
-Naturally, there are many cases in which the relationship between the contextual information and the follower about to appear next is not arbitrarily complex, but instead exhibits some structure.
-In fact, in many settings the contextual information can be predictive of the follower’s type.
-For example, in wildlife protection, tire tracks might signal the presence of poachers in a particular area of the park, and market demand trends might be indicative of which animals’ poachers are now more likely to appear.
-Similarly in airport and border security, current affairs might make it more or less likely that certain smuggler groups are active.
-Motivated by this observation, we study how the leader can learn effectively when there is additional structure between the contextual information and the follower types, thereby circumventing the previous impossibility result.
+Stackelberg games are the canonical framework for analyzing strategic interactions with commitment, where one player (the leader) commits to a strategy before another (the follower) observes this strategy and best-responds.
+While initially introduced in the context of market competition [stackelberg1934marktform], Stackelberg games have been applied across a wide range of domains including security, where a security agency commits to a patrol strategy before attackers choose vulnerable targets to attack [Sinha2018StackelbergSG], congestion control, where a city sets tolls before drivers choose routes [swamy2007effectiveness], and AI red-teaming, where AI providers deploy a model before users find adversarial use cases [shivadekar2025red, liu2025chasingmovingtargetsonline].
+Contextual Stackelberg games [harrisregret, balcan2025nearlyoptimalbanditlearningstackelberg] are a generalization of this model, in which both players’ payoffs depend on additional side information.
+This model is motivated by many applications where players have payoff-relevant information about the state of the world at the time of play, which influences their strategies (and subsequently the outcome of the game).
+For example, a wildlife park’s security might place varying levels of importance on patrolling different areas of the park depending on the time of year and animal migration patterns.
+Contextual Stackelberg games have also been used to do label-efficient fine-tuning of large language models [wang2026optimal].
+Interestingly, learning under this more general framework is not always possible.
+Specifically, harrisregret study the online setting (where the leader faces a sequence of followers with unknown utilities) and provide no-regret learning guarantees for the leader as long as either the followers or contexts are drawn i.i.d.
+from a fixed distribution.
+However, when both contextual information and follower types are chosen adversarially, the worst-case regret can grow linearly with the time horizon.
+This lower bound follows from a reduction to online classification, where an adversary can encode a mapping from contexts to follower types that is arbitrarily difficult to learn online.
+In this work, we ask how broadly this impossibility result applies and investigate the landscape of learnability in this setting.
+Specifically, we examine learnability in contextual Stackelberg games under an arbitrary, unknown mapping from contexts to follower types.
+This setting naturally captures domains where contextual data contains some signal that is predictive of the follower about to appear next.
+In wildlife protection for example, data from motion sensors and camera footage is routinely collected and used to predict the intrusion type [s18051474].
+Similarly in AI safety, different types of attack may be more or less viable depending on the domain the model is deployed in.
+We capture such applications via the structured Stackelberg game model, which we introduce.
+In this model, we study the following fundamental learnability questions:
+(1) In which structured Stackelberg games can the leader learn a utility-maximizing policy?
+(2) Are there natural dimension(s) that characterize learnability in this setting?
 1.1 Our contributions
-We revisit the framework of Harris et al.
-[27] and take a learning theoretic approach to analyzing the leader’s task in predicting the follower type from contexts.
-Importantly, we find that existing notions of complexity do not capture the difficulty of learning in this setting—we obtain (tight) Stackelberg regret and sample complexity guarantees that are parameterized by new complexity measures and that directly exploit the structure of Stackelberg games.
-Concretely, our main contributions in this work are as follows.
-Instance-optimal Stackelberg regret bounds.
-We initiate the study of online learning in Stackelberg games with side information, under the structural assumption that there exists a fixed, unknown mapping between contexts and follower types.
-In particular, we study how the leader, given access to a predefined hypothesis class, can leverage it to learn.
-We study the realizable setting, under which the class includes an optimal predictor of the followers’ types.
-Then the leader’s task is to learn an optimal policy on the fly while achieving as high cumulative utility as possible.
+Instance-optimal regret bounds.
+We initiate the study of online learning in contextual Stackelberg games under an arbitrary, unknown mapping from contexts to follower types.
+In particular, we study how the leader can learn a utility-maximizing policy, given a hypothesis class rich enough to include this underlying mapping.
 Naturally, the difficulty of the leader’s learning task scales with the complexity of the hypothesis class.
-In the learning theory literature, the (multiclass) Littlestone dimension [30, 22] is used to characterize the complexity of a class in settings with a finite set of labels (which in structured Stackelberg games corresponds to the follower types).
-Interestingly, we prove that the Littlestone dimension can give arbitrarily loose regret bounds (Theorem˜3.5).
+In the learning theory literature, the (multiclass) Littlestone dimension [littlestone1988learning, pmlr-v19-daniely11a] is used to characterize the complexity of online classification in settings with a finite set of labels (which, in structured Stackelberg games, corresponds to the set of follower types).
+We prove that the Littlestone dimension can give arbitrarily loose regret bounds in our Stackelberg game setting (Theorem˜3.6).
 Our construction leverages the fact that the Littlestone dimension is completely blind to the utility space defined by the Stackelberg game, which ultimately determines the performance of the leader.
-Based on these insights, we introduce a new notion of complexity, the Stackelberg-Littlestone dimension (Definitions˜3.6 and 3.7, which directly exploits the interaction between the hypothesis class and the utility space of the game.
-We base our definition on shattered trees, which are the workhorse of classic online learning.
-The challenge in our setting relies on having a finite set of follower types, yet a continuous utility function, which requires a much more nuanced approach to constructing the trees.
-We show that the Stackelberg-Littlestone dimension exactly characterizes the difficulty of learning in our setting (Theorem˜3.8), and in particular gives instance-optimal regret bounds.
-That is, for any hypothesis class and any Stackelberg game, the Stackelberg-Littlestone dimension identifies exactly the optimal regret.
-The proof of our upper bound Theorem˜3.9 is constructive (Algorithm˜1), which yields a complete characterization of learnability in online structured Stackelberg games.
+We introduce a new notion of complexity, the Stackelberg-Littlestone dimension (Definition˜3.9; henceforth SL dimension), designed to capture the joint complexity of the hypothesis class and the utility space of the Stackelberg game.
+Our definition is based on the notion of shattered trees, which are a classic tool in online learning.
+However in contrast to previous work, our setting involves a discrete set of follower types, yet a continuous utility function, so constructing the trees requires a more nuanced approach.
+We prove that the SL dimension tightly characterizes the complexity of learning in structured Stackelberg games.
+Specifically, we prove that for any hypothesis class and any Stackelberg game, the SL dimension both upper- and lower-bounds the optimal regret for the leader (Theorem˜3.10, Theorem˜3.11).
+The proof of our upper bound is constructive (Algorithm˜1), which yields a complete characterization of learnability in online structured Stackelberg games.
 PAC learning guarantees.
-We introduce and formalize the study of distributional learning in structured Stackelberg games (Section˜4).
-The leader has access to a set of (context, follower type) examples, where the contexts are drawn from some fixed, unknown distribution and the follower types are determined by a true hypothesis in a given hypothesis class.
-Upon the draw of a new context, the leader’s goal is to deploy a strategy such as to maximize her expected payoff.
+We also formalize the study of distributional learning in structured Stackelberg games (Section˜4).
+In this setting, the leader has access to a set of (context, follower type) examples, where the contexts are drawn from some fixed, unknown distribution and the follower types are determined by some hypothesis in the hypothesis class.
+Upon the draw of a new context, the leader’s goal is to deploy a strategy such as to maximize her expected utility.
 We show generalization guarantees with probably approximately correct (PAC)-style bounds.
-We introduce the notion of $\gamma$-Stackelberg-Natarajan dimension (Definitions˜4.9 and 4.10) and $\gamma$-Stackelberg-Graph dimension (Definitions˜4.12 and 4.13), which respectively control the lower and upper bounds on the sample complexity (Theorems˜4.11 and 4.14).
-Our definitions are inspired by analogous notions in classification and regression [31, 3, 21], yet we give a tailored approach to what constitutes a “mistake”.
-This leverages the discreteness of the set of follower types to determine how much hypotheses “differ” in a minmax sense.
-We give a simple, improper algorithm (Algorithm˜2) that with high probability achieves (almost) optimal expected utility with as many samples as our sample complexity upper bound.
-We also analyze a simpler, context-free setting, in which the training set is only over the follower types.
-Using tools from data-driven algorithm design, we show strong generalization guarantees Theorem˜4.5.
-Optimal strategy misprediction bounds.
-In the Appendix, we study structured Stackelberg games under a relaxed benchmark that seeks to minimize the total rounds in which the leader does not play the optimal strategy against the realized follower type.
-We give results for both the online (Appendix˜B) and distributional setting (Section˜B.2), and in both cases, we show stronger guarantees than those achieved by applying existing approaches from learning theory.
-These results are both conceptually and quantitatively superseded by the ones we present in the main body and we include them for completeness and further insight.
-We summarize our results in  Table˜1.
-Setting: Online Distributional
-Dimension Definitions˜3.6 and 3.7 Definitions˜4.10 and 4.13
-Lower Bound Theorem˜3.8 Theorem˜4.11
-Upper Bound Theorem˜3.9 Theorem˜4.14
-Algorithm Algorithm˜1 Algorithm˜2
-Table 1: Summary of our main results.
+We introduce the notion of the $\gamma$-SN dimension (Definitions˜4.2 and 4.3) and the $\gamma$-SG dimension (Definitions˜4.5 and 4.6), which control the lower- and upper-bounds on the sample complexity (Theorems˜4.4 and 4.7).
+Our definitions are inspired by analogous notions in classification and regression [natarajan1989learning, Attias23:regression, daniely2014:optimal], but we rely on a tailored approach to what constitutes a “mistake” in our setting.
+Our approach leverages the discreteness of the set of follower types to determine how much hypotheses differ in a minmax sense.
+We give a simple, improper algorithm (Algorithm˜2) that with high probability achieves nearly-optimal expected utility with as many samples as in our sample complexity upper bound.
+In the Appendix, we analyze a simpler, context-free offline setting, in which the training set is only over the distribution of follower types.
+Using tools from data-driven algorithm design, we show strong generalization guarantees (Theorem˜C.8).
+Moreover, all of our results directly extend to the related problems of learning to bid in auctions with side information and Bayesian persuasion with public and private states.
+We discuss this in Appendix˜A.
+Finally, computational considerations are discussed in Appendix˜D.
 1.2 Related work
 Learning in Stackelberg games.
-The problem of learning an optimal strategy for the leader with a single unknown follower type has been studied by Letchford et al.
-[29] and Blum et al.
-[16].
-These results have been extended by Peng et al.
-[32], who give faster rates and almost-matching lower bounds.
-Recently, Bacchiocchi et al.
-[4] relax some structural assumptions from previous work, including tie-breaking rules for the follower, and revise worst-case sample complexity results of learning the leader’s optimal mixed strategy.
-Our main focus is on online learning in Stackelberg games.
-Online Stackelberg games with multiple follower types were introduced in Balcan et al.
-[7].
-The authors show that when a series of unknown attacker types is chosen adversarially, playing the well-known Hedge algorithm  [25] over a carefully selected set of mixed strategies yields a no-regret algorithm.
-Our departing point is the work of Harris et al.
-[27], in which the authors study the problem of online learning in Stackelberg games with side information.
+The problem of learning an optimal strategy for the leader with a single unknown follower type has been studied by letchford2009learning and blum2014learning.
+These results have been extended by peng2019learning, who give faster rates and almost-matching lower bounds.
+Recently, bacchiocchi2024sample relax some structural assumptions from previous work, including tie-breaking rules for the follower, and revise worst-case sample complexity results of learning the leader’s optimal mixed strategy.
+Online Stackelberg games with multiple follower types were introduced in balcan2015commitment.
+The authors show that when a series of unknown attacker types is chosen adversarially, playing the Hedge algorithm  [freund1997decision] over a carefully selected set of mixed strategies is sufficient to obtain no-regret.
+Our departing point is the work of harrisregret, in which the authors study the problem of online learning in Stackelberg games with side information.
 The authors show that when either the sequence of follower types or the sequence of contexts are drawn independently from some fixed, unknown distribution, no-regret learning is indeed possible.
-Follow-up work [13] obtains better regret guarantees under bandit feedback.
+Follow-up work [balcan2025nearlyoptimalbanditlearningstackelberg] obtains better regret guarantees under bandit feedback.
 However, when both sequences are chosen adversarially, regret can scale linearly in the worst case, as the problem can be reduced from online learning linear thresholds.
-In contrast to this line of work, we circumvent this impossibility result by making an additional structural assumption about the relationship between contexts and follower types—namely that it belongs to a class of functions that the leader has access to.
+We circumvent this impossibility result by making an additional structural assumption about the relationship between contexts and follower types—namely that it belongs to a class of functions that the leader has access to.
 Under this assumption, we show that online learning is indeed possible, and our guarantees depend on an appropriate measure of complexity of this function class.
+wang2026optimal cast the problem of budget-aware supervised fine-tuning of large language models as a contextual Stackelberg game, and provide results for the setting where the leader does not observe the follower’s type by default, but can query it at a cost.
 Multiclass learnability.
-Our work can also be viewed through the lens of multiclass learnability, as it suffices for the leader to learn the hypothesis class of the function mapping the contexts to a (finite) set of follower types.
-Then the leader can compute the optimal strategy against the predicted follower.
-Multiclass prediction has been studied extensively both in the ditributional ([31, 22, 18]) and in the online setting ([22, 26]).
-Interestingly, we show that learning in Stackelberg games with contexts is at most as hard as multiclass prediction, and in some cases, can be much easier.
-For example, we give a problem instance the leader can learn to play optimally under a hypothesis class that has infinite Littlestone dimension.
-Online Learning beyond Multiclass Prediction.
-There has been tremendous interest in understanding online learnability in many settings of interest.
-Notably, in online realizable regression Daskalakis and Golowich [23] designed a proper algorithm that achieved near-optimal cumulative absolute loss with respect to the sequential fat-shattering dimension of the hypothesis class.
-Our work is more closely related to Attias et al.
-[3], who introduce the online (resp.
+Our work can also be viewed through the lens of multiclass learnability, as it suffices for the leader to (1) learn the hypothesis class of the function mapping the contexts to a (finite) set of follower types, then (2) compute the optimal strategy against the predicted follower.
+Multiclass prediction has been studied extensively both in the distributional ([natarajan1989learning, pmlr-v19-daniely11a, brukhim2022characterization]) and online ([pmlr-v19-daniely11a, hanneke2023multiclass]) settings.
+Our results show that just relying on multiclass prediction tools that do not leverage the additional structure of the problem can lead to loose regret guarantees.
+For example, it is possible to construct a problem instance in which the leader can learn to play optimally under a hypothesis class that has infinite Littlestone dimension.
+Learning theory beyond multiclass prediction.
+In online realizable regression daskalakis2022fast designed a proper algorithm that achieved near-optimal cumulative absolute loss with respect to the sequential fat-shattering dimension of the hypothesis class.
+Our work is more closely related to Attias23:regression, who introduce the online (resp.
 $\gamma$-Graph) dimension of a function class and show that it tightly characterizes online regression (resp.
 distributional learning) in the realizable setting up to a factor of 2.
-While our setting also considers utility functions with continuous image sets, the dimensions we introduce explicitly exploit the discrete structure of the Stackelberg game that arise from the finite set of follower types.
-Another recent line of work that captures incentives in online learning is online strategic classification.
-In this setting, agents can modify their features according to a directed graph over the feature space, before providing these features to the learner.
-The authors use a new combinatorial measure, the Strategic Littlestone Dimension, to achieve instance optimal mistake bounds.
-While seemingly similar, this measure of complexity is fundamentally distinct from the one we use to derive our results.
+While we also consider utility functions with continuous image sets, the dimensions we introduce explicitly exploit the discrete structure of the Stackelberg game that arises from the finite set of follower types.
+ahmadi2024strategic introduces a new combinatorial measure, the Strategic Littlestone Dimension, to achieve instance optimal mistake bounds in the online strategic classification setting.
+While both their measures of complexity and ours characterize learning in various principal-agent settings with commitment, the two are generally incomparable due to the differences in our settings.
 Learning piecewise Lipschitz functions.
-Given the piecewise linear structure of the leader’s utility function, our work also connects to the line of work on learning functions that are piecewise Lipschitz [8, 36, 10, 11, 9].
-We draw from this literature in part of our study that explores a simpler distributional setting without contexts.
-The well-structured payoff function of the leader against a fixed follower type allows us to adopt tools from Balcan et al.
-[9] and get better generalization guarantees compared to prior work that studied a similar setting [29].
-2 Problem Setting
-Notation.
-We use $[N]:=\{1,2,\ldots,N\}$ to denote the set of natural numbers up to and including $N$.
-We use $\Delta(\pazocal{A})$ to denote the probability simplex over a (finite) set $\pazocal{A}$.
-For a given string we use inclusive slice indexing such that $x[i:j]$ are the elements of string $x$ from position $i$ to $j$.
+Given the piecewise linear structure of the leader’s utility function, our work also connects to the line of work on learning functions that are piecewise Lipschitz [8555142, Sharma2019LearningPL, balcan2020semi, 10.5555/3540261.3541415, balcan2018general].
+Specifically, we draw from this literature in the part of our work that explores a simpler distributional setting without contexts (Section˜C.1).
+In particular, the well-structured payoff function of the leader against a fixed follower type allows us to adopt tools from balcan2018general and get better generalization guarantees compared to prior work that studies a similar setting [letchford2009learning].
+2 Setting and notation
+We use $N]:=\{1,2,\ldots,N\}$ to denote the set of natural numbers up to and including $N$ and $\Delta(\pazocal{A})$ to denote the probability simplex over a (finite) set $\pazocal{A}$.
+For strings, we use inclusive slice indexing such that $x[i:j]$ denotes the entries of $x$ from position $i$ to $j$.
 Dropping either $i$ or $j$ means that the slice extends from the beginning of the string or to the end, respectively.
-We use negative indexing to denote the elements of a string starting at the end i.e., $x[-1]$ is the last element of $x$.
+We also use negative indexing, so that, e.g., x[1]x[-1] is the last element of $x$.
+All proofs are deferred to the Appendix.
 We consider a Stackelberg game between a leader and a follower.
 Before the game starts, some contextual information $\mathbf{z}\in\pazocal{Z}\subseteq\mathbb{R}^{d}$ is revealed to both players.
-The leader plays first and commits to a (possibly mixed) strategy $\mathbf{x}\in\Delta(\pazocal{A})$, where $\pazocal{A}$ is the set of finite actions available to the leader.
+The leader plays first and commits to a mixed strategy $\mathbf{x}\in\Delta(\pazocal{A})$, where $\pazocal{A}$ is the set of finite actions available to the leader [^1] Our results hold for any strategy set that the leader can optimize over..
 The follower best-responds to the leader’s strategy and the context by playing some action $a_{f}\in\pazocal{A}_{f}$, where $\pazocal{A}_{f}$ is the finite set of follower actions.
-[^2] Without loss of generality the follower’s best-response will always be a pure strategy.
-We define the follower’s best-response as
-$b_{f}(\mathbf{z},\mathbf{x})\in\arg\max_{a_{f}\in\pazocal{A}_{f}}\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u_{f}(\mathbf{z},a_{l},a_{f}).$ (1)
-Here, $u_{f}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$ is the utility function of follower $f$.
-In case of ties, we assume that there is a fixed and known ordering over follower actions that determines the follower’s tie-breaking rule.
-We allow the follower to be one of $K\in\mathbb{N}_{\geq 1}$ follower types in the set $\{f^{(1)},\dots,f^{(K)}\}$.
+[^2] WLOG the follower’s best-response is a pure strategy.
+The follower’s best-response is defined as
+$b_{f}(\mathbf{z},\mathbf{x})\in\arg\max_{a_{f}\in\pazocal{A}_{f}}\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u_{f}(\mathbf{z},a_{l},a_{f}),$
+where, $u_{f}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$ is the follower’s utility function.
+In case of ties, we assume that the follower uses a fixed, known, and deterministic tie-breaking order over $\pazocal{A}_{f}$.
+We allow the follower to be one of $K$ follower types in the set $\{f^{(1)},\dots,f^{(K)}\}$.
 The type of follower characterizes their utility function, $u_{f^{(i)}}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$.
-We assume that the leader knows the set of possible follower types a priori, but not the realized type and aims to maximize her own utility, $u:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$.
-By slight abuse of notation, we sometimes use the shorthand $u(\mathbf{z},\mathbf{x},a_{f})=\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u(\mathbf{z},a_{l},a_{f}).$
+We assume that the leader knows the set of possible follower types a priori [^3] This assumption is for ease of exposition.
+Our results depend only on the leader knowing the best response function for each follower type., but not the realized type and aims to maximize her own utility, $u:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$.
+By slight abuse of notation, we sometimes use the shorthand $u(\mathbf{z},\mathbf{x},a_{f}):=\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u(\mathbf{z},a_{l},a_{f}).$
 We will denote a Stackelberg game as a tuple, $\pazocal{G}=\left(\pazocal{A},\pazocal{A}_{f},\pazocal{Z},u,u_{f^{(1)}},\ldots,u_{f^{(K)}}\right)$, consisting of the leader action set, the follower action set, the set of possible contexts, as well as the leader and followers’ utility functions.
 Remark 2.1.
-For the remainder of the paper and without loss of generality, we will consider $[K]$ to be a set of distinct follower types.
-We call any two follower types distinct, if the optimal leader’s strategy against them is not the same for every context.
-3 Online Learning in Structured Stackelberg Games
+For the remainder of the paper, and without loss of generality, we assume that the follower types in $[K]$ are distinct.
+We say that two follower types are distinct if there exists some context for which there is no single leader strategy is simultaneously optimal against both types.
+Let $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$ be a set of candidate functions (also called a hypothesis class) mapping contexts to follower types, which the leader uses to learn.
+Throughout, we focus on the realizable setting, i.e., we assume that $H$ is sufficiently rich to contain a perfect mapping $h^{*}\in\pazocal{H}$, where $h^{*}(\mathbf{z}_{t})=f_{t}$ for all $t\in[T]$.
+We use $h^{*}$ as a benchmark for the leader’s performance.
+For a single interaction with a follower at context $\mathbf{z}$, we write the leader’s loss when playing strategy $\hat{\mathbf{x}}\in\Delta(\pazocal{A})$ as
+[…]
+We let $\pi_{h^{*}}(\mathbf{z})\in\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h^{*}(\mathbf{z})}(\mathbf{z},\mathbf{x}))$ denote the optimal policy for the leader.
+For a hypothesis class $\pazocal{H}$, we denote its projection on a point as $\pazocal{H}(\mathbf{z})=\{i\in[K]:\exists h\in\pazocal{H}\;\text{s.t.}\;h(\mathbf{z})=i\}$ and on a set as $\pazocal{H}(\{\mathbf{z}_{i}\}_{i\in[n]})=\{(h(\mathbf{z}_{1}),\dots,h(\mathbf{z}_{n})):h\in\pazocal{H}\}$.
+For $i\in[K]$, we write $\pazocal{H}^{(\mathbf{z}\to i)}=\{h\in\pazocal{H}:h(\mathbf{z})=i\}$.
+3 Online learning
 In the online setting, the leader is faced with a sequence of followers.
 In each round $t\in[T]$, a context $\mathbf{z}_{t}\in\pazocal{Z}$ is revealed to both players.
-A follower of type $f_{t}$ arrives, the leader commits to a mixed strategy $\mathbf{x}_{t}\in\Delta(\pazocal{A})$, and the follower best-responds according to Equation˜1.
-We assume that the leader knows the utilities of all follower types, but the type of follower at time $t$ is not revealed to the leader until after the round is over.
-We make no assumptions about how the sequence of contexts is chosen, i.e., we allow them to be chosen adversarially and adaptively to the leader’s sequence of play.
-3.1 Leader’s objective
-Let $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$ be a set of candidate functions (also called a hypothesis class) mapping contexts to follower types, which the leader uses to learn.
-We consider the realizable setting, where there exists a true mapping $h^{*}\in\pazocal{H}$ such that $h^{*}(\mathbf{z}_{t})=f_{t}$ for all possible $(\mathbf{z}_{t},f_{t})$ pairs.
-We use $h^{*}$ as a benchmark for the leader’s performance over time, which we quantify via the notion of contextual Stackelberg regret.
+A follower of type $f_{t}$ arrives, the leader commits to a mixed strategy $\mathbf{x}_{t}\in\Delta(\pazocal{A})$, and the follower best-responds.
+We assume that the leader knows the utilities of all follower types, but the type of the follower at time $t$ is not revealed to the leader until after the round is over.
+We make no assumptions about how the sequence of contexts is chosen, i.e., they may be chosen adversarially.
+Next we present the notion of contextual Stackelberg regret, which quantifies the leader’s performance over time.
 Definition 3.1 (Contextual Stackelberg Regret).
-Given a sequence of pairs of contexts and followers $\{(\mathbf{z}_{i},f_{i})\}_{t=1}^{T}$ the leader’s regret is
-$R(T):=\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t})),$
-where $\pi_{h^{*}}(\mathbf{z})=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h^{*}(\mathbf{z})}(\mathbf{z},\mathbf{x}))$ and $\mathbf{x}_{t}\in\Delta(\pazocal{A})$ is the leader’s strategy at timestep $t$.
-The goal of the leader is to minimize her regret, that is, to achieve utility comparable to the utility she would have gotten if she could (1) optimally predict the type of follower given the context and (2) play an optimal strategy against that type of follower.
-In Section˜3 we present an algorithm that is optimal with respect to this benchmark.
+Given a sequence of pairs of contexts and followers $\{(\mathbf{z}_{t},f_{t})\}_{t=1}^{T}$ the leader’s regret is $R(T):=\sum_{t=1}^{T}r(\mathbf{z}_{t},\mathbf{x}_{t},f_{t})$ where $\mathbf{x}_{t}\in\Delta(\pazocal{A})$ is the leader’s strategy at timestep $t$.
+The goal of the leader is to minimize her regret, that is, to achieve utility comparable to the utility she would have obtained if she could correctly predict the follower’s type given the context and play optimally against it.
 Remark 3.2.
-An important part of our analysis relies on the leader being able to compute the optimal strategy for any fixed context and follower type.
-However, due to the tie-breaking rules of the follower, the leader might not be able to get an exact solution.
-Previous work  [7, 27] shows that the leader can get an arbitrarily small error in the utility by optimizing over an appropriately selected subset of the simplex.
-For ease of notation, in the remainder of this paper, we will assume that the leader can directly optimize over the entire simplex, without that affecting her regret guarantees.
-For completeness, we formally state this result in Section˜A.1.
+Our analysis relies on the leader being able to compute the optimal strategy to play for any fixed context and follower type.
+However, due to the tie-breaking rules of the follower, the leader may not be able to get an exact solution.
+Previous work [balcan2015commitment, harrisregret] shows that the leader can get an arbitrarily small error in the utility by optimizing over an appropriately selected subset of the simplex.
+For ease of exposition, we assume that the leader can directly optimize over the entire simplex, without affecting her regret guarantees.
+A formal statement of this result appears in Section˜B.1.
 3.1 Littlestone dimension
-In this section, we highlight the concepts and definitions from multiclass online learning [22] that are relevant to our work.
-We start with the notion of shattering, which captures the following intuition: the richer a set of functions is, the more complex patterns it should be able to fit.
-Definition 3.3 (Multiclass $\pazocal{H}$-Shattered Littlestone Tree).
+Given the structure induced by the finite set of follower types, a natural idea is to directly predict the follower’s type from the context, and play the corresponding optimal strategy.
+This suggests using tools from online (multiclass) classification to characterize the statistical complexity of the problem.
+In this section, we will show that this approach comes up short.
+We begin by recalling several concepts from online learning that are relevant to our work, starting with the notion of shattering, which captures the intuition that richer hypotheses classes can encode more complex label patterns.
+Definition 3.3 (Multiclass Littlestone Tree).
+A Littlestone tree is a rooted tree, in which each internal node is labeled with an instance in $\pazocal{Z}$ and each edge corresponds to a label in a finite set $[K]$.
+We say that the tree is shattered by a hypothesis class $\pazocal{H}$ if for every root-to-leaf path that traverses nodes $\mathbf{z}_{1},\dots,\mathbf{z}_{n}$, there exists a hypothesis $h\in\pazocal{H}$ such that the label of the edge in the path leaving $\mathbf{z}_{i}$ is $h(\mathbf{z}_{i})$ for all $i$.
+Example 3.4.
+Consider the class of permutations
+$\mathcal{H}_{3}=\left\{\,h:\{1,2,3\}\to\{f^{(1)},f^{(2)},f^{(3)}\}\ \middle|\ h\text{ is a bijection}\,\right\}.$
+Figure˜1 shows an example of a Littlestone Tree that is shattered by this class.
+Figure 1: $\mathcal{H}_{3}$-shattered Littlestone Tree showing a consistent hypothesis for each root-to-leaf path.
+The multiclass Littlestone dimension reflects how the ability of a function class to shatter deeper trees corresponds to greater complexity in the class.
+[^4] For binary labels and when all nodes at the same depth of a shattered tree are the same, the Littlestone dimension recovers the definition of the well-studied VC-dimension.
+This definition is due to pmlr-v19-daniely11a and it generalizes the classic definition of the Littlestone dimension for binary classification [littlestone1988learning].
+Definition 3.5 (Multiclass Littlestone Dimension).
+The Littlestone dimension of a multiclass hypothesis class $\pazocal{H}$, denoted $\textnormal{Ldim}(\pazocal{H})$, is the maximal integer $d$ such that there exists a full binary tree of depth $d$ that is shattered by $\pazocal{H}$.
+The finiteness of the Littlestone dimension of a hypothesis class is necessary for a hypothesis class to be online learnable [littlestone1988learning].
+Some examples of classes with finite Littlestone dimension are conjunctions, decision lists, linear separators with a margin, and in general anything learnable in the mistake bound model (cf.
+balcannotes:mistakebound, Proposition 17 in Alon21:partial).
+pmlr-v19-daniely11a show that the multiclass Littlestone dimension of a hypothesis class characterizes learnability in online multiclass classification, that is, it determines the number of mistakes an optimal online learner will make in the worst-case.
+In our setting, learning how to optimally predict the follower’s type given a context is an instance of multiclass prediction and is indeed sufficient for learning an optimal policy.
+Namely, given the context and the predicted follower type, the leader can compute the strategy that maximizes her utility by solving a sequence of $|\pazocal{A}_{f}|$ linear programs [conitzer2006computing].
+However, it is a priori unclear whether learning to optimally predict the follower’s type is not only sufficient but also necessary.
+We now show that this is not the case, by constructing a structured Stackelberg game instance with an infinite Littlestone dimension that can be solved without learning.
+Theorem 3.6.
+There exists a Stackelberg game $\pazocal{G}$ and a hypothesis class $\pazocal{H}$ such that $\textnormal{Ldim}(\pazocal{H})=\infty$, but the leader’s optimal policy can be determined without learning.
+Proof sketch.
+The key idea is that, although the follower types can be distinct from the leader’s perspective (Remark˜2.1) and the hypothesis class can encode arbitrarily complex type-prediction problems, the leader’s downstream utility optimization problem may be much simpler.
+Our construction uses two follower types and a hypothesis class, $\pazocal{H}$, consisting of linear threshold functions.
+Learning the true threshold can be hard ($\textnormal{Ldim}(\pazocal{H})=\infty$), but is unnecessary: near the threshold the two follower types induce the same optimal strategy for the leader.
+Figure˜2 provides an illustration of the Stackelberg game we use in the proof.
+∎
+Figure 2: Follower types as a function of the context in the construction of Theorem˜3.6.
+While $\gamma_{1}$ and $\gamma_{2}$ vary with each hypothesis, the leader’s optimal strategy only depends on whether $\mathbf{z}[2]<0.5$.
+This result shows that a finite Littlestone dimension is not necessary for online learning in our setting, as it only measures the complexity of predicting the follower types, not how those predictions affect the leader’s utilities.
+3.2 Stackelberg Littlestone dimension
+We now define the Stackeleberg Littlestone dimension, which we prove characterizes online learnability in structured Stackeleberg games.
+In the definitions that follow we consider trees with nodes labeled by elements in $\pazocal{Z}$ and edges corresponding to different labels in $[K]$.
+A depth-$d$ tree, $\mathcal{T}_{d}=\{\mathbf{z}_{s}:s\in S_{d}\}$, is indexed by sequences $S_{d}\subseteq\{\varnothing\}\cup\{s\in[K]^{\ell}:1\leq\ell\leq d\}$, where $\mathbf{z}_{\varnothing}$ denotes the root node.
+Definition 3.7 (Stackelberg Littlestone (SL) Tree).
+A SL tree, $\pazocal{T}_{d}$, is a rooted tree, in which each internal node is labeled with an instance in $\pazocal{Z}$ and each edge corresponds to a label in $[K]$.
+Each node $\mathbf{z}_{s}\in\mathcal{T}_{d}$ has a weight $\rho_{s}$, where $\rho_{s}=0$ if $\mathbf{z}_{s}$ is a leaf node and $\rho_{s}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in[K]:sj\in S_{d}}\left(r(\mathbf{z}_{s},\mathbf{x},f^{(j)}))+\rho_{sj}\right)$ otherwise.
+We say that the tree is shattered by a hypothesis class $\pazocal{H}$ under Stackelberg game $\pazocal{G}$ if for every root-to-leaf path that traverses nodes $\mathbf{z}_{\varnothing},\dots\mathbf{z}_{s\leq d}$, there exists $h\in\pazocal{H}$ such that the label of the edge in the path leaving $\mathbf{z}_{s_{\leq i}}$ is $h(\mathbf{z}_{s_{\leq i}})$ for every $i$.
+Example 3.8.
+Consider the hypothesis class of permutations
+$\mathcal{H}_{3}=\left\{\,h:\{1,2,3\}\to\{f^{(1)},f^{(2)},f^{(3)}\}\ \middle|\ h\text{ is a bijection}\,\right\}$
+and a Stackelberg game with 3 actions for both players.
+The leader’s utility matrix is $U_{L}=\mathbb{I}_{3\times 3}$ and follower $f^{(i)}$’s utility matrix is $U_{f^{(i)}}=\mathbf{1}e_{i}^{\top}$, i.e.
+she receives utility 1 if and only if she plays action iiFigure˜3 shows a corresponding SL tree.
+Figure 3: $\mathcal{H}_{3}$-shattered SL-Tree showing the consistent hypothesis for each root-to-leaf path.
+Leaf nodes have weight 0, nodes at depth 11 have weight $1/2$ and the root node has weight $1/2+2/3$.
+Definition 3.9 (Stackelberg Littlestone (SL) dimension).
+The SL dimension of a hypothesis class $\mathcal{H}$ for Stackelberg game $\pazocal{G}$ is defined as $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H}):=\sup\{\kappa\geq 0:$ there exists an SL tree shattered by $\pazocal{H}$ under $\pazocal{G}$ with root node weight $\kappa\}$.
+Note that taking the supremum is necessary since it is possible that no finite-depth tree achieves the optimal value.
+If for every $\kappa\geq 0$ there exists an SL tree with root node weight at least $\kappa$, then $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=\infty$.
+We now show that no deterministic algorithm can achieve regret lower than the SL dimension.
+Theorem 3.10.
+For any hypothesis class $\pazocal{H}$ and Stackelberg game $\pazocal{G}$, there exists an adversarial sequence of contexts and follower types such that any deterministic algorithm suffers at least $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})-\epsilon$ regret, for any $\epsilon>0$.
+Proof sketch.
+The key idea in the proof is that the weight assigned to each node of an SL tree measures how much regret the adversary can still guarantee from that point onward.
+At the root, this weight captures the minimum regret the learner must suffer no matter which strategy she plays.
+If the learner chooses the strategy that minimizes immediate regret, the adversary selects the follower type that leads into a subtree with large remaining weight, ensuring future regret.
+If the learner deviates to reduce future regret, the adversary instead extracts more regret immediately.
+By repeating this reasoning along a root-to-leaf path, the adversary can force the learner to incur regret equal to the root weight of the tree.
+Since the SL dimension is defined as the supremum root weight over all such shattered trees, no deterministic algorithm can achieve regret smaller than the SL dimension (up to an arbitrarily small additive term).
+∎
+We now present an instance-optimal algorithm that achieves regret equal to the SL dimension.
+We call Algorithm˜1 the Stackelberg Standard Optimal Algorithm (SSOA), as it shares some similarities with the Standard Optimal Algorithm (SOA) from online multiclass prediction [daniely2014:optimal].
+In each round, SSOA tracks the subset of hypotheses that are consistent with the history so far.
+The algorithm then plays the mixed strategy that minimizes the worst-case instantaneous regret, plus the SL dimension of the induced hypothesis class.
+The latter captures the difficulty of the learning task that remains after the current round ends.
+1:Input: Stackelberg game $\pazocal{G}$, hypothesis class $\pazocal{H}$
+2: Initialize version space $V_{1}=\pazocal{H}$
+3:for $t\in[T]$ do
+4:  Observe context $\mathbf{z}_{t}$
+5:  Compute optimal utility for each type $i\in V_{t}(\mathbf{z}_{t})$ $u_{*}^{(i)}=\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{f^{(i)}}(\mathbf{z}_{t},\mathbf{x}))\;$
 […]
-The notion of multiclass Littlestone dimension reflects how the ability of a function class to shatter deeper trees corresponds to greater complexity in the class.
-[^3] We note that for binary labels, the Littlestone dimension recovers the definition of the well-studied VC-dimension in the case where all nodes at the same depth of a shattered tree are the same.
-This definition is due to Daniely et al.
-[22] and it generalizes the classic definition of the Littlestone dimension for binary classification [30] to the multiclass setting.
-Definition 3.4 (Multiclass Littlestone Dimension).
-The Littlestone dimension of a multiclass hypothesis class $\pazocal{H}$, denoted as $\textnormal{Ldim}(\pazocal{H})$, is the maximal integer $d$, such that there exists a full binary tree of depth $d$ that is shattered by $\pazocal{H}$.
-Examples.
-The finiteness of the Littlestone dimension of a hypothesis class is necessary for the class to be online learnable [30].
-There are several examples of hypothesis classes that have finite Littlestone dimension and can, therefore, be (online) learned.
-These include the class of conjunctions, the class of decision lists, the class of linear separators with a margin (cf.
-Proposition 17 in [1]), and in general anything learnable in the mistake bound model.
-In Daniely et al.
-[22], the authors show that the multiclass Littlestone dimension of a hypothesis class characterizes the learnability of the class.
-In our setting, learning how to optimally predict the follower’s type given a context is an instance of multiclass prediction and is indeed sufficient for learning an optimal policy.
-Namely, given the context and the predicted follower type, the leader can deterministically find a strategy that maximizes her utility, by solving a sequence of $|\pazocal{A}_{f}|$ linear programs [20].
-However, it is a priori unclear whether learning to optimally predict the follower’s type is also necessary.
-In  Theorem˜3.5 we show that the Littlestone dimension does not characterize the complexity of online learning in Stackelberg games with side information.
-The full proof can be found in Section˜A.2.
+7:  Observe follower type $f_{t}$
+8:  Update version space $V_{t+1}=V_{t}^{(\mathbf{z}_{t}\to f_{t})}$
+9:end for
+ALGORITHM 1 SSOA
 Theorem 3.11.
-There exists a Stackelberg game, $\pazocal{G}$, and a hypothesis class, $\pazocal{H}$, such that $\textnormal{Ldim}(\pazocal{H})=\infty$, but the leader’s optimal policy can be determined without learning.
+For any Stackelberg game $\pazocal{G}$ and every hypothesis class $\pazocal{H}$, the regret of Algorithm˜1 is at most $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
 Proof sketch.
-Our proof captures the following intuition: in a structured Stackelberg game, the follower types can be completely distinct from the perspective of the leader (i.e., any one irreducible to another), and the hypothesis class can be arbitrarily hard to learn.
-Yet, to truly know how difficult the leader’s learning task is, one has to account for how the parameters of the game interact with the hypothesis class.
-Our construction uses distinct follower types and the hypothesis class indicates that there is a true threshold that splits the context domain in two parts, each of which is predictive of a different follower type.
-Learning that true threshold can be arbitrarily hard ($\textnormal{Ldim}(\pazocal{H})=\infty$), yet unnecessary if the two follower types happen to induce the same optimal strategy for the leader near the threshold.
-This result highlights that the Littlestone dimension is completely blind to how the hypothesis class interacts with the utilities of the players, which motivates a new measure of complexity.
-3.3 Stackelberg-Littlestone Dimension
-In this section we provide an instance optimal learner for structured Stackelberg games. We start with the definition of node-weighted shattered trees.
+Our proof follows a high-level idea similar to the analysis of SOA in online prediction: With each mistake, the learner’s task becomes easier, as the Littlestone dimension decreases.
+Similarly, we show that when the learner incurs instantaneous regret, the SL dimension of the induced hypothesis class decreases by at least this amount.∎
+3.3 Relationship with the Littlestone dimension
+We have so far shown that the SL dimension characterizes learnability in structured Stackelberg games. We now examine how the it relates to online multiclass prediction.
+Lemma 3.12.
+For any hypothesis class $\pazocal{H}$ and any Stackelberg game $\pazocal{G}$, $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})\leq\textnormal{Ldim}(\pazocal{H})$.
+To make the relationship between the two dimensions more concrete, consider the following example, which is a generalization of Examples 3.4 and 3.8 to $n$ dimensions.
+Example 3.13.
+Consider a setting with $n$ contexts, $n$ follower types, and the class of permutations $\pazocal{H}_{n}$.
+Note that $\textnormal{Ldim}(\pazocal{H}_{n})=n-1$, since an adversary can force a mistake in every round that she can query.
+A tree of depth $n$ cannot be shattered because once the labels of the first n1n-1 contexts have been observed, the label of the last context is fixed.
+Now consider a Stackelberg game where the leader’s utility matrix is $U=\mathbf{I}_{n\times n}$ and follower type $i$’s utility matrix is $U_{i}=\mathbf{1}_{n}e_{i}^{\top}$.
+We will show that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H}_{n})=n-H_{n}$, where $H_{n}$ is the $n$-th harmonic number.
+Consider the SL tree that witnesses the SL dimension.
+Due to symmetry, each node at depth $d$ will have ndn-d outgoing edges, one for each follower type not encountered in the path so far.
+Each node at depth n2n-2 will have weight $1/2$, since the optimal utility of the leader against any follower type is $1$, the optimal utility against any follower chosen adversarially from a set of two types is $1/2$, and any deeper nodes must have only one outgoing edge (and therefore weight 0).
+More generally, for a node $s$ at depth $d$ we have
+$\rho_{s}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{j_{1},\dots,j_{n-d}\}}\biggl[r(\mathbf{z},\mathbf{x},f^{(j)})+\sum_{j=2}^{n-d-1}1-\frac{1}{j}\biggr],$
+which simplifies to $n-d-\sum_{j=1}^{n-d}\frac{1}{j}$.
+Note that while any deterministic learner will make at least $\textnormal{Ldim}(\pazocal{H})$ mispredictions when learning $\pazocal{H}$, mispredicting the follower’s type need not lead to suboptimal utility for the leader.
+Therefore, Lemma˜3.12 does not immediately imply that the optimal multiclass prediction algorithm (SOA) would incur suboptimal regret in our setting.
+Indeed, in Theorem˜3.6 we constructed a game instance where the leader can mispredict the follower’s type at every timestep, yet still achieve optimal utility.
+We conclude this section by showing that there are indeed game instances where SOA will suffer larger regret than SSOA.
+Theorem 3.14.
+There exists a hypothesis class $\pazocal{H}$ and Stackelberg game $\pazocal{G}$ such that running SOA on $\pazocal{H}$ results in suboptimal utility for the leader.
+Proof sketch.
+The game construction exploits the mismatch between the label prediction objective of SOA and optimizing for leader utility.
+Specifically, the game is designed such that SOA focuses on parts of the hypothesis class that are combinatorially complex (large Littlestone dimension) but strategically irrelevant.
+As a result, SOA unnecessarily suffers utility loss over multiple rounds.
+∎
+4 Distributional learning
+We now shift our focus from online learning to distributional learning.
+In this setting, contexts are drawn i.i.d.
+from an unknown distribution and as in the online setting, we assume that there exists a function $h^{*}\in\pazocal{H}$ that correctly predicts each follower type given the context.
+The leader’s goal is to choose a strategy that will maximize her expected utility on a fresh sample drawn from the same distribution.
+A learning algorithm is a mapping $A:(\pazocal{Z}\times[K])^{*}\to\Delta(\pazocal{A}){Z}$.
+We consider the following notion of sample complexity, adapted to our Stackelberg setting.
+Definition 4.1 (PAC cut-off sample complexity).
+The PAC cut-off sample complexity $m(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$ of function class $\pazocal{H}$ w.r.t.
+Stackelberg game $\pazocal{G}$ is $m(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)=\inf_{A}m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$, where $m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$ is the smallest integer such that for any $m\geq m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$, every distribution $\mathcal{D}$ on $\pazocal{Z}$, and any target hypothesis $h^{*}\in\pazocal{H}$, the expected cut-off loss of algorithm $A$, $\mathbb{P}_{\mathbf{z}\sim\mathcal{D}}[r(\mathbf{z},A(\mathbf{z};S),f^{(h^{*}(\mathbf{z}))})>\gamma]$, is at most $\epsilon$ with probability at least $1-\delta$.
+While we present our results with respect to the PAC cut-off sample complexity, any bound with respect to this benchmark immediately implies a bound on the familiar PAC sample complexity [valiant1984theory] as well.
+[^5] See Lemma 1 in Attias23:regression.
+4.1 Sample complexity lower bound
+We now introduce the $\gamma$-valued Stackelberg-Natarajan dimension (henceforth SN dimension), which we show characterizes the sample complexity lower bound for distributional learning in structured Stackelberg games.
+Indeed, we prove that the finiteness of this measure is a necessary condition for learnability in our setting.
+The SN dimension is inspired by the Natarajan dimension from multiclass prediction [natarajan1989learning, daniely2014:optimal] and the $\gamma$-Natarajan dimension from realizable regression [Attias23:regression].
+We start by defining what it means for a set of contexts to be $\gamma$-SN-shattered.
+Intuitively, a set is $\gamma$-SN-shattered if the hypothesis class $\pazocal{H}$ can fit complex patterns of follower types on the set.
+Furthermore, these follower types must be sufficiently different: no leader strategy can incur loss less than $\gamma$ if the type is chosen adversarially among them.
+Definition 4.2 ($\gamma$-SN-shattered set).
+A set $\{\mathbf{z}_{1},\dots,\mathbf{z}_{n}\}\subseteq\pazocal{Z}$ is $\gamma$-SN-shattered by hypothesis class $\pazocal{H}$ with respect to Stackelberg game $\pazocal{G}$ if there exist functions $g_{0},g_{1}:\pazocal{Z}\to[K]$ such that
+  1. 1.
+    $\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{g_{0}(\mathbf{z}_{i}),g_{1}(\mathbf{z}_{i})\}}r(\mathbf{z}_{i},\mathbf{x},f^{(j)})>\gamma$ for every $\mathbf{z}_{i}\in S$, and
+  2. 2.
+    for all $b\in\{0,1\}^{n}$, there exists an $h_{b}\in\pazocal{H}$ such that
+    $h_{b}(\mathbf{z}_{i})=\begin{cases}g_{0}(\mathbf{z}_{i})\;\;\text{if}\;\;b_{i}=0\\ g_{1}(\mathbf{z}_{i})\;\;\text{if}\;\;b_{i}=1.\end{cases}$
+Definition 4.3 ($\gamma$-valued SN dimension).
+The $\gamma$-valued SN dimension, $\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$, is the cardinality of the largest set that is $\gamma$-SN-shattered by $\pazocal{H}$ with respect to $\pazocal{G}$.
+We emphasize that that the two functions, $g_{0}$ and $g_{1}$ in Definition˜4.2 must be $\gamma$-far from each other in a minmax sense: at each context the leader cannot choose a mixed strategy that performs well against both $g_{0}(\mathbf{z}_{0})$ and $g_{1}(\mathbf{z}_{0})$.
+This enables a distribution-independent lower bound that holds for every learning algorithm.
+We prove our lower bound in Theorem˜4.4 by adapting the argument of Attias23:regression to our setting.
+Theorem 4.4.
+Let $A$ be any learning algorithm and $\epsilon,\delta,\gamma\in(0,1)$ such that $\delta<\epsilon$, then
+$m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)\geq\Omega\left(\frac{\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})+\log(1/\delta)}{\epsilon}\right).$
+Proof sketch.
+The high-level idea is to reduce distributional learning in structured Stackelberg games to a hard identification problem over a $\gamma$-SN-shattered set.
+Such a set contains many contexts where the hypothesis class can realize two different follower types such that no leader strategy can simultaneously perform well against both.
+The distribution we consider on this shattered set has most probability mass on one “easy” context and small but non-negligible mass spread over many “hard” contexts.
+With limited samples, the learner is unlikely to observe enough of the hard contexts to correctly distinguish the follower type.
+Since the hypothesis class can label these unseen contexts arbitrarily while remaining consistent with the data, the learner’s strategy on them is effectively a guess.
+Because each unseen context independently forces a $\gamma$-loss with constant probability, the learner’s expected cut-off error remains large unless the number of samples scales linearly with the size of the $\gamma$-SN-shattered set.
+Standard concentration arguments then show that achieving small error with high probability requires a sample size proportional to the $\gamma$-SN dimension.
+∎
+4.2 Sample complexity upper bound
+We conclude our treatment of the distributional setting by introducing the $\gamma$-valued Stackelberg-Graph dimension (henceforth SG dimension), which provides a sufficient condition for distributional learning in structured Stackelberg games.
+The SG dimension is a natural relaxation of the $\gamma$-Natarajan dimension and an adaptation of similar dimensions from classification [natarajan1989learning, daniely2014:optimal] and regression [Attias23:regression] to our setting.
+Definition 4.5 ($\gamma$-SG-shattered set).
+A set $\{\mathbf{z}_{1},\dots,\mathbf{z}_{n}\}\subseteq\pazocal{Z}$ is $\gamma$-SG-shattered by hypothesis class $\pazocal{H}$ with respect to Stackelberg game $\pazocal{G}$ if there exists a function $g:\pazocal{Z}\to[K]$ such that for every $b\in\{0,1\}^{n}$, there exists an $h_{b}\in\pazocal{H}$ satisfying
+  1. 1.
+    $h_{b}(\mathbf{z}_{i})=g(\mathbf{z}_{i})\;\;\forall i\in[n]$ such that $b_{i}=0$, and
+  2. 2.
+    $\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{g(\mathbf{z}_{i}),h_{b}(\mathbf{z}_{i})\}}r(\mathbf{z}_{i},\mathbf{x},f^{(j)})\geq\gamma$ for all $i\in[n]$ such that $b_{i}=1$.
+Definition 4.6 ($\gamma$-valued SG dimension).
+The $\gamma$-valued SG dimension, $\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$, is the cardinality of the largest set that is $\gamma$-SG-shattered by $\pazocal{H}$ with respect to $\pazocal{G}$.
+We now present the main result of this section.
+Given a training sample $S$ of context–follower-type pairs, Algorithm˜2 first computes the subset of hypotheses that are perfectly consistent with $S$.
+Then, on a new context, the algorithm plays the leader’s strategy that is optimal against an adversarially chosen follower type predicted at that context by any hypothesis consistent with $S$.
+1:Input: Stackelberg game $\pazocal{G}$, hypothesis class $\pazocal{H}$, sample $S=\{(\mathbf{z}_{i},f_{i})\}_{i\in[n]}$, and test context $\mathbf{z}\in\pazocal{Z}$
+2: Compute $\pazocal{H}|_{S}=\{h\in\pazocal{H}:h(\mathbf{z}_{i})=f_{i}\;\forall i\in[n]\}$
+3: Compute $F=\pazocal{H}|_{S}(\mathbf{z})=\{h(\mathbf{z})\in[K]:h\in\pazocal{H}|_{S}\}$
+4: Play $\mathbf{x}^{*}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in F}r(\mathbf{z},\mathbf{x},f^{(i)})$
+ALGORITHM 2 Stackelberg Distributional Learner ($\mathfrak{L}^{*}$)
+We now show that the $\gamma$-SG dimension upper bounds the PAC sample complexity in our setting. The proof of Theorem˜4.7 follows from adapting that of Attias23:regression to our setting.
+Theorem 4.7.
+For any class $\pazocal{H}$, game, $\pazocal{G}$, and $\epsilon,\delta,\gamma\in(0,1)$, there is some constant $C_{1}$ such that Algorithm˜2 achieves
+$m_{\mathfrak{L}^{*}}^{r}(\pazocal{H},\epsilon,\delta,\gamma)\leq C_{1}\frac{\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})\log(1/\epsilon)+\log(1/\delta)}{\epsilon}.$
+Proof sketch.
+The key observation in our analysis is that the leader only needs to worry about contexts where different hypotheses would force meaningfully different leader utilities.
+Algorithm˜2 keeps all hypotheses that are consistent with the data and, given a new context, plays the strategy that is most robust to all follower types predicted by those hypotheses.
+If the learner performs poorly on a noticeable fraction of contexts, then there must exist many contexts where some hypothesis in the class disagrees with the true one and where this disagreement induces loss at least $\gamma$.
+Such contexts form a $\gamma$-SG-shattered set.
+The definition of the $\gamma$-SG dimension limits the size of these hard sets.
+Using standard uniform convergence and symmetrization arguments, we show that once the sample size exceeds this dimension, no hypothesis consistent with the data can be $\gamma$-far from the true one on more than an $\epsilon$ fraction of contexts.
+As a result, the chosen strategy achieves near-optimal expected utility with high probability.∎
+5 Conclusion and future directions
+We initiate the study of structured Stackelberg games, where contextual information is predictive of the follower’s type.
+Our main results consist of a complete characterization of instance-optimal regret bounds in the online setting, and PAC sample complexity guarantees in the distributional setting.
+There are several exciting directions for future work.
+Our main focus in this work is on the statistical complexity of learning in structured Stackelberg games.
+An interesting avenue for future research is to consider the computational complexity of learning in this setting.
+We conduct a preliminary exploration of this direction in in Appendix˜D, and find that our results open the door for efficient (albeit so far sub-optimal) learning algorithms in some settings.
+We focus on the realizable setting, where the leader is given a class of hypotheses with the promise that it includes the optimal predictor.
+The agnostic learning setting is a natural relaxation, where this assumption is not made and the goal is instead to do as well as the best hypothesis in the class.
+As such, providing a complete characterization for the agnostic case remains an important future direction.
+6 Acknowledgments
+K.F. is grateful to Brian Hu Zhang for helpful discussions on an earlier version of this paper.
+Appendix A Applications beyond Stackelberg games
+Our results also apply to the problems of learning to bid in simultaneous second-price auctions with side information and Bayesian persuasion with both public and private states.
+These settings are formally introduced in balcan2025nearlyoptimalbanditlearningstackelberg; we give an overview here.
+They are mathematically equivalent to the setting we study in the main body, with the exception of the form that the learner’s strategy space takes (detailed below).
+A.1 Simultaneous second-price auctions with side information
+This application is an extension of a setting in daskalakis2016learning to incorporate side information.
+In each round (or historical data point in the distributional setting), bidders simultaneously bid on a set of $m$ items.
+They receive the bundle of items for which they are the highest bidder, and they pay a price corresponding to the second-highest bid for each item.
+The learner in this setting is a bidder whose strategy space consists of a set of possible bids for each item.
+The bidder’s valuation depends on the entire bundle of items they receive, e.g.
+the bidder may have a low valuation for item A and item B in silos, but a high valuation for the combination of the two.
+In addition to the bundle of items, the bidder’s valuation also depends on contextual information, e.g.
+different seasons or fashion trends may influence a bidder’s valuation for a bundle of clothing items.
+Other players’ bids roughly correspond to different follower types, and we posit a hypothesis class mapping contextual information to different sets of other players’ bids.
+Our results carry over to this application as-is when the space of the other players’ bids is discrete; the only difference is that the leader’s strategy space is now the bid space $0,1]^{m}$ instead of a probability simplex.
+A.2 Bayesian persuasion with both public and private states
+In Bayesian persuasion [kamenica2011bayesian], a sender (analogous to the leader in Stackelberg games) strategically reveals payoff-relevant information about an underlying private state of the world to a receiver (analogous to the follower), who then takes a payoff-relevant action.
+This sub-section considers a generalization of Bayesian persuasion where there is some additional payoff-relevant information about the state which is common knowledge to both players.
+Receiver types map to follower types, and we posit a relationship between the common knowledge and the receiver’s type.
+As is the case in the previous sub-section, our results carry over here as-is, with the sender optimizing over a convex polytope strategy space instead of a probability simplex.
+Appendix B Details and proofs from Section˜3
+See 3.6
+Proof.
+We construct such a game explicitly.
+Let the context space be $\pazocal{Z}=\{0,1\}\times[0,1]$ and let the set of follower types be $\{f^{(1)},f^{(2)},f^{(3)},f^{(4)}\}$.
+Consider the follower types $f^{(1)}$ and $f^{(2)}$ with utility functions $u_{f^{(1)}},u_{f^{(2)}}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$ chosen such that for every context $\mathbf{z}\in\pazocal{Z}$, the leader’s optimal action differs, i.e.
+$\pi_{*}^{(1)}(\mathbf{z})\neq\pi_{*}^{(2)}(\mathbf{z})$, where $\pi_{*}^{(i)}$, refers to the optimal policy against follower $f^{(i)}$.
+Concrete payoff matrices for all follower types are shown in Table˜1 (the leader is the row player and the follower is the column player).
+$F_{1}$ $F_{2}$
+$L_{1}$ $1,0.5$ $0.5,0$
+$L_{2}$ $0.25,1$ $0.75,0$
+(a) Follower type $f^{(1)}$
+$F_{1}$ $F_{2}$
+$L_{1}$ $1,0$ $0.5,0.5$
+$L_{2}$ $0.25,0$ $0.75,1$
+(b) Follower type $f^{(2)}$
+$F_{1}$ $F_{2}$
+$L_{1}$ $1,1-\mathbf{z}[1]$ $0.5,0.5$
+$L_{2}$ $0.25,0.5$ $0.75,\mathbf{z}[1]$
+(c) Follower type $f^{(3)}$
+$F_{1}$ $F_{2}$
+$L_{1}$ $1,\mathbf{z}[1]$ $0.5,0.5$
+$L_{2}$ $0.25,0.5$ $0.75,1-\mathbf{z}[1]$
+(d) Follower type $f^{(4)}$
+Table 1: Utility tables for Theorem˜3.6
+We now construct the hypothesis class $\pazocal{H}=\{h_{\gamma_{1},\gamma_{2}}:\forall\;\;0<\gamma_{1}<0.5,\;0.5\leq\gamma_{2}<1\}$, where
 […]
-Definition 3.6 (Shattered Stackelberg-Littlestone Tree).
-[…]
-  • (Consistency) $\pazocal{H}$ is consistent with every path in $\pazocal{T}_{d}$ i.e., for any $s\in S_{d}$, there exists $h\in\pazocal{H}$ such that for all $i\in\{\varnothing,1,\dots,d-1\}$ the label of edge $(\mathbf{z}_{s[:i]},\mathbf{z}_{s[:i+1]})$ is $h(\mathbf{z}_{s[:i]})$.
-  • (Max-branch property) $\pazocal{T}_{d}$ satisfies the max-branch property, i.e., every node, $\mathbf{z}_{s}$, has children $\mathbf{z}_{si}$ for all $i\in[K]$ where $\pazocal{H}$ is consistent with path $si$.
-  • (Weights) each node, $\mathbf{z}_{s}\in\mathscr{T}_{d}$ has a weight, $\rho_{s}$ defined as follows: if $\text{len}(s)=d+1$ then $\rho_{s}=0$, else
-    $\rho_{s}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{sj\in S_{d}}\biggl{[}r(\mathbf{z},\mathbf{x},f^{(j)})+\rho_{sj}\biggr{]},$
-    where $r(\mathbf{z},\mathbf{x},j):=u(\mathbf{z},\pi_{h^{*}}(\mathbf{z}),b_{f^{(j)}}(\mathbf{z},\pi_{h^{*}}(\mathbf{z})))-u(\mathbf{z},\mathbf{x},b_{j}(\mathbf{z},\mathbf{x}))$ is the leader’s instantaneous regret, given context $\mathbf{z}$, mixed strategy $\mathbf{x}$, and follower type $j$ [^4] The infimum might not be achieved due to the followers’ tie breaking rule.
-We refer the reader to Section A.1 for further details.
-Definition 3.7 (Stackelberg-Littlestone dimension).
-The Stackelberg-Littlestone dimension of a hypothesis class $\mathscr{H}$ with respect to a Stackelberg game, $\pazocal{G}$, denoted as $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$, is the supremum root node weight value among all finite node-weighted trees that are shattered by hypothesis class $\pazocal{H}$ under Stackelberg game $\pazocal{G}$.
-Note that taking the supremum is necessary since it is possible that no finite-depth tree achieves the optimal value.
-We begin by showing that no deterministic algorithm can achieve regret lower than the Stackelberg-Littlestone dimension.
-Theorem 3.8.
-For any hypothesis class, $\pazocal{H}$, any Stackelberg game $\pazocal{G}$, and any deterministic algorithm, there exists an adversarial sequence of contexts and follower types such that the learner suffers at least $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})-\epsilon$ regret, for any $\epsilon>0$.
+Observe that by construction, for any true hypothesis $h_{*}\in\pazocal{H}$ we have that the optimal policy $\pi_{h^{*}}$, will satisfy $\pi_{h^{*}}(\mathbf{z})=\pi_{*}^{(1)}(\mathbf{z})$ if $\mathbf{z}[2]<\frac{1}{2}$ and $\pi_{h^{*}}(\mathbf{z})=\pi_{*}^{(2)}(\mathbf{z})$ otherwise.
+Therefore, no matter which hypothesis is used to predict the follower’s type, the utility of the leader will be optimal.
+On the other hand, determining $h^{*}$ requires learning the exact cutoff values, $\gamma_{1}$ and $\gamma_{2}$.
+This corresponds to learning the class of linear thresholds, which has infinite Littlestone dimension (cf.
+Example 21.4 in [uml]).
+Therefore, $\textnormal{Ldim}(\pazocal{H})=\infty$.
+∎
+See 3.10
 Proof.
 We prove by induction on the depth $d$ of shattered trees that across $d$ timesteps the learner can be forced to incur regret at least as much as the weight of the tree’s root node.
-Since the Stackelberg-Littlestone dimension is defined as the supremum root node weight over all finite-depth trees, proving this suffices to show the desired claim.
+Since the SL dimension is defined as the supremum root node weight over all finite-depth trees, proving this suffices to show the desired claim.
 In case the supremum in the definition of $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ is not achieved by any finite tree, there will always be an additive error of $\epsilon$ that will get arbitrarily small as $d\to\infty$.
-(Base Case) We consider a shattered tree of depth 111. Then
+(Base Case) We consider a shattered tree of depth $1$. Then
 $\rho_{\varnothing}$ $=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})$
@@
 (Inductive step) We assume that for any tree of depth $d$ that is shattered by $\pazocal{H}$ under $\pazocal{G}$, the weight of the root node lower bounds the learner’s cumulative regret across $d$ timesteps.
-We now consider a tree $\mathscr{T}_{d+1}$ of depth $d+1$ that is shattered by $\pazocal{H}$ under Stackelberg game $\pazocal{G}$.
+We now consider a tree $\mathcal{T}_{d+1}$ of depth $d+1$ that is shattered by $\pazocal{H}$ under Stackelberg game $\pazocal{G}$.
 We will prove that for such a tree, the adversary can guarantee regret at least $\rho_{\varnothing}$, where $\mathbf{z}_{\varnothing}$ is the root node of this tree.
-By Definition˜3.6, we know that $\rho_{\varnothing}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl{[}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})+\rho_{i}\biggr{]}$.
+By Definition˜3.7, we know that $\rho_{\varnothing}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl[r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})+\rho_{\varnothing i}\biggr]$.
 Let $\mathbf{x}_{*}\in\Delta(\pazocal{A})$ and $i_{*}\in[K]$ be the respective arginf and argmax values in the above expression.
@@
 $\implies r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})+\rho_{i_{*}}$ $\leq r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(\tilde{i})})+\rho_{\tilde{i}}$
-$\implies r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})-r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(\tilde{i})})$ $\leq\rho_{\tilde{i}}-\rho_{i_{*}}$
-We denote $\ell:=r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})-r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(\tilde{i})})$ and by definition of $\tilde{\mathbf{x}}$ we know that $\ell>0$.
-We get that $\rho_{\tilde{i}}\geq\rho_{i_{*}}+\ell$.
-Now by the boundedness of the utility function, we know that if $\rho_{i_{*}}+\ell\geq d$, then the above inequality leads to a contradiction (the learner cannot incur strictly more than $d$ regret over $d$ timesteps).
+[…]
+[…]
 Otherwise, we have a subtree with root node weight at least $\rho_{i_{*}}+\ell$ that is shattered by the induced hypothesis space $\pazocal{H}^{(\mathbf{z}_{\varnothing}\to f^{(\tilde{i})})}$.
@@
 ∎
-3.4 Stackelberg Standard Optimal Algorithm
-In this section, we present an algorithm that achieves instance-optimal regret that is equal to the Stackelberg-Littlestone dimension.
-We call Algorithm˜1 the Stackelberg Standard Optimal Algorithm (SSOA), which maintains a similar approach to the Standard Optimal Algorithm (SOA) from online multiclass prediction.
-Specifically, it keeps track of the subset of hypotheses that are consistent with the history so far (˜8) and plays the mixed strategy that minimizes the worst-case possible instantaneous regret “regularized” by the Stackelberg-Littlestone dimension of the induced hypothesis class (˜6).
-The latter works as a proxy to the difficulty of the learning task of the leader that would remain after the current round is over.
-Algorithm 1 Stackelberg Standard Optimal Algorithm (SSOA)
-1:input: time horizon, $T$, Stackelberg game $\pazocal{G}$, and hypothesis class $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$
-2:Set $V_{1}=\pazocal{H}$
-3:for $t\in[T]$ do
-4:  Observe $\mathbf{z}_{t}$
-5:  Let $u_{*}^{(i,\mathbf{z}_{t})}:=\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{f^{(i)}}(\mathbf{z}_{t},\mathbf{x}))$ for all $i\in V_{t}(\mathbf{z}_{t})$
-6:  Play $\mathbf{x}_{t}=\arg\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in V_{t}(\mathbf{z}_{t})}\left[u_{*}^{(j,\mathbf{z}_{t})}-u(\mathbf{z}_{t},\mathbf{x},b_{f^{(j)}}(\mathbf{z}_{t},\mathbf{x}))+\textnormal{SLdim}_{\pazocal{G}}\left(V_{t}^{(\mathbf{z}_{t}\to j)}\right)\right]$
-7:  Observe $f_{t}$
-8:  Set $V_{t+1}=V_{t}^{(\mathbf{z}_{t}\to f_{t})}$
-9:end
-In Theorem˜3.9, we prove the optimality of our algorithm.
-Our proof follows a high-level idea similar to that of the Standard Optimal Algorithm (SOA) from online multiclass learning: with each mistake, the learner’s task becomes easier, as the Littlestone dimension decreases.
-Similarly here we show that when the learner incurs positive instantaneous regret, the Stackelberg Littlestone dimension decreases by at least this amount.
-Theorem 3.9.
-For any Stackelberg game, $\pazocal{G}$, and every hypothesis class, $\pazocal{H}$, the regret of Algorithm˜1 is at most $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
+See 3.11
 Proof.
-We will show that at each timestep $t$, if the learner incurred instantaneous regret $r_{t}=r(\mathbf{z}_{t},\mathbf{x}_{t},f_{t})$, then the the Stackelberg-Littlestone dimension must have decreased by at least $r_{t}$.
+We will show that at each timestep $t$, if the learner incurred instantaneous regret $r_{t}=r(\mathbf{z}_{t},\mathbf{x}_{t},f_{t})$, then the the SL dimension must have decreased by at least $r_{t}$.
 In other words, $\textnormal{SLdim}_{\pazocal{G}}(V_{t+1})\leq\textnormal{SLdim}_{\pazocal{G}}(V_{t})-r_{t}$, where $V_{t}$ is the set of hypotheses consistent with the history up to timestep $t$.
@@
 where $\mathbf{x}_{*}^{(i,\mathbf{z})}$ is the leader’s optimal strategy against follower $f^{(i)}$ for context $\mathbf{z}$. ∎
-3.5 Relationship with Littlestone dimension
-We have so far shown that the Stackelberg-Littlestone dimension characterizes learnability in our setting.
-We next explore how that relates to multiclass learnability.
-While, as discussed, learning to predict the follower type from the context is sufficient to learn the leader’s optimal policy, in this section we prove that this is not a necessary condition.
-We start by showing that the Littlestone dimension upper bounds the Stackelberg Littlestone dimension.
-The proof of Lemma˜3.10 can be found in Section˜A.4.
-Lemma 3.10.
-For any hypothesis class $\pazocal{H}$ and Stackelberg game, $\pazocal{G}$, $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})\leq\textnormal{Ldim}(\pazocal{H})$.
-While it is known that any learner makes at least $\textnormal{Ldim}(\pazocal{H})$ mispredictions when learning a hypothesis class, $\pazocal{H}$, a careful reader might notice that this does not immediately show that the optimal multiclass learning algorithm (SOA) [22] achieves suboptimal regret in our setting.
-This is because mispredicting the follower type does not necessarily imply that the optimal strategy against the mispredicted and the true follower type do not yield the same utility for the leader.
-An illustrative example of this is our game construction in Theorem˜3.5, which shows that any learner can mispredict the follower type ad infinitum, yet the leader will achieve optimal cumulative utility regardless.
-In Theorem˜3.11, we prove that, indeed, there are instances where SOA will achieve smaller cumulative utility for the leader compared to our proposed algorithm.
-[^5] In fact, there are cases in which SOA yields strictly lower utility even against the algorithm that minimizes policy mispredictions.
-We prove this result in Lemma B.6.
-Theorem 3.11.
-There exists some hypothesis class $\pazocal{H}$ and a Stackelberg game, $\pazocal{G}$, such that running multiclass SOA on $\pazocal{H}$ yields suboptimal utility for the leader.
+See 3.12
 Proof.
+We assume for the sake of contradiction that there exists a hypothesis class with $\textnormal{Ldim}(\pazocal{H})=d$ and a Stackelberg game $\pazocal{G}$, such that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})>d$.
+According to Definition˜B.5, this implies that there exists a shattered tree, whose minimum-weighted root-to-leaf path has weight strictly greater than $d$.
+This implies that all paths in the tree have cumulative weight strictly greater than $d$.
+Given that the weight of each edge in the tree is at most $1$, the length of the minimum path (and therefore of all other paths in the tree) must be of length at least $d+1$.
+This implies that there exists a tree of depth $d+1$ that is shattered by $\pazocal{H}$, which leads to a contradiction.
+∎
+See 3.14
+Proof.
+Consider $\pazocal{Z}=\{1,2,3,4\}$ and followers $\{f^{(1)},f^{(2)},f^{(3)}\}$.
+We define $\pazocal{H}=\{h_{\sigma}(z):=\mathbbm{1}\{z\neq 4\}\cdot f^{(\sigma(z))}+\mathbbm{1}\{z=4\}\cdot f^{(3)}\}_{\sigma\in\Sigma}\cup\{h,h^{\prime}\}$ , where $\Sigma$ is the set of all permutations of the set $\{1,2,3\}$ and $h,h^{\prime}$ are defined as follows.
 […]
-$h(z)=\begin{cases}f^{(3)}\;\;&\text{ if }z=1\\ f^{(2)}\;\;&\text{ if }z=2\\ f^{(1)}\;\;&\text{ if }z=3\\ f^{(2)}\;\;&\text{ if }z=4\\ \end{cases}\hskip 56.9055pth^{\prime}(z)=\begin{cases}f^{(3)}\;\;&\text{ if }z=1\\ f^{(1)}\;\;&\text{ if }z=2\\ f^{(2)}\;\;&\text{ if }z=3\\ f^{(2)}\;\;&\text{ if }z=4\\ \end{cases}$
-Notice that the Littlestone dimension of this hypothesis class is at least 222. This follows from the $\pazocal{H}$-shattered tree construction shown in Figure˜1.
-Figure 1: Construction of an $\pazocal{H}$-shattered Littlestone tree. Each node represents a context in $\pazocal{Z}$ and each edge corresponds to a follower type.
+[…]
+Notice that the Littlestone dimension of this hypothesis class is at least $2$. This follows from the $\pazocal{H}$-shattered tree construction shown in Figure˜4.
 Consider now a Stackelberg game with utility functions $u_{f^{(1)}},u_{f^{(2)}},u_{f^{(3)}}$, defined such that if $z=1$ or $z=2$ then $\pi_{*}^{(1)}(z)=\pi_{*}^{(2)}(z)\neq\pi_{*}^{(3)}(z)$, if $z=3$ then $\pi_{*}^{(1)}(3)=\pi_{*}^{(2)}(3)=\pi_{*}^{(3)}(3)$, and if $z=4$, then $\pi_{*}^{(1)}(4)\neq\pi_{*}^{(2)}(4)=\pi_{*}^{(3)}(4)$.
@@
 Concretely, consider a Stackelberg game with the leader being the row player and each follower being the column player, as shown in Table˜2.
-$F_{1}$ $F_{2}$
-$L_{1}$ 1,1/21,1/21 , 1 / 2 $,0$ 0
-$L_{2}$ $,0$ 0 1,01,01 , 0
+Table 2: Utility tables for the leader (row player) and each follower type (column player)
 (a) Follower type $f^{(1)}$
 $F_{1}$ $F_{2}$
-$L_{1}$ $1,\mathbbm{1}_{\{z\neq 4\}}/2$ $,0$ 0
-$L_{2}$ $,0$ 0 $1,\mathbbm{1}_{\{z=4\}}/2$
+$L_{1}$ $1,1/2$ $0,0$
+$L_{2}$ $0,0$ $1,0$
 (b) Follower type $f^{(2)}$
 $F_{1}$ $F_{2}$
-$L_{1}$ $1,\mathbbm{1}_{\{z=3\}}/2$ $,0$ 0
-$L_{2}$ $,0$ 0 $1,\mathbbm{1}_{\{z\neq 3\}}/2$
+$L_{1}$ $1,\mathbbm{1}_{\{z\neq 4\}}/2$ $0,0$
+$L_{2}$ $0,0$ $1,\mathbbm{1}_{\{z=4\}}/2$
 (c) Follower type $f^{(3)}$
-Table 2: Utility tables for the leader (row player) and each follower type (column player)
+$F_{1}$ $F_{2}$
+$L_{1}$ $1,\mathbbm{1}_{\{z=3\}}/2$ $0,0$
+$L_{2}$ $0,0$ $1,\mathbbm{1}_{\{z\neq 3\}}/2$
 We now show that the $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=3/4$.
-We will consider any possible Stackelberg Littlestone tree.
+We will consider any possible SL tree.
 First, note that without loss of generality we can consider trees that do not include the context $z=4$ since for $z=4,\;\pi_{*}^{(2)}(4)=\pi_{*}^{(3)}(4)$, where $f^{(2)}$ and $f^{(3)}$ are the only possible follower types predicted by hypotheses in $\pazocal{H}$ for this context.
 Similarly, for $z=3$, since the leader’s optimal strategy against all follower types is identical ($\pi_{*}^{(1)}(3)=\pi_{*}^{(2)}(3)=\pi_{*}^{(3)}(3)$).
-So it suffices to consider contexts 111 and 222.
-In Figure˜2 we show the two shattered Stackelberg-Littlestone tree constructions with $z=1$ and $z=2$ as the root nodes respectively.
-We start with the tree in Figure˜2(a).
+So it suffices to consider contexts $1$ and $2$.
+In Figure˜5 we show the two shattered SL tree constructions with $z=1$ and $z=2$ as the root nodes respectively.
+We start with the tree in Figure˜5(a).
 For $z=2$, its not hard to show that $\inf_{\mathbf{x}}\max_{j\in\{1,3\}}r(\mathbf{x},2,f^{(j)})=\inf_{\mathbf{x}}\max_{j\in\{2,3\}}r(\mathbf{x},2,f^{(j)})=1/2$, achieved at $\mathbf{x}=(1/2,1/2)$.
 For the root node $z=1$, we have that $\inf_{\mathbf{x}}\max\{r(\mathbf{x},1,f^{(1)})+1/2,\;r(\mathbf{x},1,f^{(2)})+1/2,\;r(\mathbf{x},1,f^{(3)})\}=3/4$ achieved at $\mathbf{x}=(3/4,1/4)$.
-Therefore, the weight of the root node is 3/43/43 / 4.
-A similar analysis shows that the root node weight of the SL-tree in Figure˜2(b) is 3/43/43 / 4.
-Therefore, $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ (and also the worst-case cumulative regret) equals 3/43/43 / 4.
-We consider the case, where the adversary present context 111 followed by context 222. We go through the SOA outputs per timestep below.
-[…]
+Therefore, the weight of the root node is $3/4$.
+A similar analysis shows that the root node weight of the SL-tree in Figure˜5(b) is $3/4$.
+Therefore, $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ (and also the worst-case cumulative regret) equals $3/4$.
+We consider the case, where the adversary present context $1$ followed by context $2$. We go through the SOA outputs per timestep below.
+  • ($t=1$) This node has three children corresponding to follower types $f^{(1)},f^{(2)},f^{(3)}$, with $f^{(1)}$ and $f^{(2)}$ leading to a full binary subtree of maximal depth 1 each.
+To see this, notice that (1) neither $h$ nor $h^{\prime}$ are consistent with the mapping of context $z=1$ to follower type $f^{(1)}$ or $f^{(2)}$, and (2) $\pazocal{H}\setminus\{h,h^{\prime}\}$ consists of all permutations mapping contexts $\{1,2,3\}$ to follower types $\{f^{(1)},f^{(2)},f^{(3)}\}$.
+Therefore, at most two of these contexts can exist in any root-to-leaf path of any $\pazocal{H}$-shattered tree.
+The child corresponding to follower type $f^{(3)}$ leads to a full binary subtree of maximal depth at least $2$, as shown in Figure˜4.
+So, the predicted follower type is $f_{1}=f^{(3)}$ and the strategy played is $x_{1}=\pi_{*}^{(3)}(1)$.
+However, the true follower type is $f_{1}^{*}=f^{(1)}$ (instantaneous regret equals 1).
   • ($t=2$) This node has two children mapping $z=2$ to follower types $f^{(2)}$ and $f^{(3)}$.
@@
 This leads to SOA mispredicting the follower type and getting suboptimal utility in both rounds.
-Since the Stackelberg Littlestone dimension is 3/43/43 / 4 and we have shown that for the above sequence of contexts running SOA yields regret 222, this concludes our proof.
+Since the Stackelberg Littlestone dimension is $3/4$ and we have shown that for the above sequence of contexts running SOA yields regret $2$, this concludes our proof.
+∎
+Figure 4: Example construction of an $\pazocal{H}$-shattered Littlestone tree. Each node represents a context in $\pazocal{Z}=\{1,2,3,4\}$ and each edge corresponds to a follower type.
 Figure 5: $\pazocal{H}$-shattered SL-Tree with (a) $z=1$ as the root node and (b) $z=2$ as the root node.
-4 Distributional Learning in Structured Stackelberg Games
-We now shift our focus from online learning to distributional learning.
-Consider an unknown, joint distribution over contexts and follower types, from which the leader receives samples drawn i.i.d.
-The leader’s objective is to design her strategy in such a way to maximize the utility she will get on a new sample drawn from the same distribution.
-We first study the simpler setting, without contexts, in which we give strong generalization guarantees by exploiting the structure of the leader’s payoff function.
-When contexts come into play in Section˜4.2, we prove necessary and sufficient conditions for PAC learnability.
-4.1 Learning without Contexts
-Suppose there is a fixed, unknown distribution, $\mathscr{D}$, over the set of possible follower types.
+B.1 A technical subtlety
+An important component of our analysis is that the leader must be able to compute the optimal strategy for a given context and follower type.
+For this, we state that for any context $\mathbf{z}\in\pazocal{Z}$ the leader incurs negligible regret by restricting herself to a subset $\mathcal{E}_{\mathbf{z}}\subseteq\Delta(\pazocal{A})$ which is finite and computable.
+To formally state our result we present the notion of a contextual best-response region, which was first introduced in [harrisregret] and which we repeat here for completeness.
+Definition B.1 (Contextual Follower Best-Response Region).
+For context $\mathbf{z}$, follower type $f^{(i)}$, and follower action $a_{f}\in\pazocal{A}_{f}$, let $X_{\mathbf{z}}(f^{(i)},a_{f})\subseteq\pazocal{X}$ denote the leader’s mixed strategies such that a follower of type $f^{(i)}$ best-responds to all $x\in X_{\mathbf{z}}(f^{(i)},a_{f})$ by playing action $a_{f}$, that is, $X_{\mathbf{z}}(f^{(i)},a_{f})=\{\mathbf{x}\in\pazocal{X}:b_{f^{(i)}}(\mathbf{z},\mathbf{x})=a_{f}\}$
+Notice that any possible follower best-response region is (1) a subset of $\Delta(\pazocal{A})$ and (2) the intersection of finitely-many halfspaces.
+Therefore, every $X_{\mathbf{z}}(f^{(i)},a_{f})$ is convex and bounded.
+However, the leader may not be able to optimize over the extreme points of a best-response region, as it may not be closed due to the followers’ tie-breaking.
+Definition B.2 ($\delta$-approximate extreme points).
+[…]
+Notice that for sufficiently small $\delta$ i.e.
+$\delta=\mathcal{O}(\frac{1}{T})$, we get that the additional regret of the leader by choosing actions over $\mathcal{E}^{(i)}$ rather than $X(f^{(i)},a_{f})$ is negligible.
+Therefore, we use the shorthand $\mathcal{E}^{(i)}:=\mathcal{E}^{(i)}(\delta)$ for the rest of the section.
+The following lemma from [harrisregret] formalizes this notion.
+Lemma B.3.
+For any sequence of followers $f_{1},\ldots f_{T}$ and any leader policy $\pi$, there exists a policy $\pi^{(\mathcal{E})}:\pazocal{Z}\rightarrow\cup_{\mathbf{z}\in\pazocal{Z}}\mathcal{E}_{\mathbf{z}}$ that, when given context $\mathbf{z}$, plays a mixed strategy in $\mathcal{E}_{\mathbf{z}}$ and guarantees that $\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\pi^{(\mathcal{E})}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi^{(\mathcal{E})}(\mathbf{z}_{t})))\leq 1$.
+Moreover, the same result holds in expectation over any distribution over follower types.
+Given Lemma˜B.3 and for ease of notation, in this paper we assume that the leader will always set the parameter $\delta$, such that it does not affect our regret bounds.
+B.2 Alternate definition of the Stackelberg Littlestone dimension
+In this section we provide an alternate definition of the Stackelberg Littlestone dimension that can be defined for infinite trees.
+In Theorem˜B.6 we prove that for finite trees, this is equivalent to Definition˜3.9.
+Definition B.4 ((Edge-weighted) SL tree).
+An edge-weighted SL tree, $\pazocal{T}$, is a rooted tree, in which each internal node is labeled with an instance in $\pazocal{Z}$ and each edge corresponds to a label in $[K]$.
+For each node, $\mathbf{z}_{s}\in\pazocal{T}$ with children $\{\mathbf{z}_{si}\}_{i\in M\subseteq[K]}$ the outgoing edges have weights, $\{\nu_{s}^{(i)}\}_{i\in M}$, such that for every $\mathbf{x}\in\Delta(\pazocal{A})$ there exists $i\in M$ such that $r(\mathbf{z}_{s},\mathbf{x},f^{(i)})\geq\nu_{s}^{(i)}$.
+We say that the tree is shattered by a hypothesis class $\pazocal{H}$ under Stackelberg game $\pazocal{G}$ if for every path that traverses nodes $\mathbf{z}_{\varnothing},\dots\mathbf{z}_{s_{\leq d}}$, there exists $h\in\pazocal{H}$ such that the label of the edge in the path leaving $\mathbf{z}_{s\leq i}$ is $h(\mathbf{z}_{s_{\leq i}})$ for every $i$.
+Definition B.5 ((Edge-weighted) SL dimension).
+The SL dimension of a hypothesis class $\mathcal{H}$ with respect to a Stackelberg game, $\pazocal{G}$, is defined as $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=\sup\{\kappa\in\mathbb{R}_{\geq 0}:$ exists finite edge-weighted SL tree shattered by $\pazocal{H}$ under $\pazocal{G}$ where all root-to-leaf paths have cumulative weight at least $\kappa$}.
+Theorem B.6.
+For finite trees, Definition˜3.9 and Definition˜B.5 coincide.
+Proof.
+It suffices to show that for any finite depth $d$ edge-weighted tree with all root-to-leaf paths having cumulative weight at least $\kappa$ can be shattered under Definition˜B.4 if and only if there exists a tree with root node weight $\kappa$ that can be shattered under Definition˜3.7.
+We proceed by induction on the depth.
+(Base Case) Consider a depth-1 tree, $\pazocal{T}_{1}$, with root node $\mathbf{z}_{\varnothing}$.
+In the node-weight formulation, $\rho_{\varnothing}=\inf_{\mathbf{x}\in\Delta(\mathcal{A})}\max_{i\in\mathcal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})$.
+In the edge-weight formulation, the valid edge-weight assignments are $\mathcal{N}_{\mathbf{z}_{\varnothing}}=\Bigl\{N=\{\nu_{\varnothing}^{(i)}\}_{i\in\mathcal{H}(\mathbf{z}_{\varnothing})}:\;\forall\mathbf{x}\;\exists i\;r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\nu_{\varnothing}^{(i)}\Bigr\}$.
+Their minimal path weight is $\max_{N\in\mathcal{N}_{\mathbf{z}_{\varnothing}}}\min_{i\in\mathcal{H}(\mathbf{z}_{\varnothing})}\nu_{\varnothing}^{(i)}$.
+We will first show that $\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\max_{N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}}\min_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\{\nu_{\varnothing}^{(i)}:\nu_{\varnothing}^{(i)}\in N\}$.
+Fix some $N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}$.
+Then for every $\mathbf{x}\in\Delta(\pazocal{A})$, $\exists i\in\pazocal{H}(\mathbf{z}_{\varnothing})$ such that $r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\nu_{\varnothing}^{(i)}$.
+This means that
+$\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\min_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\{\nu_{\varnothing}^{(i)}:\nu_{\varnothing}^{(i)}\in N\}\;\;\;\forall\mathbf{x}\in\Delta(\pazocal{A})\;\;\forall N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}.$
+Taking the infimum over $\mathbf{x}$ and the maximum over $N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}$ yields the inequality.
+For the reverse inequality, let $\mathbf{x}^{*}\in\arg\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})$ and $N^{*}=\{\nu_{\varnothing}^{(i)}:=\arg\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x}^{*},f^{(i)})\}_{i\in[K]}$.
+We argue that $N^{*}\in\pazocal{N}_{\mathbf{z}_{\varnothing}}$, i.e., that
+$\forall\mathbf{x}\in\Delta(\pazocal{A})\;\;\exists i\in\pazocal{H}(\mathbf{z}_{\varnothing})\;\;\text{such that}\;\;r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\arg\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z},\mathbf{x}^{*},f^{(i)})$
+which holds since $\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})$.
+(Inductive Step) Assume that the equivalence holds for all trees of depth d1d-1.
+Consider such a depth-$d$ tree, $\pazocal{T}_{d}$, with root node weight $\mathbf{z}_{\varnothing}$ and children $\mathbf{z}_{\varnothing i}$ for $i\in\pazocal{H}(\mathbf{z}_{\varnothing})$.
+We want to show that the root node weight $\rho_{\varnothing}$ defined as in Definition˜3.7 is equal to the weight of the minimum path in the tree, where the edge weights are defined as in Definition˜B.4.
+Concretely, we want to show that
+$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl[r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(j)})+\rho_{\varnothing j}\biggr]$ $=\sup_{\pazocal{N}_{d}}\inf_{s\in S_{d}}\Biggl[\sum_{i=0}^{d}\nu_{s[:i]}^{(s[i+1])}\Biggr],$
+where $\pazocal{N}_{d}$ is the collection of sets of all possible edge weights, one for each node in the tree of depth $d$. This holds if and only if
+$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl[r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(j)})+\rho_{\varnothing j}\biggr]$ $=\max_{\pazocal{N}_{0}}\min_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\Biggl[\nu_{\varnothing}^{(s[1])}+\max_{\pazocal{N}_{d-1}}\inf_{s\in S_{d-1}}\sum_{i=1}^{d}\nu_{s[:i]}^{(s[i+1])}\Biggr]$
+We know that for any $j\in\pazocal{H}(\mathbf{z}_{\varnothing})$ and $s\in S_{d}$ such that $s[1]=j$, $\rho_{\varnothing j}=\max_{\pazocal{N}_{d-1}}\inf_{s\in S_{d-1}}\sum_{i=1}^{d}\nu_{s[:i]}^{(s[i+1])}$ by the inductive hypothesis.
+Therefore, using an argument similar to the base case, the equation above holds.
+∎
+Appendix C Details and proofs from Section˜4
+In our work, we consider two variants of sample complexity, PAC cut-off sample complexity (Definition˜4.1) and PAC sample complexity defined as follows.
+Definition C.1 (PAC sample complexity).
+The PAC sample complexity of a function class $\pazocal{H}$ w.r.t.
+a Stackelberg game $\pazocal{G}$ is defined as $m(\pazocal{H},\pazocal{G};\epsilon,\delta)=\inf_{A}m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta)$, where the infimum is over all learning algorithms.
+$m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta)$ is the smallest integer such that for any $m\geq m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta)$, every distribution $\mathcal{D}$ on $\pazocal{Z}$, and any true hypothesis $h^{*}\in\pazocal{H}$, the expected loss of $A$, $\mathop{\mathbb{E}}_{\mathbf{z}\sim\mathcal{D}}\left[r(\mathbf{z},A(\mathbf{z};S),f^{(h^{*}(\mathbf{z}))})\right]$, is at most $\epsilon$ with probability at least $1-\delta$, where $S=\{(\mathbf{z}_{i},f^{(h^{*}(\mathbf{z}_{i}))})\}_{i\in[m]}$ and $\mathbf{z}_{i}\sim\mathcal{D}$.
+According to the following lemma, any bound with respect to the PAC cut-off complexity immediately implies a bound on the PAC sample complexity as well.
+Lemma C.2 (Lemma 1 in [Attias23:regression]).
+For every $\epsilon,\delta\in(0,1)^{2}$ and every $\pazocal{H}\subseteq[0,1]^{\pazocal{Z}}$, it holds that
+$m(\pazocal{H};\sqrt{\epsilon},\delta,\sqrt{\epsilon})\leq m(\pazocal{H};\epsilon,\delta)\leq m(\pazocal{H},\epsilon/2,\delta,\epsilon/2).$
+See 4.4
+Proof.
+Let $n_{0}=\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$.
+Then we know that for any $1\leq n\leq n_{0}$ there exists a set $S_{n}\in\pazocal{Z}^{n}$ and functions $g_{0},g_{1}:\pazocal{Z}\to\Delta(\pazocal{A})$ such that for any set $T\subseteq S$ there exists an $h$ that agrees with $g_{0}$ on all $\mathbf{z}\in T$ and with $g_{1}$ for any $\mathbf{z}\in S\setminus T$.
+We will show that for a given learning algorithm $A:(\pazocal{Z}\times[K])^{*}\to\Delta(\pazocal{A}){Z}$, there exists a distribution $\mathcal{D}$ on $\pazocal{Z}$ such that the algorithm requires at least $C_{0}\frac{\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})+\log(\frac{1}{\delta})}{\epsilon}$ samples.
+To do that, we will construct a distribution with support the $n_{0}$-sized $\gamma$-shattered set $S$.
+We take $\delta<1/15$ and define the distribution $\mathcal{D}$ on $S$ as follows.
+$\mathbb{P}[\mathbf{z}_{1}]=1-16\epsilon,\;\;\;\;\mathbb{P}[\mathbf{z}_{i}]=\frac{16\epsilon}{d-1}\;\forall i\in\{2,\dots,d-1\}.$
+We draw a set $X$ of $m=\frac{d-1}{64\epsilon}$ i.i.d.
+samples from this distribution.
+We let $A$ be an algorithm that takes $X$ as the input and then produces a mapping $A(X):\pazocal{Z}\to\Delta(\pazocal{A})$.
+We define
+$\text{er}_{\mathcal{D},\gamma,h}(A(X)):=\mathbb{P}_{\mathbf{z}\sim\mathcal{D}}[r(\mathbf{z},A(X)(\mathbf{z}),f^{(h(\mathbf{z}))})>\gamma].$
+It suffices to establish that there exists $h^{*}\in\pazocal{H}$ such that $\mathbb{P}_{X\sim\mathcal{D}^{m}}\left[\text{er}_{\mathcal{D},\gamma,h^{*}}(A(X))>\epsilon\right]>1/15$.
+First consider the event
+$B:|\{\mathbf{x}\in X:\mathbf{x}\in\{\mathbf{z}_{2},\dots,\mathbf{z}_{n}\}\}|\leq(d-1)/2.$
+Using Markov’s inequality, it is easy to show that $\mathbb{P}[B]\geq 1/2$.
+Let $U=\{\mathbf{z}_{2},\dots,\mathbf{z}_{n}\}\setminus X$ be the points excluding $\mathbf{z}_{1}$ that are unseen by the algorithm through the training set.
+Given $B$ we know that $|U|\geq\frac{d-1}{2}$.
+For the remainder of this section we will focus on
+$\text{er}^{\prime}_{\mathcal{D},\gamma,h}(A(X))$ $:=\mathbb{P}_{\mathbf{z}\sim\mathcal{D}}[r(\mathbf{z},A(X)(\mathbf{z}),f^{(h(\mathbf{z}))})$
+$>\gamma\land\mathbf{z}\in\{\mathbf{z}_{2},\dots,\mathbf{z}_{n_{0}}\}],$
+which lower bounds $\text{er}_{\mathcal{D},\gamma,h}(A(X))$.
+Also, we consider the following way to choose a true hypothesis $h$.
+First, we choose vector $b$ uniformly at random from the set $\{0,1\}^{n_{0}}$.
+We then set $h(\mathbf{z}_{i})=g_{0}(\mathbf{z}_{i})$ for all $b_{i}=0$ and $h(\mathbf{z}_{i})=g_{1}(\mathbf{z}_{i})$ for all $b_{i}=1$.
+We note that choosing a random $h$ in this way is equivalent to flipping a fair coin for each point in $S$ to determine its label.
+Since for any point $\mathbf{z}_{i}\in U$ both labels $g_{0}(\mathbf{z}_{i})$ and $g_{1}(\mathbf{z}_{i})$ are consistent with the training set $X$ (by definition of an SN-shattered set), we have that $A(X)$ is independent of the labeling of points in $U$.
+We also know that if we condition on the point drawn from $\mathcal{D}$ being some unseen point $\mathbf{u}\in U$, we get that for any strategy $A(X)(\mathbf{u})\in\Delta(\pazocal{A})$ of the learner $\text{er}^{\prime}_{\mathcal{D},\gamma,h}(A(X))\geq 1$ if $h(\mathbf{u})=\arg\max_{j\in\{g_{0}(\mathbf{u}),g_{1}(\mathbf{u})\}}r(\mathbf{u},A(X)(\mathbf{u}),f^{(j)})$ and 0 otherwise.
+Since we are randomly choosing $h(\mathbf{z})$ to be $g_{0}(\mathbf{z}_{i})$ or $g_{1}(\mathbf{z}_{i})$, we have that the two conditions above happen with probability $1/2$ each.
+Given that there are at least (d1)/2(d-1)/2 unseen points, each sampled with probability $\frac{16\epsilon}{d-1}$, and each forcing the learner to make a $\gamma$-error with probability $1/2$, we get that
+$\mathop{\mathbb{E}}_{h,X}\left[\text{er}^{\prime}_{\mathcal{D},\gamma,h}(A(X))\;|\;B\right]>\frac{1}{2}\cdot\frac{16\epsilon}{d-1}\cdot\frac{(d-1)}{2}=4\epsilon.$
+From this analysis and the probability of $B$, we get that
+$\mathop{\mathbb{E}}_{h,X}\left[\text{er}^{\prime}_{\mathcal{D},\gamma,h}(A(X))\right]$ $=\mathop{\mathbb{E}}_{h,X}\left[\text{er}^{\prime}_{\mathcal{D},\gamma,h}(A(X))\;|\;B\right]\cdot\mathbb{P}[B]$
+$>4\epsilon\cdot\frac{1}{2}=2\epsilon.$
+So there must exists some $h^{*}$ such that $\mathop{\mathbb{E}}_{X}\left[\text{er}^{\prime}_{\mathcal{D},\gamma,h^{*}}(A(X))\right]>2\epsilon$.
+We take $h^{*}$ as the target concept and show that $A$ is likely to produce high error rate.
+We know that
+$\mathbb{P}_{X}\left[\text{er}_{\mathcal{D},\gamma,h^{*}}(A(X))>\epsilon\right]\cdot\mathop{\mathbb{E}}_{X}\left[\text{er}_{\mathcal{D},\gamma,h^{*}}(A(X))\;|\;\text{er}_{\mathcal{D},\gamma,h}(A(X))>\epsilon\right]$
+$+(1-\mathbb{P}_{X}\left[\text{er}_{\mathcal{D},\gamma,h^{*}}(A(X))>\epsilon\right])\cdot\mathop{\mathbb{E}}_{X}\left[\text{er}_{\mathcal{D},\gamma,h^{*}}(A(X))\;|\;\text{er}_{\mathcal{D},\gamma,h^{*}}(A(X))\leq\epsilon\right]>2\epsilon$
+Given that for any algorithm, $A$, it holds that $\text{er}^{\prime}_{\mathcal{D},\gamma,h}(A(X))\leq\mathbb{P}[\mathbf{z}\in\{\mathbf{z}_{2},\dots,\mathbf{z}_{n}\}]=16\epsilon$ we get
+$\mathop{\mathbb{E}}_{X}\left[\text{er}^{\prime}_{\mathcal{D},\gamma,h}(A(X))\;|\;\text{er}_{\mathcal{D},\gamma,h}(A(X))>\epsilon\right]\leq 16\epsilon\;\;\text{for any}\;\;h$
+which implies that for $h^{*}$
+$\mathbb{P}_{X}\left[\text{er}^{\prime}_{\mathcal{D},\gamma,h^{*}}(A(X))>\epsilon\right]\cdot 16\epsilon+(1-\mathbb{P}_{X}\left[\text{er}^{\prime}_{\mathcal{D},\gamma,h^{*}}(A(X))>\epsilon\right])\cdot\epsilon>2\epsilon$
+$\iff\mathbb{P}_{X}\left[\text{er}^{\prime}_{\mathcal{D},\gamma,h^{*}}(A(X))>\epsilon\right]>1/15.$
+Therefore, the learner needs at least $m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)\geq\frac{d-1}{64\epsilon}$ samples.
+We now consider the case, where $X$ contains only $\mathbf{z}_{1}$.
+This occurs with probability $(1-16\epsilon)^{m}\geq e^{-32m\epsilon}$, which is greater than $\delta$, when $m\leq\frac{\log(1/\delta)}{32\epsilon}$.
+Therefore we get that for any algorithm $A$:
+$m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)\geq\max\biggl\{\frac{d-1}{64\epsilon},\frac{\log(1/\delta)}{16\epsilon}\biggr\}=C_{0}\frac{\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})+\log(\frac{1}{\delta})}{\epsilon}$
+Lastly, we consider the case, where $\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})=\infty$.
+Then for any shattered set of size $n\in\mathbb{N}$, we can repeat the argument above, which would show that the cut-off sample complexity of any algorithm is at least $\Omega\left(\frac{n+\log(\frac{1}{\delta})}{\epsilon}\right)$, thus establishing that $m(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)=\infty$.
+∎
+See 4.7
+Proof.
+We work with the cut-off loss problem for some parameters $(\epsilon,\delta,\gamma)\in(0,1)^{3}$. We define the cut-off loss
+$\text{er}_{\mathcal{D},\gamma}(h):=\mathbb{P}_{\mathbf{z}\sim\mathcal{D}}\left[\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}),h^{*}(\mathbf{z})\}}r(\mathbf{z},\mathbf{x},f^{(j)})>\gamma\right]$
+and the empirical cut-off loss for a dataset $S\in(\pazocal{Z}\times[K])^{n}$
+$\hat{\text{er}}_{S,\gamma}(h):=\frac{1}{n}\sum_{(\mathbf{z}_{i},f_{i})\in S}\mathbbm{1}\biggl\{\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}),h^{*}(\mathbf{z})\}}r(\mathbf{z},\mathbf{x},f^{(j)})>\gamma\biggr\}$
+and the empirical loss
+$\hat{\text{er}}_{S}(h):=\frac{1}{n}\sum_{(\mathbf{z}_{i},f_{i})\in S}\mathbbm{1}\biggl\{h(\mathbf{z}_{i})\neq h^{\prime}(\mathbf{z}_{i})\biggr\}.$
+For $\pazocal{M}=\pazocal{Z}\times[K]$ we are interested in the following set
+$Q_{\epsilon,\gamma}=\{M\in\pazocal{M}^{n}:\exists h\in\pazocal{H}:\hat{\text{er}}_{M}(h)=0,\text{er}_{\mathcal{D},\gamma}(h)>\epsilon\}$
+Now using a classic symmetrization argument [uml, Attias23:regression], we can upper bound the probability of $Q_{\epsilon,\gamma}$ with that of the following set
+$R_{\epsilon,\gamma}=\{(N,S)\in\pazocal{M}^{n}\times\pazocal{M}^{n}:\exists h\in\pazocal{H}:\hat{\text{er}}_{N}(h)=0,\hat{\text{er}}_{S,\gamma}(h)\geq\epsilon/2\}.$
+The idea here is to bound the probability that there exists a hypothesis yielding maximum utility for the leader over the sample, but is $\gamma$-far from the optimal hypothesis, by the probability that this hypothesis has poor performance on another sample.
+More concretely, it can be shown that $\mathcal{D}^{n}(Q_{\epsilon,\gamma})\leq 2\mathcal{D}^{2n}(R_{\epsilon,\gamma})$ for $n\geq c/\epsilon$, for some constant $c$.
+To see this, we first write
+$\mathcal{D}^{2n}(R_{\epsilon,\gamma})=\int_{Q_{\epsilon,\gamma}}\mathcal{D}^{n}(\{S\in\pazocal{M}^{n}:\exists h\in\pazocal{H}:\hat{\text{er}}_{N}(h)=0,\hat{\text{er}}_{S,\gamma}(h)\geq\epsilon/2\})d\mathcal{D}^{n}(N)$
+For some $N\in Q_{\epsilon,\gamma}$ consider some $h_{N}\in\pazocal{H}$ that satisfies $\hat{\text{er}}_{N,\gamma}(h_{N})=0$ and $\text{er}_{\mathcal{D},\gamma}(h_{N})>\epsilon$.
+We show that $\mathcal{D}(\hat{\text{er}}_{S,\gamma}(h_{S})\geq\epsilon/2)\geq 1/2$, from which the claim follows straightforwardly.
+We know that $\text{er}_{\mathcal{D},\gamma}(h_{N})>\epsilon$, and therefore $n\cdot\hat{\text{er}}_{S,\gamma}(h_{N})\leq n\cdot\hat{\text{er}}_{S}(h_{N})$ follows a binomial distribution with probability of success at least $\epsilon$.
+We note that the inequality above follows since mispredicting the follower type is a necessary condition for suboptimal utility.
+Using a Chernoff bound, we can show that
+$\mathcal{D}^{n}(S:\hat{\text{er}}_{S,\gamma}(h_{N})<\epsilon/2)\leq e^{-n\epsilon/8}$
+and for $n=c/\epsilon$, we get the desired result.
+The next step is to use a random swap argument to upper bound $\mathcal{D}^{2n}(R_{\epsilon,\gamma})$ with something that is independent of $\mathcal{D}$.
+We let $\Gamma_{n}$ denote the set of all permutations on $\{1,\dots,2n\}$ that potentially swap the $i$’th and $n+i$’th elements for all $i\in[n]$.
+So for all $\sigma\in\Gamma_{n}$ and $i\in[n]$, we have that either $\sigma(i)=i,\sigma(n+i)=n+i$ or $\sigma(i)=n+i,\sigma(n+i)=i$.
+From Lemma 3 in Attias23:regression (cf.
+also [Anthony_Bartlett_1999]), we get that for some $\pazocal{M}=\pazocal{Z}\times[K]$ and distribution $\mathcal{D}$ on $V$:
+$\mathcal{D}^{2n}(R_{\epsilon,\gamma})=\mathop{\mathbb{E}}_{Z\sim\mathcal{D}^{2n}}\mathbb{P}_{\sigma\sim\text{Unif}(\Gamma_{n})}[\sigma(V)\in R_{\epsilon,\gamma}]\leq\max_{Z\in\pazocal{Z}^{2n}}\mathbb{P}_{\sigma\sim\text{Unif}(\Gamma_{n})}[\sigma(V)\in R_{\epsilon,\gamma}].$
+The last step of the proof, requires reducing to learning a partial binary hypothesis class and using a disambiguation argument to turn the partial class into a total one, as in Attias23:regression.
+Concretely, we start by fixing a set $V\in\pazocal{M}^{2n}$, where $v_{i}=(\mathbf{z}_{i},h^{*}(\mathbf{z}_{i}))$ and a set $S=\{\mathbf{z}_{1},\dots,\mathbf{z}_{2n}\}$.
+For the projection, $\pazocal{H}(S)$ of the hypothesis class on the set, one can define the partial binary hypothesis class
+$\pazocal{H}:=\left\{h^{\prime}\in\{0,1,\star\}^{2n}:\exists h\in\pazocal{H}|_{S}:\forall i\in[2n]\;\;h^{\prime}(\mathbf{z}_{i})=\begin{cases}0\;\;\text{if}\;\;h(\mathbf{z}_{i})=h^{*}(\mathbf{z}_{i})\\ 1\;\;\text{if}\;\;\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}),h^{*}(\mathbf{z})\}}r(\mathbf{z},\mathbf{x},f^{(j)})>\gamma\\ \star\;\;\text{if}\;\;\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}),h^{*}(\mathbf{z})\}}r(\mathbf{z},\mathbf{x},f^{(j)})\leq\gamma\end{cases}\right\}$
+It is easy to verify that $\textnormal{VCdim}(\pazocal{H})\leq\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$.
+We can now turn this partial concept class into a total concept class of size $n^{O(\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})\log(n))}$ using the following disambiguation argument.
+Lemma C.3 ([Alon21:partial]).
+Let $\pazocal{H}^{\prime}$ be a partial concept class on a finite space $S$, with $\textnormal{VCdim}(\pazocal{H}^{\prime})=d$.
+Then there exists a disambiguation $\tilde{\pazocal{H}}^{\prime}$ of $\pazocal{H}^{\prime}$ of size $|\tilde{\pazocal{H}}^{\prime}|=|S|^{O(d\log|S|)}$.
+We have that $\sigma(V)\in R_{\epsilon,\gamma}$ iff there exists some $h\in\pazocal{H}$ such that $h(\mathbf{z}_{\sigma(i)})=h^{*}(\mathbf{z}_{\sigma(i)})\;\;\;\forall\;i\in[n]$ and
+$\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\biggl\{\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}_{\sigma(n+i)}),h^{*}(\mathbf{z}_{\sigma(n+i)})\}}r(\mathbf{z}_{\sigma(n+i)},\mathbf{x},f^{(j)})>\gamma\biggr\}\geq\epsilon/2.$
+[…]
+C.1 Learning without contexts
+Suppose there is a fixed, unknown distribution, $\mathcal{D}$, over the set of possible follower types.
 We denote with $u(\mathbf{x},b_{f^{(i)}}(\mathbf{x}))$ the leader’s payoff when committing to a (possibly mixed) strategy $\mathbf{x}\in\Delta(\pazocal{A})$ against a follower of type $f^{(i)}$.
-We use the notation $u_{\mathscr{D}}(\mathbf{x})=\mathbb{E}_{f\sim\mathscr{D}}[u(\mathbf{x},b_{f}(\mathbf{x}))]$ and given a sample $S=\{f_{1},\dots,f_{m}\}$ of follower types, we write $u_{S}(\mathbf{x})=\frac{1}{m}\sum_{f\in S}u(\mathbf{x},b_{f}(\mathbf{x}))$.
-Given such a sample, the goal of the leader is to choose a mixed strategy that fares well in expectation against any new sample from $\mathscr{D}$.
+We use the notation $u_{\mathcal{D}}(\mathbf{x})=\mathbb{E}_{f\sim\mathcal{D}}[u(\mathbf{x},b_{f}(\mathbf{x}))]$ and given a sample $S=\{f_{1},\dots,f_{m}\}$ of follower types, we write $u_{S}(\mathbf{x})=\frac{1}{m}\sum_{f\in S}u(\mathbf{x},b_{f}(\mathbf{x}))$.
+Given such a sample, the goal of the leader is to choose a mixed strategy that fares well in expectation against any new sample from $\mathcal{D}$.
 Concretely, we want to choose $\hat{\mathbf{x}}\in\Delta(\pazocal{A})$ such that
-$\mathbb{P}_{S}\left[\sup_{\mathbf{x}}u_{\mathscr{D}}(\mathbf{x})-u_{\mathscr{D}}(\hat{\mathbf{x}})>\epsilon\right]\leq\delta.$
+$\mathbb{P}_{S}\left[\sup_{\mathbf{x}}u_{\mathcal{D}}(\mathbf{x})-u_{\mathcal{D}}(\hat{\mathbf{x}})>\epsilon\right]\leq\delta.$
 To achieve this, it suffices to show a generalization guarantee i.e., a bound on the difference between the leader’s expected payoff and the empirical payoff over a set of follower types for any possible mixed strategy.
-Definition 4.1 (Generalization guarantee).
-A generalization guarantee for the leader in a Stackelberg game, $\pazocal{G}$, is a function $\epsilon_{\pazocal{G}}:\mathbb{Z}_{\geq 1}\times(0,1)\to\mathbb{R}_{\geq 0}$ such that for any $\delta\in(0,1)$, any $m\in\mathbb{Z}_{\geq 1}$, and any distribution $\mathscr{D}$ over the follower types, with probability at least $1-\delta$ over the draw of a set $S\sim\mathscr{D}^{m}$ for any $\mathbf{x}\in\Delta(\pazocal{A})$, the difference between the average payoff of $\mathbf{x}$ over $S$ and the expected payoff of $\mathbf{x}$ over $\mathscr{D}$ is at most $\epsilon_{\pazocal{G}}(m,\delta)$:
-$\mathbb{P}_{S\sim\mathscr{D}^{m}}\left[\exists\mathbf{x}\in\Delta(\pazocal{A})\;\;\text{such that}\;\;|u_{\mathscr{D}}(\mathbf{x})-u_{S}(\mathbf{x})|>\epsilon_{\pazocal{G}}(m,\delta)\right]\leq\delta$
+Definition C.4 (Generalization guarantee).
+A generalization guarantee for the leader in a Stackelberg game, $\pazocal{G}$, is a function $\epsilon_{\pazocal{G}}:\mathbb{Z}_{\geq 1}\times(0,1)\to\mathbb{R}_{\geq 0}$ such that for any $\delta\in(0,1)$, any $m\in\mathbb{Z}_{\geq 1}$, and any distribution $\mathcal{D}$ over the follower types, with probability at least $1-\delta$ over the draw of a set $S\sim\mathcal{D}^{m}$ for any $\mathbf{x}\in\Delta(\pazocal{A})$, the difference between the average payoff of $\mathbf{x}$ over $S$ and the expected payoff of $\mathbf{x}$ over $\mathcal{D}$ is at most $\epsilon_{\pazocal{G}}(m,\delta)$:
+$\mathbb{P}_{S\sim\mathcal{D}^{m}}\left[\exists\mathbf{x}\in\Delta(\pazocal{A})\;\;\text{s.t.}\;\;|u_{\mathcal{D}}(\mathbf{x})-u_{S}(\mathbf{x})|>\epsilon_{\pazocal{G}}(m,\delta)\right]\leq\delta$
 This implies a relationship between the utility of the leader’s strategy maximizing the average payoff over the training data to the expected utility of the optimal strategy.
-Formally, for $\hat{\mathbf{x}}\in\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u_{S}(\mathbf{x})$ and $\mathbf{x}^{*}\in\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u_{\mathscr{D}}(\mathbf{x})$
-$\mathbb{P}_{S\sim\mathscr{D}^{m}}\left[|u_{\mathscr{D}}(\mathbf{x}^{*})-u_{S}(\hat{\mathbf{x}})|>2\epsilon_{\pazocal{G}}(m,\delta)\right]\leq\delta.$
+Formally, for $\hat{\mathbf{x}}\in\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u_{S}(\mathbf{x})$ and $\mathbf{x}^{*}\in\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u_{\mathcal{D}}(\mathbf{x})$
+$\mathbb{P}_{S\sim\mathcal{D}^{m}}\left[|u_{\mathcal{D}}(\mathbf{x}^{*})-u_{S}(\hat{\mathbf{x}})|>2\epsilon_{\pazocal{G}}(m,\delta)\right]\leq\delta.$
 We let $\pazocal{P}=\{u_{\mathbf{x}}:[K]\to\mathbb{R}\;\;\text{such that}\;\;u_{\mathbf{x}}(f)=u(\mathbf{x},b_{f}(\mathbf{x}))\}$ be the set of payoff functions for the leader.
@@
 We also define the dual class $\tilde{\pazocal{P}}=\{\tilde{u}_{f}:\Delta(\pazocal{A})\to\mathbb{R}\;\;\text{such that}\;\;\tilde{u}_{f}(\mathbf{x})=u(\mathbf{x},b_{f}(\mathbf{x}))\}$.
-Consider the following definition, introduced by Balcan et al.
-[9] and adapted to our setting.
-Definition 4.2.
-[…]
+Consider the following definition, introduced by balcan2018general and adapted to our setting.
+Definition C.5.
+A class $\pazocal{H}$ of hypotheses, $h:[K]\to\mathbb{R}$, is $(d,t)$-delineable if
   • The class $\pazocal{H}$ consists of hypotheses that are parameterized by vectors $\mathbf{x}$ from a set $\pazocal{X}\subseteq\mathbb{R}^{d}$,
-  • For any $f\in[K]$, there is a set $\mathscr{J}$ of $t$ hyperplanes such that for any connected component $\pazocal{X}^{\prime}=\pazocal{X}\setminus\mathscr{J}$, the dual function $h_{f}(\mathbf{x})$ is linear over $\pazocal{X}^{\prime}$.
+  • For any $f\in[K]$, there is a set $\mathcal{J}$ of $t$ hyperplanes such that for any connected component $\pazocal{X}^{\prime}=\pazocal{X}\setminus\mathcal{J}$, the dual function $h_{f}(\mathbf{x})$ is linear over $\pazocal{X}^{\prime}$.
 We can now show that the primal class $\pazocal{P}$ is deliniable, which is a property we will use to show a generalization guarantee.
-Lemma 4.3.
+Lemma C.6.
 The class of leader payoff functions, $\pazocal{P}$ that are parameterized by the leader’s mixed strategy and take as input the follower’s type is $(|\pazocal{A}|,\binom{|\pazocal{A}_{f}|}{2})-$delineable.
@@
 ∎
-Next, we present the definition of pseudo-dimension [34], which is complexity measure that characterizes a hypothesis class and in our setting helps us reason about how difficult the learning task of the leader is.
+Next, we present the definition of pseudo-dimension [pollard2012convergence], which is complexity measure that characterizes a hypothesis class and in our setting helps us reason about how difficult the learning task of the leader is.
 Informally speaking, the more complex the hypothesis class, $\pazocal{P}$, the harder it is for the leader to distinguish between strategies, $\mathbf{x}$, which parameterize the class, and therefore, more samples are needed to learn.
-Definition 4.4 (Pseudo-dimension).
+Definition C.7 (Pseudo-dimension).
 Let $\pazocal{F}$ be a class of functions $f:A\to\mathbb{R}$, with an abstract domain, $A$.
@@
 We then say that the pseudo-dimension of $\pazocal{F}$, denoted $\textnormal{Pdim}(\pazocal{F})$, is the size of the largest set that can be shattered by $\pazocal{F}$.
-Using Lemma˜4.3 and results from the data-driven algorithm design literature [6, 12, 14], we immediately get the following bound on the pseudo-dimension.
-Theorem 4.5 (Adaptation of Lemma 3.10 in Balcan et al. [9]).
+Using Lemma˜C.6 and results from the data-driven algorithm design literature [Balcan_2021, 10.1145/3676278, balcan2025generalization], we immediately get the following bound on the pseudo-dimension.
+Theorem C.8 (Adaptation of Lemma 3.10 in balcan2018general).
 $\textnormal{Pdim}(\pazocal{P})\leq 9|\pazocal{A}|\log\left(4|\pazocal{A}||\pazocal{A}_{f}|^{2}\right)$.
-It is known that the finiteness of the pseudo-dimension of a function class is a sufficient condition to show a generalization guarantee for the class [34, 24].
+It is known that the finiteness of the pseudo-dimension of a function class is a sufficient condition to show a generalization guarantee for the class [pollard2012convergence, dudley2010universal].
 In our case, the above bound implies a generalization guarantee with function
@@
 Informally, it asserts that for any distribution over the follower types, if the leader has so far encountered a number of followers that scales linearly with the size of her action set and logarithmically with the follower’s action set, it suffices for her to play a strategy that maximizes her average utility over the followers seen so far.
-This result holds even for an infinite number of follower types and extends prior work on learning in Stackelberg games without contextual information [29].
-4.2 Learning with Contexts: Problem Setting and Definitions
-In this section, we introduce the study of the statistical complexity of distributional learning in structured Stackelberg games.
-The leader has access to a data set consisting of tuples of contexts and follower types drawn from a fixed and unknown distribution.
-Given new contextual information drawn from the same distribution, the leader aims to choose a strategy that maximizes her payoff against a follower whose type is unknown.
-As in the online setting, this comes with the promise that there exists a function, $h^{*}$, in a class, $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$, that perfectly predicts each follower type given the context.
-We denote a learning algorithm for distributional learning in structured Stackelberg games as a mapping $A:(\pazocal{Z}\times[K])^{*}\to\Delta(\pazocal{A}){Z}$.
-Given a context, $\mathbf{z}$, we write the leader’s loss with respect to $h^{*}$, when playing strategy $\hat{\mathbf{x}}\in\Delta(\pazocal{A})$ as
-$r(\mathbf{z},\hat{\mathbf{x}},f^{(h^{*}(\mathbf{z}))})=\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u\left(\mathbf{z},\mathbf{x},b_{f^{(h^{*}(\mathbf{z}))}}(\mathbf{z},\mathbf{x})\right)-u\left(\mathbf{z},\hat{\mathbf{x}},b_{f^{(h^{*}(\mathbf{z}))}}(\mathbf{z},\hat{\mathbf{x}})\right)$
-We handle two variants of the notion of sample complexity, which will be central in showing our results. We begin with the classic PAC sample complexity, which we adapt to our setting.
-Definition 4.6 (PAC Sample complexity).
-The PAC sample complexity $m(\pazocal{H},\pazocal{G};\epsilon,\delta)$ of a function class $\pazocal{H}$ with respect to a Stackelberg game $\pazocal{G}$ is defined as $m(\pazocal{H},\pazocal{G};\epsilon,\delta)=\inf_{A}m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta)$, where the infimum is over all possible learning algorithms.
-With $m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta)$ we denote the minimal integer such that for any $m\geq m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta)$, every distribution $\mathscr{D}$ on $\pazocal{Z}$, and any true target hypothesis $h^{*}\in\pazocal{H}$, the expected loss of $A$,
-$\mathop{\mathbb{E}}_{\mathbf{z}\sim\mathscr{D}}\left[r(\mathbf{z},A(S)(\mathbf{z}),f^{(h^{*}(\mathbf{z}))})\right],$
-is at most $\epsilon$ with probability $1-\delta$ over the set $S=\{(\mathbf{z}_{i},f^{(h^{*}(\mathbf{z}_{i}))})\}_{i\in[m]}$, where $\forall i\in[m]\;\mathbf{z}_{i}\sim\mathscr{D}$.
-Definition 4.7 (PAC Cut-off sample complexity).
-The PAC sample complexity $m(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$ of a function class $\pazocal{H}$ with respect to a Stackelberg game $\pazocal{G}$ is defined as $m(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)=\inf_{A}m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$, where $m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$ is the minimal integer such that for any $m\geq m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)$, every distribution $\mathscr{D}$ on $\pazocal{Z}$, and any true target hypothesis $h^{*}\in\pazocal{H}$, the expected cut-off loss of algorithm $A$,
-$\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}[r(\mathbf{z},A(S)(\mathbf{z}),f^{(h^{*}(\mathbf{z}))})>\gamma],$
-is at most $\epsilon$ with probability $1-\delta$ over the set $S=\{(\mathbf{z}_{i},f^{(h^{*}(\mathbf{z}_{i}))})\}_{i\in[m]}$, where $\forall i\in[m]\;\mathbf{z}_{i}\sim\mathscr{D}$.
-For the remainder of this section our results will be with respect to the PAC cut-off sample complexity.
-Any bound with respect to this benchmark immediately implies a bound on the PAC sample complexity as well.
-It is not hard to generalize the result of Lemma˜4.8 to our setting, so we omit the proof.
-Lemma 4.8 (Lemma 1 in [3]).
-For every $\epsilon,\delta\in(0,1)^{2}$ and every $\pazocal{H}\subseteq[0,1]^{\pazocal{Z}}$, it holds that
-$m(\pazocal{H};\sqrt{\epsilon},\delta,\sqrt{\epsilon})\leq m(\pazocal{H};\epsilon,\delta)\leq m(\pazocal{H},\epsilon/2,\delta,\epsilon/2).$
-4.3 Lower bound
-In this section, we introduce the $\gamma$-valued Stackelberg-Natarajan dimension, a new measure of complexity that characterizes a structured Stackelberg game and, as we will see, controls the lower bound of the sample complexity.
-Indeed, we prove that the finiteness of this measure is a necessary condition for learnability in our setting.
-This definition is inspired by the classic Natarajan dimension from multiclass prediction [31, 21], and more recently, the $\gamma$-Natarajan dimension from realizable regression [3].
-We start with introducing the notion of a $\gamma$-SN-shattered set.
-Informally, a set of contexts is $\gamma$-SN-shattered if $\pazocal{H}$ can fit complex patterns of follower types on the set.
-Further, these follower types must be sufficiently different, such that if the type is chosen adversarially among them, the leader incurs loss at least $\gamma$ regardless of her strategy.
-Definition 4.9 ($\gamma$-SN-Shattered Set).
-Let $\pazocal{H}:[K]^{\pazocal{Z}}$ and $\pazocal{G}$ be a Stackelberg game.
-We say that a set $S\in\pazocal{Z}^{n}$ is $\gamma$-SN-shattered by $\pazocal{H}$ with respect to $\pazocal{G}$ if there exist $g_{0},g_{1}:\pazocal{Z}\to[K]$ such that
-$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{g_{0}(\mathbf{z}_{i}),g_{1}(\mathbf{z}_{i})\}}r(\mathbf{z}_{i},\mathbf{x},f^{(j)})>\gamma\;\;\forall\mathbf{z}_{i}\in S$
-and for all $b\in\{0,1\}^{n}$ there exists an $h_{b}\in\pazocal{H}$ such that $h_{b}(\mathbf{z}_{i})=\begin{cases}g_{0}(\mathbf{z}_{i})\;\;\text{if}\;\;b_{i}=0\\ g_{1}(\mathbf{z}_{i})\;\;\text{if}\;\;b_{i}=1\end{cases}$.
-Definition 4.10 ($\gamma$-valued Stackelberg-Natarajan Dimension).
-The $\gamma$-valued Stackelberg-Natarajan dimension, denoted as $\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$, is the cardinality of the largest set that is $\gamma$-SN-shattered by $\pazocal{H}$ with respect to $\pazocal{G}$.
-Before stating our main result of this section, we point out that it is important that the two functions, $g_{0}$ and $g_{1}$ be $\gamma$-far from each other in a minmax sense.
-This allows us to prove a lower bound that holds for any learning algorithm.
-Theorem 4.11.
-Let $A$ be any learning algorithm and $\epsilon,\delta,\gamma\in(0,1)^{3}$ such that $\delta<\epsilon$, then
-$m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)\geq\Omega\left(\frac{\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})+\log(1/\delta)}{\epsilon}\right).$
-Proof.
-Let $n_{0}=\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$.
-Then we know that for any $1\leq n\leq n_{0}$ there exists a set $S_{n}\in\pazocal{Z}^{n}$ and functions $g_{0},g_{1}:\pazocal{Z}\to\Delta(\pazocal{A})$ such that for any set $T\subseteq S$ there exists an $h$ that agrees with $g_{0}$ on all $\mathbf{z}\in T$ and with $g_{1}$ for any $\mathbf{z}\in S\setminus T$.
-We will show that for a given learning algorithm $A:(\pazocal{Z}\times[K])^{*}\to\Delta(\pazocal{A}){Z}$, there exists a distribution $\mathscr{D}$ on $\pazocal{Z}$ such that the algorithm requires at least $C_{0}\frac{\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})+\log(\frac{1}{\delta})}{\epsilon}$ samples.
-To do that, we will construct a distribution with support the $n_{0}$-sized $\gamma$-shattered set $S$.
-We take $\delta<1/15$ and define the distribution $\mathscr{D}$ on $S$ as follows.
-$\mathbb{P}[\mathbf{z}_{1}]=1-16\epsilon,\;\;\;\;\mathbb{P}[\mathbf{z}_{i}]=\frac{16\epsilon}{d-1}\;\forall i\in\{2,\dots,d-1\}.$
-We draw a set $X$ of $m=\frac{d-1}{64\epsilon}$ i.i.d.
-samples from this distribution.
-We let $A$ be an algorithm that takes $X$ as the input and then produces a mapping $A(X):\pazocal{Z}\to\Delta(\pazocal{A})$.
-We define
-$\text{er}_{\mathscr{D},\gamma,h}(A(X)):=\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}[r(\mathbf{z},A(X)(\mathbf{z}),f^{(h(\mathbf{z}))})>\gamma].$
-It suffices to establish that there exists $h^{*}\in\pazocal{H}$ such that $\mathbb{P}_{X\sim\mathscr{D}^{m}}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right]>1/15$.
-First consider the event
-$B:|\{\mathbf{x}\in X:\mathbf{x}\in\{\mathbf{z}_{2},\dots,\mathbf{z}_{n}\}\}|\leq(d-1)/2.$
-Using Markov’s inequality, it is easy to show that $\mathbb{P}[B]\geq 1/2$.
-Let $U=\{\mathbf{z}_{2},\dots,\mathbf{z}_{n}\}\setminus X$ be the points excluding $\mathbf{z}_{1}$ that are unseen by the algorithm through the training set.
-Given $B$ we know that $|U|\geq\frac{d-1}{2}$.
-For the remainder of this section we will focus on
-$\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X)):=\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}[r(\mathbf{z},A(X)(\mathbf{z}),f^{(h(\mathbf{z}))})>\gamma\land\mathbf{z}\in\{\mathbf{z}_{2},\dots,\mathbf{z}_{n_{0}}\}],$
-which trivially lower bounds $\text{er}_{\mathscr{D},\gamma,h}(A(X))$.
-Also, we consider the following way to choose a true hypothesis $h$.
-First, we choose vector $b$ uniformly at random from the set $\{0,1\}^{n_{0}}$.
-We then set $h(\mathbf{z}_{i})=g_{0}(\mathbf{z}_{i})$ for all $b_{i}=0$ and $h(\mathbf{z}_{i})=g_{1}(\mathbf{z}_{i})$ for all $b_{i}=1$.
-We note that choosing a random $h$ in this way is equivalent to flipping a fair coin for each point in $S$ to determine its label.
-Since for any point $\mathbf{z}_{i}\in U$ both labels $g_{0}(\mathbf{z}_{i})$ and $g_{1}(\mathbf{z}_{i})$ are consistent with the training set $X$ (by definition of an SN-shattered set), we have that $A(X)$ is independent of the labeling of points in $U$.
-We also know that if we condition on the point drawn from $\mathscr{D}$ being some unseen point $\mathbf{u}\in U$, we get that for any strategy $A(X)(\mathbf{u})\in\Delta(\pazocal{A})$ of the learner
-$\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\geq\begin{cases}1\;\;\text{if}\;h(\mathbf{u})=\arg\max_{j\in\{g_{0}(\mathbf{u}),g_{1}(\mathbf{u})\}}r(\mathbf{u},A(X)(\mathbf{u}),f^{(j)})\\ 0\;\;\text{otherwise}.\end{cases}$
-Since we are randomly choosing $h(\mathbf{z})$ to be $g_{0}(\mathbf{z}_{i})$ or $g_{1}(\mathbf{z}_{i})$, we have that the two conditions above happen with probability 1/21/21 / 2 each.
-Given that there are at least $(d-1)/2$ unseen points, each sampled with probability $\frac{16\epsilon}{d-1}$, and each forcing the learner to make a $\gamma$-error with probability 1/21/21 / 2, we get that
-$\mathop{\mathbb{E}}_{h,X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\;|\;B\right]>\frac{1}{2}\cdot\frac{16\epsilon}{d-1}\cdot\frac{(d-1)}{2}=4\epsilon.$
-From this analysis and the probability of $B$, we get that
-$\mathop{\mathbb{E}}_{h,X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\right]$ $=\mathop{\mathbb{E}}_{h,X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\;|\;B\right]\cdot\mathbb{P}[B]$
-$>4\epsilon\cdot\frac{1}{2}=2\epsilon.$
-So there must exists some $h^{*}$ such that $\mathop{\mathbb{E}}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h^{*}}(A(X))\right]>2\epsilon$.
-We take $h^{*}$ as the target concept and show that $A$ is likely to produce high error rate.
-We know that
-$\mathbb{P}_{X}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right]\cdot\mathop{\mathbb{E}}_{X}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))\;|\;\text{er}_{\mathscr{D},\gamma,h}(A(X))>\epsilon\right]$
-$+(1-\mathbb{P}_{X}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right])\cdot\mathop{\mathbb{E}}_{X}\left[\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))\;|\;\text{er}_{\mathscr{D},\gamma,h^{*}}(A(X))\leq\epsilon\right]>2\epsilon$
-Given that for any algorithm, $A$, it holds that $\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\leq\mathbb{P}[\mathbf{z}\in\{\mathbf{z}_{2},\dots,\mathbf{z}_{n}\}]=16\epsilon$ we get
-$\mathop{\mathbb{E}}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h}(A(X))\;|\;\text{er}_{\mathscr{D},\gamma,h}(A(X))>\epsilon\right]\leq 16\epsilon\;\;\text{for any}\;\;h$
-which implies that for $h^{*}$
-$\mathbb{P}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right]\cdot 16\epsilon+(1-\mathbb{P}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right])\cdot\epsilon>2\epsilon$
-$\iff\mathbb{P}_{X}\left[\text{er}^{\prime}_{\mathscr{D},\gamma,h^{*}}(A(X))>\epsilon\right]>1/15.$
-Therefore, the learner needs at least $m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)\geq\frac{d-1}{64\epsilon}$ samples.
-We now consider the case, where $X$ contains only $\mathbf{z}_{1}$.
-This occurs with probability $(1-16\epsilon)^{m}\geq e^{-32m\epsilon}$, which is greater than $\delta$, when $m\leq\frac{\log(1/\delta)}{32\epsilon}$.
-Therefore we get that for any algorithm $A$:
-$m_{A}(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)\geq\max\biggl{\{}\frac{d-1}{64\epsilon},\frac{\log(1/\delta)}{16\epsilon}\biggr{\}}=C_{0}\frac{\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})+\log(\frac{1}{\delta})}{\epsilon}$
-Lastly, we consider the case, where $\textnormal{SNdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})=\infty$.
-Then for any shattered set of size $n\in\mathbb{N}$, we can repeat the argument above, which would show that the cut-off sample complexity of any algorithm is at least $\Omega\left(\frac{n+\log(\frac{1}{\delta})}{\epsilon}\right)$, thus establishing that $m(\pazocal{H},\pazocal{G};\epsilon,\delta,\gamma)=\infty$.
-∎
-4.4 Upper bound
-In this section, we introduce the $\gamma$-valued Stackelberg-Graph dimension, which, as we will see, is a sufficient condition for distributional learning in structured Stackelberg games.
-It is also a natural relaxation of the $\gamma$-Natarajan dimension and an adaptation of similar dimensions from classification [31, 21] and regression [3] to our setting.
-Definition 4.12 ($\gamma$-SG-Shattered Set).
-Let $\pazocal{H}:[K]^{\pazocal{Z}}$ and $\pazocal{G}$ be a Stackelberg game.
-We say that a sample $S\in\pazocal{Z}^{n}$ is shattered by $\pazocal{H}$ with respect to $\pazocal{G}$ if there exists a function $g:\pazocal{Z}\to[K]$ such that for all $b\in\{0,1\}^{n}$ there exists an $h_{b}\in\pazocal{H}$ such that:
-  • $h_{b}(\mathbf{z}_{i})=g(i)\;\;\forall i\in[n]$ such that $b_{i}=0$, and
-  • $\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{g(\mathbf{z}_{i}),h_{b}(\mathbf{z}_{i})\}}r(\mathbf{z}_{i},\mathbf{x},f^{(j)})\geq\gamma\;\;\forall i\in[n]$ such that $b_{i}=1$.
-Definition 4.13 ($\gamma$-valued Stackelberg-Graph Dimension).
-The $\gamma$-valued Stackelberg-Graph dimension, denoted as $\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$ is the cardinality of the largest set that is $\gamma$-shattered by $\pazocal{H}$ with respect to $\pazocal{G}$.
-We are not ready to show our main result in this section, which shows that the $\gamma$-Graph dimension upper bounds the PAC sample complexity in our setting.
-We design the following simple, improper algorithm that achieves the corresponding bound.
-Algorithm 2 Stackelberg Distributional Learner ($\mathfrak{L}^{*}$)
-1:input: sample $S=\{(\mathbf{z}_{1},f_{1}),\dots,(\mathbf{z}_{n},f_{n})\}$ that is realizable by $\pazocal{H}$ and test point $\mathbf{z}\in\pazocal{Z}$
-2:Consider the subset of hypotheses consistent with the sample, $\pazocal{H}_{c}=\pazocal{H}|_{S}$.
-3:Consider the set of followers predicted by $\pazocal{H}_{c}$ on test point $\mathbf{z}$, $F=\pazocal{H}_{c}(\mathbf{z})$
-4:Compute and return $\mathbf{x}^{*}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in F}r(\mathbf{z},\mathbf{x},f^{(i)})$
-5:end
-Theorem 4.14.
-For any class $\pazocal{H}:[K]^{\pazocal{Z}}$, any Stackelberg game, $\pazocal{G}$, and for any $\epsilon,\delta,\gamma\in(0,1)$, there exists some constant $C_{1}$ such that Algorithm˜2 achieves:
-$m_{\mathfrak{L}^{*}}^{r}(\pazocal{H},\epsilon,\delta,\gamma)\leq C_{1}\frac{\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})\log(1/\epsilon)+\log(1/\delta)}{\epsilon}.$
-Proof.
-We work with the cut-off loss problem for some parameters $(\epsilon,\delta,\gamma)\in(0,1)^{3}$. We define the cut-off loss
-$\text{er}_{\mathscr{D},\gamma}(h):=\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}\left[\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}),h^{*}(\mathbf{z})\}}r(\mathbf{z},\mathbf{x},f^{(j)})>\gamma\right]$
-and the empirical cut-off loss for a dataset $S\in(\pazocal{Z}\times[K])^{n}$
-$\hat{\text{er}}_{S,\gamma}(h):=\frac{1}{n}\sum_{(\mathbf{z}_{i},f_{i})\in S}\mathbbm{1}\biggl{\{}\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}),h^{*}(\mathbf{z})\}}r(\mathbf{z},\mathbf{x},f^{(j)})>\gamma\biggr{\}}$
-and the empirical loss
-$\hat{\text{er}}_{S}(h):=\frac{1}{n}\sum_{(\mathbf{z}_{i},f_{i})\in S}\mathbbm{1}\biggl{\{}h(\mathbf{z}_{i})\neq h^{\prime}(\mathbf{z}_{i})\biggr{\}}.$
-For $\pazocal{M}=\pazocal{Z}\times[K]$ we are interested in the following set
-$Q_{\epsilon,\gamma}=\{M\in\pazocal{M}^{n}:\exists h\in\pazocal{H}:\hat{\text{er}}_{M}(h)=0,\text{er}_{\mathscr{D},\gamma}(h)>\epsilon\}$
-Now using a classic symmetrization argument [35, 3], we can upper bound the probability of $Q_{\epsilon,\gamma}$ with that of the following set
-$R_{\epsilon,\gamma}=\{(N,S)\in\pazocal{M}^{n}\times\pazocal{M}^{n}:\exists h\in\pazocal{H}:\hat{\text{er}}_{N}(h)=0,\hat{\text{er}}_{S,\gamma}(h)\geq\epsilon/2\}.$
-The idea here is to bound the probability that there exists a hypothesis yielding maximum utility for the leader over the sample, but is $\gamma$-far from the optimal hypothesis, by the probability that this hypothesis has poor performance on another sample.
-More concretely, it can be shown that $\mathscr{D}^{n}(Q_{\epsilon,\gamma})\leq 2\mathscr{D}^{2n}(R_{\epsilon,\gamma})$ for $n\geq c/\epsilon$, for some constant $c$.
-To see this, we first write
-$\mathscr{D}^{2n}(R_{\epsilon,\gamma})=\int_{Q_{\epsilon,\gamma}}\mathscr{D}^{n}(\{S\in\pazocal{M}^{n}:\exists h\in\pazocal{H}:\hat{\text{er}}_{N}(h)=0,\hat{\text{er}}_{S,\gamma}(h)\geq\epsilon/2\})d\mathscr{D}^{n}(N)$
-For some $N\in Q_{\epsilon,\gamma}$ consider some $h_{N}\in\pazocal{H}$ that satisfies $\hat{\text{er}}_{N,\gamma}(h_{N})=0$ and $\text{er}_{\mathscr{D},\gamma}(h_{N})>\epsilon$.
-We show that $\mathscr{D}(\hat{\text{er}}_{S,\gamma}(h_{S})\geq\epsilon/2)\geq 1/2$, from which the claim follows straightforwardly.
-We know that $\text{er}_{\mathscr{D},\gamma}(h_{N})>\epsilon$, and therefore $n\cdot\hat{\text{er}}_{S,\gamma}(h_{N})\leq n\cdot\hat{\text{er}}_{S}(h_{N})$ follows a binomial distribution with probability of success at least $\epsilon$.
-We note that the inequality above follows since mispredicting the follower type is a necessary condition for suboptimal utility.
-Using a Chernoff bound, we can show that
-$\mathscr{D}^{n}(S:\hat{\text{er}}_{S,\gamma}(h_{N})<\epsilon/2)\leq e^{-n\epsilon/8}$
-and for $n=c/\epsilon$, we get the desired result.
-The next step is to use a random swap argument to upper bound $\mathscr{D}^{2n}(R_{\epsilon,\gamma})$ with something that is independent of $\mathscr{D}$.
-We let $\Gamma_{n}$ denote the set of all permutations on $\{1,\dots,2n\}$ that potentially swap the $i$’th and $n+i$’th elements for all $i\in[n]$.
-So for all $\sigma\in\Gamma_{n}$ and $i\in[n]$, we have that either $\sigma(i)=i,\sigma(n+i)=n+i$ or $\sigma(i)=n+i,\sigma(n+i)=i$.
-From Lemma 3 in Attias et al.
-[3] (cf.
-also [2]), we get that for some $\pazocal{M}=\pazocal{Z}\times[K]$ and distribution $\mathscr{D}$ on $V$:
-$\mathscr{D}^{2n}(R_{\epsilon,\gamma})=\mathop{\mathbb{E}}_{Z\sim\mathscr{D}^{2n}}\mathbb{P}_{\sigma\sim\text{Unif}(\Gamma_{n})}[\sigma(V)\in R_{\epsilon,\gamma}]\leq\max_{Z\in\pazocal{Z}^{2n}}\mathbb{P}_{\sigma\sim\text{Unif}(\Gamma_{n})}[\sigma(V)\in R_{\epsilon,\gamma}].$
-The last step of the proof, requires reducing to learning a partial binary hypothesis class and using a disambiguation argument to turn the partial class into a total one, as in Attias et al.
-[3].
-Concretely, we start by fixing a set $V\in\pazocal{M}^{2n}$, where $v_{i}=(\mathbf{z}_{i},h^{*}(\mathbf{z}_{i}))$ and a set $S=\{\mathbf{z}_{1},\dots,\mathbf{z}_{2n}\}$.
-For the projection, $\pazocal{H}(S)$ of the hypothesis class on the set, one can define the partial binary hypothesis class
-$\pazocal{H}:=\left\{h^{\prime}\in\{0,1,\star\}^{2n}:\exists h\in\pazocal{H}|_{S}:\forall i\in[2n]\;\;h^{\prime}(\mathbf{z}_{i})=\begin{cases}0\;\;\text{if}\;\;h(\mathbf{z}_{i})=h^{*}(\mathbf{z}_{i})\\ 1\;\;\text{if}\;\;\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}),h^{*}(\mathbf{z})\}}r(\mathbf{z},\mathbf{x},f^{(j)})>\gamma\\ \star\;\;\text{if}\;\;\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}),h^{*}(\mathbf{z})\}}r(\mathbf{z},\mathbf{x},f^{(j)})\leq\gamma\end{cases}\right\}$
-It is easy to verify that $\textnormal{VCdim}(\pazocal{H})\leq\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})$.
-We can now turn this partial concept class into a total concept class of size $n^{O(\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})\log(n))}$ using the following disambiguation argument.
-Lemma 4.15 ([1]).
-Let $\pazocal{H}^{\prime}$ be a partial concept class on a finite space $S$, with $\textnormal{VCdim}(\pazocal{H}^{\prime})=d$.
-Then there exists a disambiguation $\tilde{\pazocal{H}}^{\prime}$ of $\pazocal{H}^{\prime}$ of size $|\tilde{\pazocal{H}}^{\prime}|=|S|^{O(d\log|S|)}$.
-We have that $\sigma(V)\in R_{\epsilon,\gamma}$ iff there exists some $h\in\pazocal{H}$ such that $h(\mathbf{z}_{\sigma(i)})=h^{*}(\mathbf{z}_{\sigma(i)})\;\;\;\forall\;i\in[n]$ and
-$\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\biggl{\{}\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in\{h(\mathbf{z}_{\sigma(n+i)}),h^{*}(\mathbf{z}_{\sigma(n+i)})\}}r(\mathbf{z}_{\sigma(n+i)},\mathbf{x},f^{(j)})>\gamma\biggr{\}}\geq\epsilon/2.$
-We know that if there exists such an $h$ that is correct on the first $n$ points but $\gamma$-far on at least $\epsilon n/2$ of the last $n$ points, there exists a function in $\tilde{\pazocal{H}}^{\prime}$, that is 0 on the first $n$ points and 111 on the points on which $h$ is $\gamma$-far.
-We also know that if for some $\sigma\in\Gamma_{n}$, $\tilde{h}^{\prime}\in\pazocal{H}^{\prime}$ is a witness that $\sigma(V)\in R_{\epsilon,\gamma}$, then at least one of $\tilde{h}^{\prime}(i),\tilde{h}^{\prime}(i)$ must be 0, and at least $n\epsilon/2$ elements must be non-zero.
-Thus for a permutation drawn uniformly at random, the probability that all 1’s are on the last n elements is at most $2^{-n\epsilon/2}$.
-Also since $|\tilde{\pazocal{H}}^{\prime}|$ is bounded we an use a union bound that yields $\mathbb{P}_{\sigma\sim\text{Unif}(\Gamma_{n})}[\sigma(V)\in R_{\epsilon,\gamma}]\leq n^{O(\textnormal{SGdim}^{(\gamma)}_{\pazocal{G}}(\pazocal{H})\log(n))}2^{-n\epsilon/2}$.
-Given that $V\in\pazocal{M}^{2n}$, we can take a $\max$ over all $V\in\pazocal{M}^{2n}$, which gives the desired result.
-∎
-Remark 4.16.
+This result holds even for an infinite number of follower types and extends prior work on learning in Stackelberg games without contextual information [letchford2009learning].
+Remark C.9.
 It is worth noting that even when there is only one context (i.e.
-$\pazocal{Z}$ is a singleton), our results in this section are incomparable with those in Section˜4.1.
+$\pazocal{Z}$ is a singleton), our results in this section are incomparable with those in Section˜4.
 This is because in this section, the goal of the leader is to learn given a deterministic mapping from contexts to follower types.
 So, if the contexts are the same between all examples in the training set, the follower types corresponding to these examples will also be the same.
-In contrast, the context-free setting we consider in Section˜4.1 allows for a distribution over follower types.
+In contrast, the context-free setting we consider in Section˜C.1 allows for a distribution over follower types.
 Thus the two bounds are generally incomparable.
-5 Computational Complexity Considerations
+Appendix D Computational complexity considerations
 Computational complexity has been a longstanding concern in the analysis of Stackelberg games.
-In the simplest setting, where a leader interacts with a single known follower type Conitzer and Sandholm [20] have shown that computing the optimal mixed strategy for the leader can be achieved in polynomial time using linear programming.
-However, in more complex Stackelberg games such as Bayesian Stackelberg games, Stackelberg security games, and even 3-player Stackelberg games computing the leader’s optimal strategy has been shown to be NP-hard [20, 28].
+In the simplest setting, where a leader interacts with a single known follower type conitzer2006computing have shown that computing the optimal mixed strategy for the leader can be achieved in polynomial time using linear programming.
+However, in more complex Stackelberg games such as Bayesian Stackelberg games, Stackelberg security games, and even 3-player Stackelberg games computing the leader’s optimal strategy has been shown to be NP-hard [conitzer2006computing, korzhyk10:stackelberg].
 Therefore computational considerations have so far been largely ignored.
 In our setting, there are several interesting cases of structured Stackelberg games for which the leader’s optimal strategy can be learned in polynomial time with provable (albeit not optimal) regret guarantees.
-Observation.
 In a structured Stackelberg game with hypothesis class $\pazocal{H}$ the leader’s optimal policy can be learned in polynomial time if $\pazocal{H}$ is efficiently online learnable.
@@
 The leader can use this algorithm to predict the follower’s type and based on that compute the optimal mixed strategy against that type as if her prediction were correct.
-As shown by Conitzer and Sandholm [20], for a fixed and known follower type, computing the leader’s optimal mixed strategy can be achieved in polynomial time using linear programming.
+As shown by conitzer2006computing, for a fixed and known follower type, computing the leader’s optimal mixed strategy can be achieved in polynomial time using linear programming.
 The learner will make a finite number of follower type mispredictions before finding $h^{*}$, and at this point will play optimally against every follower type that appears in every subsequent round.
@@
 We now present one such example of a hypothesis class that can be learned efficiently when used to predict the followers’ types.
-Example.
-[…]
-Theorem 5.1.
+Example D.1.
+Consider the hypothesis class, $\pazocal{H}_{\text{MDL}}$ of multiclass decision lists.
+Each decision list is a sequence of “if-then” rules, specifically of the form “if $\ell_{1}$ then $k_{1}$, else if $\ell_{2}$ then $k_{2}$, else if $\ell_{3}$ then $k_{3}$, …, else $k_{\text{final}}$”.
+Each decision list $h\in\pazocal{H}$ can be represented by a list $(\ell_{1},k_{1}),\ldots,(\ell_{n},k_{n}),(k_{\text{final}})]$ that consists of condition-value pairs $(\ell_{i},k_{i})$ and a final default value $(k_{\text{final}})$.
+Each condition $\ell_{i}$ is a literal (of either a variable in a set of input variables, $\pazocal{Z}=\{x_{1},\ldots,x_{n}\}$, or its negation) and each value $k_{i}$ belongs to a set of output labels, $[K]$.
+A decision list is evaluated sequentially.
+It takes as input an assignment of variables in $\pazocal{Z}$ i.e.
+$\mathbf{z}\in\{0,1\}^{n}$ and its output is the value corresponding to the first condition that is satisfied by the input, $\mathbf{z}$.
+Theorem D.2.
 The class of multiclass decision lists is efficiently online learnable.
-We present the full proof of Theorem˜5.1 and the corresponding algorithm in Section˜A.5.
-6 Conclusion and Future Directions
-We initiate the study of structured Stackelberg games, where contextual information is predictive of the follower’s type.
-Our main results include a complete characterization of instance-optimal regret bounds in the online setting and PAC sample complexity guarantees in the distributional setting.
-In this section, we discuss concrete open problems and future directions.
-Computationally efficient algorithms.
-An interesting avenue for future research is to consider the computational complexity of structured Stackelberg game algorithms.
-Although we have shown how to achieve optimal regret bounds, our proposed algorithm is not computationally efficient.
-As we discuss in Section˜5, our model opens the door for efficient, yet so far sub-optimal learning algorithms in certain settings.
-It would be interesting to further explore whether the structure introduced in our work can lead to both optimal and efficient algorithms or if there is an underlying tradeoff between computational complexity and tight regret guarantees.
-Regret guarantees for the agnostic setting.
-Our work focuses on structured Stackelberg games in the realizable setting, where the leader is given a class of hypotheses with the promise that it includes the optimal predictor.
-The agnostic learning setting is a natural relaxation of this assumption.
-While we provide some results (cf.
-Section˜B.1.4) for this setting, our bounds are not instance optimal.
-Providing a complete characterization for the agnostic case, analogously to the realizable one we study here, remains an avenue for future research.
-It is plausible that an entirely new complexity measure will be required here, which would show a clear separation between our setting and multiclass classification, where the Littlestone dimension characterizes both the realizable and the agnostic setting.
-Beyond worst-case analysis frameworks.
-Another direction for future research is to analyze structured Stackelberg games in the beyond worst-case analysis framework.
-It has recently been shown that smooth online learning is as easy as distributional learning, with regret bounds scaling with the VC dimension of the hypothesis class.
-Extending these findings to multiclass learning and structured Stackelberg games might lead to improved results both from a statistical and computational point of view.
-Appendix A Details and Omitted Proofs
-A.1 A technical subtlety
-An important component of our analysis is that the leader must be able to compute the optimal strategy for a given context and follower type.
-For this, we state that for any context $\mathbf{z}\in\pazocal{Z}$ the leader incurs negligible regret by restricting herself to a subset $\mathscr{E}_{\mathbf{z}}\subseteq\Delta(\pazocal{A})$ which is finite and computable.
-To formally state our result we present the notion of a contextual best-response region, which was first introduced in [27] and which we repeat here for completeness.
-Definition A.1 (Contextual Follower Best-Response Region).
-For context $\mathbf{z}$, follower type $f^{(i)}$, and follower action $a_{f}\in\pazocal{A}_{f}$, let $X_{\mathbf{z}}(f^{(i)},a_{f})\subseteq\pazocal{X}$ denote the leader’s mixed strategies such that a follower of type $f^{(i)}$ best-responds to all $x\in X_{\mathbf{z}}(f^{(i)},a_{f})$ by playing action $a_{f}$, that is, $X_{\mathbf{z}}(f^{(i)},a_{f})=\{\mathbf{x}\in\pazocal{X}:b_{f^{(i)}}(\mathbf{z},\mathbf{x})=a_{f}\}$
-Notice that any possible follower best-response region is (1) a subset of $\Delta(\pazocal{A})$ and (2) the intersection of finitely-many halfspaces.
-Therefore, every $X_{\mathbf{z}}(f^{(i)},a_{f})$ is convex and bounded.
-However, the leader may not be able to optimize over the extreme points of a best-response region, as it may not be closed due to the followers’ tie-breaking.
-Definition A.2 ($\delta$-approximate extreme points).
-[…]
-Notice that for sufficiently small $\delta$ i.e.
-$\delta=\mathscr{O}(\frac{1}{T})$, we get that the additional regret of the leader by choosing actions over $\mathscr{E}^{(i)}$ rather than $X(f^{(i)},a_{f})$ is negligible.
-Therefore, we use the shorthand $\mathscr{E}^{(i)}:=\mathscr{E}^{(i)}(\delta)$ for the rest of the section.
-The following lemma from [27] formalizes this notion.
-Lemma A.3.
-For any sequence of followers $f_{1},\ldots f_{T}$ and any leader policy $\pi$, there exists a policy $\pi^{(\mathscr{E})}:\pazocal{Z}\rightarrow\cup_{\mathbf{z}\in\pazocal{Z}}\mathscr{E}_{\mathbf{z}}$ that, when given context $\mathbf{z}$, plays a mixed strategy in $\mathscr{E}_{\mathbf{z}}$ and guarantees that $\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\pi^{(\mathscr{E})}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi^{(\mathscr{E})}(\mathbf{z}_{t})))\leq 1$.
-Moreover, the same result holds in expectation over any distribution over follower types.
-Given Lemma˜A.3 and for ease of notation, in this paper we assume that the leader will always set the parameter $\delta$, such that it does not affect our regret bounds.
-A.2 Proof of Theorem˜3.5
-See 3.5
 Proof.
-Consider the setting where the context space is $\pazocal{Z}=\{1,2\}\times[0,1]$ and we have 4 follower types $\{f^{(1)},f^{(2)},f^{(3)},f^{(4)}\}$.
-Follower types $f^{(1)}$ and $f^{(2)}$ have fixed utility functions $u_{f^{(1)}},u_{f^{(2)}}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$ such that for all $\mathbf{z}\in\pazocal{Z}$, we have that $\pi_{*}^{(1)}(\mathbf{z})\neq\pi_{*}^{(2)}(\mathbf{z})$, where $\pi_{*}^{(i)}$, refers to the optimal policy against follower $f^{(i)}$.
-Let follower type $f^{(3)}$ have utility function $u_{f^{(3)}}(\mathbf{z},a_{l},a_{f})=\mathbbm{1}\{\mathbf{z}[1]=1\}\cdot u_{f^{(1)}}(\mathbf{z},a_{l},a_{f})+\mathbbm{1}\{\mathbf{z}[1]=2\}\cdot u_{f^{(2)}}(\mathbf{z},a_{l},a_{f})$ and follower type $f^{(4)}$ have utility function $u_{f^{(4)}}(\mathbf{z},a_{l},a_{f})=\mathbbm{1}\{\mathbf{z}[1]=2\}\cdot u_{f^{(1)}}(\mathbf{z},a_{l},a_{f})+\mathbbm{1}\{\mathbf{z}[1]=1\}\cdot u_{f^{(2)}}(\mathbf{z},a_{l},a_{f})$.
-Concretely, one can consider followers with the following utility functions.
-The leader in this case is the row player and the follower is the column player.
-$F_{1}$ $F_{2}$
-$L_{1}$ 1,0.51,0.51 , 0.5 0.5,00.5,00.5 , 0
-$L_{2}$ 0.25,10.25,10.25 , 1 0.75,00.75,00.75 , 0
-(a) Follower type $f^{(1)}$
-$F_{1}$ $F_{2}$
-$L_{1}$ 1,01,01 , 0 $,0.5$ 0.5
-$L_{2}$ 0.25,00.25,00.25 , 0 0.75,10.75,10.75 , 1
-(b) Follower type $f^{(2)}$
-$F_{1}$ $F_{2}$
-$L_{1}$ $1,|\mathbf{z}[1]-2|$ $,0.5$ 0.5
-$L_{2}$ 0.25,0.50.25,0.50.25 , 0.5 $0.75,|\mathbf{z}[1]-1|$
-(c) Follower type $f^{(3)}$
-$F_{1}$ $F_{2}$
-$L_{1}$ $1,|\mathbf{z}[1]-1|$ $,0.5$ 0.5
-$L_{2}$ 0.25,0.50.25,0.50.25 , 0.5 $0.75,|\mathbf{z}[1]-2|$
-(d) Follower type $f^{(4)}$
-Table 3: Utility tables for Theorem˜3.5
-Let
-$h_{\gamma_{1},\gamma_{2}}(\mathbf{z})=\begin{cases}f^{(1)}\;\;&\text{ if }0\leq\mathbf{z}[2]<\gamma_{1}\\ f^{(2)}\;\;&\text{ if }\gamma_{2}<\mathbf{z}[2]\leq 1\\ f^{(3)}\;\;&\text{ if }(\mathbf{z}[1]=1\text{ and }\gamma_{1}\leq\mathbf{z}[2]<0.5)\text{ or }(\mathbf{z}[1]=2\text{ and }\gamma_{2}\leq\mathbf{z}[2]\leq 0.5)\\ f^{(4)}\;\;&\text{ if }(\mathbf{z}[1]=2\text{ and }\gamma_{1}\leq\mathbf{z}[2]<0.5)\text{ or }(\mathbf{z}[1]=1\text{ and }\gamma_{2}\leq\mathbf{z}[2]\leq 0.5)\\ \end{cases}$
-[…]
-Figure 3: Follower type as a function of the context in the construction of Theorem˜3.5.
-While $\gamma_{1}$ and $\gamma_{2}$ vary with each hypothesis, the leader’s optimal strategy only depends on whether $\mathbf{z}[2]<0.5$.
-A.3 Alternate definition of the Stackelberg Littlestone dimension
-In this section we provide an alternate definition of the Stackelberg Littlestone dimension that can be defined for infinite trees.
-In Theorem˜A.6 we prove that for finite trees, this is equivalent to Definition˜3.7.
-Definition A.4 ($\kappa$-shattered edge-weighted tree).
-[…]
-  1. 1.
-    (Consistency) $\pazocal{H}$ is consistent with every path in $\pazocal{T}_{d}$ i.e., for any $s\in S_{d}$, there exists $h\in\pazocal{H}$ such that for all $i\in\{\varnothing,1,\dots,d-1\}$ the label of edge $(\mathbf{z}_{s[:i]},\mathbf{z}_{s[:i+1]})$ is $h(\mathbf{z}_{s[:i]})$.
-  2. 2.
-    (Max-branch property) $\pazocal{T}_{d}$ satisfies the max-branch property, i.e., every node, $\mathbf{z}_{s}$, has children $\mathbf{z}_{si}$ for all $i\in[K]$ where $\pazocal{H}$ is consistent with path $si$.
-  3. 3.
-    (Weights) for each node, $\mathbf{z}_{s}\in\pazocal{Z}$ with children $\{\mathbf{z}_{si}\}_{i\in M\subseteq[K]}$ the set of node-to-child edge weights, $\{\nu_{s}^{(i)}\}_{i\in M\subseteq[K]}$, is such that for every $\mathbf{x}\in\Delta(\pazocal{A})$ there exists a follower type $i$ such that $r(\mathbf{z},\mathbf{x},f^{(i)})\geq\nu_{s}^{(i)}$.
-  4. 4.
-    $\inf_{s\in S_{d}}\biggl{[}\sum_{i=0}^{d}\nu_{s[:i]}^{(s[i+1])}\biggr{]}=\kappa$
-Definition A.5 (Stackelberg-Littlestone dimension).
-The Stackelberg-Littlestone dimension of a hypothesis class $\mathscr{H}$ with respect to a Stackelberg game, $\pazocal{G}$, denoted as $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$, is the supremum value of $\kappa$ among all edge-weighted finite trees that are $\kappa$-shattered by hypothesis class $\pazocal{H}$ under Stackelberg game $\pazocal{G}$.
-Theorem A.6.
-Definition˜3.7 and Definition˜A.5 are equivalent for finite trees.
-Proof.
-It suffices to show that any fixed finite-depth tree can be $\kappa$-shattered under Definition˜A.4 if and only if it can be shattered under Definition˜3.6 with root node weight $\kappa$.
-We proceed by induction on the depth of shattered trees.
-(Base Case) Consider a tree, $\pazocal{T}_{1}$, of depth 111 with root node $\mathbf{z}_{\varnothing}$.
-We will argue that the root node weight specified by Definition˜3.6 is equal to the minimum edge weight of the maximum set of valid edge weights under Definition˜A.4 and Definition˜A.5.
-It suffices to show that
-$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})=\max_{N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}}\min_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\{\nu_{\varnothing}^{(i)}:\nu_{\varnothing}^{(i)}\in N\}$
-where $\pazocal{N}_{\mathbf{z}_{\varnothing}}$ is the collection of all sets $N=\{\nu_{\varnothing}^{(i)}\}_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}$ with the property that for all $\mathbf{x}\in\Delta(\pazocal{A})$, there exists an $i\in\pazocal{H}(\mathbf{z}_{\varnothing})$ such that $r(\mathbf{z},\mathbf{x},f^{(i)})\geq\nu_{\varnothing}^{(i)}$.
-We will first show that $\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\max_{N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}}\min_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\{\nu_{\varnothing}^{(i)}:\nu_{\varnothing}^{(i)}\in N\}$.
-Consider any set of valid edge-weights $N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}$.
-By definition we know that for every $\mathbf{x}\in\Delta(\pazocal{A})$, $\exists i\in\pazocal{H}(\mathbf{z}_{\varnothing})$ such that $r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\nu_{\varnothing}^{(i)}$.
-This means that
-$\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\min_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\{\nu_{\varnothing}^{(i)}:\nu_{\varnothing}^{(i)}\in N\}\;\;\;\forall\mathbf{x}\in\Delta(\pazocal{A})\;\;\forall N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}.$
-Considering the $\mathbf{x}\in\Delta(\pazocal{A})$ that minimizes the LHS and an $N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}$ that maximizes the RHS, we get
-$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\max_{N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}}\min_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\{\nu_{\varnothing}^{(i)}:\nu_{\varnothing}^{(i)}\in N\}.$
-We will now show that $\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\leq\max_{N\in\pazocal{N}_{\mathbf{z}_{\varnothing}}}\min_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\{\nu_{\varnothing}^{(i)}:\nu_{\varnothing}^{(i)}\in N\}$.
-Let us denote $\mathbf{x}^{*}=\arg\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})$, and $i^{*}=\arg\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing}})r(\mathbf{z}_{\varnothing},\mathbf{x}^{*},f^{(i)})$.
-Let $N^{*}=\{\nu_{\varnothing}^{(i)}:=r(\mathbf{z}_{\varnothing},x^{*},f^{(i^{*})})\}_{i\in[K]}$.
-We argue that $N^{*}\in\pazocal{N}_{\mathbf{z}_{\varnothing}}$, i.e., that
-$\forall\mathbf{x}\in\Delta(\pazocal{A})\;\;\exists i\in\pazocal{H}(\mathbf{z}_{\varnothing})\;\;\text{such that}\;\;r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq r(\mathbf{z},\mathbf{x}^{*},f^{(i^{*})})$
-and it suffices to argue that $\forall\mathbf{x}\in\Delta(\pazocal{A})$ if $i_{\mathbf{x}}=\textnormal{argmax}_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})$, then $\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})\geq\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})$, which holds trivially.
-(Inductive Step) Say that the two dimensions are equivalent for any tree of depth $d-1$.
-We will show that they are also equivalent for any tree of depth $d$.
-Consider such a tree, $\pazocal{T}_{d}$, with root node weight $\mathbf{z}_{\varnothing}$ and children $\{\mathbf{z}_{0i}\}_{i\in[K]}$.
-We want to show that the root node weight $\rho_{\varnothing}$ defined as in Definition˜3.6 is equal to the weight of the minimum path in the tree, where the edge weights are defined as in Definition˜A.4.
-Concretely, we want to show that
-$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in[K]}\biggl{[}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(j)})+\rho_{\varnothing j}\biggr{]}$ $=\sup_{\pazocal{N}_{d}}\inf_{s\in S_{d}}\Biggl{[}\sum_{i=0}^{d}\nu_{s[:i]}^{(s[i+1])}\Biggr{]},$
-where $\pazocal{N}_{d}$ is the collection of sets of all possible edge weights, one for each node in the tree of depth $d$. This holds if and only if
-$\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in[K]}\biggl{[}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(j)})+\rho_{\varnothing j}\biggr{]}$ $=\max_{\pazocal{N}_{0}}\min_{i\in[K]}\Biggl{[}\nu_{\varnothing}^{(s[1])}+\max_{\pazocal{N}_{d-1}}\inf_{s\in S_{d-1}}\sum_{i=1}^{d}\nu_{s[:i]}^{(s[i+1])}\Biggr{]}$
-We know that for any $j\in[K]$ and $s\in S_{d}$ such that $s[1]=j$, $\rho_{\varnothing j}=\max_{\pazocal{N}_{d-1}}\inf_{s\in S_{d-1}}\sum_{i=1}^{d}\nu_{s[:i]}^{(s[i+1])}$ by the inductive hypothesis.
-Therefore, using an argument similar to the base case, the equation above holds.
+Consider a true multiclass decision list $h^{*}:\{0,1\}^{n}\to[K]$.
+We show that Algorithm˜3 makes at most $O(nKL)$ mistakes in the worst case, where $n=|\pazocal{Z}|$ is the number of variables, $L$ is the length of the target decision list and $K$ is the number of possible labels.
+We make the following three observations: (1) At every timestep in which the algorithm makes a mistake, at least one rule is moved down by one level.
+(2) Any rule will be moved down a level at most $L+1$ times, where $L$ is the length of the target decision list.
+(3) By induction, the $i$-th rule of the target hypothesis will never be moved below the $i$-th level of $\tilde{h}$.
+Therefore the number of mistake that the algorithm makes is upper bounded the by $(L+1)(2nK+2)$, where $2nK+2$ is the number of all possible rules.
 ∎
-A.4 Proof of Lemma˜3.10
-See 3.10
-Proof.
-We assume for the sake of contradiction that there exists a hypothesis class with $\textnormal{Ldim}(\pazocal{H})=d$ and a Stackelberg game $\pazocal{G}$, such that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})>d$.
-According to Definition˜A.5, this implies that there exists a shattered tree, whose minimum-weighted root-to-leaf path has weight strictly greater than $d$.
-This implies that all paths in the tree have cumulative weight strictly greater than $d$.
-Given that the weight of each edge in the tree is at most 111, the length of the minimum path (and therefore of all other paths in the tree) must be of length at least $d+1$.
-This implies that there exists a tree of depth $d+1$ that is shattered by $\pazocal{H}$, which leads to a contradiction.
-∎
-A.5 Proof of Theorem˜5.1
-Algorithm 3 Multiclass Decision Lists Learning Algorithm
-[…]
+1:input: input variable space, $\pazocal{Z}$, and output label space, $[K]$.
 2: Define the set of rules $R=\{(\ell_{i},k_{i})\;\text{such that}\;\ell_{i}\in\{x_{i},\bar{x}_{i}\},x_{i}\in\pazocal{Z},\;k_{i}\in[K]\}$
@@
 14:     Break
-15:   else
-16:     Continue
-17:end
-See 5.1
-Proof.
-[…]
-Appendix B Optimal Strategy Misprediction Bounds
-In the main body of our work we include examples that highlight that existing tools and techniques from multiclass prediction would yield loose regret bounds in our setting (cf.
-proofs of Theorems˜3.5 and 3.11).
-A common theme that emerges from these is that for two distinct follower types, $i,j\in[K]$, the optimal policies $\pi_{*}^{(i)},\pi_{*}^{(j)}:\pazocal{Z}\to\Delta(\pazocal{A})$ for each follower type respectively might coincide for a subset of the context space, that is, they yield the same optimal strategy for the leader.
-Naturally, this prompts us to consider the goal of minimizing the number of timesteps in which the leader mispredicts the optimal strategy.
-Would we get tighter regret bounds if the leader played optimally against this benchmark?
-We answer this in the affirmative, by giving a full characterization of learning in both the online and distributional learning settings.
-While our approaches here perform better than using off-the-shelf classification algorithms, the results we present in the main body strictly supersede them.
-We find this to be for two reasons: (1) what counts as a “misprediction” of the optimal strategy is sensitive to the leader’s tie-breaking rule, and (2) among the suboptimal strategies, some yield higher utility than others, yet each misprediction counts as either adding loss of 1 or 0 to the total.
-B.1 Online Learning
-In this section, we present no-regret learning algorithms for online learning in structured Stackelberg games, that are designed to minimize the number of optimal strategy mispredictions that the leader makes in a time horizon $T$.
-In general, if the regret of an algorithm is sublinear in the time horizon $T$, i.e., $R(T)=o(T)$, we say that the algorithm is a no-regret algorithm.
-We first introduce the notion of a Mistake Stackelberg Littlestone- (MSL-)Tree, which is a natural extension of the Littlestone tree to our setting.
-Definition B.1 ($\pazocal{H}$-Shattered MSL Tree).
-For a Stackelberg game, $\pazocal{G}$, let a MSL tree, $T$, be a tree whose internal nodes are instances in $\pazocal{Z}$ and where for each node, the node-to-child edges correspond to different vectors in $\Delta(\pazocal{A})$.
-We say that $T$ is shattered by a hypothesis class $\pazocal{H}$ with respect to $\pazocal{G}$ if for every root-to-leaf path in the tree that traverses the nodes $\mathbf{z}_{1},\ldots,\mathbf{z}_{n}$, there exists a hypothesis $h\in\mathscr{H}$, such that the label of every edge $(\mathbf{z}_{i},\mathbf{z}_{i+1})$ is $\mathbf{x}_{i}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{i},\mathbf{x},b_{h(\mathbf{z}_{i})}(\mathbf{z}_{i},\mathbf{x}))$.
-In the case of ties, we assume that the leader follows a fixed and arbitrary tie-breaking rule.
-Equipped with this definition, we are now ready to introduce the MSL dimension, which characterizes the complexity of online learning when the leader only cares about minimizing the number of mistakes.
-Definition B.2 (MSL Dimension).
-The MSL dimension of a hypothesis class $\mathscr{H}$ with respect to a Stackelberg game, $\pazocal{G}$, denoted as $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$, is the maximal integer $d$, such that there exists a full binary MSL tree of depth $d$ that is shattered by $\pazocal{H}$.
-B.1.1 Lower bound
-We begin by showing that for any online Stackelberg game $\mathscr{G}$, no algorithm can make fewer than $\Omega(\textnormal{MSLdim}_{\pazocal{G}}(\mathscr{H}))$ mistakes.
-Theorem B.3.
-There exists a Stackelberg game, $\pazocal{G}$, such that for any hypothesis class $\pazocal{H}$ any deterministic algorithm will make at least $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes.
-Furthermore, any randomized algorithm will make at least $\frac{1}{2}\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes in expectation.
-Proof.
-Consider a Stackelberg game, in which for any fixed context, playing the optimal strategy against the true follower type yields utility 111, whereas playing any other strategy yields utility 0.
-In other words, the leader gets utility 1 if and only if she correctly predicts the follower type and plays optimally against them.
-Concretely, consider the following game.
-Let $\beta_{1},\beta_{2},\ldots,\beta_{K}$ be distinct, positive real numbers and consider the following payoff matrix, where the leader is the row player and the column player is follower type $f^{(i)}$.
-The leader has 2 pure actions and the follower has 3 pure actions.
-Let $\epsilon\in(0,\min_{i}\beta_{i})$ be a small real number.
-Notice that if the leader puts probability $p>\beta_{i}+\epsilon$ on action $L_{1}$, then the follower will play action $F_{1}$ and the leader gets payoff 0.
-Similarly, if the leader puts probability $p<\beta_{i}-\epsilon$ on action $L_{1}$ then the follower plays action $F_{3}$ and the leader gets payoff 0.
-To guarantee that the leader gets utility 1, the leader must play action $L_{1}$ with probability $p\in(\beta_{i}-\epsilon,\beta_{i}+\epsilon)$.
-Then it is always beneficial for the follower to play $F_{2}$ (this holds no matter how the follower tie-breaks).
-So as long as the intervals $(\beta_{1}-\epsilon,\beta_{1}+\epsilon),(\beta_{2}-\epsilon,\beta_{2}+\epsilon),\ldots,(\beta_{K}-\epsilon,\beta_{K}+\epsilon)$ do not overlap, the leader will get utility 1 if and only if she correctly predicts the follower type she is playing against and plays optimally against them.
-$F_{1}$ $F_{2}$ $F_{3}$
-$L_{1}$ $0,\frac{1}{\beta_{i}+\epsilon}$ $,1$ 1 $,0$ 0
-$L_{2}$ $,0$ 0 $,1$ 1 $0,\frac{1}{1-\beta_{i}+\epsilon}$
-Table 4: Game construction for Theorem˜B.3
-(Claim 1.) An adversary can force the learner to make $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes by using the following strategy: (1) select $\mathbf{z}_{t}$ as the next input from an MSL-Tree of depth $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ that is shattered by $\pazocal{H}$ and (2) set the optimal mixed strategy, $\mathbf{x}_{t}^{*}$ to be the edge in the shattered tree not chosen by the leader, that is $\mathbf{x}_{t}^{*}\neq\widehat{\mathbf{x}}_{t}$, where $\widehat{\mathbf{x}}_{t}$ is the strategy played by the learner at time $t$.
-Since the tree is shattered by $\pazocal{H}$, there must exist some $h\in\pazocal{H}$ that is consistent with $(\mathbf{z}_{t},f_{t})_{t=1}^{\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})}$.
-Therefore, any deterministic learner will make at least $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes.
-(Claim 2.) When the learner is randomized, at each timestep $t$, the adversary can select an $\mathbf{x}_{t}^{*}$ such that it minimizes the probability that the leader chooses $\widehat{\mathbf{x}}_{t}=\mathbf{x}_{t}^{*}$.
-In this case, the probability that the learner will make a mistake is at least 0.5, and equality holds when the leader selects uniformly among two branches of the tree.
-Summing over each layer of the tree gives us the lower bound on the learner’s expected number of mistakes.
-∎
-B.1.2 Upper bounds
-Having established that the MSL dimension lower-bounds the number of mispredictions of the optimal strategy for any learner, in this section we present an algorithm and show that it achieves the optimal mistake bound.
-Warm-up: context-free utilities.
-As a warm-up, consider a simpler setting, in which both the leader and the follower utility functions do not take the context as an input (i.e.
-they are context-free).
-This means that the context at each timestep is now only predictive of the follower’s type at that timestep.
-In Theorem˜B.4, we show that we can reduce the problem of learning the optimal policy of the leader to online multiclass prediction.
-Additionally, the known optimal algorithm in online multiclass prediction (the Standard Optimal Algorithm (SOA)) meets the lower bound shown in Theorem˜B.3.
-Theorem B.4.
-Consider a Stackelberg game, $\pazocal{G}$, such that the utilities of both the leader and the follower are context-free.
-Then for any finite class, $\pazocal{H}$, of hypotheses $h:\pazocal{Z}\to[K]$, there exists a reduction of this setting to multiclass prediction that runs in $O(\text{poly}(K,|\pazocal{A}_{f}|,|\pazocal{A}|,|\pazocal{H}|))$ time and makes at most $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes.
-Proof.
-The reduction is as follows.
-First, compute the optimal strategy against each type of follower offline.
-This can be achieved using linear programming and requires $O(\text{poly}(K,|\pazocal{A}_{f}|,|\pazocal{A}|))$ runtime [20].
-Then create a (possibly many-to-one) look-up table mapping each follower type $f^{(i)}$ to an optimal strategy, $\mathbf{x}_{*}^{(i)}$, for the leader against that follower.
-Compute $\pazocal{H}^{\prime}=\{h^{\prime}(\mathbf{z}):=\sum_{i=1}^{K}\mathbf{x}_{*}^{(i)}\cdot\mathbbm{1}\{h(\mathbf{z})=f^{(i)}\}\}_{h\in\pazocal{H}}$, which takes at most $O(\text{poly}(|\pazocal{H}|))$ time.
-Finally, we run the multiclass Standard Optimal Algorithm (see [22]) on $\pazocal{H}^{\prime}$ considering every optimal strategy for the leader in the look-up table as one label.
-It remains to show that $\textnormal{Ldim}(\pazocal{H}^{\prime})=\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$.
-Notice that $\pazocal{H}^{\prime}$ is the set of all hypotheses, mapping each context to an optimal strategy for the leader against every follower type that is consistent with $\pazocal{H}$ for that context.
-Consider any MSL tree and any Littlestone tree such that the context at each corresponding node is the same among them.
-If the MSL tree is shattered by $\pazocal{H}$ then for every root-to-leaf path there must exist a hypothesis in $\pazocal{H}$ that is consistent with the labels along that path.
-But by definition of $\pazocal{H}^{\prime}$ for that corresponding path in the Littlestone tree there must exist an $h^{\prime}\in\pazocal{H}^{\prime}$ that is consistent with the labelings of that path.
-With the same argument, the converse is also true.
-Therefore, the MSL tree with maximal depth that is shattered by $\pazocal{H}$ is precisely the generalized Littlestone tree with maximal depth that is shattered by $\pazocal{H}^{\prime}$.
-∎
-We note that the algorithm described above coincides with Algorithm˜4 (presented below), which we prove to be optimal in the general contextual Stackelberg game setting.
-Context-dependent utilities.
-We now let the utilities of both players to be functions of the context.
-We present our main algorithm (Algorithm˜4), which we show achieves regret that matches the lower bound in Theorem˜B.3 for any Stackelberg game.
-Our approach is inspired by the Standard Optimal Algorithm for online multiclass prediction, with the main difference that our algorithm uses a hypothesis class, $\pazocal{H}$, to optimize directly over leader strategies in the simplex.
-Intuitively, our algorithm keeps track of all the “surviving” hypotheses in $\pazocal{H}$, that is, the hypotheses that are consistent with the data seen so far.
-At each timestep the algorithm observes the context and considers the predicted follower types of all surviving hypotheses.
-For every such follower type the algorithm computes the optimal leader strategy against them.
-Then it plays the strategy that maximizes the MSL dimension of the set of surviving hypotheses were that to be the optimal strategy for that timestep.
-If the algorithm makes a mistake, then the set of surviving hypotheses will have strictly smaller MSL dimension than before.
-In other words, the algorithm learns from this mistake.
-Notice that the number of possible strategies in the simplex is infinite, yet our algorithm only considers a bounded number of them at each timestep, at most one for each follower type.
-Algorithm 4 Mistake-Stackelberg Standard Optimal Algorithm (MSSOA)
-1:Input: time horizon, $T$, Stackelberg game $\pazocal{G}$, and hypothesis class $\pazocal{H}$, where $h:\pazocal{Z}\to[K]$ for every $h\in\pazocal{H}$.
-2:Set $V_{0}=\pazocal{H}$
-3:for $t\in[T]$ do
-4:  Observe $\mathbf{z}_{t}$
-5:  Let $\mathbf{x}^{(h,t)}=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{h(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))\quad\forall h\in\pazocal{H}$
-6:  Let $X_{t}=\{\mathbf{x}^{(h)}:h\in V_{t-1}\}$triangleright Note that $|X_{t}|\leq K$
-7:  Let $V_{t}^{(\mathbf{x})}=\{h\in V_{t-1}:\mathbf{x}^{(h,t)}=\mathbf{x}\}\quad\forall\mathbf{x}\in X_{t}$
-8:  Play action $\mathbf{x}_{t}=\arg\max_{\mathbf{x}\in X_{t}}\textnormal{MSLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})$
-9:  Observe $f_{t}$
-10:  Set $V_{t}=\{h\in V_{t-1}:h(\mathbf{z}_{t})=f_{t}\}$
-Theorem B.5.
-For any Stackelberg game, $\pazocal{G}$ and every hypothesis class, $\pazocal{H}$, Algorithm˜4 makes at most $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes.
-Proof.
-Let $T_{e}\subseteq[T]$ be the set of time steps on which MSSOA errs and let $T_{c}=[T]\setminus T_{e}$ be the set of time steps on which MSSOA correctly predicts the optimal strategy against the true follower type.
-Recall from Remark˜3.2 that we can make $\sum_{t\in T_{c}}u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$ arbitrarily small.
-Now from the boundedness of the utility function, we also know that $\sum_{t\in T_{e}}u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))\leq|T_{e}|$.
-It suffices to show that for every time step $t\in T_{e}$ that MSSOA makes a mistake, we have that $\textnormal{MSLdim}_{\pazocal{G}}(V_{t+1})\leq\textnormal{MSLdim}_{\pazocal{G}}(V_{t})-1$.
-For the sake of contradiction, assume that $\textnormal{MSLdim}_{\pazocal{G}}(V_{t})=\textnormal{MSLdim}_{\pazocal{G}}(V_{t+1})$.
-Given that MSSOA chooses a strategy $\mathbf{x}$ that maximizes $\textnormal{MSLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})$ and that MSSOA made a mistake at this timestep, there must have been at least two distinct strategies $\mathbf{x},\mathbf{x}^{\prime}$ such that $\textnormal{MSLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})=\textnormal{MSLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x}^{\prime})})=\textnormal{MSLdim}_{\pazocal{G}}(V_{t})$.
-However, this would mean that we could construct a full binary MSL tree that is shattered by $\pazocal{H}$ and that has depth $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})+1$, which leads to a contradiction.
-B.1.3 Using multiclass SOA as a black-box algorithm
-We consider the following algorithm that uses SOA [22] to predict the follower’s type given the context.
-Then proceeds to play the optimal strategy against that follower type as if the prediction were correct.
-Algorithm 5 Learning with full-feedback and realizable follower types
-1:for $t\in[T]$ do
-2:  Observe $\mathbf{z}_{t}$
-3:  Evaluate $\hat{f}_{t}=h_{t}(\mathbf{z}_{t})\leftarrow\text{SOA}(\mathbf{z}_{t})$
-4:  Commit to $\mathbf{x}_{t}=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{\hat{f}_{t}}(\mathbf{z}_{t},\mathbf{x}))$
-5:  Observe follower type $f_{t}$
-6:end
-Lemma B.6.
-There exists a hypothesis class $\pazocal{H}$ and a Stackelberg game $\pazocal{G}$ such that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})<\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})<\textnormal{Ldim}(\pazocal{H})$.
-Further, SOA achieves strictly larger regret than MSSOA, which in turn achieves strictly larger regret than SSOA.
-Proof.
-We revisit the game presented in the proof of Theorem˜3.11.
-We have shown that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=3/4$ and $\textnormal{Ldim}(\pazocal{H})=2$.
-We will now show that $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})=1$.
-We consider any possible MSL tree.
-First, note that the context $z=4$ cannot be a node in an $\pazocal{H}$-shattered MSL-Tree since for $z=4,\;\pi_{*}^{(2)}(4)=\pi_{*}^{(3)}(4)$, where $f^{(2)}$ and $f^{(3)}$ are the only possible follower types predicted by hypotheses in $\pazocal{H}$ for this context.
-Similarly, $z=3$ cannot be a node in an $\pazocal{H}$-shattered SL tree, since the leader’s optimal strategy against all follower types is identical ($\pi_{*}^{(1)}(3)=\pi_{*}^{(2)}(3)=\pi_{*}^{(3)}(3)$).
-So it suffices to consider contexts 111 and 222.
-In Figure˜4 we show the two $\pazocal{H}$-shattered MSL tree constructions with $z=1$ and $z=2$ as the root nodes respectively.
-Notice that the maximal depth of a full binary tree is 111, as desired.
-We now prove the regret values.
-First, since $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ tightly characterizes the regret of SSOA, we know that the regret of SSOA is 3/43/43 / 4.
-Also, from Theorem˜3.11 we immediately get that the cumulative regret of SOA equals 2 when the sequence of contexts is 111, followed by 222.
-We now proceed with the regret of MSSOA for the same sequence.
-  • ($t=1$) This root node has two branches: $\pi_{*}^{(1)}(1)=\pi_{*}^{(2)}(1)$ and $\pi_{*}^{(3)}(1)$.
-Given the arguments used to prove that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=1$, the former branch leads to a full binary subtree of depth 111, while the second branch cannot lead to a full binary tree.
-So SSOA chooses $x_{1}=\pi_{*}^{(1)}(1)=\pi_{*}^{(2)}(1)$ and the true follower type is $f_{1}^{*}=f^{(1)}$ (no instantaneous regret).
-  • ($t=2$) The context $z=2$ leads to two branches: $\pi_{*}^{(2)}(2)$ and $\pi_{*}^{(3)}(2)$, none of which leads to a complete binary tree.
-SSOA tie breaks, in which case an adversary can always force a mistake by choosing the correct follower type to be the one not predicted by MSSOA (thereby incurring instantaneous regret 1).
-Figure 4: $\pazocal{H}$-shattered MSL-Tree with (a) $z=1$ and (b) $z=2$ as the root node.
-Theorem B.7.
-For every online contextual Stackelberg game with $K$ distinct follower types and every hypothesis class $\pazocal{H}$, there exists a hypothesis class $\Pi$ such that running multiclass SOA on $\Pi$ incurs $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ mistakes.
-Moreover, this algorithm coincides with MSSOA.
-Proof.
-Consider a mapping $\zeta:\pazocal{H}\to\Pi$ that maps every $h\in\pazocal{H}$ to a function $\pi_{h}:\pazocal{Z}\to\Delta(\pazocal{A})$ such that $\pi_{h}(\mathbf{z}):=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x})).$
-Notice that $\zeta$ is a bijection given that each follower type is distinct (cf.
-Remark˜2.1).
-Also note that given that learning $\pazocal{H}$ is a realizable problem, so is learning $\Pi$.
-This is because if there exists $h^{*}\in\pazocal{H}$ such that $\forall t\in[T]\quad h^{*}(\mathbf{z}_{t})=f_{t}$, then by definition there must exist $\pi^{*}\in\Pi$ such that $\forall t\in[T]\quad\pi^{*}(\mathbf{z}_{t})=\mathbf{x}^{*}_{t}$, since $\mathbf{x}^{*}_{t}=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{h^{*}(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))$.
-We know that multiclass SOA makes at most $\textnormal{Ldim}(\Pi)$ mistakes for any hypothesis class $\Pi$ with infinite labels (cf.
-Theorem 1 in [26]).
-Therefore, it remains to show that $\textnormal{Ldim}(\Pi)=\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$.
-Consider any MSL-Tree for $\pazocal{H}$.
-Notice that all nodes are contexts and all edges represent optimal leader strategies, so this is a valid Littlestone tree for $\Pi$ as well.
-Also, for any root-to-leaf path on this tree that is consistent with a function $h\in\pazocal{H}$, there exists $\pi_{h}\in\Pi$ that is consistent with it as well, by definition of $\pi_{h}$.
-The converse is also true: in any Littlestone tree for $\Pi$ consider any root-to-leaf path, that is consistent with some $\pi_{h}\in\Pi$.
-Then by $\zeta$ being a bijection, there exists an $h\in\pazocal{H}$ that is also consistent with it.
-Therefore, a tree is MSL-shattered by $\pazocal{H}$ if and only if it is shattered by $\Pi$.
-This holds for all trees, so it will hold for the maximal-depth trees as well.
-B.1.4 No-regret learning in the agnostic setting
-In this section we relax the realizability assumption, and we consider the setting where the sequence of follower types $f_{1},\cdots,f_{T}$ is not necessarily consistent with some hypothesis $h\in\mathscr{H}$.
-Unlike the realizable setting, here we will not have an absolute benchmark such as the mistake bound.
-Instead, we measure the performance of our algorithms relative to the best in hindsight hypothesis in the class.
-We define $h^{*}\in\arg\max_{h\in\mathscr{H}}\sum_{t\in[T]}\mathbbm{1}\{h(\mathbf{z}_{t})=y_{t}\}$ and generalize the notion of regret:
-Definition B.8 ((Agnostic Setting) Contextual Stackelberg Regret).
-Given a sequence of followers $f_{1},\dots,f_{T}$ and contexts $\mathbf{z}_{1},\dots\mathbf{z}_{T}$, the leader’s contextual Stackelberg regret in the agnostic setting is given by:
-$R(T):=\max_{h\in\mathscr{H}}\sum_{t\in[T]}u(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t})),$
-where $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$ is the sequence of mixed strategies played and $\pi_{h}(\mathbf{z}):=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x}))$.
-Algorithmic procedure.
-Consider the “learning with expert advice” setting where the leader at each timestep $t$ would get to observe not only the context $\mathbf{z}_{t}$ but also $N$ advices, $\{\phi_{t}^{(1)},\cdots,\phi_{t}^{(N)}\}\in\Delta(\pazocal{A})^{N}$.
-Then we can denote the cumulative loss of expert $i$ at timestep $T$ as:
-$L_{T,i}=-\sum_{t=1}^{T}u(\mathbf{z}_{t},\phi_{t}^{(i)},b_{f_{t}}(\mathbf{z}_{t},\phi_{t}^{(i)})).$
-For an appropriately defined set of experts we can formalize the following algorithm.
-Algorithm 6 Learning in the full-feedback, agnostic setting
-1:input: time horizon $T$
-2:Set $\eta=\sqrt{8\ln(N)/T}$
-3:for $t\in[T]$ do
-4:  Observe $\mathbf{z}_{t}$
-5:  Receive advices $\{\phi_{t}^{(1)},\cdots,\phi_{t}^{(N)}\}\in\Delta(A)^{N}$
-6:  Commit to $\mathbf{x}_{t}=\phi_{t}^{(i)}$ with probability proportional to $\exp(-\eta L_{t-1,i})$
-7:  Receive utility $u(\mathbf{z}_{t},a_{l},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$, where $a_{l}$ is drawn from $\mathbf{x}_{t}$
-8:  Observe follower type $f_{t}$
-9:end
-Combining our MSSOA algorithm with the above results and tools from multiclass prediction [22], we get the following regret bound.
-Theorem B.9.
-Consider a Stackelberg game, $\pazocal{G}$, and a hypothesis class, $\pazocal{H}$.
-In the agnostic setting, there exists an algorithm that achieves regret $R(T)\leq\sqrt{\frac{1}{2}\textnormal{MSLdim}_{\pazocal{G}}(\mathscr{H})T\log{(TK)}}$.
-Proof.
-First consider the case where we construct an expert for every $h\in\mathscr{H}$ and keep track of their cumulative loss $L_{t,h}$ at each timestep $t$.
-In our case, for each $h\in\pazocal{H}$, we can construct an expert $i$ that will predict the type, $f_{t}^{(i)}=h(\mathbf{z}_{t})$ at that timestep and based on that evaluate
-$\phi_{t}^{(i)}=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u(\mathbf{z}_{t},a_{l},b_{f_{t}^{(i)}}(\mathbf{z}_{t},\mathbf{x})).$
-From the regret analysis of the Hedge algorithm (see [19] Chapter 2), we know that
-$\sum_{t\in[T]}\mathbb{E}\left[u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))\right]\geq\max_{i\in[N]}\sum_{t=1}^{T}u(\mathbf{z}_{t},\phi_{t}^{(i)},b_{f_{t}}(\mathbf{z}_{t},\phi_{t}^{i}))-\sqrt{\frac{1}{2}\ln(|\mathscr{H}|)T},$
-where the expectation is taken over the randomness in the Hedge algorithm.
-Now it remains to construct a smaller set of experts such that the upper bound on the regret depends on the $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$, instead of $\ln(|\pazocal{H}|)$.
-For some time horizon $T$, we define $S_{T}=\{S\subset[T]:|S|\leq\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})\}$.
-For each $S\in S_{T}$ and $g:S\rightarrow\Delta(\pazocal{A})$, we construct an expert, $E_{S,g}$.
-Expert $E_{S,g}$ immitates an instance of the MSSOA algorithm, where MSSOA errs at all timesteps $t\in S$ and for each timestep $t\in S$, the optimal strategy for the leader at the timestep is given by $g(t)$.
-We show the implementation of each expert in Algorithm˜7 below.
-Algorithm 7 Implementation of expert $E_{S,g}$
-1:input: time horizon $T$
-2:Set $V_{0}=\pazocal{H}$
-3:for $t\in[T]$ do
-4:  Observe $\mathbf{z}_{t}$
-5:  Let $\mathbf{x}^{(h,t)}=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{h(\mathbf{z}_{t})}(\mathbf{z}_{t},\mathbf{x}))\quad\forall h\in\pazocal{H}$
-6:  Let $X_{t}=\{\mathbf{x}^{(h)}:h\in V_{t-1}\}$triangleright $|X_{t}|\leq K$
-7:  Let $V_{t}^{(\mathbf{x})}=\{h\in V_{t-1}:\mathbf{x}^{(h,t)}=\mathbf{x}\}\quad\forall\mathbf{x}\in X_{t}$
-8:  Play action $l_{t}=\arg\max_{\mathbf{x}\in X_{t}}\textnormal{MSLdim}_{\pazocal{G}}(V_{t}^{(\mathbf{x})})$
-9:  Set $l_{t}=\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}\textnormal{MSLdim}_{\pazocal{G}}(V_{t}^{\mathbf{z}_{t}\rightarrow k})$
-10:  if $t\in S$ then
-11:   Predict $g(t)$
-12:   $V_{t+1}\leftarrow V_{t}^{\mathbf{z}_{t}\rightarrow g(t)}$
-13:  else
-14:   Predict $l_{t}$
-15:   $V_{t+1}\leftarrow V_{t}^{\mathbf{z}_{t}\rightarrow l_{t}}$
-16:end
-Notice that for each possible timestep $t\in T$, a single context $\mathbf{z}_{t}$ is presented, so there are at most $K$ possible optimal strategies for the leader, one for each follower type.
-So, the number of experts that can be constructed in this way is $\sum_{i=0}^{\textnormal{MSLdim}_{\pazocal{G}}(\mathscr{H})}\binom{T}{i}(K-1)^{i}\leq(TK)^{\textnormal{MSLdim}_{\pazocal{G}}(\mathscr{H})}$.
-So using the above regret upper bound with this set of experts we get that Algorithm˜6 achieves utility at least $\max_{S,g}-L_{S,g,T}-\sqrt{\frac{1}{2}\textnormal{MSLdim}_{\pazocal{G}}(\mathscr{H})T\ln(TK)}$, where $-L_{S,g,T}$ is the accumulated utility of expert $E_{S,g}$ up to timestep $T$.
-It remains to show that $\max_{S,g}-L_{S,g,T}\geq\max_{h\in\mathscr{H}}-L_{h,T}$.
-Notice that for any $h\in\mathscr{H}$ and for any set of timesteps $S$ that MSSOA errs in, there exists an expert in the constructed set of experts such that $L_{S,g,T}=L_{h,T}$ when executed on the sequence $(\mathbf{z}_{1},h(\mathbf{z}_{1})),\cdots,(\mathbf{z}_{T},h(\mathbf{z}_{T}))$.
-We use a similar approach to Lemma 14 from [15] to get a (nearly-)matching lower bound.
-Theorem B.10.
-Let $\pazocal{H}$ be any hypothesis class.
-There exists a Stackelberg game, $\pazocal{G}$, such that $\pazocal{H}$ has a finite MSL dimension $\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$ and such that no algorithm in the agnostic, full-feedback setting can achieve regret smaller than $\Omega\left(\sqrt{\frac{1}{8}\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})T}\right)$.
-Proof.
-We let $d=\textnormal{MSLdim}_{\pazocal{G}}(\pazocal{H})$.
-For simplicity assume that $T$ is an integer multiple of $d$, say $T=kd$ for $k\in\mathbb{N}$.
-Consider a tree, $(\mathbf{v}_{1},\mathbf{v}_{1},\ldots,\mathbf{v}_{2^{d}-1})$, of depth, $d$, that is SL-shattered by $\pazocal{H}$ such that each node $\mathbf{v}_{i}\in\pazocal{Z}$ has two edges (each leading to a child) corresponding to two strategies for the leader $\mathbf{x}_{i,1},\mathbf{x}_{i,2}\in\Delta(\pazocal{A})$, such that $\mathbf{x}_{i,1}\neq\mathbf{x}_{i,2}$.
-By definition, we have that for each $i$, there exist at least two functions $h_{i,1},h_{i,2}\in\pazocal{H}$ such that at context $\mathbf{v}_{i}$, they predict follower types $f_{1}=h_{i,1}(\mathbf{v}_{i})$ and $f_{2}=h_{i,2}(\mathbf{v}_{i})$, against which the strategies $\mathbf{x}_{i,1}$ and respectively $\mathbf{x}_{i,2}$ are optimal.
-In other words, we construct a Stackelberg game, playing the optimal strategy against the mispredicted follower type gives utility 0, whereas playing the optimal strategy against the correct follower type gives utility 111 (for an explicit construction see proof of Theorem˜B.3).
-Then, we have that for every $i$
-$u(\mathbf{v}_{i},\mathbf{x}_{i,1},b_{f_{1}}(\mathbf{v}_{i},\mathbf{x}_{i,1}))-u(\mathbf{v}_{i},\mathbf{x}_{i,2},b_{f_{1}}(\mathbf{v}_{i},\mathbf{x}_{i,2}))=1$
-and similarly,
-$u(\mathbf{v}_{i},\mathbf{x}_{i,2},b_{f_{2}}(\mathbf{v}_{i},\mathbf{x}_{i,2}))-u(\mathbf{v}_{i},\mathbf{x}_{i,1},b_{f_{2}}(\mathbf{v}_{i},\mathbf{x}_{i,1}))=1.$
-We will now construct a sequence $(\mathbf{z}_{1},\mathbf{x}_{1}),(\mathbf{z}_{2},\mathbf{x}_{2}),\ldots,(\mathbf{z}_{T},\mathbf{x}_{T})$, which we will show forces any algorithm to meet the lower bound in the theorem statement.
-From the MSL-shattered tree, we follow a root-to-leaf path, say $(\tilde{\mathbf{z}}_{1},\tilde{\mathbf{x}}_{1}),\ldots,(\tilde{\mathbf{z}}_{d},\tilde{\mathbf{x}}_{d})$.
-For each node $\tilde{\mathbf{z}}_{i}$ on the path, where $i\in[d]$, we construct a set of $k$ pairs of contexts and follower types $\{(\mathbf{z}_{(i-1)k+1}^{a},\mathbf{x}_{(i-1)k+1}^{a}),\ldots,(\mathbf{z}_{ik}^{a},\mathbf{x}_{ik}^{a})\}$ as follows.
-  • We set $\mathbf{z}_{(i-1)k+1}^{a}=\ldots=\mathbf{z}_{ik}^{a}=\tilde{\mathbf{z}}_{i}$.
-  • We draw each $\mathbf{x}_{t}^{a}$ for $t\in\{(i-1)k+1,\ldots,ik\}$, independently and uniformly at random from the two optimal strategies corresponding to the edges leading to the two children of $\mathbf{z}_{i}$.
-We denote the follower type and optimal strategy for each child, $f_{A,i},\mathbf{x}_{A,i}$ and $f_{B,i},\mathbf{x}_{B,i}$ respectively.
-For each set of $k$ examples, let $T_{i}=\{(i-1)k+1,\ldots,ik\}$ be the set of indeces of the examples corresponding to $\tilde{\mathbf{z}}_{i}$. It suffices to lower bound:
-$\max_{h\in\pazocal{H}}\sum_{t\in T_{i}}u(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$
-For ease of notation, we denote $h^{*}=\arg\max_{h\in\pazocal{H}}\sum_{t\in T_{i}}u(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h}(\mathbf{z}_{t})))$, $\;u_{h^{*},t}=u(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{h^{*}}(\mathbf{z}_{t})))$, $u_{t}=u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$, and $u_{t}^{*}=u(\mathbf{z}_{t},\pi_{f_{t}}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi_{f_{t}}(\mathbf{z}_{t})))$, where $\pi_{f_{t}}(\mathbf{z}_{t})$ corresponds to the optimal policy against the true follower type and context at time $t$.
-So equivalently, we want to lower bound
-$\sum_{t\in T_{i}}u_{h^{*},t}-u_{t}=\sum_{t\in T_{i}}(u_{t}^{*}-u_{t})-(u_{t}^{*}-u_{h^{*},t}).$
-Let $r_{i}$ be the number of timesteps in $T_{i}$, where $f_{A,i}$ appears in the example. Then
-$u_{t}^{*}-u_{h^{*},t}=\begin{cases}k-r\quad\text{if}\;r\geq k/2\\ r\hskip 28.45274pt\text{if}\;r<k/2\end{cases}$
-Therefore, we get that $k/2-\sum_{t\in T_{i}}u_{t}^{*}-u_{h^{*},t}=|r-k/2|$. By taking an expectation over the randomness in the $f_{t}$’s, we get that
-$k/2-\mathbb{E}\left[\sum_{t\in T_{i}}u_{t}^{*}-u_{h^{*},t}\right]=\mathbb{E}\left[|r-k/2|\right]\geq\sqrt{k/8},$
-where the last inequality holds by Khinchine’s inequality (see Lemma A.9 in [19]). Now it suffices to sum over all $i\in[d]$ to get
-$\frac{dk}{2}-\mathbb{E}\left[\sum_{t\in[T]}u_{t}^{*}-u_{h^{*},t}\right]=\sum_{i\in[d]}\left(\frac{k}{2}-\mathbb{E}\left[\sum_{t\in T_{i}}u_{t}^{*}-u_{h^{*},t}\right]\right)\geq d\sqrt{k/8}$
-Finally, notice that $dk/2=T/2=\sum_{t\in[T]}(u_{t}^{*}-u_{t})$ for any algorithm, which concludes the proof. ∎
-B.2 Distributional Learning
-We now give similar results in the distributional setting, in which the leader has access to a batch of training data. We start with some key definitions.
-Consider a distribution $\mathscr{D}$ over $\pazocal{Z}\times[K]$. We define the leader’s payoff as a function of a policy $\pi:\pazocal{Z}\to\Delta(\pazocal{A})$ with respect to $\mathscr{D}$ as
-$u_{\mathscr{D}}(\pi)=\mathop{\mathbb{E}}_{(\mathbf{z},f^{(i)})\sim\mathscr{D}}\left[u(\mathbf{z},\pi(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi(\mathbf{z})))\right]$
-and the leader’s average payoff on a sample $S_{m}=\{(\mathbf{z}_{1},f_{1}),\ldots,(\mathbf{z}_{m},f_{m})\}\in(\pazocal{Z}\times[K])^{m}$ as
-$u_{S_{m}}(\pi)=\frac{1}{m}\sum_{i=1}^{m}\left[u(\mathbf{z}_{i},\pi(\mathbf{z}_{i}),b_{f_{i}}(\mathbf{z}_{i},\pi(\mathbf{z}_{i})))\right].$
-The goal of the learner is to deploy an algorithm, $\textsc{Alg}:\cup_{m=0}^{\infty}(\pazocal{Z}\times[K])^{m}\to\Delta(A)^{\pazocal{Z}}$, such that given a training sequence $S_{m}=\{(\mathbf{z}_{1},f_{1}),\cdots,(\mathbf{z}_{m},f_{m})\}$, the algorithm outputs a policy $\pi_{\textsc{Alg}}:\pazocal{Z}\to\Delta(A)$ that maximizes $u_{\mathscr{D}}(\pi)$.
-Suppose that the leader has access to a set of hypotheses, $\pazocal{H}$, such that each hypothesis in $\pazocal{H}$, maps contexts to follower types ($h:\pazocal{Z}\to[K]$).
-We define the true error function of a hypothesis $h$ with respect to a distribution, $\mathscr{D}$, as
-$\text{Err}_{\mathscr{D}}(h)=\mathbb{P}_{(\mathbf{z},f)\sim\mathscr{D}}\left[h(\mathbf{z})\neq f\right].$
-Similarly, we define the empirical error of a hypothesis $h$ on a sample $S_{m}$ as
-$\text{Err}(h)=\frac{1}{m}|\{i\in[m]:h(\mathbf{z}_{i})\neq f_{i}\}|$
-and the approximation error of a hypothesis class $\pazocal{H}$ on a distribution $\mathscr{D}$ to be
-$\text{Err}_{\mathscr{D}}(\pazocal{H}):=\inf_{h\in\pazocal{H}}\text{Err}_{\mathscr{D}}(h).$
-Remark B.11.
-For the remainder of this section, we will assume that for any context, $\mathbf{z}$, and follower type $f^{(i)}$ there exists a fixed and known tie-breaking rule over the leader’s optimal strategy against $f^{(i)}$ given $\mathbf{z}$.
-Consider an optimal hypothesis, $h^{*}\in\pazocal{H}$, such that given any context, $h^{*}$ always predicts the follower type correctly, i.e.
-$\text{Err}_{\mathscr{D}}(h^{*})=0$.
-Since we are in the realizable setting, for the remainder of this section we will overload notation and use $\mathscr{D}$ to refer to the marginal distribution over the context space $\pazocal{Z}$.
-We now define the Natarajan dimension, a measure of complexity of a multiclass hypothesis class (introduced by  Natarajan [31]), which will be useful in our setting.
-Informally, the Natarajan dimension measures the ability of a class of functions, whose image is a finite set, to fit complex patterns.
-[^6] Note that for binary labels, the definition of Natarajan dimension recovers the well-studied VC-dimension of a hypothesis class.
-Definition B.12 (Natarajan Dimension).
-Let $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$ be a hypothesis class and let $S\subseteq\pazocal{Z}$.
-We say that $\pazocal{H}$ N-shatters $S$ if there exist $\phi_{1},\phi_{2}:S\to[K]$ such that $\forall\mathbf{z}\in\pazocal{Z}\;\;\phi_{1}(\mathbf{z})\neq\phi_{2}(\mathbf{z})$ and for any $T\subseteq S$ there exists a $g\in\pazocal{H}$ such that
-$\forall\mathbf{z}\in T,\;g(\mathbf{z})=\phi_{1}(\mathbf{z})\quad\text{and}\quad\forall\mathbf{z}\in S\setminus T,\;g(\mathbf{z})=\phi_{2}(\mathbf{z}).$
-The Natarajan dimension of $\pazocal{H}$, denoted $d_{N}(\pazocal{H})$, is the maximal cardinality of a set that is shattered by $\pazocal{H}$.
-We define the realizable sample complexity of a Stackelberg-learning algorithm Alg for a hypothesis class $\pazocal{H}$ to be the function $m_{\textsc{Alg},\pazocal{H}}^{r}$ such that for every $\epsilon,\delta>0$, $m_{\textsc{Alg},\pazocal{H}}^{r}$ is the minimal integer such that for every $m\geq m_{\textsc{Alg},\pazocal{H}}^{r}$ and every distribution $\mathscr{D}$ over $\pazocal{Z}$, we have
-$\mathbb{P}_{S_{m}}\left(u_{\mathscr{D}}(\textsc{Alg}(S_{m}))<u_{\mathscr{D}}(\pi_{h^{*}})-\epsilon\right)\leq\delta.$ (2)
-where $\pi_{h^{*}}(\mathbf{z}):=\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h^{*}(\mathbf{z})}(\mathbf{z},\mathbf{x}))$.
-B.2.1 Vanilla ERM learner
-We define an Empirical Risk Minimization (ERM) learner for the hypothesis class $\pazocal{H}$ as an algorithm, such that for any sample $S_{m}$ returns a function that minimizes the empirical error of that function relative to any other function in $\pazocal{H}$.
-Formally, we say
-$h_{ERM}$ $\in\arg\min_{h\in\pazocal{H}}\frac{1}{m}\sum_{i=1}^{m}\mathbbm{1}\{h(\mathbf{z}_{i})\neq f_{i}\}.$ (3)
-We now show that the finiteness of the Natarajan dimension is a sufficient condition for learning the optimal policy for the leader in structured Stackelberg games.
-This can be achieved by using a simple ERM learner as in (3).
-Proposition B.13.
-Consider an algorithm $\textsc{Alg}_{\text{ERM}}:\cup_{n=0}^{\infty}(\pazocal{Z}\times[K])^{n}\to\Delta(\pazocal{A})^{\pazocal{Z}}$ that outputs the policy
-$\pi_{\text{ERM}}(\mathbf{z})$ $\in\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h_{ERM}(\mathbf{z})}(\mathbf{z},\mathbf{x})),$
-where $h_{ERM}$ satisfies Equation˜3. Then for any such $h_{\text{ERM}}$, there exists constant $C>0$, such that for every $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$,
-$m^{r}_{\pazocal{A}_{\text{ERM}}}\leq C\left(\frac{d_{N}(\pazocal{H})\ln(K)\ln(\frac{1}{\epsilon})+\ln(\frac{1}{\delta})}{\epsilon}\right)$
-Proof.
-We have that
-$\mathbb{P}_{S_{m}}\left[u_{\mathscr{D}}(\pi_{h^{*}})-u_{\mathscr{D}}(\pi_{\text{ERM}})\geq\epsilon\right]$
-$\leq$ $\mathbb{P}_{S_{m}}\left[\mathop{\mathbb{E}}_{\mathbf{z}\sim\mathscr{D}}\left[u(\mathbf{z},\pi_{h^{*}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{h^{*}}(\mathbf{z})))-u(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z})))\right]>\epsilon\right]$
-By the boundedness of the utility function we have that the expectation above lies within [1,1][-1,1][ - 1 , 1 ]. Therefore, we get
-$\leq$ $\mathbb{P}_{S_{m}}\left[\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}\left[u(\mathbf{z},\pi_{h^{*}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{h^{*}}(\mathbf{z})))-u(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z}),b_{f^{(i)}}(\mathbf{z},\pi_{\text{ERM}}(\mathbf{z})))>0\right]>\epsilon\right]$
-But getting less utility than the optimal predictor in the class, implies that our prediction of the follower’s type is incorrect. So,
-$\leq$ $\mathbb{P}_{S_{m}}\left[\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}\left[h_{*}(\mathbf{z})\neq h_{\text{ERM}}(\mathbf{z})\right]>\epsilon\right]$
-But from Theorem 6 in [22], we know that for any $\epsilon,\delta>0$, there exists constant $C_{2}$ such that for any $m\geq C_{2}\left(\frac{d_{N}(\pazocal{H})\ln(K)\ln(\frac{1}{\epsilon})+\ln(\frac{1}{\delta})}{\epsilon}\right)$, we have that for any $h_{\text{ERM}}$ that satisfies Equation˜3
-$\mathbb{P}_{S_{m}}\left(\text{Err}_{\mathscr{D}}(h_{\text{ERM}})>\text{Err}_{\mathscr{D}}(\pazocal{H})+\epsilon\right)\leq\delta$
-Since we are in the realizable setting, we know that $\text{Err}_{\mathscr{D}}(\pazocal{H})=0$.
-By definition, we have that $\text{Err}_{\mathscr{D}}(h_{\text{ERM}})>\epsilon\iff\mathbb{P}_{\mathbf{z}\sim\mathscr{D}}\left[h_{\text{ERM}}(\mathbf{z})\neq h^{*}(\mathbf{z})\right]>\epsilon$, which concludes our proof.
-∎
-B.2.2 Stackelberg ERM learner
-In this section, we show how to improve upon the sample complexity bound shown in Proposition˜B.13, by leveraging the interaction between the Stackelberg game and the hypothesis class.
-Consider the following hypothesis class of possible leader policies
-$\Pi_{\pazocal{H},\pazocal{G}}=\Bigl{\{}\pi_{h}:\pazocal{Z}\to\Delta(\pazocal{A})|h\in\pazocal{H},\pi_{h}(\mathbf{z})\in\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h(\mathbf{z})}(\mathbf{z},\mathbf{x}))\Bigl{\}}$
-and the following variation of an ERM rule.
-Definition B.14 (Stackelberg ERM).
-Consider a hypothesis class $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$ and a Stackelberg game, $\pazocal{G}$.
-We define the Stackelberg ERM learning algorithm, $\textsc{Alg}_{\text{S-ERM}}:\cup_{n=0}^{\infty}(\pazocal{Z}\times[K])^{n}\to\Delta(\pazocal{A})^{\pazocal{Z}}$, such that given a training sequence $S_{m}=\{(\mathbf{z}_{1},f_{1}),\cdots,(\mathbf{z}_{m},f_{m})\}$, outputs the policy $\pi_{\text{S-ERM}}:\pazocal{Z}^{m}\to\Delta(\pazocal{A})$ that satisfies
-$\pi_{\text{S-ERM}}\in\arg\max_{\pi\in\Pi_{\pazocal{H},\pazocal{G}}}\frac{1}{m}\sum_{i=1}^{m}u(\mathbf{z}_{i},\pi(\mathbf{z}_{i}),b_{f_{i}}(\mathbf{z}_{i},\pi(\mathbf{z}_{i})).$ (4)
-We define the following measure of complexity of a hypothesis class with respect to a Stackelberg game.
-Definition B.15 (Stackelberg-Natarajan Dimension).
-Consider a hypothesis class $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$, a Stackelberg game, $\pazocal{G}$, and let $S\subseteq\pazocal{Z}$.
-We say that $\pazocal{H}$ SN-shatters $S$ with respect to game $\pazocal{G}$, if there exist $\phi_{1},\phi_{2}:S\to\Delta(\pazocal{A})$ such that $\forall\mathbf{z}\in\pazocal{Z}\;\phi_{1}(\mathbf{z})\neq\phi_{2}(\mathbf{z})$ and for any $T\subseteq S$ there exists a $g\in\pazocal{H}$ such that
-$\forall\mathbf{z}\in T,\;\pi_{g}(\mathbf{z})=\phi_{1}(\mathbf{z})\quad\text{and}\quad\forall\mathbf{z}\in S\setminus T,\;\pi_{g}(\mathbf{z})=\phi_{2}(\mathbf{z}).$
-The Stackelberg-Natarajan dimension of $\pazocal{H}$ with respect to game $\pazocal{G}$, denoted as $d_{SN,\pazocal{G}}(\pazocal{H})$, is the maximal cardinality of a set that is SN-shattered by $\pazocal{H}$.
-Theorem B.16.
-For any hypothesis class $\pazocal{H}$, $d_{SN,\pazocal{G}}(\pazocal{H})\leq d_{N}(\pazocal{H})$.
-Additionally, there exists a hypothesis class and a Stackelberg game with distinct follower types for which the strict inequality holds.
-Proof.
-(Claim 1.) Consider a hypothesis class, $\pazocal{H}$.
-It suffices to show that for any set $S\subseteq\pazocal{Z}$ SN-shattered by $\pazocal{H}$, $S$ is also N-shattered by $\pazocal{H}$.
-This holds because for any pair of points $\mathbf{z}_{1},\mathbf{z}_{2}\in S$, for which there are two different leader strategies consistent with some $g\in\pazocal{H}$, (and assuming a fixed tie-breaking rule) then it must be that $g(\mathbf{z}_{1})\neq g(\mathbf{z}_{2})$, i.e.
-that function predicts two different follower types for these contexts.
-(Claim 2.) Consider the set $\pazocal{Z}=\{\mathbf{z}_{1},\mathbf{z}_{2}\}$ and the set of follower types $\{f^{(0)},f^{(1)},f^{(2)}\}$.
-Also consider the following hypothesis class $\pazocal{H}=\{h_{i}\}_{i\in[4]}$, where $h_{1}(\mathbf{z}_{1})=h_{4}(\mathbf{z}_{1})=f^{(0)}$, $h_{2}(\mathbf{z}_{1})=h_{3}(\mathbf{z}_{1})=f^{(1)}$, $h_{1}(\mathbf{z}_{2})=h_{3}(\mathbf{z}_{2})=f^{(2)}$, and $h_{2}(\mathbf{z}_{2})=h_{4}(\mathbf{z}_{2})=f^{(1)}$.
-It is easy to verify that for $\phi_{1},\phi_{2}$, where $\phi_{1}(\mathbf{z}_{1})=f^{(0)},\;\phi_{1}(\mathbf{z}_{2})=f^{(1)},\;\phi_{2}(\mathbf{z}_{1})=f^{(1)}$, and $\phi_{2}(\mathbf{z}_{2})=f^{(2)}$, $\pazocal{H}$ shatters the set $\{\mathbf{z}_{1},\mathbf{z}_{2}\}$, therefore, $d_{N}(\pazocal{H})=2$.
-Now consider a Stackelberg game, where for all contexts $\mathbf{z}$, $\pi_{*}^{(0)}(\mathbf{z})\neq\pi_{*}^{(2)}(\mathbf{z})$, while $\pi_{*}^{(1)}(\mathbf{z}_{1})=\pi_{*}^{(0)}(\mathbf{z}_{1})$, and $\pi_{*}^{(1)}(\mathbf{z}_{2})=\pi_{*}^{(2)}(\mathbf{z}_{2})$, where $\pi_{*}^{(i)}:\pazocal{Z}\to\Delta(\pazocal{A})$, denotes the optimal policy of the leader against a follower type $f^{(i)}$.
-Then notice that we always have $\pi_{h^{*}}(\mathbf{z}_{1})=\pi_{*}^{(0)}(\mathbf{z}_{1})$ and $\pi_{h^{*}}(\mathbf{z}_{2})=\pi_{*}^{(2)}(\mathbf{z}_{2})$.
-Therefore, $d_{SN,\pazocal{G}}(\pazocal{H})=0$.
-∎
-As in the online setting, depending on the hypothesis class and the Stackelberg game at hand, the gap between the Stackelberg-Natarajan dimension and the Natarajan dimension can be arbitrarily large.
-We illustrate this via the following example.
-Example B.17.
-Consider the hypothesis class of conjunctions, where each hypothesis is defined over a set of Boolean variables, $\pazocal{Z}=\{z_{1},\cdots,z_{n}\}$, such that $\pazocal{H}_{\text{conj}}=\{h:\{0,1\}^{n}\to\{0,1\}|h(z)=\land_{i\in S}z_{i}\land_{j\in T}z_{j},S,T\subseteq[n],S\cap T=\emptyset\}$.
-Since each hypothesis outputs in binary, this hypothesis class could naturally be considered in a Stackelberg game with two follower types, $f^{(0)},f^{(1)}$.
-The Natarajan dimension in this case reduces to the well-studied notion of the VC-dimension, which is known to be equal to $n$ for $\pazocal{H}_{\text{conj}}$.
-At the same time, the $d_{SN,\pazocal{G}}(\pazocal{H}_{\text{conj}})$ will be upper bounded by the size of the set $\{z\in\{0,1\}^{n}|\pi_{*}^{(0)}(z)\neq\pi_{*}^{(1)}(z)\}$, which could be as small as 111 for two distinct follower types.
-Theorem B.18.
-Consider the Stackelberg ERM policy, $\pi_{\text{S-ERM}}$, as in Definition˜B.15 and let $h_{\text{S-ERM}}$ be the corresponding hypothesis in $\pazocal{H}$.
-Then universally for any such policy, there exists constant $C$, such that for every $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$ and distribution $\mathscr{D}$,
-$m^{r}_{\pazocal{A}_{\text{ERM}}}\leq C\frac{1}{\epsilon}\left(d_{SN}(\pazocal{H})\log\left(\frac{d_{SN}(\pazocal{H})\cdot K}{\epsilon}\right)+\log\left(\frac{1}{\delta}\right)\right)$
-Proof Sketch: The proof follows from a similar argument as that of the sample complexity bounds of multiclass learning [22].
-The only part of the proof that does not apply here is Natarajan’s lemma.
-We prove an appropriate generalization in Lemma˜B.24.
-Proof.
-Note that the following proof follows parts of Balcan [5], which we reproduce for completeness.
-By definition, it suffices to show that for any $\epsilon,\delta\in(0,1)$ if we draw a sample $S$ from $\mathscr{D}$ of size:
-$m>C\cdot\frac{1}{\epsilon}\Bigr{[}d_{SN,\pazocal{G}}(\pazocal{H})\log_{2}\Bigl{(}\frac{d_{SN,\pazocal{G}}(\pazocal{H})\cdot K}{\epsilon}\Bigl{)}+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}\Bigr{]}$
-for some large enough constant $C$, then with probability at least $1-\delta$,
-$\mathbb{P}_{S_{m}}\left(u_{\mathscr{D}}(\pi_{\text{S-ERM}})\geq u_{\mathscr{D}}(\pi_{h^{*}})-\epsilon\right).$ (5)
-In other words, it suffices to prove that all the policies in $\Pi_{\pazocal{H},\pazocal{G}}$ with $u_{\mathscr{D}}(\pi_{h})<u_{\mathscr{D}}(\pi_{h^{*}})-\epsilon$ have low utility in the training sample too, that is, $u_{S}(\pi_{h})<u_{S}(\pi_{h^{*}})$.
-It suffices then to bound the following bad event:
-$B:\exists\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}\text{ with }u_{S}(\pi_{h})=u_{S}(\pi_{h^{*}})\text{ but }u_{D}(\pi_{h})<u_{D}(\pi_{h^{*}})-\epsilon.$
-Let us denote the training sample by $S=\{\mathbf{z}_{1},\mathbf{z}_{2},\ldots,\mathbf{z}_{m}\}$.
-Now suppose $S^{\prime}=\{\mathbf{z}^{\prime}_{1},\mathbf{z}^{\prime}_{2},\ldots,\mathbf{z}^{\prime}_{m}\}$ is another sample drawn i.i.d.
-from $\mathscr{D}$ (a ghost sample).
-We define the event:
-$B^{\prime}:\exists\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}\text{ with }u_{S}(\pi_{h})=u_{S}(\pi_{h^{*}})\text{ but }u_{S^{\prime}}(\pi_{h})<u_{S^{\prime}}(\pi_{h^{*}})-\epsilon/2.$
-Lemma B.19.
-If $m>\frac{8}{\epsilon}$, then $\mathbb{P}[B]\leq 2\mathbb{P}[B^{\prime}]$.
-Proof.
-Suppose that we know that there exists an $\pi_{h}$ that is consistent with $S$ but $L_{\mathscr{D}}(\pi_{h}):=u_{D}(\pi_{h^{*}})-u_{D}(\pi_{h})>\epsilon$.
-Since $S^{\prime}$ is drawn i.i.d.
-from $\mathscr{D}$, $\mathbb{E}[L_{S^{\prime}}(\pi_{h})]\geq\epsilon m$.
-By a Chernoff bound, $\mathbb{P}[L_{S^{\prime}}(\pi_{h})\leq\epsilon m/2]<e^{-m\epsilon/8}\leq\frac{1}{2}$ for $m>\frac{8}{\epsilon}$.
-This implies that:
-$\frac{\mathbb{P}[B^{\prime}]}{\mathbb{P}[B]}\geq\frac{\mathbb{P}[B^{\prime}\cap B]}{\mathbb{P}[B]}=\mathbb{P}[B^{\prime}|B]\geq\frac{1}{2},$
-so $\mathbb{P}[B]\leq 2\mathbb{P}[B^{\prime}]$. ∎
-Thus, it suffices to bound $\mathbb{P}[B^{\prime}]$ (this probability is over the choices of $S$ and $S^{\prime}$). Given two samples $S$ and $S^{\prime}$, consider the random process SwapR:
-  • For $i$ from 111 to $m$, flip a fair coin.
-  • If heads, swap $\mathbf{z}_{i}$ and $\mathbf{z}^{\prime}_{i}$, else do nothing.
-Let $T$ and $T^{\prime}$ denote the resulting collections.
-Lemma B.20.
-Suppose $S$ and $S^{\prime}$ are chosen according to $\mathscr{D}$ and SwapR is applied. Then $T$ and $T^{\prime}$ are identically distributed to $S$ and $S^{\prime}$.
-Now, we define the event:
-$B^{\prime\prime}:\exists\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}\text{ with }L_{T}(\pi_{h})=0\text{ but }L_{T^{\prime}}(\pi_{h})>\epsilon/2.$
-By Lemma˜B.20, $\mathbb{P}[B^{\prime\prime}]=\mathbb{P}[B^{\prime}]$ so, it suffices to upper bound $\mathbb{P}[B^{\prime\prime}]$.
-Notice that the first probability is over the choice of $S$, $S^{\prime}$, and the randomness of SwapR, while the second is over the choice of $S$ and $S^{\prime}$.
-Lemma B.21.
-Fix $h\in\pazocal{H}$. Then $\mathbb{P}[L_{T}(\pi_{h})=0\land L_{T^{\prime}}(\pi_{h})>\epsilon/2|S,S^{\prime}]\leq 2^{-\epsilon m/2}.$
-Proof.
-We consider
-$\pi_{h}(\mathbf{z}_{1}),\pi_{h}(\mathbf{z}_{2}),\ldots,\pi_{h}(\mathbf{z}_{m})$
-$\pi_{h}(\mathbf{z}^{\prime}_{1}),\pi_{h}(\mathbf{z}^{\prime}_{2}),\ldots,\pi_{h}(\mathbf{z}^{\prime}_{m}).$
-Notice the following: (1.) If there is a column, with both predictions wrong, then $L_{T}(\pi_{h})=0$ can never happen, so we are done.
-(2.) Similarly, if more than $(1-\epsilon/2)m$ columns have both predictions correct, $L_{T^{\prime}}(\pi_{h})>\epsilon/2$ cannot happen.
-So, there must be at least $r\geq\epsilon m/2$ columns with one correct and one wrong prediction.
-In this case, SwapR must always ensure that the correct prediction remains on top, which happens with probability at most $2^{-r}\leq 2^{-\epsilon m/2}$.
-∎
-We are now ready to upper bound $\mathbb{P}[B^{\prime\prime}]$.
-Lemma B.22.
-$\mathbb{P}[B^{\prime\prime}]\leq\Pi_{\pazocal{H},\pazocal{G}}[2m]\cdot 2^{-\epsilon m/2}$.
-Proof.
-By definition, we have that
-$\mathbb{P}[B^{\prime\prime}]$ $=\mathop{\mathbb{E}}_{S,S^{\prime}}[\mathbb{P}_{\texttt{SwapR}}[\exists\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}},\;L_{T}(h)=0\land L_{T^{\prime}}(h)\geq m\epsilon/2|S,S^{\prime}]]$
-$=\mathop{\mathbb{E}}_{S,S^{\prime}}[\mathbb{P}_{\texttt{SwapR}}[\exists\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}[S\cup S^{\prime}],\;L_{T}(h)=0\land L_{T^{\prime}}(h)\geq m\epsilon/2|S,S^{\prime}]].$
-Using the union bound, we get that
-$\mathbb{P}[B^{\prime\prime}]$ $\leq\mathop{\mathbb{E}}_{S,S^{\prime}}\Bigr{[}\sum_{\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}[S\cup S^{\prime}]}\mathbb{P}_{\texttt{SwapR}}[\;L_{T}(\pi_{h})=0\land L_{T^{\prime}}(\pi_{h})\geq m\epsilon/2|S,S^{\prime}]\Bigr{]}$
-$\leq\Pi_{\pazocal{H},\pazocal{G}}[2m]2^{-\epsilon m/2}$
-Combining all these, we get that $\mathbb{P}[B]\leq\delta$, whenever $2\Pi_{\pazocal{H},\pazocal{G}}[2m]2^{-\epsilon m/2}\leq\delta$, which proves that
-$m>\frac{2}{\epsilon}\Bigr{[}\log_{2}(2\Pi_{\pazocal{H},\pazocal{G}}[2m])+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}\Bigr{]}$
-It remains to show that $\Pi_{\pazocal{H},\pazocal{G}}\leq d_{SN,\pazocal{G}}(\pazocal{H})\ln(K)$.
-We prove this using Lemma˜B.24, which states that $\Pi_{\pazocal{H},\pazocal{G}}[m]\leq m^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})}$.
-So we get that
-$m$ $>\frac{2}{\epsilon}\Bigr{[}\log_{2}(2\cdot(2m)^{d_{SN,\pazocal{G}}(\pazocal{H})}\cdot K^{2d_{SN,\pazocal{G}}(\pazocal{H})})+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}\Bigr{]}$
-$=\frac{2}{\epsilon}\Bigr{[}1+d_{SN,\pazocal{G}}(\pazocal{H})\log_{2}(m)+d_{SN,\pazocal{G}}(\pazocal{H})+2d_{SN,\pazocal{G}}(\pazocal{H})\log(K)+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}\Bigr{]}$
-Rearranging, we obtain the requirement that
-$m>\frac{2d_{SN,\pazocal{G}}(\pazocal{H})}{\epsilon}\Bigr{[}\log_{2}(m)\Bigr{]}+\frac{2}{\epsilon}\Bigr{[}2d_{SN,\pazocal{G}}(\pazocal{H})\log(K)+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}+d_{SN,\pazocal{G}}(\pazocal{H})+1\Bigr{]}$
-Using Lemma˜B.23, a sufficient condition for the above to hold is that
-$m>\frac{8d_{SN,\pazocal{G}}(\pazocal{H})}{\epsilon}\Bigr{[}\log_{2}\Bigl{(}\frac{4d_{SN,\pazocal{G}}(\pazocal{H})}{\epsilon}\Bigl{)}\Bigr{]}+\frac{4}{\epsilon}\Bigr{[}2d_{SN,\pazocal{G}}(\pazocal{H})\log(K)+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}+d_{SN,\pazocal{G}}(\pazocal{H})+1\Bigr{]}$
-And a sufficient condition for the above is
-$m>\frac{8d_{SN,\pazocal{G}}(\pazocal{H})}{\epsilon}\Bigr{[}\log_{2}\Bigl{(}\frac{4d_{SN,\pazocal{G}}(\pazocal{H})\cdot K}{\epsilon}\Bigl{)}\Bigr{]}+\frac{4}{\epsilon}\Bigr{[}\log\Bigl{(}\frac{1}{\delta}\Bigl{)}+d_{SN,\pazocal{G}}(\pazocal{H})+1\Bigr{]}$
-The above holds if there exists a constant $C$, large enough such that
-$m>C\cdot\frac{1}{\epsilon}\Bigr{[}d_{SN,\pazocal{G}}(\pazocal{H})\log_{2}\Bigl{(}\frac{d_{SN,\pazocal{G}}(\pazocal{H})\cdot K}{\epsilon}\Bigl{)}+\log\Bigl{(}\frac{1}{\delta}\Bigl{)}\Bigr{]}.$
-Lemma B.23 (Shalev-Shwartz and Ben-David [35], Lemma A.2).
-Let $a\geq 1$ and $b>0$. Then $x\geq 4a\log(2a)+2b$ implies that $x\geq a\log(x)+b$.
-Lemma B.24.
-For a given class $\pazocal{H}$ and Stackelberg game $\pazocal{G}$, let $d_{SN,\pazocal{G}}(\pazocal{H})=d$.
-Then for a set of contexts $S=\{\mathbf{z}_{1},\dots,\mathbf{z}_{|S|}\}$, let $\Pi_{\pazocal{H},\pazocal{G}}(S)=\{(\pi(\mathbf{z}_{1}),\dots,\pi(\mathbf{z}_{|S|}))\;\;|\;\;\pi\in\Pi_{\pazocal{H}},\pazocal{G}\}$.
-It holds that
-$|\Pi_{\pazocal{H},\pazocal{G}}(S)|\leq|S|^{d}\cdot K^{2d}.$ (6)
-Proof.
-[…]
-We will now show Equation˜6 for $|S|=m+1$.
-First we fix a set $S=\{\mathbf{z}_{0},\mathbf{z}_{1},\cdots,\mathbf{z}_{m}\}$.
-For simplicity and without loss of generality we write $\Pi_{\pazocal{H},\pazocal{G}}(S)$ as $\Pi_{\pazocal{H},\pazocal{G}}$.
-We now define the following
-$\Lambda^{(i)}=\{\pi_{h}\in\Pi_{\pazocal{H},\pazocal{G}}|\pi_{h}(\mathbf{z}_{0})=\mathbf{x}_{*}^{(i)}\}$
-$\Lambda^{(i,j)}=\{\pi_{h}\in\Lambda^{(i)}|\exists\pi_{h}^{\prime}\in\Lambda^{(j)}\;\text{s.t.}\;\pi_{h}(\mathbf{z})=\pi_{h}^{\prime}(\mathbf{z})\;\forall\mathbf{z}\in\pazocal{Z}\},$
-where $\mathbf{x}_{*}^{(i)}$ is the leader’s optimal strategy against $f^{(i)}$, given context $\mathbf{z}_{0}$. We also define
-$\Pi_{0}=\Pi_{\pazocal{H},\pazocal{G}}\setminus\left(\cup_{i\neq j}\Lambda^{(i,j)}\right).$
-Then it immediately follows that
-$|\Pi_{\pazocal{H},\pazocal{G}}(S)|\leq|\Pi_{0}(S)|+\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}|\Lambda^{(i,j)}(S)|.$ (7)
-By definition we know that each policy in $\Pi_{0}$, produces a distinct set of optimal strategies on the contexts in $S\setminus\{\mathbf{z}_{0}\}$.
-Therefore, $|\Pi_{0}(S)|=|\Pi_{0}(S\setminus\{\mathbf{z}_{0}\})|$.
-Also, $\Pi_{0}\subset\Pi_{\pazocal{H}},\pazocal{G}$ so $\Pi_{0}$ cannot SN-shatter a set of size greater than $d$.
-Therefore, from the inductive hypothesis we get that
-$|\Pi_{0}(S)|=|\Pi_{0}(S\setminus\{\mathbf{z}_{0}\})|\leq m^{d}\cdot K^{2d}.$ (8)
-Next, we argue that each $\Lambda^{(i,j)}$ cannot SN-shatter a set of size greater than $d-1$.
-To see this, we let $L=\{\tilde{\mathbf{z}}_{1},\dots,\tilde{\mathbf{z}}_{|}L|\}$ be a set that $\Lambda^{(i,j)}$ can SN-shatter.
-Then $L$ cannot include $\mathbf{z}_{0}$, since $\pi(\mathbf{z}_{0})=\mathbf{x}_{*}^{(i)}$ for all $\pi\in\Lambda^{(i,j)}$.
-However, $\Pi_{\pazocal{H},\pazocal{G}}$ can SN-shatter $L\cup\{\mathbf{z}_{0}\}$.
-So $|L|\leq d-1$.
-Therefore, by the inductive hypothesis we get
-$|\Lambda^{(i,j)}(S)|\leq m^{d-1}\cdot K^{2(d-1)}$ (9)
-Combining inequalities (7), (8), and (9), we get that
-$|\Pi_{\pazocal{H},\pazocal{G}}(S)|$ $\leq m^{d}\cdot K^{2d}+\sum_{i=1}^{K}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{K}m^{d-1}\cdot K^{2(d-1)}$
-$\leq m^{d}\cdot K^{2d}+m^{d-1}\cdot K^{2d}$
-$=m^{d-1}K^{2d}(m+1)$
-$\leq(m+1)^{d}K^{2d}$
+15:    end if
+16:   else
+17:    Continue
+18:   end if
+19:  end for
+20:end for
+ALGORITHM 3 Multiclass Decision Lists Learning Algorithm